A Complete Bibliography of ACM Transactions on Intelligent Systems and Technology (TIST)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

04 June 2024 Version 1.45

Title word cross-reference

3 [BDP12, JLH19, LLS^{+}22, MCEG23, TDVC13, WGF^{+}23, ZWH^{+}22]. 6 [FDE15]. ^{2} [FGL17, SZL^{+}23]. K [GQY^{+}19, EFMRK^{+}20, PKCC18, THY^{+}11, WLWC23, ZLZ^{+}17]. N [DOTD16, MG16, EK15]. P [SSV19].

-Facility [WLWC23]. -Means [PKCC18]. -Nearest [GQY^{+}19]. -Values [SSV19].

1 [WZFL21, YTL^{+}22a]. 19 [BZX^{+}22, SBS^{+}23].

2 [WZFL22, YTL^{+}22b]. 2018 [ZSLC19]. 2019 [ADM^{+}21]. 2nd [SY12].

Abandonment [POM20]. Abduction [SSS11, SDS12]. abnormality [SSZ^{+}13]. Abstraction [ABB^{+}15, TRH12]. abstractive [CL13]. Abuse [YL17]. Academic [YLC^{+}19]. Accelerated [BC19, STA22]. Accelerators [KG23]. Access [HTSC^{+}17, SCLZ17, YTH17]. Accessible [WLH10]. Accidents [HLF^{+}21]. Accounting [Hec19]. Accuracy [JV20, GJ13]. Accurate [HCJM15, JJ15, LCY^{+}22, TC19]. ACM [ADM^{+}21, AJL18, CKW19, CL15, LN10, WZFL21, Yan10, ZSLC19, ZLB^{+}16, ZWGW17]. ACML [SY12]. ACP [WW13]. Acquisition [JPS^{+}16, BPS13]. across [DCM15, FC15, Min16, TS17, YDZ20, dMFA^{+}13]. Action [DHF22, EHG21, ZCG15, ZHLL21, Zhu19].

Active [BN21a, ERR13, LLLC19, LCC^{+}20, RV18, SZT12, WH11b, YCZY21, YCY23, ZCWZ18]. Activities [ADJ^{+}20, HL19, RYS10, ZNYH11]. Activity [FGL17, HRCT16, KHNB15, LZY^{+}16, PEK^{+}16, QWC^{+}23, SMX15, SS11, YDZ20, HLJ11, RC13, WLG11, ZPY11]. Ad [SBS^{+}23, WWL^{+}22, XWW^{+}21, YLY^{+}23]. Ad-Hoc [SBS^{+}23, YLY^{+}23]. Adam [CSHL21]. Adaptation [BC19, LCN^{+}21, MMS17, PS11, TPM23, WCF^{+}20, WC20, WYNW20]. adaptative [LGZ^{+}21]. Adaptive [AHJB20, AKR^{+}18, BD23, CDLV13, CCZ^{+}15, CYC^{+}23, GKG^{+}11, HS19, HHJ22, LCM^{+}12, LLS^{+}22, LGJ^{+}22, LYKS23, LFWY23, MZY^{+}22, QWC^{+}23, RXK^{+}17, ZAK^{+}23, ZFH^{+}22, OY13]. Additive [QCZ^{+}21, WYC^{+}17]. Addressing [HGE17]. Ads [BB15]. Adult [OLY^{+}17]. Advanced [PMR^{+}17, LWW^{+}23]. Advances [QCL15, ZGL^{+}17, ZZH^{+}22]. Adversarial [BZX^{+}22, DLLT23, DZY^{+}22, GXS^{+}22, LBT23, LXJ^{+}20, QTM23, SC22, SSG^{+}20, SDS12, WLL^{+}21, WGL^{+}22, XHZ^{+}23, YCZ^{+}23, ZSAL20, ZFQX20]. Adverse [SGJC18, YY15]. Advertisement [CWCK15]. Advertising [ABG^{+}11, CMR15, SA15, WYG^{+}22, ZWXZ12]. advisors [ZC13]. AEGIS [EBG^{+}12]. Aerial [CCZ^{+}23]. Affiliation [VNL^{+}11]. Affinity [LCN^{+}16]. Against [CWCY22, CBPG22, LLPS20]. Age [LGZ^{+}17]. Agent [CRRH11, GKG^{+}11, GDC19, RVRJ11, SS11, XSJW23, Zhu19, CABD13, CAB^{+}13, EvdHW13, HXC^{+}23, NOZ20, ZAK^{+}23]. Agent-Based [CRRH11, RVRJ11]. Agents [ZRX^{+}22, OSM^{+}13]. AggEnhance [OSW^{+}22]. Aggregate [MG16]. Aggregation [CVCL22, JJKZ22, ODL^{+}20, OSW^{+}22, ZHW^{+}21]. Aggregation-based [CVCL22]. AggregatoR [KMH22].

Aggressive [LSZH18]. aging [CDS13]. AI [CZKJ22, LN10, LWF^{+}23]. aided [GJS23]. AIOps [NCG21]. Air [ASSR18, GME17, LLL^{+}18, BKB10]. Aircraft [SSLM21]. Airline [ACPS17, GG15]. Airport [HMG^{+}23]. alarm [DL13]. ALERA [BC19]. Alert [SSG^{+}20]. Algorithm [JJ15, MGS17b, SSV19, WCBL18]. Algorithmic [ZZZ^{+}22, ZZC^{+}22]. Algorithmic-Cryptographic [ZZC^{+}22]. Algorithms [AHJB20, BSRSS16, BFC^{+}17, HQY^{+}22, LZW^{+}23, PFS17, RXL^{+}23, SHB^{+}12, YTL^{+}22a, YTL^{+}22b, ZFH^{+}18, GPSB11, OSM^{+}13]. Alignment [FLY^{+}23, LSW23, LWWL11, RSCOVCMM17, SZL^{+}23, ZLZW23, BMV13]. All-Pay [LDTX16]. Allocation [DPC16, HXC^{+}23, NOZ20, RK15, WZY^{+}18, LZCS11]. Alternating [CYYL18]. Alternative [RKH14]. Alternatives [ARGK15]. Ambiguity [JSL^{+}19]. American [ZZZ20a]. Analogical [LCY^{+}18]. Analysis [BTL20, BCGJ11, CKP^{+}22, CWR^{+}16, CCW^{+}19, CDW^{+}19, DWKP16, EMF12, FSS15, FLLX18, GOB^{+}12, HM19, HWCL17, KZL^{+}21, LS16, LGL^{+}16, LCLN18, LZCQ12, LZP^{+}12, LYWH20, MPA13, MMDY15, NAPI14, ODL^{+}20, PT12, PHL^{+}20, PCL18, RHD^{+}12, STP^{+}18, XZH^{+}17, YY15, YL17, YCGH12, ZSLC19, ZDL^{+}12, ZWL^{+}15, ZWL^{+}19, ZWZZ23, GXZ^{+}11, OY13, ZNYH11]. Analysts [GJSC16, GJC17]. Analytics [CKW19, HASS22, JGL^{+}15, JLX^{+}17, NZW^{+}17, PHL^{+}20, RXL^{+}23, XZXM19, ZW19, ZLSY22a, ZLSY22b]. Analyzing [ADJ^{+}20, CCWS17, DJNC21, HTSC^{+}17, KHNB15, SS22, dMFA^{+}13]. animal [LHJ^{+}11]. Animation [YZL^{+}19]. Anisotropic [LGZ^{+}17]. annotated [HY11]. Annotation [CLBM15, HDPH16, HYC^{+}16, KAH^{+}16, LCN^{+}16, TTLG17, TS17, WFJY12, SZC11, THY^{+}11, WH11b, YCP^{+}13]. Anomaly [DC21, HKMN20, KW17, MWS^{+}18, RXL^{+}23, ZPL^{+}20, LWW^{+}23]. Anonymous [WMR17, XZW^{+}15]. Answering [GH18, TPG^{+}19]. Anti [TMZ^{+}20, BKB10]. anti-air [BKB10]. Anti-Spoofing [TMZ^{+}20]. Anticipation [DHF22]. Anytime [BSRSS16]. App [FLLX18, LLZ^{+}23, XLF^{+}20]. Appearance [FSW^{+}20, ZYSL12, LHS^{+}13]. Application [DWKP16, DBDM16, LLLC19, Min16, XZH^{+}17, YFJ^{+}18, ZWZS16, ZZZ^{+}19, CKS10, YJHL11]. Applications [AKR^{+}18, BCGJ11, LCX^{+}23, PFS17, RXL^{+}23, SYHB17, WZS^{+}15, WZYM19, XKW^{+}16, YLCT19, YTL^{+}22a, YTL^{+}22b, YWS^{+}23, ZFH^{+}18, ZCWY14b, Che10, Gin13, KG11, Lin11]. Applying [JCH14, PCC10]. Approach [ASSR18, BTL20, BCR21, CGZ18, DWKP16, DSS^{+}22, DHF22, FDE15, GBC^{+}22, GRR^{+}15, GMX^{+}21, HSJ^{+}22, HL17, HBL16, HWT17, KKG18, KZL^{+}17, KPF18, KSL^{+}15, LYW^{+}19, LYWW18, LRD^{+}22, LDTX16, LYLX23, MJVL16, MRJ16, ODF17, ODP^{+}17, PCF^{+}19, PPPM18, SLC23, WGL^{+}22, WHW^{+}21, XXL^{+}17, XHZ^{+}23, ZYC^{+}22, ZDL^{+}12, ZC15, ZFWL17, ZLZ^{+}22, ZWL^{+}19, BBMP13, BVCH13, BSW^{+}13, BGMS13a, CL13, LHG11, LHC^{+}13, RYS10, TDVC13, WW13]. Approaches [LC15, LCLG19, RZS^{+}15, ZSLC19, PMSR11]. Approximate [PDR23]. approximation [TNSP13]. Apps [LLZ^{+}23, WMR17]. Architecture [AdCK^{+}22, CLL23, HLC^{+}21, JYL^{+}23, ZSY^{+}12]. Architectures [DMGSD23, HTSC^{+}17]. Area [NZS^{+}22, SSZ^{+}13]. Argument [MP23]. Arranging [LH22]. array [RFI^{+}11]. Arrival [BYK^{+}21, WLW^{+}23]. Arrivals [DPC16]. Arriving [CGZ18]. ART [EvdHW13]. Articles [SRB15]. Artificial [CC12]. As-You-Type [LCV17]. Asia [SY12]. Aspect [FTE21, LFW23].

Aspect-Aware [FTE21]. Assess

[PPPM18]. Assessment [HMS^{+}14, MRJ16]. Assignment [TTFS18]. assistance [BGPYS11]. Assisted [SMGMC^{+}15, TRZ^{+}19]. Associating [UAS15]. Association [GLL^{+}17, HCCY15, YY15]. Association-Based [GLL^{+}17, HCCY15]. Asymmetrical [ZZL^{+}23]. Attack [DLLT23, GME17, HNA20, RDX22, WGL^{+}22, WMA20, YCZ^{+}23]. Attacks [CWCY22, CBPG22, GSM23, LLPS20, ODP^{+}17, QTM23, XLG^{+}23, ZSAL20, ZLL^{+}22]. Attention [CMPR21, CDW^{+}21, GWL^{+}23, HLH^{+}21, LFY^{+}22, LFW23, LYWW18, LLX^{+}22, MGJW20, SLZ^{+}23, WJY^{+}18, WLL^{+}21, YFJ^{+}18, YCZ^{+}23, ZZL^{+}23, ZZS^{+}21, ZLG^{+}20, WLL^{+}22, WYNW20, ZCZ^{+}23]. Attention-Based [LYWW18, WJY^{+}18, MGJW20]. Attention-enhanced [GWL^{+}23]. Attention-guided [YCZ^{+}23]. Attentive [CMPR21, ZFQX20, ZHW^{+}21]. Attitude [NAPI14]. Attribute [DLLT21, GTM^{+}14, HYHFV22, XLB23]. Attributes [ASK^{+}21, NZW^{+}17, WLH17]. Auction [LDTX16]. Auctions [QCL15]. Audio [DCM15, RSCOVCMM17, SHZ13, WCBK11]. Audio-to-Score [RSCOVCMM17]. augment [ZDC^{+}13]. Augmentation [KCSW23, ZB20]. Augmented [CSN^{+}17, HQW22, ZZD^{+}17]. Authentication [ZYH^{+}20]. Authoring [LMAP16, WFJY12]. Authorities [BR15]. Authority [CCWS17]. Auto [DB16, EFMRK^{+}20, LNYV22, LCY^{+}15, SZZ^{+}21]. Auto-Encoders [SZZ^{+}21]. Auto-Experimentation [DB16]. Auto-Generated [LCY^{+}15]. Auto-tuned [EFMRK^{+}20]. Auto-weighted [LNYV22]. Autoencoder [HQW22, YWZ^{+}23, ZZL^{+}23, ZWC23, ZCL^{+}18]. AutoLCA [HMS^{+}14].

Automated [DGJ^{+}23, DB16, EBG^{+}12, JD15, SDHS15, Che10, WHJ^{+}11]. Automatic [CLBM15, KW17, LGZ^{+}17, LCLN18, Pac17, TLWZ11, ZSLC19, ZZZ^{+}19, TDVC13]. Automotive [WM21, ZWL^{+}19]. Autonomous [CCZ^{+}23, CYC^{+}21, JLJ^{+}20, KL23, ZZZ^{+}22]. Autonomy [BCR21]. Auxiliary [CWC^{+}20, CASR22, VNL^{+}11, XYS^{+}23, ZBZX12, PG13]. Availability [RYC22]. Average [CDGZ16, SDXG16]. Aware [AC15, CLH^{+}22, DCF^{+}18, FXHM16, FTE21, GXYZ21, GMX^{+}21, HXY^{+}22, JYT^{+}12, LCV17, LJC^{+}11, MMPS23, PEK^{+}16, TL23, XLF^{+}20, YZL^{+}19, YTH17, ZSL^{+}15, ZFWL17, ZCX^{+}15, BGMS13a, CKS10, DSS^{+}22, FLY^{+}23, HLL^{+}23, LHZ22a, LHZ13, SLH13, TC19, WFX^{+}21, WZM^{+}22, WPL13, XXZ^{+}21, XLZ^{+}22, YZEC13, ZFQX20, ZLD^{+}23, ZMH^{+}22, ZLL^{+}22, ZLG^{+}20, ZRX^{+}22, ZXY^{+}23]. Awareness [CCZ^{+}23, DGK^{+}22, ZPP^{+}21].

Backdoor [CBPG22]. Background [CCH15]. Badminton [WCP^{+}23]. Bagging [HYHFV22]. Bagging-Based [HYHFV22]. Balanced [LYW^{+}19, SJCM23]. Balancing [HS19, WLWC23]. Ballet [KSL^{+}15]. BAMB [LYW^{+}19]. Bargaining [HDTG15, KAH12]. Base [JV20]. Based [ACPS17, ABO17, ASSR18, BBM17, BDP12, Bha21, BN21b, CGZ18, CUG^{+}12, CZP^{+}14, CLH^{+}22, CYKL16, CRRH11, CDW^{+}19, DJI^{+}16, DOTD16, DTL15, EK15, EL14, FDE15, FZX15, GLL^{+}17, GRR^{+}15, GZ21, HCCY15, HDPH16, HMG^{+}23, HYHFV22, HKMN20, HCRF21, HLL14, HMCW15, HTL^{+}20, JJ14, JJ15, JGL^{+}15, JHK^{+}22, KW17, KPF18, LCD17, LHS18, LCD18, LLZ^{+}23, LBP19, LH22, LZP^{+}12, LTS^{+}15, LYWW18, LLPS20, LWWX20, MFI19, MRJ16, NYBG17, PMR^{+}17, Pat15, QHH^{+}21, RHD^{+}12, SKF^{+}14, ST19, SZZ^{+}21, SRJP12, TJL^{+}21, TL23, WSGM14, WJY^{+}18, WMR17, XLG^{+}23, XTW17, YZQ16, YCH^{+}22, YMC16, YTH17, YCZY21, ZLBZ23, ZYSL12, ZWH17, ZL19, ZZZ^{+}22, ZZZ^{+}11, ZCG15, ZZKT20, ADM^{+}21, AKA^{+}21, BYK^{+}21, BZW^{+}22, CFG13, CDS13, CVCL22, GJ13, GQY^{+}19, GJS23, HQY^{+}22, HSJ^{+}22, HM19, HG21, HLT11, HHL^{+}22, HKO13, JJKZ22, JZG22, LYZ^{+}23, LSQ11, LCC^{+}20, LAS20, MGJW20, MLJZ21, ME13, NOZ20]. based [QTM23, RVRJ11, RC13, RYS10, RP23, SLZ^{+}23, SGTK20, SGD13, TDVC13, THY^{+}11, TAL^{+}19, TNSP13, WCS^{+}20, WM21, XJS^{+}21, YCY23, YJHL11, ZYC^{+}22, ZAK^{+}23, ZC15, ZFLD22, ZCZ^{+}23, LWW^{+}23, MLSK23, LHC^{+}13]. Baseline [WTK^{+}19]. basis [Sin13]. Basket [XYS^{+}23]. Basketball [CTC^{+}22]. Batch [GPSB11, RV18, SZT12]. Batch-Mode [RV18]. BATS [WHW^{+}21]. Bayes [LAS20]. Bayesian [AHJB20, HRCT16, HYHFV22, JWJC16, KPF18, SLR^{+}16, VKA^{+}19, XLZ21]. BayesPiles [VKA^{+}19]. Be [LLL23]. Beats [WYM17]. Behavior [AC15, DSB^{+}18, HDPH16, HL17, LAsO^{+}19, LZ18, LWH^{+}20, LLZ^{+}23, NYBG17, NZS^{+}22, ODP^{+}17, PYD^{+}17, SCC^{+}23, VPD^{+}22, WZCJ21, WYP22, WWL^{+}22, YZQ16, YLC^{+}19, YKTL14, ZYW^{+}15, ZLH18, OY13, ZC13, dMFA^{+}13]. Behavior-Based [HDPH16]. Behavioral [LCLN18, YNS13]. behavioral-cultural [YNS13]. Behaviors [CHP17, WLW^{+}22, WLW^{+}23]. Behaviour [GDC19]. belief [TDVC13]. Ben [BI17]. Ben-Israel [BI17]. Benchmark [HLL^{+}22, VDL^{+}19]. Benchmarking [CSTZ16, DJS16, DBDM16]. Benefits [CCL15, WLZ^{+}23]. Berkeley [FLF^{+}20]. Bernstein [ZS18]. Better [AT15, LJLZ19, MBM21]. between [CRYT12, GME17, LZC23, ODP^{+}17, RKH14]. Beyond [BBS^{+}16]. Bézier [ZS18].

Bi [CHY15]. Bi-Histogram [CHY15]. Bias [YGY^{+}23, ZJSY21, BNS13]. Biased [RK15]. Bibliographic [WXZ^{+}16]. Biclustering [WHW^{+}21]. Bicycle [CLL^{+}21]. Bicycle-sharing [CLL^{+}21]. Bid [BB15]. bids [ZC13]. Big [JLX^{+}17, JCW^{+}22]. Bike [EL14, QHH^{+}21]. Bilingual [ZHW^{+}21]. Billboard [WYG^{+}22]. Binarization [MFI19]. Binarized [GWDJ15]. Binary [SDXG16]. BiNeTClus [BN21b]. Biobehavioral [YLD^{+}22]. biological [BVK10]. biomass [CGMC11]. Biometric [ZYH^{+}20]. Bipartite [BN21b, UAS15]. BISTRO [FLF^{+}20]. Black [CGZ23, DLLT23, XLG^{+}23]. Black-Box [CGZ23, DLLT23, XLG^{+}23]. Blackmarket [ADJ^{+}20]. Blanket [LYW^{+}19]. Blau [JCH14, JCH14]. Blockmodel [TY14]. board [BCR21]. Bookmarking [DJS16]. Bookmarks [TRDD12]. Boosted [MFI19]. boosting [DSM^{+}11]. Bootstrapping [BL16, HB12, HLNL18]. Both [CGZ18, Zhu19]. Bottlenecks [LBC^{+}22]. Bound [ODF17]. Bounded [LSZH18]. Bounds [LG16, SDXG16]. Box [CGZ23, WGL^{+}22, DLLT23, BRSG20, XLG^{+}23]. BOXREC [BRSG20]. Bracelets [YLS15]. Brainwaves [ZYH^{+}20]. branch [WLL^{+}22]. Breast [TWC^{+}23]. Breathing [WYM17]. Bridges [DSB^{+}18]. Bridging [GME17]. Bring [HCWH22]. Broadband [LZY^{+}16]. broadcast [TLWZ11]. browse [JPL13]. Browsing [AC15, NYBG17, TRH12]. Budget [NVDMFD22, ZSL^{+}15]. Building [EvdHW13]. Buildings [GRR^{+}15]. Bulk [ZZC^{+}20]. Buses [WFX^{+}21]. Business [BCGJ11, PSRL12, TCK20, VDL^{+}19, Lin11, MMC^{+}13]. Buying [BB15, ZC13].

Cab [VPD^{+}22]. CAFE [ZFH^{+}22]. calculus [MMC^{+}13]. calendaring [BGPYS11]. Calibrated [LC15]. Camera [LXM^{+}18, TTL^{+}21, ZLT15]. Cameras [MZL12]. Campaign [LCCS13]. Campus [YLC^{+}19]. Campuses [WMR17]. Can [HCWH22, SSG^{+}20]. Cancer [RP23, TWC^{+}23]. Candidate [RKH14, RK15]. Candidates [CCH15]. Capsule [CCGP22]. Capture [LCKY14]. CapVis [LJLZ19]. Car [NTM^{+}16, NZS^{+}22]. Car-hailing [NZS^{+}22]. cardiopulmonary [BCC^{+}13]. Carefulness [FXR^{+}17]. Caregivers [CCL15]. Carved [LC15]. Cascade [QCZ^{+}21]. Cascaded [YLY^{+}23]. Case [EL14, KAH^{+}16, MOC^{+}11, PHL^{+}20, Dha11, EvdHW13]. Cash [ZZZ^{+}19]. Cash-Out [ZZZ^{+}19]. Categorical [CYC^{+}23]. Categories [SSLM21, FPVC13]. Categorization [GGC21, TDVC13]. Category [SP16]. Causal [CDGZ16, FNS16, HM19, HBL16, KBM^{+}21, LLL^{+}16, LCN^{+}21, LLX^{+}20, LC16, LCLG19, QCZ^{+}21, SDXG16, WC12, YGY^{+}23, ZWZS16, ZLB^{+}16, ZZGH19]. Causality [FZ16, ZNYH11]. Cause [Hec19, LCLG19]. Causes [Hec19, WYY^{+}23]. CAVE [KSL^{+}15]. CDSM [YLY^{+}23]. Center [FZH^{+}21, SLZ^{+}23, ZZL^{+}19]. Center-dark [ZZL^{+}19]. Centered [GBC^{+}22]. Centers [HWCL17, PMSR11]. Centric [KKG18, ZZZ^{+}11, CCW^{+}19]. Chains [LAsO^{+}19]. Challenge [ADM^{+}21, ZSLC19]. Challenges [HGE17, RXL^{+}23, XWW^{+}21, ZFH^{+}18]. Change [WPA^{+}12]. Changeable [TBW21]. changes [SKOM13]. Channel [ZZL^{+}19, DPB20]. Characterization [ABB^{+}15, HLL^{+}22]. Characterizing [YFJ^{+}18]. Charging [MBR^{+}14, WZS^{+}20, WFX^{+}21]. chatting [HLJ11]. Cheating [LSW^{+}20]. Check [HWT17, JCH14, LX14, SMX15, SCC^{+}23, YKTL14]. Check-In [HWT17, LX14, SMX15, YKTL14, SCC^{+}23]. Check-ins [JCH14]. Children [YLS15]. chiller [PMSR11]. Chinese [ZDC^{+}13]. Choice [HTSC^{+}17, LXJ^{+}20, OY13]. Choosing [RKH14, RK15]. CIM [CZP^{+}14].

Circular [LH22]. Citation [KSKÇ15]. Cities [ABO17]. Citizen [YMLM16]. Citizen-Sourcing [YMLM16]. City [ABTS15, JCH14, MFLP14, YMLM16, ZQP^{+}15, ZLH18, LGZ^{+}21, YMLM16]. City-Scale [MFLP14, ZQP^{+}15]. Citywide [JCW^{+}22]. Class [AKR^{+}18, BLAK19, JSL^{+}19, OSW^{+}22, WHC13]. Classes [YCL^{+}21]. Classi [SSLM21]. Classi-Fly [SSLM21]. Classification [BTTT19, BLAK19, CZG^{+}23, CLL23, DTL15, GZZY17, HYHFV22, HWL^{+}17, KCS18, KMH22, LPL^{+}22, LCY^{+}18, MRW^{+}12, MG16, RAK23, SLM^{+}21, SKF^{+}14, TWC^{+}23, WMH18, XHZ^{+}23, YCL^{+}21, YCY23, ZLZ^{+}17, GPSB11, SHZ13, WHC13]. Classifier [GQY^{+}19, HLT11]. Classifiers [JV20, TCK20]. classifying [BVCH13]. CLC [ZYC^{+}22]. Cleaning [FWYX22]. Clearing [NOZ20]. Clearing-based [NOZ20]. CLEaVER [LNO^{+}18]. Click [AC15, CWCK15]. clinical [LKD13]. Clinically [DC21]. Clique [GGY^{+}23]. Closer [Wid17]. Cloud [CUG^{+}12, CCGP22, HTM15, JPS^{+}16, XLZ^{+}22, ZBW^{+}22]. Cloud-Edge [XLZ^{+}22]. Cluster [JJ14, LAsO^{+}19, ZXY^{+}23]. Cluster-aware [ZXY^{+}23]. Cluster-Based [JJ14]. Clustering [AWSF21, BN21b, EL14, HQY^{+}22, HCTC12, KLLL20, KPF18, LYLX23, TRDD12, WYY^{+}23, XJS^{+}21, YCY23, ZL12, ZWZ^{+}19, LMWS13]. Clustering-based [YCY23]. Clusters [Pat15]. CNN [SGTK20, WJY^{+}18]. CNN-based [SGTK20]. CNNs [ST19]. Co [HSJ^{+}22, KCS18, KLLL20, LLF^{+}19, ZFH^{+}18, ZZC^{+}22]. Co-Clustering [KLLL20]. Co-design [ZZC^{+}22]. Co-Saliency [ZFH^{+}18, LLF^{+}19]. Co-Training [HSJ^{+}22, KCS18]. coaching [LLWC13]. Coalition [BFC^{+}17, MFB^{+}20]. Coalitions [PMR^{+}17, FT10]. Coarse [OLY^{+}17]. CoClustering [HMCW15]. Code [RBG22, SJCM23]. Coding [FWZ17, HG21, ZYSL12]. CoFi [LPM20]. CoFi-points [LPM20]. Cognitive [LWC^{+}18]. Cold [LHS18]. Cold-Start [LHS18]. Collaboration [MOC^{+}11, SRB15]. Collaborative [CSN^{+}17, FS17, JJ14, LPM20, LFG^{+}23, ST20, TB22, WSGM14, ZWZZ23, ZHLL21, BCD^{+}13, ERR13, LHZ13]. Collection [HB12, YJHL11]. Collection-based [YJHL11]. Collections [TRH12]. Collective [ACC21, GST12, WFZ^{+}18, YZQ16]. College [WMR17, YLC^{+}19, BSW^{+}13]. Collocation [LWWL11]. Colluding [QTM23]. Collusive [ADJ^{+}20, DJNC21]. colocation [WCBK11]. Colonography [MMDY15]. Color [CCZ^{+}15, WH11a]. Color-Guided [CCZ^{+}15]. colorblindness [WLH10]. COM [RC13]. Combating [SQJ^{+}19]. Combination [HYL^{+}18]. Combined [FDE15]. Combining [LLS^{+}21, LFL^{+}20, SPDR15]. COMET [LFG^{+}23]. Comfort [ASSR18, VPD^{+}22]. Comfort-Based [ASSR18]. commentary [WW13]. Comments [GW17]. Commentsphere [PSLB12]. Commercial [ZZZ20a]. Commitment [POM20, BBMP13]. Commodities [ZZC^{+}20]. Commodity [WYM17]. Common [AWSF21, Hec19]. Commonsense [HB12]. Communication [CHP17, MZY^{+}22, CABD13, CAB^{+}13]. Communication-Efficient [MZY^{+}22]. Communities [BR15, CCWS17, TY14, YL14]. Community [BBM17, BN21b, CZP^{+}14, CBP13, GGY^{+}23, JLJ^{+}20, KLL17, KLLL20, PBvL14, TPG^{+}19, WFZ^{+}18, WLH17, YCGH12, ZNWC14, ZMH^{+}15, Goo10]. Community-Based [CZP^{+}14]. Commuter [VPD^{+}22]. companion [TZY^{+}13]. Comparative [DMGSD23]. Comparing [FSW^{+}20]. Comparison [LZW^{+}23, VDL^{+}19, BCD^{+}13].

Comparisons [JV20]. Competing [EBS^{+}22]. Complementarity [WZCJ21]. Complementary [EHG21]. Completion [ZWH17]. Complex [ASW^{+}19, JPS^{+}16, ZSY^{+}12]. compliant [JTP^{+}21]. Complier [CDGZ16]. Component [LGL^{+}16, LYWH20]. Composing [WL23]. Compound [XSJW23]. Comprehension [DSS^{+}22, LXJ^{+}20]. Comprehensive [BCR21, DT16, KZL^{+}17, RAK23, WWD^{+}21, XWW^{+}21]. Compression [DLGT19, YZL^{+}19, CL13]. Compressive [LYKS23]. Computation [NTM^{+}16, YCP^{+}13]. Computational [GST12, LWF^{+}23, YLD^{+}22, GY11, WCBK11]. Computationally [FTCP^{+}13]. Computations [RHF16]. Computer [GJS23, MFB^{+}20]. Computer-aided [GJS23]. Computing [DB16, HTM15, HLC^{+}21, KP17, NDW^{+}19, SLC23, SRMW19, Ten23, XLZ^{+}22, ZCWY14a, ZCWY14b, ZDW19, LN10, YNS13, YZEC13]. Con [Wid17]. Conceal [WMWR22]. Concept [CYC^{+}23, JTZ^{+}11, JLW^{+}23, LWH12, LJC^{+}11, WJY^{+}18, XLZ21, YLCT19, CZLS13, SLWW13]. Concepts [ZCWY14b]. Conceptual [HAAM12]. concurrent [HLJ11]. Condition [DLY^{+}21]. Conditional [GWD^{+}21, WYNW20, ZZGH19]. Conditioning [ASSR18]. Conditions [SGJC18]. Cone [TTLG17]. Conference [SY12, CXW^{+}13]. Confidence [ODF17, WZH16]. Confidence-Weighted [WZH16]. Configurations [HTM15, CCG^{+}13]. Configure [GWL^{+}23]. conformant [TNSP13]. Confounding [QCZ^{+}21]. Congestion [WYY^{+}23]. Connected [SRMW19, ZCL^{+}21]. Connecting [CXW^{+}13, KBM^{+}21]. Connections [LGJ^{+}22]. Conquer [PKCC18, WMH18]. Conscious [WLC^{+}20]. Consensus [ZYC^{+}22]. Consensus-based [ZYC^{+}22]. consider [ZC13]. Considering [CGZ18]. Consistency [LJLZ19, LWLG22, TTL^{+}21, ZCS^{+}12]. Consistent [FWZ17, FDE15, ZLZW23]. Constitutive [BBMP13]. Constrained [BFC^{+}17, LFWY23, SZZ^{+}21]. Constrains [WLL^{+}20]. Constraint [ODF17, SHX^{+}23, WM21, RYS10]. Constraint-based [WM21, RYS10]. Constraints [LCD18, ZS18, ZSL^{+}15, MMC^{+}13, PCC10]. Constraints-Based [LCD18]. Construction [EFMRK^{+}20, KW17, PKCC18]. Consumer [YY15, ZT11]. Consumption [LAsO^{+}19, TBW21]. Contact [FZH^{+}21]. Content [BLL^{+}14, CRYT12, CCC^{+}12, EMF12, SDD^{+}16, TL23, YWZ^{+}17, ZZZ^{+}11, CCL13]. Content-Aware [TL23]. Content-Centric [ZZZ^{+}11]. content-driven [CCL13]. Context [CASR22, FXHM16, HQW22, JYT^{+}12, LCCT12, LCKY14, LJC^{+}11, XLF^{+}20, ZLG^{+}20, ZCX^{+}15, BGMS13a, LHZ13, ME13, SBD13, SRM^{+}13, SLH13]. Context-Aware [FXHM16, JYT^{+}12, LJC^{+}11, XLF^{+}20, ZCX^{+}15, ZLG^{+}20, BGMS13a, SLH13]. Context-Free [LCKY14]. Context-Sensing [LCCT12]. Contextual [ABG^{+}11, CHHH18, ZPL^{+}20, SLH13]. Contextualized [SC17]. Continuation [ZSLC19]. Continuous [DWKP16, GBBD22, HKMN20, LG16, ZLZ^{+}22]. Contracts [CWCK15]. contrarian [HLT11]. Contrastive [XXL^{+}23, ZWX^{+}22]. Contributed [YY15]. Contribution [LCY^{+}22, ZFLD22, BMV13]. Control [BYK^{+}21, BC19, HDPH16, HTSC^{+}17, KW17, PMR^{+}17, QTM23, YTH17, PCC10, RVRJ11]. Controlling [SSV19]. Controversial [SRB15]. Convergence [STA22, WZZ^{+}21]. conversation [WCBK11]. Conveying [HNV14].

Convolution [JYL^{+}23, LGJ^{+}22, ZDW19]. Convolutional [CZG^{+}23, FLY^{+}23, JLH19, LLS^{+}22, LWZ^{+}23, LFG^{+}23, LCJ^{+}19, QHH^{+}21, WTL20, WZM^{+}22, ZLG^{+}20]. Convoys [YBZ^{+}20]. Cooperative [LFW23, YHF21]. Coordinated [GXT^{+}23]. Coordinates [CDS12]. CoPhy [EBS^{+}22]. CoPhy-PGNN [EBS^{+}22]. Copula [HKMN20]. Copula-Based [HKMN20]. CORALS [BKB10]. Coranking [WXZ^{+}16]. coreference [CST13]. Cores [DJS16]. CORN [LHG11]. Corpora [CDS12, GOB^{+}12]. Corpus [TRDD12]. Correction [WWD^{+}21, ZYC^{+}22]. Correlated [BLAK19]. Correlation [CPHL15, DLLT21, LHG11, LSW^{+}20, SST^{+}15, TZC^{+}20, WZM^{+}22]. Correlation-aware [WZM^{+}22]. Correlation-driven [LHG11]. Correlations [WZY^{+}18, YSN^{+}17, ZCL^{+}21, TEP11]. Correspondence [PS11]. Corrupted [LNYV22]. Cost [BL16, HWCL17, SLZ^{+}23, WZY^{+}18, ZZZ^{+}20b, BD11, LHC^{+}13]. Cost-Effective [WZY^{+}18]. Cost-Optimized [HWCL17]. Cost-sensitive [SLZ^{+}23, LHC^{+}13]. costly [WH10]. Count [EL14]. Counterfactual [GSM23, HM19]. counterinsurgency [HKO13]. Counterterrorism [SDHS15]. Counting [MZL12]. Coupled [AWSF21, RGC^{+}22]. Courier [WLW^{+}22, WLW^{+}23]. Couriers [WLW^{+}23]. cover [MGB^{+}11]. Coverage [WLC^{+}16, YMC16]. Covert [WMWR22]. COVID [BZX^{+}22, SBS^{+}23]. COVID-19 [BZX^{+}22, SBS^{+}23]. COVID-GAN [BZX^{+}22]. CpA [TNSP13]. CRADLE [MGS17b]. crater [DSM^{+}11]. Crime [LWLG22]. CrimeTensor [LWLG22]. criminal [TEP11]. crisis [Goo10]. criteria [LKD13]. Criterion [ZCWZ18]. CRM [NTM^{+}16]. Cross [DCWP22, GB22, HWCL17, KLL22, KL23, LHS18, LBT23, PS11, STA22, VDL^{+}19, WZZ^{+}16, WLL^{+}22, WWL^{+}22, ZWC23, ZNWC14, ZZC^{+}20, PG13]. Cross-benchmark [VDL^{+}19]. Cross-branch [WLL^{+}22]. Cross-Domain [LHS18, ZZC^{+}20]. cross-level [PG13]. Cross-Lingual [PS11]. Cross-Media [HWCL17, WZZ^{+}16]. Cross-Modal [ZWC23]. Cross-platform [WWL^{+}22]. Cross-Region [GB22, ZNWC14]. Cross-Silo [DCWP22, STA22]. Cross-view [KLL22]. Crowd [ABO17, BL16, CALK16, DB16, GCY^{+}15, JCW^{+}22, KCTT16, LFY^{+}22, SNL^{+}16, SLR^{+}16]. Crowd-Mobility [ABO17]. Crowd-Powered [SNL^{+}16]. Crowdedness [DCF^{+}18]. Crowdedness-Aware [DCF^{+}18]. Crowds [KF18, FK13]. Crowdsensing [WZY^{+}18, XZW^{+}15]. Crowdsource [MRJ16]. Crowdsource-Based [MRJ16]. Crowdsourced [ZLH18]. Crowdsourcing [CWR^{+}16, CCK^{+}18, FS17, HDPH16, HTL^{+}20, KAH^{+}16, LDTX16, MIS20, MIRS23, RFJ16, SLR^{+}16, TTFS18, BPS13, RBK^{+}13]. Cryptographic [ZZC^{+}22]. CSL [ACC21]. CSM [JPS^{+}16]. CSOC [SGJC18]. CTR [WWL^{+}22]. CUDIA [PG13]. Cultural [YZQ16, Bai10, LN10, YNS13]. Cultures [GKG^{+}11, KF18]. Current [ZRX^{+}22]. Curriculum [GYT19]. Curve [PHL^{+}20]. Curves [ZS18]. Customer [FZH^{+}21, PSRL12, TBW21, WYZ23, GXZ^{+}11]. Customized [LMWS13]. Cut [MTC^{+}20]. Cut-n-Reveal [MTC^{+}20]. Cyber [GME17, HGE17, MWS^{+}18, ODP^{+}17, SSG^{+}20, TAL^{+}19, ZDW19]. Cyber-Alert [SSG^{+}20]. Cyber-Attack [GME17]. Cyber-Physical [MWS^{+}18, TAL^{+}19]. Cybersecurity [GJSC16, GJC17].

D [FDE15, BDP12, CCW^{+}19, FDE15, HYZ15, JLH19, LLS^{+}22, MCEG23, TDVC13, WGF^{+}23, ZYT^{+}15, ZSS^{+}15, ZLT15, ZWH^{+}22]. D-Guided [WGF^{+}23]. D-Map [CCW^{+}19]. Daehr [XZH^{+}17]. DAG [ZAK^{+}23]. Daily [LCLN18]. Dance [KSL^{+}15]. Dancing [YLWX20]. dark [ZZL^{+}19]. Data [ACPS17, ABO17, BYK^{+}21, BD23, BMTT16, BTTT19, BLNN20, CKP^{+}22, CXW^{+}19, CGZ23, CYC^{+}23, CTY^{+}19, DC21, DCWP22, DYQ^{+}23, DOTD16, DGZ15, FWYX22, FE15, FNS16, GXS^{+}22, GB22, Hec19, HTM15, HKMN20, HLL14, HWCL17, HYL^{+}21, HCWH22, HWT17, JLX^{+}17, JSL^{+}19, KCSW23, LWH12, LNYV22, LLL23, LLX^{+}20, LH22, LM11, LC16, LZY^{+}16, MMPS23, Min16, OSW^{+}22, ODP^{+}17, PFS17, PCF^{+}19, PPPM18, PCC17, RDX22, RYC22, RV18, SCLZ17, SLZ^{+}23, SZT12, SS15, SSLM21, TRZ^{+}19, TZC^{+}20, WC12, WTK^{+}19, WCS^{+}20, WZFL21, WZCJ21, WYG^{+}22, WZFL22, WWL^{+}22, WLZ^{+}23, XZXM19, XZH^{+}17, XLZ21, YBZ^{+}20, YLD^{+}22, YY15, YZQ16, YCY23, ZWL^{+}15, ZYH^{+}17, ZLH18, ZW19, ZWL^{+}19, Zhe15, ZBZX12, ZZC^{+}20, ZWGW17, AAX13, BVCH13, BK11, CDK^{+}13, KDC13, LHJ^{+}11, LZCS11, MGB^{+}11, PMSR11, TZY^{+}13, YCP^{+}13]. Data-Aware [MMPS23]. Data-Driven [ACPS17, ODP^{+}17, SCLZ17, ZWGW17, PCF^{+}19, WYG^{+}22]. Data-Mined [CGZ23]. Database [KAH^{+}16]. Databases [MMDY15]. Dataset [ADM^{+}21, DBDM16, YWS^{+}23]. Datasets [FE15, WYP22, XZS20, YDZ20]. DBSCAN [XJS^{+}21]. DClusterE [ZL12]. DDNAS [CLL23]. Dealers [HYL^{+}21]. Dealing [WHR13, YL17]. Deanonymization [FZX15]. Death [SDXG16]. Debiased [ZZL^{+}23]. Decay [Pai16]. Decentralized [HBK^{+}16, WLZ^{+}23]. Decision [BWC15, Bha21, WLW^{+}22, YZZ23, ZZZ^{+}20b, KDC13]. Declarative [MMPS23]. Decoding [ZYH^{+}20]. Decomposition [BWC15, HBK^{+}16, RSCOVCMM17, WYM17]. decoupled [BBMP13]. Deduplication [LLLC19]. Deep [BQF^{+}23, CCGP22, CDW^{+}19, DLGT19, FMdA^{+}23, GXYZ21, GMX^{+}21, HHL^{+}22, JCW^{+}22, JHK^{+}22, KZL^{+}21, KG23, LLL21, LXC^{+}21, LLS^{+}21, LLS^{+}22, LPL^{+}22, LYWH20, LWWX20, LGJ^{+}21, LRD^{+}22, NDW^{+}19, OLY^{+}17, QTM23, RAK23, TLW^{+}15, TLB^{+}21, WZFL21, WZFL22, WYY^{+}23, WC20, WL23, WCFH23, XW23, YLY^{+}23, YLWX20, ZGP^{+}18, ZZZ^{+}19, ZSAL20, ZYH^{+}20, ZMH^{+}22, ZBW^{+}22, ZZC^{+}22]. Deep-learning [ZSAL20]. DeepApp [XLF^{+}20]. DeepExpress [RGC^{+}22]. DeepKey [ZYH^{+}20]. DeepRoute [WLW^{+}22]. DeepTracker [LCJ^{+}19]. Defending [CBPG22, LLPS20]. Defense [WMA20, BKB10]. Defensive [CTC^{+}22]. Definition [MPS23]. Degradation [BC19]. Delays [BMTT16]. Delivery [HCRF21, RGC^{+}22]. Demand [HS19, QHH^{+}21, ZMH^{+}22]. Demographic [LLL21]. Demonstration [ZSY^{+}12]. Density [HTL^{+}20, LC15, XJS^{+}21, LSQ11, SSZ^{+}13]. Density-based [XJS^{+}21, LSQ11]. Dependence [GDC19]. Dependencies [WCP^{+}23]. Dependent [WZZ^{+}16, RCN10]. Deployment [SCLZ17, XLZ^{+}22]. Depth [CCZ^{+}15, ZCG15]. Depth-Based [ZCG15]. Depth-Selective [CCZ^{+}15]. Describe [LH12]. Describing [LHO23]. Description [PBvL14, OY13]. Description-Driven [PBvL14]. Descriptive [FZX15]. Descriptor [GWDJ15]. Design [HDPH16, LDTX16, MGSK19, MBR^{+}14, SMGMC^{+}15, CCG^{+}13, ZZC^{+}22]. Designing [MJVL16]. Detect [YY15, MSSV11]. Detecting [ADJ^{+}20, CCWS17, DJNC21, HLF^{+}21, HM19, HQW22, SKOM13, TY14, LMC^{+}15]. Detection [BBM17, BN21b, CCH15, CDW^{+}21, DC21, DGL^{+}22, FMdA^{+}23, FXR^{+}17, GGY^{+}23, HNA20, HXY^{+}22, HHJ22, JLJ^{+}20, KZL^{+}17, KLL17, KLLL20, KW17, LLF^{+}19, LGL^{+}22, LC15, LZW^{+}23, LSW^{+}20, MGJW20, MRW^{+}12, MLJZ21, MFI19, MWS^{+}18, POM20, PBvL14, WPA^{+}12, WLH17, XZH^{+}17, XLZ21, YCZY21, ZQP^{+}15, ZLT15, ZLY^{+}18, ZFH^{+}18, ZZZ^{+}19, ZPL^{+}20, ZZL^{+}19, ZCZ^{+}23, LWW^{+}23, CDS13, CBP13, SSZ^{+}13]. Detector [HG21, JYT^{+}12, SGTK20]. Determinant [LYLX23]. Development [CRRH11, CDR19]. Developments [ZGP^{+}18]. Device [GWDJ15, ZBW^{+}22]. Devices [TAL^{+}19, ZZZ20a]. dhCM [GGC21]. DHPA [PHL^{+}20]. Diagnose [TWC^{+}23]. Diagnosis [AGP17, LWC^{+}18, RP23, SLC23, SWZ^{+}21]. Diagnostic [GJS23]. Diagram [SPDR15]. Dialogue [FTE21, LFW23, MP23, ZFQX20]. Difference [LCY^{+}15]. Different [ABB^{+}15, GKG^{+}11, KF18, TCK20]. Differentiable [CLL23]. Differential [CP23, KLL17, YCH^{+}22]. Differentially [JZG22, MLSK23]. Difficulty [TPG^{+}19]. DiffQue [TPG^{+}19]. Diffusion [CCW^{+}19, STP^{+}18, WMA20, SKOM13]. Digg [PT12]. Digital [MFLP14]. DILSA [KZL^{+}21]. Dimension [JZG22, LFG^{+}23, XTW17]. Dimensional [HYHFV22, HKMN20, PCC17, PKCC18, BLNN20]. Dimensionality [CSA^{+}23]. Direct [KKG18, LG16]. Directed [WYC^{+}22]. Directions [MPS23, QCL15, CAB^{+}13]. Directly [XXL^{+}17]. Dirichlet [LLX^{+}20, LZCS11, TY14]. Disaggregation [LZ18, WCBL18]. Disambiguation [LFL^{+}23]. Disasters [SZS^{+}17]. Disclosure [ARGK15, HTM15]. Discover [ZW19]. Discovering [FGP11, WYY^{+}23, XZS20, ZZKT20]. Discovery [CWLZ15, FSW^{+}20, GGC21, LYW^{+}19, LC16, Pat15, QCZ^{+}21, SSV19, YCGH12, ZLB^{+}16, ZZGH19, ZMH^{+}15, DSM^{+}11, Edi13, LHJ^{+}11, TZY^{+}13]. Discrete [GZ23, LC16, POM20, SSL^{+}18]. Discretized [CLL23]. Discriminant [GG23, XZH^{+}17]. Discriminative [DLGT19, LCD18, TZC^{+}20, TMZ^{+}20, YWZ^{+}23, ZLG^{+}20]. Disease [HMG^{+}23, UAS15]. Diseases [GJS23, UAS15]. Disentangled [CYW^{+}21]. disjuncts [Dha11]. Disorders [XZH^{+}17]. Dispatching [YTH17]. Dispersal [KZL^{+}21]. Display [CMR15]. Dissimilar [DPSS19]. Dissipative [LLZW17]. Distance [HJTZ12, WHJ^{+}11, WYNW20, WHC13]. Distant [SP16, YHF21]. Distillation [LXC^{+}21]. Distinguishing [BDP12]. Distributed [EFMRK^{+}20, HWCL17, JJKZ22, MLJZ21, MMDY15, NVDMFD22, ODF17, TB22, XLL^{+}22, ZZZ^{+}19, ZBW^{+}22]. Distribution [HWCL17, SHX^{+}23, TTL^{+}21, WCF^{+}20, ZYH^{+}17, ME13]. Distributional [Mar13]. Diversification [XXL^{+}17]. Diversified [LXC^{+}21, WLC^{+}16]. Diversifying [KSKÇ15]. Diversity [BBS^{+}16, CWLZ15, JV20, LCC^{+}20, MG16, XXL^{+}17]. Divide [WMH18]. Divide-and-Conquer [WMH18]. DMAD [SCLZ17]. Dockless [CLL^{+}21]. Doctors [TWC^{+}23]. Document [HWZL20, RHD^{+}12, WHW^{+}21, ZL12]. Documentation [RBG22]. documents [ESNN13]. Doing [GB22, ZZZ^{+}11]. Domain [LHS18, LCN^{+}21, LBT23, MFI19, Min16, QWC^{+}23, SGTK20, TTL^{+}21, TPM23, WC20, WYNW20, WHR13, ZZC^{+}20]. Domain-attention [WYNW20]. Domains [Min16, MGS17b]. Domestic [PMR^{+}17]. Double [DLGT19, PKCC18, ZCL^{+}18]. Double-Sided [PKCC18]. down [CST13]. Downstream [BN21a]. Drawn [JTZ^{+}11]. Drift [XLZ21]. Drifts [LWH12]. drinking [BSW^{+}13]. Driven [ACPS17, BC19, DLY^{+}21, LSW23, ODP^{+}17, PBvL14, RSCOVCMM17, SCLZ17, YWZ^{+}17, ZSL^{+}15, ZWGW17, CCL13, LHG11, PCF^{+}19, WYG^{+}22]. Driver [PHL^{+}20, VPD^{+}22, ZLT15]. Driving [CYC^{+}21, NTM^{+}16, VPD^{+}22, ZLBZ23, ZZZ^{+}22]. Drug [HYL^{+}21, YY15, YL17]. Dual [CDW^{+}19, JYL^{+}23, SLZ^{+}23, WLL^{+}22, XLG^{+}23, HLT11]. Dual-attention [WLL^{+}22]. dual-classifier [HLT11]. Dual-stage [SLZ^{+}23]. DUCT [ODF17]. Durability [ZWL^{+}19]. During [VPD^{+}22]. Dutch [SDD^{+}16]. Duty [LCCT12]. Dynamic [CLL^{+}21, CCWS17, DLY^{+}21, DPC16, GJSC16, GGC21, HLGW13, HXY^{+}22, JJKZ22, LGL^{+}22, NOZ20, PHL^{+}20, QHH^{+}21, RSCOVCMM17, SGJC18, SS11, TY14, WPA^{+}12, WCF^{+}20, WYC^{+}22, WYZ23, XSJW23, ZZZ^{+}20b, ZMH^{+}22, ZRX^{+}22, BNS13]. Dynamic-Aware [HXY^{+}22]. Dynamics [DCF^{+}18, GRR^{+}15, JCW^{+}22, LH12, PEK^{+}16, ZLZ^{+}22, Gin13]. Dystemo [SP16].

EachWiki [WFJY12]. Early [BCC^{+}13, KZL^{+}17, LCC^{+}20, XZH^{+}17]. Earth [HSJ^{+}22, JHK^{+}22]. Echoing [YLWX20]. Economic [PMR^{+}17, ZLH18]. Ecosystem [JTP^{+}21]. Edge [CCGP22, JBLW21, KMH22, NDW^{+}19, SLC23, SSV19, TAL^{+}19, XLZ^{+}22, YCL^{+}21, ZBW^{+}22, ZDW19]. Edge-specific [SSV19]. Editing [LLDT16, SPDR15]. Editor [SRB15]. EEG [DPB20]. EEMC [XZW^{+}15]. Effect [CDGZ16, Hec19, PCL18]. Effective [CCG^{+}13, DSS^{+}22, FZX15, LLS^{+}21, MRJ16, ODL^{+}20, TBW21, WZY^{+}18, ZYSL12, ZLY^{+}18]. Effectively [BN21a, RKH14]. Effectiveness [SGJC18, WLWC23]. Effects [GW17, LCLG19, LG16, SDXG16]. Efficiency [WLWC23, ZKC^{+}23]. Efficient [BBM17, CWCY22, CYYL18, CTY^{+}19, DSS^{+}22, KL23, LCCT12, LLS^{+}21, LYZ^{+}23, LM11, LCY^{+}22, MZY^{+}22, RK15, RZS^{+}15, STA22, TRZ^{+}19, WCBL18, XLG^{+}23, XLL^{+}22, XKW^{+}16, XZW^{+}15, ZFWL17, ZYY^{+}23, CCG^{+}13, FTCP^{+}13, TAL^{+}19]. Effort [RFJ16]. Ego [CCW^{+}19]. Ego-centric [CCW^{+}19]. Egocentric [CXW^{+}19]. EHR [CYC^{+}23]. Eigenvalue [EBS^{+}22]. Elderly [LCLN18]. Elders [SMGMC^{+}15]. Electric [MBR^{+}14, WZS^{+}20, WFX^{+}21]. Electricity [PMR^{+}17, TBW21]. Electricity-Based [PMR^{+}17]. Electro [BC19]. Electro-Mechanical [BC19]. Electromyogram [ZZD^{+}17]. Electronic [DLWF22, XZH^{+}17, BVCH13, ZT11, ZC13]. elicitation [ERR13]. eligibility [LKD13]. Elimination [JV20]. Elusive [ZCL^{+}21]. Embedded [FMdA^{+}23]. Embedding [CYC^{+}23, DMGSD23, WFZ^{+}18, WZCJ21]. Embeddings [LHZ22a, LLL23]. Emergency [YTH17]. emerging [CDS13]. Emilia [CGMC11]. Emilia-Romagna [CGMC11]. EMMA [KAH^{+}16]. Emotion [KAH^{+}16, SP16, YC12]. Empathetic [KAH^{+}16]. Empirical [GJS23, TCK20, BCD^{+}13]. Empowering [CCL15]. Enable [GDC19]. Enabling [XZW^{+}15]. Encoded [BC19]. Encoders [SZZ^{+}21]. Encoding [TCK20, ZCL^{+}18]. Encoding-Layer [ZCL^{+}18]. End [CDR19, THL^{+}15, WLL^{+}20, ZBW^{+}22]. End-to-End [THL^{+}15, WLL^{+}20]. End-User [CDR19]. Endpoint [LG16]. Energy [LLL21, LCCT12, LZ18, PMR^{+}17, STA22, TAL^{+}19, XZW^{+}15, ZSL^{+}15, ZFWL17, RVRJ11]. Energy-Efficient [LCCT12, STA22, XZW^{+}15, ZFWL17, TAL^{+}19]. Engine [LCM^{+}12]. Engineering [BTL20]. Enhanced [BZX^{+}22, DJI^{+}16, KZL^{+}21, LWZ^{+}23, GWL^{+}23]. Enhancement [OSW^{+}22]. Enhancing [MIRS23, ZPP^{+}21]. Enough [WLW^{+}23]. Ensemble [JV20, JSL^{+}19, KCS18, LXC^{+}21, SLM^{+}21, TCK20, YCZY21, ZSY^{+}12].

Ensembles [BFHP12, LNO^{+}18]. Entities [DJNC21, KXZG15]. Entity [HWZL20, LLY12, LYZ^{+}23, LSW23, MPS23, LWZZ13]. Entity-Relationship [LLY12]. Enumerating [SRMW19]. Environment [GXYZ21, HTM15, KSL^{+}15, LXBW20]. Environment-Aware [GXYZ21]. Environmentally [ZGP^{+}18]. Environments [DZY^{+}22, KHNB15, POM20, VKLY18, dMFA^{+}13]. Ephemeral [CVCL22]. Epidemic [YXL^{+}23]. Episode [ASW^{+}19]. Equation [RGH19]. Equations [CP23]. Era [JLX^{+}17]. Error [BC19, CSHL21, WWD^{+}21]. Espressione [Wid17]. Essence [FSW^{+}20, Wid17]. Estimate [LPR19, WLW^{+}23]. Estimating [BZX^{+}22, SSV19, TPG^{+}19, UAS15, XZR12, XXZ^{+}21]. Estimation [BD23, HTL^{+}20, JYL^{+}23, LGZ^{+}17, ST19, SS15, WTK^{+}19, ZWZS16, ZFLD22]. Ethnicity [BDP12]. Ethnicity-Based [BDP12]. Euler [ZWZZ23]. Evacuation [IVS^{+}16]. Evaluating [MG16, ZL12]. Evaluation [CVCL22, LCY^{+}22, PCF^{+}19, SSG^{+}20, SS15, SHB^{+}12, XXL^{+}17, ZQP^{+}15]. Event [ASW^{+}19, CCW^{+}19, CDW^{+}21, DGL^{+}22, GZZY17, GGC21, Min16, TCK20, ZQP^{+}15, ZSL^{+}15, ZLY^{+}18, ZL19, CST13, MMC^{+}13]. Event-Based [ZL19]. Event-centric [CCW^{+}19]. Event-Driven [ZSL^{+}15]. Events [ABTS15, DC21, HQW22, JCW^{+}22, KZL^{+}17, KZL^{+}21, LC12]. Evolution [RCN10, WZS^{+}20]. Evolutionary [JTP^{+}21, RV18]. Examinee [LWC^{+}18]. example [HASS22]. Excitation [ZHW^{+}21]. Exclusive [LYLX23]. Expansion [WFX^{+}21, BGMS13b]. Expect [AT15]. Experience [FZH^{+}21, KCTT16]. Experimental [ODL^{+}20]. Experimentation [DB16]. expertise [SLWW13]. Explainable [BZW^{+}22, JLW^{+}23, MP23, ZFS^{+}19]. explaining [MPA13]. Explanation [GSM23, LZC23]. Explanations [GSM23, LZC23, MTC^{+}20, ZLZW23, CLSL13]. Explicit [LYZ^{+}23, MKL11]. Explicitly [BBS^{+}16]. Exploiting [BLL^{+}14, JLL18, SFX17, WWZ^{+}16, WZY^{+}18, ZZL^{+}19]. Exploration [CDS12, CCW^{+}19, CCZ^{+}23, CCG^{+}13]. Explorations [LLS^{+}21]. Exploratory [MGS17b]. Exploring [CHP17, FL20, LSW^{+}20, NZS^{+}22, SST^{+}15, WPL13, YZY^{+}17, ZZD^{+}17]. Express [RGC^{+}22]. Expression [RP23, XTW17, ZWH^{+}22, WH10]. Expressional [ZWH^{+}22]. Expressions [JLH19]. expressive [LHC^{+}13]. Extensible [THL^{+}15]. Extension [CDW^{+}19, ZY12]. Extensions [LC16]. Extracting [ABTS15]. Extraction [CUG^{+}12, LCLN18, LYZ^{+}23, LFW23, LWWL11, Min16, YLY^{+}19, YHF21, ZLZ15, ZB20, LCCS13]. Extreme [HYHFV22, HQW22, ZWZ^{+}19].

Face [BDP12, DT16, HXY^{+}22, LGL^{+}16, LLDT16, LTS^{+}15, TMZ^{+}20, WGF^{+}23]. Faceted [TRH12]. Facial [BDP12, DLLT21, JLH19, XTW17, ZWH^{+}22]. Facilitating [WFJY12]. Facility [WLC^{+}20, WLWC23]. Factor [PCL18, BSW^{+}13, LLL^{+}18]. Factorization [CWCY22, CWC^{+}20, CZJL15, CYYL18, LLL21, LCD17, LCD18, PKH^{+}17, Ren12, TTLG17, SLH13]. factors [QSRGDAJD13]. Failure [NCG21]. Fair [VS11]. Fairness [LHZ22a, LCX^{+}23]. Fairness-aware [LHZ22a]. FairSR [LHZ22a]. Faithful [ZLZW23]. Fake [DGL^{+}22, HNA20, LGL^{+}22, SQJ^{+}19, YCZY21]. False [SSV19]. Fashion [SNL^{+}16]. Fast [CYC^{+}21, CZJL15, EFMRK^{+}20, MWY^{+}23, DL13]. Fatigue [ZLT15]. FCL [SLC23]. Feature [BTL20, DMGSD23, EK15, FC15, GG23, HHJ22, JYT^{+}12, JLL18, LYW^{+}19, MLJZ21, MMDY15, PKH^{+}17, QWC^{+}23, RHD^{+}12, SLZ^{+}23, WPA^{+}12, YGY^{+}23, YLY^{+}19, ZYT^{+}15, ZLZ15, LWW^{+}23, FC15]. Feature-Based [EK15, RHD^{+}12]. Feature-Rich [FC15]. Feature-Space [FC15]. Features [BZX^{+}22, BDP12, HCCY15, KZL^{+}21, SGTK20, TS17, WZS^{+}15, WWZ^{+}16, YGU15, ZCG15, SHZ13]. FedBERT [TWL^{+}22]. FedCTR [WWL^{+}22]. FedCVT [KLL22]. Federated [AdCK^{+}22, CWCY22, CKP^{+}22, CBPG22, DGK^{+}22, DLWF22, DCWP22, GWL^{+}23, GYL^{+}22, GGY^{+}23, HXY^{+}22, HLL^{+}22, HLC^{+}21, JTP^{+}21, JTS^{+}21, JJKZ22, JZG22, KLL22, LNYV22, LHZ^{+}22b, LYF^{+}22, LCY^{+}22, MZY^{+}22, NVDMFD22, OSW^{+}22, RDX22, RYC22, STA22, TWL^{+}22, WWL^{+}22, XLL^{+}22, XLZ^{+}22, YLCT19, YTL^{+}22a, YTL^{+}22b, ZYC^{+}22, ZFLD22, ZLD^{+}23, ZGF^{+}23, ZKC^{+}23, ZBW^{+}22]. Federation [GWL^{+}23]. FEED [YMLM16]. Feedback [CSHL21, HLL^{+}23, NYBG17, ZYH^{+}17]. Feedback-aware [HLL^{+}23]. Few [LCY^{+}18, LZW^{+}23, MPS23]. Few-Shot [LCY^{+}18, LZW^{+}23, MPS23]. Fi [SCLZ17, ZZZ20a]. field [TLWZ11]. fields [SGD13]. Filtering [CSN^{+}17, JJ14, LPM20, LFG^{+}23, ST20, TB22, WSGM14, BCD^{+}13, ERR13]. financial [Dha11]. Find [LCY^{+}15, SS22]. Find-the-Difference [LCY^{+}15]. Finding [DPSS19, GLL^{+}17, LCLG19]. Fine [BN21a, LWLG22, OLY^{+}17, TWC^{+}23, WWZ^{+}16, WLL^{+}21, YTH17]. Fine-Grained [TWC^{+}23, WWZ^{+}16, WLL^{+}21, YTH17]. Fine-Scale [LWLG22]. Fine-to-Coarse [OLY^{+}17]. Fine-Tuning [BN21a]. FL [JZG22]. FLAG [HLL^{+}23]. Flashback [DYQ^{+}23]. Flattening [KLL17]. Flatter [MBM21]. FLEE [ZBW^{+}22]. Fleet [ZMH^{+}22, DGK^{+}22]. Flexible [ZLC^{+}20]. Flight [BMTT16]. Flow [KZL^{+}17, LFY^{+}22, LLS^{+}22, LGJ^{+}22]. Fly [SSLM21]. Fog [TRZ^{+}19]. Fog-assisted [TRZ^{+}19]. folder [BD11]. FolderPredictor [BD11]. Folksonomy [CUG^{+}12, FSS15, SHB^{+}12]. Folksonomy-Based [CUG^{+}12]. Following [SZS^{+}17]. Footprints [MFLP14]. Foraging [DZY^{+}22]. Force [GRR^{+}15]. Forecasting [AAX13, CP23, HYL^{+}18, JJ15, LGJ^{+}22, QHH^{+}21, SLZ^{+}23, TBW21, ZZC^{+}20]. Foreground [CCH15]. Forensics [DLGT19]. Forest [WAL18, ZZZ^{+}19, MGB^{+}11]. Forgery [HXY^{+}22, LC15]. Formalizing [GS13]. Formation [BFC^{+}17, MFB^{+}20, WH10]. Foundations [LCX^{+}23]. Foursquare [JCH14]. Framework [CCZ^{+}15, CYKL16, DBDM16, FWYX22, FLLX18, GG23, HCCY15, HRCT16, HMS^{+}14, HMCW15, KLL17, KLLL20, LFWY23, MG16, PHL^{+}20, RKH14, SZC^{+}14, TTFS18, WSGM14, WLC^{+}16, XLG^{+}23, XLL^{+}22, XZH^{+}17, YLD^{+}22, ZL12, ZSL^{+}15, ZLZ15, ZBW^{+}22, CLSL13, CCL13, LKD13, TZY^{+}13, ZT11, ZC13]. Fraud [ZZZ^{+}19]. Free [LCKY14, ZGF^{+}23, TPM23]. Frequency [ACPS17, SGTK20]. Frequency-Based [ACPS17]. Frequent [ASW^{+}19, PDR23]. friendship [MPA13]. Frontal [WGF^{+}23]. FROST [WLC^{+}20]. Fruit [FMdA^{+}23]. FSR [FC15]. Fully [CZG^{+}23, SSZ^{+}13]. Function [WWZ^{+}16]. Functional [ZWZS16]. Functions [EBS^{+}22, TDVC13, HLY^{+}14]. Fundamentals [ZFH^{+}18]. Furniture [HWL^{+}17]. Furtherance [MIRS23]. Fused [BD23, XKW^{+}16, ZZL^{+}23]. Fusing [CWC^{+}20, ZCG15]. Fusion [GWDJ15, HLF^{+}21, HYL^{+}21, HMCW15, PFS17, QWC^{+}23, SZC^{+}13, WYZ23, WLL^{+}22, XSJW23, YZZ23, ZWZ^{+}19, ZZS^{+}21, ZZC^{+}20, CZLS13]. Future [QCL15, SGY^{+}22, Ten23, WXZ^{+}16, Goo10]. Fuzzy [DJI^{+}16, KAH12, LWC^{+}18, YWZ^{+}23].

Gait [WWZ^{+}16]. Gaits [ZYH^{+}20]. Game [Bha21, HB12, JD15, LCY^{+}15, MRJ16, SSG^{+}20]. Game-Theory [MRJ16]. Game-Theory-Based [Bha21]. Games [MFB^{+}20, SRMW19, WMA20, FT10]. Gamesourced [SS15]. Gamification [MIRS23]. GAN [BZX^{+}22, DDZ^{+}21]. Gap [GME17]. Gaps [SS22, SSS11]. Gated [CLL^{+}21, LGJ^{+}22]. Gathering [KZL^{+}17, KG11]. Gaussian [FNS16, LGZ^{+}17]. Gaze [HLNL18]. GDPR [JTP^{+}21]. GDPR-compliant [JTP^{+}21]. Gen [BI17]. Gene [UAS15]. Gene-Disease [UAS15]. Genealogy [FE15]. General [GMX^{+}21, LS16, THL^{+}15, WLF^{+}18, ZWZS16]. General-Purpose [LS16]. Generalizable [TMZ^{+}20]. Generalization [QWC^{+}23]. Generalized [XKW^{+}16, CFG13]. Generate [WLC^{+}16]. Generated [CRYT12, CCC^{+}12, LCY^{+}15, SDD^{+}16, ZX11]. Generating [MD13, YMC16, ZDC^{+}13, CLSL13]. Generation [CUG^{+}12, CTC^{+}22, CASR22, DB16, FTE21, GWD^{+}21, MCEG23, SDHS15, WGF^{+}23, ZFQX20, ZFS^{+}19, Mar13, YSJ13]. Generative [BZX^{+}22, GXS^{+}22, LXJ^{+}20, RDX22, WYC^{+}17, WLL^{+}21]. Generic [ZDL^{+}12]. Genetic [AHJB20, UAS15]. Genres [TS17]. Geo [HWCL17, JGL^{+}15, ZWH17, ZLY^{+}18]. Geo-Distributed [HWCL17]. Geo-Social [JGL^{+}15]. Geo-Tagged [ZWH17, ZLY^{+}18]. GeoBurst [ZLY^{+}18]. GeoCloud [ZWL^{+}15]. Geographic [CRRH11, JSL^{+}19, YLT13]. geographic-temporal-semantic [YLT13]. Geographical [HL19]. geolocating [CCL13]. Geolocation [SZC^{+}14]. Geosocial [KLLL20]. Geospatial [SSS11, SDS12, Siz12]. Geotagged [ZLH18, ZZC12, SSHL13]. Geotemporal [HWT17]. Gestures [LMAP16]. Getting [Wid17]. Global [HLL^{+}23, LFL^{+}23, SBS^{+}23, WLZ^{+}23, YLH^{+}23, ZCS^{+}12, ZHLL21, MGB^{+}11]. GNN [LWW^{+}23]. GNN-based [LWW^{+}23]. Go [LMAP16]. Goal [MGSK19]. governing [Sin13]. GPS [FWYX22, ZX11]. GPSClean [FWYX22]. Gradient [CZJL15, LLS^{+}21]. Gradients [WCFH23]. Grained [TWC^{+}23, WWZ^{+}16, WLL^{+}21, YTH17]. Grammars [LCKY14]. Grammatical [WWD^{+}21]. Granularity [HLY^{+}14]. Graph [BFC^{+}17, CZG^{+}23, CSA^{+}23, CLL^{+}21, CP23, CDW^{+}21, EFMRK^{+}20, FLY^{+}23, FZX15, HSJ^{+}22, HQW22, HLH^{+}21, JBLW21, JJKZ22, JYL^{+}23, KMH22, LS16, LCN^{+}16, LCD17, LLF^{+}19, LFY^{+}22, LHZ22a, LWZ^{+}23, LFW23, LHZ^{+}22b, LYF^{+}22, LDFL23, LAS20, LLX^{+}22, LGJ^{+}22, LFWY23, PKH^{+}17, RXL^{+}23, RAZE18, SZZ^{+}21, SRMW19, WXLY12, WZM^{+}22, YLH^{+}23, YXL^{+}23, YHF21, ZLZW23, ZZH^{+}22, ZXY^{+}23, SZC11, THY^{+}11]. Graph-based [LAS20, THY^{+}11]. Graph-Constrained [BFC^{+}17]. Graph-Mining [LS16]. Graph-Restricted [SRMW19]. Graphical [LGL^{+}22]. Graphs [BD23, GYT19, KLL17, LWZ^{+}23, OOD^{+}17, UAS15, WFZ^{+}18, WYC^{+}22, YLY^{+}23, CST13, SLWW13]. Gravity [ZC15]. Gravity-based [ZC15]. Gray [WGL^{+}22]. Gray-Box [WGL^{+}22]. green [RVRJ11]. grid [RVRJ11]. GRNN [RDX22]. Ground [FSW^{+}20, KP17, YL14]. Ground-Truth [YL14]. Group [BB15, CVCL22, MMDY15, TWL11, QSRGDAJD13, YJHL11]. Group-Buying [BB15]. Grouping [MLJZ21]. Grouping-based [MLJZ21]. Groups [CVCL22]. Growing [LNO^{+}18]. GTAE [SZZ^{+}21]. GTG [LCY^{+}22]. GTG-Shapley [LCY^{+}22]. Guidance [CTY^{+}19]. Guided [CCZ^{+}15, HSJ^{+}22, TZC^{+}20, WGF^{+}23, BTL20, EBS^{+}22, TTL^{+}21, YCZ^{+}23].

H [TNSP13]. hailing [HHL^{+}22, NZS^{+}22]. Hand [JTZ^{+}11, TLW^{+}15]. Hand-Drawn [JTZ^{+}11]. Handling [CCK^{+}18, CLSL13]. Handwriting [ZLZ15]. Happiness [PCL18]. Hardware [KG23]. Harmonious [GWD^{+}21]. Harnessing [SR17]. Hashing [CDW^{+}19, EFMRK^{+}20, SSL^{+}18, ZWC23, ZLC^{+}20]. Hawkes [LZ18]. Hazy [CHY15]. Health [AKR^{+}18, DC21, DLWF22, KXZG15, PEK^{+}16, TRZ^{+}19, XZH^{+}17, YY15, BVCH13, RC13, RY13]. Health-Consumer-Contributed [YY15]. Healthcare [AdCK^{+}22, CCL15, CL15, TY12, WHR13]. Heating [ASSR18, PMR^{+}17]. Heavy [LCCT12]. Heavy-Duty [LCCT12]. helpfulness [ZT11]. HERA [LFL^{+}20]. Heterogeneous [AHJB20, CXW^{+}19, FC15, HLL^{+}23, JSL^{+}19, LGL^{+}16, LH22, LFL^{+}20, RGC^{+}22, WYY^{+}19, XZXM19, YGU15, YDZ20, ZAK^{+}23]. Heuristics [Ten23]. Hidden [Hec19, SRJP12, TY14, YLY^{+}19]. Hierarchical [BWC15, CDS12, CCZ^{+}23, GWL^{+}23, GGC21, HXC^{+}23, HLH^{+}21, LCKY14, YFJ^{+}18, YLH^{+}23, YMC16, ZBW^{+}22]. Hierarchies [ASW^{+}19]. High [BYK^{+}21, HYHFV22, HKMN20, KG23, LXM^{+}18, PCC17, WH18, WCS^{+}20, SSZ^{+}13]. high-density [SSZ^{+}13]. High-Dimensional [HYHFV22, HKMN20, PCC17]. High-performance [KG23]. High-Precision [LXM^{+}18]. High-utility [WCS^{+}20]. highest [BCC^{+}13]. Highly [KL23, LRD^{+}22]. HiGRN [YLH^{+}23]. Histogram [CHY15]. Historical [CLH^{+}22]. History [LX14, WLC^{+}20]. History-Conscious [WLC^{+}20]. HMPSoCs [TAL^{+}19]. Hoc [SBS^{+}23, XWW^{+}21, YLY^{+}23]. Hoeffding [BFHP12]. Holistic [DHF22]. Home [DC21, KHNB15, MND14, RC13]. home-based [RC13]. homeostatic [RVRJ11]. Homes [CPHL15]. Homogeneity [AGP17]. Horizontal [DCWP22, ZFLD22]. Hotel [ADM^{+}21]. House [GBC^{+}22, LPR19]. HTF [AHJB20]. HTF-MPR [AHJB20]. Human [BZX^{+}22, BL16, CKS10, DHF22, FE15, GWD^{+}21, HAAM12, HDPH16, LMAP16, MFB^{+}20, ODP^{+}17, PHL^{+}20, PPPM18, RHF16, SZS^{+}17, SS15, SRJP12, TJL^{+}21, TRH16, UAS15, WST^{+}15, YSN^{+}17, ZHZ18, ZLH18, FGP11, RC13, TBK^{+}10]. Human-aware [CKS10]. Human-computer [MFB^{+}20]. Human-Like [LMAP16]. Human-Machine [GWD^{+}21]. Human-Robot [HAAM12, TBK^{+}10]. Human-Thing [YSN^{+}17]. Humor [GYL^{+}22]. Hybrid [CCH15, HLC^{+}21, LSW23, SRJP12, YCSH23, YZZ23, ZCZ^{+}23, BCD^{+}13, HMCW15]. Hyper [LYLX23]. Hyper-Laplacian [LYLX23]. Hypergraph [FWZ17, XZXM19]. Hyperlocal [TTFS18]. Hypotheses [TS17].

ICS [LWW^{+}23]. ID [NVDMFD22].

Identification

[DMGSD23, FGL17, LYWW18, MND14, SQJ^{+}19, SLR^{+}16, TLLS17, WPA^{+}12, BSW^{+}13, LCC^{+}20, TTL^{+}21]. Identify [POM20, WXLY12]. Identifying [BR15, HYL^{+}21, RHT^{+}18, SRB15, YGU15]. Identities [WLWJ21, WMWR22]. Identity [TTL^{+}21, TMZ^{+}20]. Identity-Discriminative [TMZ^{+}20]. Identity-guided [TTL^{+}21]. IF [CDR19]. II [HTDJ12, ZLSY22b]. iid [FNS16, OSW^{+}22]. illiad [MWS^{+}18]. Illicit [HYL^{+}21, YL17]. Image [CHY15, DDZ^{+}21, DTL15, GHZ^{+}17, HJCK20, JHK^{+}22, KKG18, LCN^{+}16, LLL23, LLZW17, LCY^{+}18, LWWX20, LGJ^{+}21, OLY^{+}17, RAK23, SLM^{+}21, SGTK20, THY^{+}11, TTLG17, TC19, WH11a, WLL^{+}20, WLL^{+}21, ZYT^{+}15, ZWH17, ZYY^{+}23, ZZS^{+}21, ZHW^{+}21, SZC11, WLH10, WHC13]. Image-Centric [KKG18]. image-to-class [WHC13]. Imagery [HCTC12, HSJ^{+}22, LXW^{+}23, MIS20, RHF16]. Images [CWR^{+}16, HHJ22, LPR19, LC15, ZSS^{+}15, THY^{+}11]. Imbalance [BLAK19]. Imitation [HAAM12]. Impact [FSS15, VPD^{+}22, ZZD^{+}17]. Impacts [YWS^{+}23]. Imperceptible [DLLT23]. Imperfect [HXC^{+}23, JHK^{+}22]. Implications [HLL^{+}22, WCBK11]. Implicit [LLZW17, MKL11]. Improve [HCTC12, KCTT16, RHF16, SMGMC^{+}15, WCFH23, RBK^{+}13]. Improved [LZC23, LC15, TRDD12, YCZY21, MD13]. Improving [CDK^{+}13, EHG21, GJ13, RYC22]. imputation [PG13]. In-App [FLLX18]. In-Network [LXC^{+}21]. Incentive [LDTX16, YTH17]. Incentive-Aware [YTH17]. Incentives [FS17, MIRS23, RFJ16]. Incomplete [BLAK19, HWT17, WHRGC22, Zhu19]. incorporating [SLWW13]. Incorporation [CGZ23]. Incremental [WH18, YMC16, BCC^{+}13]. Incubation [CZKJ22]. Independence [FNS16, ZZGH19]. Independent [ZZGH19]. index [DL13]. Indexing [LCM^{+}12]. Indirect [LG16]. Indirectly [DC21]. Individual [DZY^{+}22, HMG^{+}23, MG16, SHX^{+}23, PG13, RYS10]. Indoor [FDE15, YZY^{+}17, SGD13]. Induction [SHB^{+}12]. Inductive [DJI^{+}16]. Industrial [CWR^{+}16, DGJ^{+}23, JTS^{+}21, KW17, WZS^{+}15]. Industry [Bha21, WM21]. Infer [LCCT12, WZHL14]. Inference [BD23, CWCY22, CDGZ16, DHF22, FNS16, GTM^{+}14, GH18, GZ23, LLX^{+}20, RYC22, SLM^{+}23, ZLB^{+}16, ME13, SZC11]. inference-based [ME13]. Inferencing [KG23]. Inferring [Hec19, HL19, HWT17, SSLM21, WCBK11, WCP^{+}23]. Influence [BBM17, CZP^{+}14, CCWS17, LBP19, Pat15, TY12, WXZ^{+}16, WCP^{+}23, ZC15, ZFLD22]. Influence-Based [BBM17, Pat15, ZFLD22]. influencing [HKO13]. Informatics [CL15, RY13]. Information [ARGK15, CHHH18, CCW^{+}19, CWC^{+}20, CRRH11, FZX15, GG23, GLJ^{+}14, HXC^{+}23, HMCW15, JLL18, KAH12, LMC^{+}15, LFW23, Pai16, PSLB12, RK15, SFX17, SR17, STP^{+}18, WLH17, YLWX20, ZYH^{+}17, ZW19, CBP13, KG11, PG13, SKOM13, WHJ^{+}11]. Informed [SBS^{+}23]. Infrastructure [HDPH16, MBR^{+}14]. Inland [HSBRM22]. INN [CZKJ22]. Input [SDHS15]. ins [JCH14]. Insect [FMdA^{+}23]. Insights [HWCL17, KDC13]. Inspection [SSG^{+}20, ZPL^{+}20]. Inspired [WAL18]. Instagram [HYL^{+}21, YL17]. Instance [ZCZ^{+}23]. Instant [CTC^{+}22]. Insurance [NTM^{+}16]. Integrate [PKCC18]. Integrated [FLF^{+}20, HL17, PKH^{+}17]. Integrating [CBP13, PKCC18, ZZZ^{+}22]. Integration [CKP^{+}22, CYC^{+}23, YCGH12, LWW^{+}23]. integrative [WW13]. Intelligence [AJL18, CC12, GST12, GCY^{+}15, KAH^{+}16, ZWGW17]. Intelligent [Bha21, CALK16, CL15, FGL17, HGE17, HJTZ12, HTDJ12, IVS^{+}16, KCSW23, LCM^{+}12, LZCQ12, MMS17, NAPI14, NDW^{+}19, SYHB17, SLC23, TAL^{+}19, WW13, ZLSY22a, ZLSY22b, ZRX^{+}22, Edi13, FKSS13, HTDJ11, LKD13, LLWC13, RY13, YSJ13, YZEC13, ZPY11, MWS^{+}18, Ten23]. interacting [SZC^{+}13]. Interaction [GWD^{+}21, HAAM12, LYZ^{+}23, LFG^{+}23, YWZ^{+}17, ZRX^{+}22, LMWS13]. Interaction-aware [ZRX^{+}22]. Interaction-Driven [YWZ^{+}17]. Interactions [HLNL18, YSN^{+}17]. Interactive [CCW^{+}19, LZP^{+}12, RAZE18, SWZ^{+}21, WH11a, PCC10, YSJ13]. Interdata [WZY^{+}18]. Interest [CYKL16, MIS20, YGU15, YKTL14].

Interesting [XZS20, ZNWC14]. Interests [LXW^{+}23, WZHL14]. Interface [NAPI14]. Interfaces [DPSS19, LZCQ12]. Interior [OSW^{+}22, SHX^{+}23]. Intermittent [GBBD22, PMR^{+}17]. International [RFI^{+}11]. Interpretability [CGZ23]. Interpretable [CZKJ22]. Interpreter [TL23]. Intersections [KL23]. Interval [WCS^{+}20]. Interval-based [WCS^{+}20]. Intervention [HM19, KPF18]. Intervention-Based [KPF18, HM19]. interventions [BVK10]. Intradata [WZY^{+}18]. Intrinsic [ZFLD22]. Introduction [BBGG13, BTVY17, CWLZ15, CALK16, CC12, CABD13, CL15, CCC^{+}12, CSTZ16, Edi13, FS13, GST12, GY11, GCY^{+}15, GCZ11, GCZ13, HLY^{+}14, HJTZ12, HYZ15, Hsu11, HTDJ11, HTDJ12, JLX^{+}17, KN13, LLWC13, Lin11, LN10, LN11, LZCQ12, RY13, SYHB17, SA15, SY12, WDSZ13, WZFL22, Yan10, YNS13, YTL^{+}22a, YZEC13, ZPY11, ZCWY14a, ZLSY22a, ZLSY22b]. Intrusion [MFI19]. Invariant [DT16, HG21, MWS^{+}18, WGF^{+}23]. Investigation [TCK20]. Involved [ADJ^{+}20]. IoT [LLLC19, YTH17]. IoT-Based [YTH17]. IoV [XLZ^{+}22]. Irregular [LFY^{+}22]. Irrelevance [ZYH^{+}17]. Isoform [RP23]. Isolated [GME17]. Israel [BI17]. Issue [AJL18, BTVY17, CKW19, CWLZ15, CALK16, CL15, CSTZ16, GSM23, GCY^{+}15, HLY^{+}14, HYZ15, JLX^{+}17, SA15, WZFL21, WZFL22, YTL^{+}22a, YMLM16, ZLB^{+}16, ZLSY22a, ZLSY22b, ZWGW17, BBGG13, Che10, GY11, Hsu11, HTDJ11, Lin11, LN10, ZPY11]. Italian [CGMC11]. Item [CYW^{+}21, WYC^{+}17, WYD^{+}18, XYS^{+}23, GJ13]. Item-Set [LPM20]. item-specific [GJ13]. Items [EK15]. Itemset [WZCJ21]. Iteratively [WMH18].

Joint [BD23, GTM^{+}14, LAsO^{+}19, LWH^{+}20, LYZ^{+}23, LC15, XTW17, HLGW13, LSQ11]. Jointly [ZWH^{+}22]. Joyful [Pac17]. JPEG [DLGT19, LSQ11, LC15]. Just [ABG^{+}11]. Just-in-Time [ABG^{+}11].

Kernel [LSZH18, MMDY15, SLM^{+}21]. Key [HG21, LLLC19, XWW^{+}21]. Key-point [HG21]. Keyword [CWCK15]. Keywords [ZWXZ12]. Kinect [CCZ^{+}15, KSL^{+}15, SZX15]. KinectFusion [FDE15]. Kinematic [LMAP16]. KNN [MLSK23]. KNN-Based [MLSK23]. Knockout [VS11]. Knots [KBM^{+}21]. Knowledge [CGZ23, DJI^{+}16, DGL^{+}22, GH18, HCWH22, JLW^{+}23, LXC^{+}21, LGZ^{+}21, LDFL23, LFWY23, Min16, OOD^{+}17, SRJP12, WYY^{+}19, ZFQX20, Edi13]. Knowledge-aware [ZFQX20]. Knowledge-Based [SRJP12]. Knowledge-Leverage-Based [DJI^{+}16]. KOMPOS [KBM^{+}21]. krypta [FPVC13].

L21 [LYLX23]. Label [BD23, FLLX18, LFL^{+}20, LFWY23, LFL^{+}23, RV18, SKF^{+}14, ZYC^{+}22, BLAK19, THY^{+}11, YJHL11]. Labeled [LWH12]. Labeling [ZSS^{+}15, TLWZ11]. Labels [HSJ^{+}22, JHK^{+}22, YDZ20]. Laboratory [DGJ^{+}23]. Land [XHZ^{+}23]. Landmark [GWDJ15, SSHL13]. Landmarking [WPA^{+}12]. Language [BLNN20, LHO23, SZX15, ZSAL20, ESNN13, ZZZ20a]. Laplacian [CSA^{+}23, LYLX23]. Large [ASW^{+}19, BWC15, CQZ^{+}12, GOB^{+}12, HDTG15, HKMN20, HYL^{+}21, KBM^{+}21, LJC^{+}11, PCC17, PCL18, TTLG17, WTK^{+}19, WFX^{+}21, FGP11, Hsu11, SSZ^{+}13]. Large-Scale [ASW^{+}19, CQZ^{+}12, HDTG15, HKMN20, LJC^{+}11, PCC17, PCL18, HYL^{+}21, WTK^{+}19, WFX^{+}21, FGP11, Hsu11]. laser [SZC^{+}13]. Lasso [XKW^{+}16]. Latent [BD23, GRR^{+}15, LT20, RGH19, Siz12, SZX15, YCGH12, LZCS11]. Latent-Force- [GRR^{+}15]. Layer [KLL17, ZCL^{+}18]. LBSNs [HYC^{+}16]. LDA [TRH16, YCL^{+}21]. Leaders [WMWR22]. Leakage [RDX22]. Learning [AdCK^{+}22, BC19, BQF^{+}23, BN21a, BLL^{+}14, BLNN20, CSA^{+}23, CCL15, CKP^{+}22, CYYL18, DGK^{+}22, DLWF22, DCWP22, DJI^{+}16, DGZ15, EBS^{+}22, EHG21, EMF12, FC15, FZ16, FMdA^{+}23, FLLX18, FL20, GJSC16, GHZ^{+}17, GBC^{+}22, GWL^{+}23, GG23, GYT19, GXT^{+}23, GXYZ21, GYL^{+}22, HY11, HAAM12, HCTC12, HYHFV22, HJTZ12, HCRF21, HXY^{+}22, HLL^{+}22, HCWH22, HWZL20, HLC^{+}21, HHL^{+}22, JV20, JSL^{+}19, JTP^{+}21, JCW^{+}22, JZG22, JHK^{+}22, JLH19, KLL22, KG23, LC12, LXBW20, LCCT12, LCKY14, LPM20, LXC^{+}21, LLS^{+}21, LNYV22, LHZ22a, LPL^{+}22, LBT23, LLL23, LXW^{+}23, LZC23, LWZ^{+}23, LWLG22, LLZW17, LCY^{+}18, LCLG19, LCC^{+}20, LGJ^{+}21, LHZ^{+}22b, LCY^{+}22, LRD^{+}22, LSZH18, LYKS23, LNO^{+}18, LFL^{+}20, LFWY23, LFL^{+}23, MSSV11, MKL11, MZY^{+}22, Min16, MNSB15, NVDMFD22, NZW^{+}17, NDW^{+}19, OSW^{+}22, PYD^{+}17, PHL^{+}20, PCF^{+}19, PKH^{+}17, PCC17, PS11, QTM23, RDX22, RYC22, RXL^{+}23, RV18, RAZE18, SC17, SLM^{+}21, SZT12, STA22, SY12, SWA23]. Learning [TJL^{+}21, TLB^{+}21, TZC^{+}20, TWL^{+}22, TMZ^{+}20, VKA^{+}19, WC12, WHC13, WZS^{+}15, WZH16, WJY^{+}18, WFZ^{+}18, WZYM19, WCF^{+}20, WZFL21, WLFY21, WGL^{+}22, WZFL22, WYC^{+}22, WLW^{+}23, WMH18, WYY^{+}19, WHRGC22, WLZ^{+}23, WL23, XLF^{+}20, XZXM19, XLL^{+}22, XZR12, XXZ^{+}21, XLZ^{+}22, XXL^{+}23, XW23, XSJW23, XHZ^{+}23, YLCT19, YTL^{+}22a, YTL^{+}22b, YWZ^{+}23, YDZ20, YCZY21, YCY23, YXL^{+}23, YHF21, ZYC^{+}22, ZYT^{+}15, ZAK^{+}23, ZCS^{+}12, ZSY^{+}12, ZY12, ZLZ^{+}17, ZGP^{+}18, ZHZ18, ZCWZ18, ZWZ^{+}19, ZZZ^{+}22, ZFLD22, ZLD^{+}23, ZGF^{+}23, ZKC^{+}23, ZX11, ZPL^{+}20, ZMH^{+}22, ZBW^{+}22, ZBZX12, ZJSY21, ZWX^{+}22, ZWZZ23, ZHLL21, ZCZ^{+}23, ZXY^{+}23, ZCL^{+}18, BSW^{+}13, BPS13, CBPG22, DSM^{+}11, ERR13, GPSB11, GXZ^{+}11, Hsu11, HLT11, KDC13, LHG11, Lin11, LGZ^{+}21, SSZ^{+}13, SZC^{+}13, WH11b, WHJ^{+}11, YCSH23, ZSAL20, ZLZ^{+}22]. Learning-based [HHL^{+}22, QTM23, ZAK^{+}23, HLT11]. Least [LHS18]. lecture [YSJ13]. Lending [ZGL^{+}17]. Length [WHRGC22, YMC16]. Lesion [SLC23]. Less [GB22]. Let [LBC^{+}22]. Letter [BI17]. Level [HSBRM22, LZY^{+}16, SGJC18, PG13]. Levels [ABB^{+}15]. Leverage [DJI^{+}16]. Leveraging [FXR^{+}17, LXW^{+}23, RHF16, TRDD12, YLY^{+}23, ZBZX12, ZZC^{+}20]. Lexical [HCB13]. libFM [Ren12]. Libraries [MGSK19]. Library [LS16, CL11]. LIBSVM [CL11]. Life [LM11]. Lifespan [FE15]. Lights [ZLBZ23]. Lightweight [ZDW19]. Like [LMAP16, SCC^{+}23, TWC^{+}23]. Limited [HSJ^{+}22, LWH12]. limiting [ZC13]. Linear [CZG^{+}23, WZZ^{+}21]. Lingual [PS11]. Linguistic [SZZ^{+}21]. Linguistic-Constrained [SZZ^{+}21]. Link [BTL20, GTM^{+}14, WZHL14, FTCP^{+}13]. Linkage [GSM23]. Linked [DOTD16]. Linking [HLY^{+}14, WLWJ21]. Links [JJ14]. List [UAS15]. Living [SMGMC^{+}15]. Load [TBW21, WCBL18]. Local [GQY^{+}19, HLL^{+}23, JYT^{+}12, JBLW21, JZG22, LTS^{+}15, LCLG19, SMX15, ZCS^{+}12, ZLY^{+}18, ZLD^{+}23, ZPP^{+}21, ZHLL21, ZCZ^{+}23]. Local-Global [ZHLL21]. Locality [EFMRK^{+}20]. Locality-sensitive [EFMRK^{+}20]. Localization [HKMN20, LXM^{+}18, SGD13]. Locally [ZNWC14]. Localness [HWT17]. LocateMe [SGD13]. Location [AC15, ABO17, CLH^{+}22, CYKL16, DYQ^{+}23, FXHM16, GBC^{+}22, HQY^{+}22, HCRF21, HLL14, JGL^{+}15, JWJC16, LX14, LBP19, MND14, SFX17, WMR17, YZQ16, ZC15, ZFWL17, ZWH17, CGMC11, YLT13]. Location- [AC15]. Location-Aware [ZFWL17]. Location-Based [CYKL16, HLL14, LBP19, WMR17, YZQ16, ZWH17, HQY^{+}22]. Location-Centered [GBC^{+}22]. Location-Specific [LBP19]. Locations [SGY^{+}22, FGP11]. Locks [HSBRM22]. Log [FZH^{+}21, LYLX23]. Log-Determinant [LYLX23]. Logic [ACC21, KAH12]. Logs [LJC^{+}11, TCK20, JPL13]. LONET [YSJ13]. Long [SLM^{+}23, TJL^{+}21, WZS^{+}15, WZS^{+}20, WCP^{+}23, ZPP^{+}21]. Long-Tail [WZS^{+}15]. Long-Tailed [SLM^{+}23]. Long-Term [TJL^{+}21, WZS^{+}20]. Longitudinal [WZS^{+}20]. Look [LPR19, SCC^{+}23]. looping [HY11]. Loss [EBS^{+}22, LFL^{+}20, SLM^{+}21]. Low [AWSF21, BL16, CSA^{+}23, HBK^{+}16, HCWH22, LCN^{+}16, LFL^{+}20, ZLZ15, ZMH^{+}15, HHJ22, SHZ13]. Low-Cost [BL16]. Low-Rank [HBK^{+}16, LCN^{+}16, LFL^{+}20, ZLZ15, ZMH^{+}15, CSA^{+}23, SHZ13]. Low-shot [HCWH22]. LSTM [RP23, SLZ^{+}23, ZCL^{+}21]. luminance [HHJ22]. Lunch [ZGF^{+}23].

Machine [GWD^{+}21, HYHFV22, HCWH22, LXJ^{+}20, NVDMFD22, PCF^{+}19, SY12, SZX15, YC12, YLCT19, ZWZ^{+}19, BSW^{+}13, BPS13, Hsu11, KDC13, Lin11, Mar13]. Machines [AKR^{+}18, CYYL18, Ren12, CL11]. Macrosociological [JCH14]. Magnetic [SGD13]. Magnetic-fields-based [SGD13]. Maj [BI17]. Make [LGJ^{+}22]. Making [Bha21, SBS^{+}23, ZZZ^{+}20b]. malicious [MSSV11]. Manage [CCL15]. Management [IVS^{+}16, NCG21, YMLM16, ZMH^{+}22, Edi13, PMSR11]. MaNIACS [PDR23]. manifesta [FPVC13]. Manifesto [Wid17]. Manifold [PKCC18, TZC^{+}20].

Manipulating [JTZ^{+}11]. Manipulation [SGTK20]. Manufacturing [CZKJ22]. Map [WH11a, ZZZ^{+}11, CCW^{+}19]. Map-Based [ZZZ^{+}11]. Mapper [AHJB20]. Mapping [HRCT16, MIS20, YZQ16, TLWZ11]. Maps [JTZ^{+}11, Goo10]. maps/crisis [Goo10]. Marked [GBBD22]. Market [JD15, NOZ20, YLWX20]. Marketplace [JPS^{+}16]. marketplaces [ZC13]. Markets [KAH12, Dha11]. Markov [BWC15, Gin13, LAsO^{+}19, LYW^{+}19, SRJP12]. Masking [KCSW23]. Massive [LZY^{+}16, WLWJ21]. Matchability [XYS^{+}23]. Matches [WCP^{+}23]. Matching [DCM15, LLF^{+}19, YLY^{+}23, ZNWC14, OSM^{+}13]. matchmaking [LKD13]. Mathematical [OY13]. Matrix [CWCY22, CWC^{+}20, CZJL15, HBK^{+}16, LCD17, LCD18, PKH^{+}17, TTLG17, ZCL^{+}21, ZMH^{+}15, SHZ13, SLH13]. Maximization [ACPS17, CZP^{+}14, YMC16]. MC [SZL^{+}23]. Mean [GQY^{+}19]. Meaning [KCSW23, ME13]. Meaning-Sensitive [KCSW23]. Meaningful [DC21]. Means [PKCC18]. Measurement [WZS^{+}20]. Measurements [MGS17a]. Measures [LAS20, XXL^{+}17]. Measuring [HLL14, SCLZ17, TRH16, ZZGH19]. Mechanical [BC19, LCN^{+}21]. Mechanism [LHS18, LCN^{+}21, LLX^{+}22, LDTX16, ZLD^{+}23]. Media [BTVY17, CCW^{+}19, GGC21, HWCL17, LCM^{+}12, LGL^{+}22, PT12, RHT^{+}18, Siz12, STP^{+}18, TY12, TY14, TLLS17, WZZ^{+}16, ZJSY21, ZMH^{+}15, CZLS13, Goo10, HCB13, LCCS13]. Mediated [ST20]. Mediation [TB22]. Medical [ABB^{+}15, CKP^{+}22, CYC^{+}23, GH18, HXC^{+}23, LFW23, LWWX20, MMDY15, SJCM23, HBSC13, KDC13]. medicine [WW13]. Meets [JGL^{+}15, LAS20, TWL^{+}22, WLC^{+}16]. Members [WMWR22]. Memory [CZJL15, MWY^{+}23, TL23]. Mental [XZH^{+}17]. mention [CST13]. MER [EBG^{+}12]. Merging [LWWX20]. Mesh [JLH19, YZL^{+}19]. Meta [CBPG22, DGL^{+}22, GWL^{+}23, LGZ^{+}21, ZLZ^{+}22, ABB^{+}15]. Meta-Learning [GWL^{+}23, CBPG22, LGZ^{+}21, ZLZ^{+}22]. MetaDetector [DGL^{+}22]. Metaphor [HNV14]. MetaStore [LGZ^{+}21]. Metastrategies [HDTG15]. Metering [WCBL18]. Method [CCH15, CZJL15, CYYL18, LWWL11, SP16, SRJP12, TTL^{+}21, ZYH^{+}17, RC13]. Methodologies [ZCWY14b]. Methods [BBM17, FSW^{+}20, LLS^{+}21, LCX^{+}23, LSZH18, NCG21, VDL^{+}19, WZYM19, WTL20]. Metric [BQF^{+}23, HCTC12, HJTZ12, LRD^{+}22, XZR12, ZCS^{+}12, ZY12, WHC13, WHJ^{+}11]. metrics [WLG11]. Metropolises [YZY^{+}17]. Microblog [CDW^{+}19, HWCL17, CCL13]. Microbloggers [HL17]. microblogging [KN13]. Microblogs [FXR^{+}17, GZZY17, LJL^{+}17, LYWW18]. microprocessor [CCG^{+}13]. Microscopic [RXK^{+}17]. Microtask [MIRS23]. Microtopic [LJL^{+}17]. Might [SCC^{+}23]. Mined [CGZ23]. Minimal [MJVL16]. Minimization [LYLX23]. Minimize [SLR^{+}16]. Minimizing [GJSC16, GJC17]. Mining [ASW^{+}19, BYK^{+}21, BMTT16, BCGJ11, CPHL15, CCC^{+}12, EL14, FE15, FZH^{+}21, HCCY15, JGL^{+}15, JPL13, LS16, LHJ^{+}11, LWH12, LLL^{+}16, LX14, LJC^{+}11, LYWW18, MMPS23, PFS17, PPPM18, PSRL12, PDR23, RAZE18, SMX15, SLH13, TBW21, WH18, WCS^{+}20, WYP22, XLZ21, YY15, YSN^{+}17, YLT13, YKTL14, ZSS^{+}15, ZZC12, Zhe15, ZCX^{+}15, BVCH13, BK11, BBGG13, MGB^{+}11, PMSR11, RC13]. Misinformation [LYWW18]. Missing [CDGZ16, DCM15]. Mitigation [AGP17, SQJ^{+}19]. Mixture [LAsO^{+}19, LLX^{+}20, LC16]. MKEL [SLM^{+}21]. Mobile [CHP17, CCK^{+}18, GWDJ15, GME17, JLX^{+}17, LZY^{+}16, NZW^{+}17, SFX17, WMR17, XZW^{+}15, YLD^{+}22, YWZ^{+}17, ZS18, ZFWL17, ZDW19, ZCX^{+}15, BGMS13a, CKS10, Edi13]. Mobility [ABO17, BZX^{+}22, FGL17, HMG^{+}23, HS19, LCLN18, PPPM18, SLM^{+}23, SZS^{+}17, TJL^{+}21, WFZ^{+}18, WZS^{+}20, WLWJ21, WYC^{+}22, YCP^{+}13, ZYW^{+}15, ZHZ18]. Mobility-on-Demand [HS19]. Modal [GYT19, ZWC23, ZZS^{+}21, GG23, ZLC^{+}20]. Modalities [WYZ23]. Modality [WZZ^{+}16]. Modality-Dependent [WZZ^{+}16]. Mode [RV18, SZT12]. Model [CRRH11, DLY^{+}21, EL14, GLL^{+}17, GRR^{+}15, HMG^{+}23, HXC^{+}23, HYC^{+}16, HLL^{+}23, HLNL18, HTL^{+}20, JWJC16, KP17, LLL21, LAsO^{+}19, LWH^{+}20, LGL^{+}22, LYZ^{+}23, LLX^{+}20, LC16, LLZW17, LGZ^{+}21, MGJW20, MFI19, NYBG17, Pai16, PCC17, QHH^{+}21, RP23, RAK23, SRJP12, SWZ^{+}21, WYC^{+}17, WLF^{+}18, WYD^{+}18, XXZ^{+}21, YCZY21, YLWX20, ZYSL12, ZWZS16, ZCWZ18, ZLD^{+}23, ZLL^{+}22, CDS13, CZLS13, HLGW13, HLT11, ME13]. Model-Based [EL14, GRR^{+}15]. Modeling [AC15, CCL15, CHHH18, CASR22, DWKP16, DJI^{+}16, DCF^{+}18, FZH^{+}21, GRR^{+}15, GOB^{+}12, GZ21, GBBD22, GZ23, HL17, JTS^{+}21, JLJ^{+}20, LZ18, LYWH20, RGC^{+}22, SZX15, TJL^{+}21, WZCJ21, WLW^{+}22, WHW^{+}21, XYS^{+}23, YLD^{+}22, YCGH12, ZC15, ZHZ18, ZPP^{+}21, LN10, YNS13, ZT11, ZC13]. Modelling [LWC^{+}18]. Models [CMPR21, EK15, GST12, HM19, KW17, LH12, MNSB15, PFS17, QCZ^{+}21, RGH19, SLR^{+}16, XXL^{+}23, ZSS^{+}15, ZWZS16, ZSAL20, ZWH^{+}22, Zhu19, ZZKT20, Bai10, FGP11, Gin13, HLJ11, LHS^{+}13]. Modification [CHY15]. modified [CLSL13]. Moment [DYQ^{+}23, TZC^{+}20]. Moment-Guided [TZC^{+}20]. Mondrian [DL13]. Monitoring [AKR^{+}18, MGB^{+}11, MMC^{+}13, NTM^{+}16, SBS^{+}23, VDL^{+}19, WYM17, WCBL18, RC13]. Monolingual [LWWL11, RBK^{+}13]. Monotonic [LH22]. MOOCs [JLW^{+}23]. Mood [KAH^{+}16]. Motion [SRJP12, YZL^{+}19, ZZZ^{+}22, LMWS13]. Motion-Aware [YZL^{+}19]. mouth [ZT11]. Movement [CCK^{+}18, WWZ^{+}16, WLC^{+}20, LHJ^{+}11]. Movements [LMAP16]. MoveMine [LHJ^{+}11]. Movie [DBDM16, SBD13, LHZ13, SLH13]. Moving [HCJM15, Pat15, LHJ^{+}11]. Moving-Object [HCJM15]. MPR [AHJB20]. MS [KSL^{+}15]. Multi [BLAK19, CKP^{+}22, CWCK15, CWC^{+}20, CYC^{+}23, CDW^{+}21, DPB20, DCWP22, DLGT19, DLLT21, FLY^{+}23, FWZ17, FLLX18, GBC^{+}22, GG23, GDC19, GYT19, HXC^{+}23, HWZL20, KLL17, KW17, LHZ22a, LFW23, LLL^{+}18, LGZ^{+}21, LHZ^{+}22b, LYLX23, MFI19, NOZ20, PCL18, RV18, SKF^{+}14, ST20, SP16, SS11, TTL^{+}21, WJY^{+}18, WZCJ21, WZM^{+}22, WYNW20, XLF^{+}20, XXZ^{+}21, XLB23, YHF21, ZAK^{+}23, ZHZ18, ZWZ^{+}19, ZZS^{+}21, ZPL^{+}20, ZLC^{+}20, Zhu19]. Multi-Agent [GDC19, SS11, Zhu19, HXC^{+}23, NOZ20, ZAK^{+}23]. Multi-aspect [LFW23]. Multi-Attribute [XLB23]. Multi-Auxiliary [CWC^{+}20]. Multi-Category [SP16]. Multi-channel [DPB20]. Multi-city [LGZ^{+}21]. Multi-Click [CWCK15]. Multi-Domain [MFI19]. Multi-Factor [PCL18, LLL^{+}18]. Multi-Graph [YHF21]. Multi-Hypergraph [FWZ17]. Multi-Keyword [CWCK15]. Multi-Label [FLLX18, RV18, SKF^{+}14, BLAK19]. Multi-Layer [KLL17]. Multi-Modal [GYT19, ZZS^{+}21, GG23, ZLC^{+}20]. Multi-Objective [SKF^{+}14]. Multi-Party [ST20]. Multi-relational [CYC^{+}23, FLY^{+}23]. Multi-scale [DLGT19, WZM^{+}22]. Multi-Source [ZZS^{+}21, WYNW20]. Multi-Stage [ZZS^{+}21]. Multi-target [TTL^{+}21]. Multi-Task [GBC^{+}22, HWZL20, LHZ22a, XLF^{+}20, ZPL^{+}20, DLLT21, LHZ^{+}22b, XXZ^{+}21]. Multi-Threaded [KW17]. Multi-Tier [DCWP22]. Multi-Type [WZCJ21]. Multi-View [CDW^{+}21, FLLX18, LYLX23, WJY^{+}18, ZHZ18, ZWZ^{+}19, CKP^{+}22]. Multiagent [CGZ18, DPC16, JD15, BNS13, FS13, ZC13]. Multiagent-Based [CGZ18]. Multiclass [YCL^{+}21]. Multicontext [OLY^{+}17]. Multidimensional [ACC21]. Multiexpert [SDHS15]. Multifeature [GWDJ15, LGZ^{+}17]. Multifocal [GXZ^{+}11]. Multigroup [HMCW15]. Multilabel [JLL18]. Multimedia [HTDJ12, JGL^{+}15, JLX^{+}17, LTW^{+}16, NZW^{+}17, PCL18, SSL^{+}18, ZLC^{+}20, BK11, HTDJ11, WH11b]. Multimodal [FTE21, HYL^{+}21, SC22, YL17, ZYH^{+}20]. Multimodular [SDD^{+}16]. Multiobjective [RZS^{+}15]. Multiobjects [WXZ^{+}16]. multipartite [SLWW13]. Multiperson [WYM17]. Multiple [ARGK15, GJS23, JV20, LXJ^{+}20, MZL12, SLM^{+}21, SGTK20, SZL^{+}23, WLWJ21, WYZ23, ZS18, ZCG15, ZRX^{+}22, ZCZ^{+}23, LMWS13, SZC^{+}13]. Multiple-Choice [LXJ^{+}20]. Multiresolution [CDS12, DTL15]. Multitask [LCN^{+}16, SJCM23]. Multitechnique [BMV13]. Multivariate [WC12, WZM^{+}22]. Multiview [AWSF21, SSL^{+}18, XTW17, ZYT^{+}15, ZCS^{+}12]. Music [OOD^{+}17, SC17, SYHB17, SR17, TS17, Wid17, YC12, ZSLC19]. Music-Related [SR17]. Mutual [LGL^{+}16, WXZ^{+}16]. MVGAN [CDW^{+}21]. Myerson [SRMW19]. MySpace [PT12].

n [MTC^{+}20]. Named

[KXZG15, LWZZ13, MPS23]. Naming [LX14]. Narrative [PCC10]. Native [WWL^{+}22]. Natural [LHO23, SZS^{+}17, ZSAL20]. Navigate [LXBW20]. navigation [WHR13]. NEAR [KMH22]. Nearest [GQY^{+}19]. Negative [JJ14, YCSH23, CLSL13, TTLG17]. Negotiating [GKG^{+}11]. Negotiation [AKA^{+}21]. Neighbor [GQY^{+}19, YLY^{+}23]. Neighborhood [KMH22, LYWH20, MLSK23]. Neighboring [LSQ11, LC15]. Neighbourly [GDC19]. Nested [MFI19]. Net [CYC^{+}21, HLF^{+}21]. Network [ASK^{+}21, BBM17, BZW^{+}22, BCGJ11, BN21b, CLL^{+}21, CZKJ22, CYC^{+}21, CDW^{+}21, DLLT21, FLY^{+}23, GWL^{+}23, GTM^{+}14, GXT^{+}23, GXYZ21, HLF^{+}21, HLH^{+}21, JLH19, LS16, LCN^{+}21, LXC^{+}21, LFY^{+}22, LWZ^{+}23, LSW23, LWWX20, LYF^{+}22, LLX^{+}22, LGJ^{+}22, LSW^{+}20, MGS17a, MWY^{+}23, MFI19, OLY^{+}17, PEK^{+}16, QHH^{+}21, RDX22, SC17, SJCM23, SZL^{+}23, TL23, VKA^{+}19, WXZ^{+}16, WLL^{+}21, YLH^{+}23, ZZS^{+}21, ZLZW23, ZBW^{+}22, ZLG^{+}20, ZJSY21, ZZC^{+}22, MPA13, YSJ13]. Network-Based [TL23]. Network-Oblivious [BBM17]. Networked [SZT12]. Networking [LWH^{+}20, ZWGW17]. Networks [ABO17, BZX^{+}22, CZG^{+}23, CZP^{+}14, CHP17, CYKL16, CCWS17, CCGP22, DLY^{+}21, DCWP22, DLGT19, DCM15, EBS^{+}22, FLY^{+}23, FL20, GXS^{+}22, GMX^{+}21, GME17, HQY^{+}22, HS19, HTSC^{+}17, HHJ22, JJ14, JJKZ22, LCKY14, LLS^{+}22, LBP19, LFW23, LCJ^{+}19, LXJ^{+}20, MGJW20, RAK23, SC22, SRB15, SLC23, TLW^{+}15, VNL^{+}11, WAL18, WTL20, WZS^{+}20, WLFY21, WZM^{+}22, WMWR22, WLH17, WL23, WCFH23, XWW^{+}21, YL14, YZQ16, YFJ^{+}18, ZL19, ZZL^{+}23, ZDW19, ZZH^{+}22, BVK10, CBP13, FTCP^{+}13, HKO13, SKOM13, SLWW13, WCBK11]. Neural [BZW^{+}22, CLL^{+}21, CZKJ22, CLL23, CP23, DMGSD23, DLLT21, EBS^{+}22, GMX^{+}21, GZ23, JJKZ22, JLH19, LWH^{+}20, LLS^{+}22, LCJ^{+}19, LYF^{+}22, LLX^{+}22, MGJW20, QHH^{+}21, RDX22, RAK23, SC17, TLW^{+}15, WAL18, WTL20, WL23, ZLZW23, ZBW^{+}22, ZDW19, ZLG^{+}20, ZZH^{+}22, ZZC^{+}22]. Neutrality [MGS17a]. News [CQZ^{+}12, DGL^{+}22, GW17, HNA20, LTW^{+}16, LGL^{+}22, SQJ^{+}19]. Newton [CYYL18, WTL20]. Next [SCC^{+}23, TCK20]. Neyman [GPSB11]. Night [LGJ^{+}21]. NN [EFMRK^{+}20, THY^{+}11, ZLZ^{+}17]. NN-sparse [THY^{+}11]. No [ZGF^{+}23]. NoC [TAL^{+}19]. Node [JLJ^{+}20]. Nodes [ASK^{+}21]. Noise [DWKP16, JLL18, MJVL16, QCZ^{+}21, ZLD^{+}23]. Noise-aware [ZLD^{+}23]. Noise-Minimal [MJVL16]. Noise-Resilient [JLL18]. noisily [THY^{+}11]. Noisy [DGZ15, KXZG15]. Non [FNS16, KBM^{+}21, LH22, NVDMFD22, OSW^{+}22, RYC22, TTLG17]. Non-ID [NVDMFD22]. Non-iid [FNS16, OSW^{+}22]. Non-Monotonic [LH22]. Non-negative [TTLG17]. Non-overlapping [RYC22]. Non-Parametric [KBM^{+}21]. nonconvex [GPSB11]. Nonhomogeneous [LAsO^{+}19]. Nonignorable [CDGZ16]. Nonintrusive [WCBL18]. Nonlinear [BC19, KBM^{+}21, QCZ^{+}21, WMH18, ZWZS16]. Nonnegative [DTL15, LCD17, LCD18, PKH^{+}17]. Nonparametric [XLZ21, LHG11]. Nontrivial [SSHL13]. Normalization [SDD^{+}16, HCB13]. Norms [AKA^{+}21, Sin13]. Nova [AKA^{+}21]. Novel [BTL20, DWKP16, DLLT21, JJ15, KKG18, KLL17, KLLL20, MG16, WSGM14]. Novelty [BBS^{+}16, CRYT12]. Nowcasting [LC12].

OARF [HLL^{+}22]. obesity [MPA13].

Obfuscating [YWS^{+}23]. Object [HCJM15, HHJ22, LZW^{+}23, MWY^{+}23, SST^{+}15, TLC^{+}14, WLL^{+}21, WLL^{+}22, YLX^{+}20, YCZ^{+}23, ZZL^{+}19, LHJ^{+}11, LHS^{+}13]. Object-Attention [WLL^{+}21].

Object-Oriented [TLC^{+}14]. Objective [SKF^{+}14]. Objects [CCL15]. Oblivious [BBM17]. Observable [JD15]. Observational [Hec19, LLL^{+}16, LLX^{+}20]. Observations [GBBD22]. Obtained [CCL15]. Ode [Pac17]. Odometry [FDE15]. Off [ZKC^{+}23]. offense [TEP11]. Offline [HL19, ZLH18]. Offloading [NDW^{+}19]. Ohmage [THL^{+}15]. On-board [BCR21]. On-Device [GWDJ15]. Onboard [MRW^{+}12]. One [AKR^{+}18]. One-Class [AKR^{+}18]. ONION [ZWC23]. Online [AKR^{+}18, BWC15, BRSG20, BLL^{+}14, BR15, DGK^{+}22, FE15, GHZ^{+}17, HTSC^{+}17, HL19, HWT17, HYL^{+}18, LSZH18, MGS17b, RSCOVCMM17, SA15, SLR^{+}16, WXLY12, WYY^{+}19, WLZ^{+}23, ZWL^{+}15, ZLH18, ZWC23, ZZZ^{+}20b, GPSB11, SSZ^{+}13, SZC^{+}13, ZDC^{+}13]. ontological [KDC13]. Ontologies [CYC^{+}23]. Ontology [ZLBZ23]. Ontology-Based [ZLBZ23]. Open [DOTD16, GDC19, SSLM21, ZZC^{+}20, TBK^{+}10]. Operational [SGJC18]. operations [BKB10, RFI^{+}11]. Opinion [HCCY15, WH10, ZT11]. Opportunity [EBG^{+}12]. Optimal [GJC17, LGZ^{+}21, POM20]. Optimization [AHJB20, BB15, BLL^{+}14, GXT^{+}23, GXYZ21, HSBRM22, HKO13, ODF17, SGJC18, SKF^{+}14, VKLY18, XLL^{+}22, ZLBZ23, FLF^{+}20]. Optimization-based [HKO13]. Optimize [XXL^{+}17]. Optimized [HWCL17, KG11]. Optimizer [HJCK20, XLB23]. Optimizing [GG15, HTSC^{+}17, ZWH^{+}22]. Optimum [ZS18]. Option [CWCK15]. Orchestration [Pac17]. Ordering [TRH12]. Ordinal [TZC^{+}20]. Organization [ZL19]. Oriented [LXJ^{+}20, TLC^{+}14]. Orienteering [VKLY18]. Origin [CLH^{+}22]. Origin-Aware [CLH^{+}22]. Orthogonal [LGZ^{+}17, LCD18]. Other [WLZ^{+}23]. Outcome [CLBM15, SDXG16]. Outcomes [CDGZ16]. Outdoor [WYG^{+}22]. Outfits [BRSG20]. Outlier [MLJZ21]. Overcoming [GB22]. Overlapping [GGY^{+}23, JLJ^{+}20, RYC22]. Overlaps [YL14]. Overview [ZGP^{+}18, Zhe15]. Own [WLZ^{+}23].

P2P [ZGL^{+}17]. Package

[TLC^{+}14, WLW^{+}22, WLW^{+}23]. Page [ABG^{+}11, HDPH16, TRDD12]. Pages [ZWXZ12]. Paid [FS17, MIS20, MIRS23]. Paint [WM21]. Pair [LCC^{+}20, ZNYH11]. pair-activities [ZNYH11]. Pair-based [LCC^{+}20]. Paradigm [HAAM12]. Parallel [CDS12, CZJL15, DLY^{+}21, LZCS11, ZWH17, ZCL^{+}21]. Parameter [LLS^{+}21, XW23]. Parameterized [Pai16]. Parameters [MCEG23]. Parametric [KBM^{+}21]. Paraphrase [BPS13, BMV13, Mar13, ME13]. paraphrases [BMV13, MD13]. paraphrasing [RBK^{+}13, WDSZ13]. Pareto [RZS^{+}15]. Pareto-Efficient [RZS^{+}15]. Paris [EL14]. Parkinsonian [WWZ^{+}16]. PARP [DLY^{+}21]. Part [HTDJ12, ADM^{+}21, WZFL21, WZFL22, YTL^{+}22a, YTL^{+}22b, ZLSY22a, ZLSY22b]. Partial [LHS18, LFL^{+}20, LFWY23, ZCZ^{+}23]. Partially [JD15, TRDD12]. Participant [LCY^{+}22, ZFLD22]. Participants [XZW^{+}15]. Participatory [CTY^{+}19, GCY^{+}15, THL^{+}15, YZQ16, ZSL^{+}15]. Particle [HJCK20]. Partitioning [DCWP22, JBLW21]. Partners [RKH14]. parts [TDVC13]. parts-based [TDVC13]. Party [ST20]. Passenger [DCF^{+}18, WYC^{+}22]. Passive [LSZH18]. Passive-Aggressive [LSZH18]. Password [ZZD^{+}17]. Past [ZRX^{+}22]. Paste [ZPL^{+}20]. Path [DOTD16, YSJ13]. Path-Based [DOTD16]. Paths [LCLG19, MNSB15]. Patient [KXZG15, HBSC13, KDC13, LMWS13]. Patient-Related [KXZG15]. Patients [CCL15]. Pattern [CPHL15, DCM15, WWZ^{+}16, WH18, ZWZZ23, BVCH13, WPL13]. pattern-aware [WPL13]. Patterns [ABO17, BYK^{+}21, CCW^{+}19, FE15, LCLN18, LXM^{+}18, PDR23, WCS^{+}20, WZS^{+}20, WYP22, ZZC12, LHJ^{+}11, RC13, SKOM13, YLT13]. Pay [LDTX16, WYP22]. Payment [WYP22]. PC [SSV19]. Peacock [WZS^{+}15]. Pearson [GPSB11]. Pedestrian [KF18]. Peer [RFJ16]. Peers [CCL15, DPSS19]. People [ARGK15, GKG^{+}11, HWT17, MZL12, CXW^{+}13]. Percentile [MBM21]. Perception [LGJ^{+}21, TRH16]. Perceptual [FZ16]. Percolation [GGY^{+}23]. Perform [RV18]. Performance [CVCL22, DGK^{+}22, HBL16, LZC23, LWC^{+}18, LZW^{+}23, PCF^{+}19, WLG11, YLC^{+}19, ZFLD22, KG23]. Performing [RP23]. Periodic [WFZ^{+}18]. Periodicity [TJL^{+}21]. Person [DMGSD23, LCC^{+}20, TTL^{+}21]. Personalized [BGPYS11, CFG13, CDS13, CSN^{+}17, ESNN13, FXHM16, GWL^{+}23, GYL^{+}22, HLNL18, LJL^{+}17, LX14, LCY^{+}15, LLL^{+}18, WYD^{+}18, XLF^{+}20, ZNWC14, ZFS^{+}19, ZCX^{+}15]. Personnel [YTH17]. Perspective [LLL^{+}18, LWF^{+}23, VPD^{+}22, WFZ^{+}18, YLC^{+}19, ZYW^{+}15, ZLH18, ERR13]. Perspectives [CDLV13]. Perturbation [LLS^{+}21]. Pervasive [SRM^{+}13]. Pests [FMdA^{+}23]. Pets [PCL18]. PG [LLS^{+}21]. PGNN [EBS^{+}22]. PhC [CDS12]. Phenotype [UAS15]. Phenotype-Gene [UAS15]. Phones [GME17]. photo [WHJ^{+}11]. Photography [WST^{+}15]. Photos [ZZC12, SSHL13, YJHL11]. phrasal [Mar13]. Physical [MWS^{+}18, PEK^{+}16, ZDW19, CXW^{+}13, TAL^{+}19]. Physics [EBS^{+}22, HSJ^{+}22]. Physics-Guided [HSJ^{+}22, EBS^{+}22]. Pick [WLW^{+}22, WLW^{+}23]. Pick-up [WLW^{+}22, WLW^{+}23]. Pilot [YMLM16]. pinpointing [BMV13]. Pipeline [HCRF21, LZCS11]. Placement [LGZ^{+}21, LZCS11]. Plan [MGS17b, MGSK19, POM20]. planner [BKB10, TNSP13]. Planning [BWC15, BVK10, CLL^{+}21, DLY^{+}21, KL23, MP23, RFI^{+}11, TBK^{+}10, ZZZ^{+}22, Che10, CKS10, LHC^{+}13, PCC10]. Plans [Zhu19, ZZKT20]. plant [CGMC11]. Platform [LGL^{+}22, THL^{+}15, Goo10, WWL^{+}22]. Platforms [HHL^{+}22, LWH^{+}20, ZGL^{+}17, ZT11]. Play [SSG^{+}20, LCY^{+}15]. Player [PCF^{+}19, TLWZ11]. PlayeRank [PCF^{+}19]. Playlist [ZSLC19]. PLDA [LZCS11]. POI [CZW^{+}20, CSN^{+}17, GB22]. Point [CYKL16, GBBD22, LXW^{+}23, SCLZ17, WZZ^{+}21, YKTL14, HG21]. Point-of-Interest [CYKL16, YKTL14]. Point-of-Interests [LXW^{+}23]. Points [HQY^{+}22, MIS20, OSW^{+}22, YGU15, LPM20]. Pointwise [LPM20]. POIs [HYC^{+}16, XHZ^{+}23]. Poisoning [CBPG22]. Policies [SDHS15, WL23]. Policy [LLS^{+}21, SBS^{+}23, XSJW23]. POLLA [ZPP^{+}21]. Polyline [JHK^{+}22]. POMDP [FKSS13]. Popularities [JLJ^{+}20]. Popularity [CRYT12, JJ15]. Population [FE15, SHX^{+}23]. Portfolio [HYL^{+}18, LHG11]. Pose [DT16, SS15, WGF^{+}23]. Pose-Invariant [DT16, WGF^{+}23]. Position [WST^{+}15]. Positive [JJ14]. Possible [SS22]. Post [ZWZS16]. Post-Nonlinear [ZWZS16]. Posts [GW17]. Posture [TLW^{+}15]. Potential [CHP17]. Power [SLZ^{+}23, CGMC11]. Powered [SNL^{+}16]. PP [LLS^{+}21]. PP-PG [LLS^{+}21]. PPLib [DB16]. Practical [CZW^{+}20, GME17]. Pre [TWL^{+}22]. Pre-training [TWL^{+}22]. Precision [LXM^{+}18]. Predict [LH12]. Predicting [ASK^{+}21, BMTT16, JCW^{+}22, KZL^{+}21, SGY^{+}22, TJL^{+}21, XLF^{+}20, YLC^{+}19]. Prediction [CMR15, CLH^{+}22, DGK^{+}22, DYQ^{+}23, Dha11, DLLT21, GBC^{+}22, GTM^{+}14, HLH^{+}21, JJ14, JWJC16, JD15, LLL21, LWH^{+}20, LFY^{+}22, LLS^{+}22, LWLG22, LLL^{+}18, LLX^{+}22, LZY^{+}16, MNSB15, RGC^{+}22, SC22, SZS^{+}17, SJCM23, SCC^{+}23, TCK20, TC19, VDL^{+}19, WZM^{+}22, WYC^{+}22, WYZ23, WLW^{+}22, WWL^{+}22, YLH^{+}23, YXL^{+}23, ZFH^{+}22, ZLZ^{+}22, ZYY^{+}23, ZCL^{+}21, ZRX^{+}22, BCC^{+}13, BSW^{+}13, FTCP^{+}13, LMWS13, YNS13, YLT13]. Predictive [SMGMC^{+}15, WW13]. Preemption [DPC16]. Preface [Che10, YTL^{+}22b, ZLB^{+}16]. Preference [BLL^{+}14, LPM20, LHZ22a, PHL^{+}20]. Preferences [LCKY14, TL23, WLW^{+}22, ZCX^{+}15, GJ13, RCN10]. Preferred [BRSG20]. Prerequisite [CLBM15]. Presence [YGY^{+}23]. Presentation [CCL15]. Preservation [GLJ^{+}14, YCH^{+}22]. Preserving [CZW^{+}20, LLL23, GGY^{+}23, TRZ^{+}19, TB22, ZLL^{+}22, ZZC^{+}22]. Price [GBC^{+}22, ZZC^{+}20]. Prices [LPR19]. Pricing [HS19, HHL^{+}22, JD15, WFX^{+}21]. Pricing-aware [WFX^{+}21]. Principles [TS17]. Prior [LFWY23, XSJW23, ZZL^{+}19]. Prioritization [PSRL12]. PRISM [TLLS17]. Privacy [CZW^{+}20, GSM23, GGY^{+}23, GLJ^{+}14, HTL^{+}20, LLL23, PPPM18, TRZ^{+}19, TB22, TC19, XLZ^{+}22, YCH^{+}22, ZZD^{+}17, ZKC^{+}23, ZLL^{+}22, ZZC^{+}22, HBSC13, WCBK11]. Privacy-aware [TC19, XLZ^{+}22]. Privacy-Preserving [LLL23, GGY^{+}23, TRZ^{+}19, TB22, ZLL^{+}22, ZZC^{+}22]. privacy-sensitive [WCBK11]. Private [CKP^{+}22, JZG22, MLSK23, WYZ23]. Probabilistic [BK11, HLJ11, LCKY14, LGL^{+}22, PG13, WLF^{+}18, FGP11]. Problem [LH22, WLWC23, ZSY^{+}12, GXZ^{+}11]. Problem-Solving [ZSY^{+}12]. Problems [EBS^{+}22, ODF17, SSS11, SDS12, VKLY18, WZZ^{+}21]. Process [DB16, FZH^{+}21, LGZ^{+}17, LCJ^{+}19, MMPS23, MMS17, TCK20, TY14, VDL^{+}19]. Processes [BWC15, FNS16, GBBD22, LZ18]. Processing [BTVY17, ZSAL20, DL13, LZCS11]. Products [HMS^{+}14, ZWL^{+}19]. Prof. [BI17]. Profession [TLLS17]. Profile [ZS18]. Profiles [NTM^{+}16]. Profiling [FGL17, LLZ^{+}23, TWL11]. Profit [ACPS17]. Programs [DB16]. Progressive [HASS22, LGL^{+}22]. Projection [PKCC18, ZLZ15]. Promoting [BBS^{+}16, LCC^{+}20]. Propagation [PEK^{+}16, THY^{+}11, YJHL11]. Proposal [AdCK^{+}22, ZHLL21]. Prospects [ZGL^{+}17]. Protection [HTL^{+}20]. protocol [GS13]. Protocols [ST20, BBMP13]. Providers [BSRSS16]. proximity [CXW^{+}13]. Pruning [LNO^{+}18]. PSDF [XLZ^{+}22]. Pseudo [TPM23]. Psychological [XZR12]. PTIME [BGPYS11]. PU [YCSH23]. PU-learning [YCSH23]. Public [CLL^{+}21, LH22]. Publication [GLJ^{+}14, YCH^{+}22]. Published [LLL23]. Pulse [WMR17, ZYW^{+}15]. Purchase [DSB^{+}18, GG15]. Purpose [HB12, LS16]. Purposes [WLF^{+}18].

QoI [ZSL^{+}15]. QoI-Aware [ZSL^{+}15]. Qrowdsmith [MIRS23]. qualitative [FK13]. Quality [HDPH16]. Quantification [LBP19]. Quantitative [SWA23]. Quantization [GZ21, MZY^{+}22, RAK23]. Quantized [CSHL21]. Queries [CHHH18, LLY12, XLB23]. Query [AC15, LJC^{+}11, XLG^{+}23, XLB23, BGMS13b]. Query-Aware [AC15]. Query-Efficient [XLG^{+}23]. Querying [YBZ^{+}20]. Question [GH18, TPG^{+}19, WJY^{+}18]. Questions [RHT^{+}18, TPG^{+}19]. Quick [HLNL18].

Random [CST13, LWZ^{+}23, WAL18, YKTL14, CLSL13]. Random-Forest-Inspired [WAL18]. Range [XLB23]. Rank [AWSF21, DGZ15, HBK^{+}16, LCN^{+}16, LZC23, LFL^{+}20, ODL^{+}20, ZLZ15, ZBZX12, ZMH^{+}15, CSA^{+}23, SHZ13]. Ranked [UAS15]. Ranking [DOTD16, HWZL20, KCTT16, LNO^{+}18, PYD^{+}17, PCF^{+}19, RP23, TY12, WSGM14, WMH18, CDK^{+}13, LHZ13]. Ranking-Based [WSGM14, RP23]. Rapid [BL16, LMAP16]. Rate [SSV19, WL23]. Rating [DBDM16, GW17, VPD^{+}22, ERR13]. ratings [CLSL13, ZDC^{+}13]. Rationale [ZLZW23]. RCMC [ABO17]. Re [DMGSD23, LCC^{+}20, TTL^{+}21]. Re-Identification [DMGSD23, LCC^{+}20, TTL^{+}21]. reaching [BD11]. Reactions [YY15]. Reactor [DWKP16]. reading [ESNN13]. Real [BB15, BFC^{+}17, FSS15, GXT^{+}23, HLNL18, LM11, MWY^{+}23, TLW^{+}15, TTFS18, WFX^{+}21, WCFH23, ZLT15, ZLY^{+}18, ZHZ18, BKB10]. Real-Life [LM11]. Real-Time [BB15, MWY^{+}23, TLW^{+}15, TTFS18, ZLT15, ZLY^{+}18, ZHZ18, GXT^{+}23, WFX^{+}21, BKB10]. Real-Use [HLNL18]. Real-World [FSS15, WCFH23]. Reality [KSL^{+}15, ZZD^{+}17]. Reason [ZFS^{+}19]. Reasoning [CXW^{+}19, FK13]. Recalibrated [SJCM23]. recency [CDK^{+}13]. Recognition [BDP12, DT16, EHG21, GWDJ15, GYL^{+}22, HRCT16, JGL^{+}15, JTP^{+}21, KAH^{+}16, KHNB15, KXZG15, LGL^{+}16, LTS^{+}15, LLZW17, MGS17b, MGSK19, MPS23, OLY^{+}17, QWC^{+}23, SP16, SRJP12, SS11, SZX15, TLW^{+}15, WGF^{+}23, XTW17, YC12, ZGP^{+}18, ZZZ20a, ZCG15, HLJ11, LWZZ13, WLG11, ZPY11]. Recognizing [ABO17, WWZ^{+}16, ZNYH11, Zhu19]. Recombination [DB16]. recommend [MKL11]. Recommendation [BBS^{+}16, BSRSS16, CZW^{+}20, CWC^{+}20, CYKL16, CSN^{+}17, DPSS19, EK15, FXHM16, FSS15, GB22, HLL^{+}23, HMCW15, HWZL20, JLW^{+}23, LJL^{+}17, LHS18, LHZ22a, LCX^{+}23, LZC23, LM11, LYF^{+}22, LAS20, OOD^{+}17, SMX15, TLC^{+}14, TC19, VNL^{+}11, WST^{+}15, WYC^{+}17, WYD^{+}18, WYG^{+}22, XYS^{+}23, YCSH23, ZWXZ12, ZZL^{+}23, ZZZ^{+}11, ZFS^{+}19, ZWX^{+}22, ZCX^{+}15, BGMS13a, CFG13, GJ13, HBSC13, LHZ13, SBD13, SLH13, SSHL13]. Recommendations [ADM^{+}21, DOTD16, DCF^{+}18, JJ15, KSKÇ15, LT20, LBT23, MG16, MLSK23, WLC^{+}16, YKTL14, ZC15, ZNWC14, ZDC^{+}13, ZX11]. Recommender [ADM^{+}21, AT15, CWLZ15, CVCL22, CSTZ16, CYW^{+}21, DJS16, MBM21, ODL^{+}20, RZS^{+}15, TL23, GCZ13, QSRGDAJD13]. recommenders [BCD^{+}13]. Recommending [BRSG20, CDR19, HLL14]. Reconstruction [YLY^{+}19]. Reconstructions [FDE15]. Record [XZH^{+}17, BVCH13]. Records [DLWF22, TEP11]. Recovery [CCZ^{+}15, ZLZ15, ZMH^{+}15]. RecRules [CDR19]. RecSys [ZSLC19]. Recurrent [YBZ^{+}20, YFJ^{+}18, YLH^{+}23]. Recurring [LWH12]. Recursive [MGJW20]. RECYCLE [HY11]. Redesign [HMS^{+}14]. Reduce [HSBRM22]. Reducing [BD11]. Reduction [CSA^{+}23, CSN^{+}17, XTW17]. Redundant [LFL^{+}23]. Referring [ZHW^{+}21]. Refined [LCD17]. Refined-Graph [LCD17]. refinements [GS13]. Refining [ZXY^{+}23]. Refueling [ZYW^{+}15]. Region [CCZ^{+}15, GB22, LWWX20, SS22, ZNWC14, CGMC11]. Region-Adaptive [CCZ^{+}15]. Regions [LFY^{+}22]. registration [TLWZ11]. Regression [HYHFV22, KBM^{+}21, LHS18, RDX22]. Regret [ZZZ^{+}20b]. Regularization [LCD17, LYLX23, LFL^{+}20, WZHL14, SZC11]. Regularization-Based [LCD17]. Regularized [LYLX23, XTW17]. regulative [BBMP13]. Reinforced [HXC^{+}23, JLW^{+}23, YCSH23]. Reinforcement [BC19, GJSC16, GXT^{+}23, GXYZ21, HHL^{+}22, LLS^{+}21, LYKS23, NDW^{+}19, QTM23, SWA23, TJL^{+}21, WXZ^{+}16, XW23, ZAK^{+}23, ZMH^{+}22]. Reinforcement-Learning-Based [TJL^{+}21]. Related [KXZG15, SR17]. Relation [FLY^{+}23, LYZ^{+}23, YHF21]. Relation-aware [FLY^{+}23]. Relational [SLWW13, CYC^{+}23, FLY^{+}23]. RelationLines [CXW^{+}19]. Relations [CXW^{+}19, RHF16, WC12, MKL11]. Relationship [CRYT12, LLY12, LZC23, ODP^{+}17, ZLL^{+}22]. Relationships [HM19]. Relative [TPG^{+}19]. Relaxing [RYC22]. Relevance [BBS^{+}16, MRJ16, WLC^{+}16, FPVC13]. Relevant [GLL^{+}17]. Reliable [HBSC13, MZL12, SLR^{+}16]. Relocation [WLC^{+}20, WLWC23]. Remaining [VDL^{+}19]. Remapping [FC15]. Remote [HCTC12, MRW^{+}12, ZZS^{+}21]. Removal [HJCK20, ZWH^{+}22, CZLS13]. Rendering [KKG18]. Rendezvous [SS22]. Reorder [SWZ^{+}13]. Repairing [FWYX22]. Repetitive [LXM^{+}18]. Replacing [DCM15]. repository [KDC13]. Representation [AWSF21, CYW^{+}21, DTL15, FL20, GG23, GQY^{+}19, HJCK20, LWZ^{+}23, LTS^{+}15, LFL^{+}23, TLB^{+}21, WLFY21, WYC^{+}22, WHRGC22, XHZ^{+}23, YWZ^{+}23, YZZ23, YLY^{+}19, ZLZ15, ZWZZ23, ZXY^{+}23, ZCL^{+}18, SHZ13]. Representation-Based [DTL15, GQY^{+}19]. Representations [GH18, LXW^{+}23, TMZ^{+}20, XXL^{+}23, FKSS13]. Reputation [MOC^{+}11]. Research [CAB^{+}13, MPS23, SBS^{+}23, WW13]. Residual [ZWH^{+}22]. Residuals [SGTK20, ZZGH19]. Resilient [JLL18].

Resists [ZLL^{+}22]. Resolution [DDZ^{+}21, CST13]. Resource [DPC16, HXC^{+}23, FT10]. Resources [CLBM15, PMR^{+}17, CXW^{+}13]. respiratory [LMWS13]. Response [CMR15, DZY^{+}22, FTE21, ZFQX20]. Responses [BZX^{+}22]. Restoration [CHY15]. Restricted [BFHP12, SRMW19]. Result [KCTT16, XXL^{+}17]. Results [AGP17, KZL^{+}17, ZWZS16]. Resume [ZW19]. ResumeVis [ZW19]. Ret [BI17]. Retrieval [CHHH18, Pai16, PSLB12, SR17, SST^{+}15, WJY^{+}18, WZZ^{+}16, YWZ^{+}17, ZYH^{+}17, ZWH17, ZB20, ZWC23, ZLC^{+}20, CZLS13, WH11b]. Return [HCTC12]. Retweet [LMC^{+}15]. Retweeting [ADJ^{+}20]. Reuse [MLSK23]. ReuseKNN [MLSK23]. Reveal [MTC^{+}20]. Revealing [FE15]. Reversion [HYL^{+}18]. Review [AdCK^{+}22, BTTT19, GJS23, IVS^{+}16, RBG22, WXLY12, YC12, ZFH^{+}18]. Reviewing [LLPS20]. Reviews [YCZY21, ZDC^{+}13]. Revisit [WZZ^{+}21]. Reward [XSJW23]. RFID [FGL17]. RGB [FDE15, HYZ15, ZYT^{+}15, ZSS^{+}15, ZLT15]. RGB-D [FDE15, HYZ15, ZYT^{+}15, ZSS^{+}15, ZLT15]. Rhetorical [RHT^{+}18]. RHUPS [BYK^{+}21]. Rhythms [YLD^{+}22]. Rich [FC15, Min16]. Ride [HHL^{+}22]. Ride-hailing [HHL^{+}22]. Rider [HLF^{+}21]. Rides [VPD^{+}22]. Right [DYQ^{+}23, BD11]. Risk [DZY^{+}22, GJSC16, GJC17, NYBG17, PPPM18, VKLY18, BSW^{+}13, OY13, RCN10]. Risk-Scoring [NYBG17]. Risk-Sensitive [VKLY18]. Risky [NZS^{+}22]. Road [DLY^{+}21, FL20, GXT^{+}23, RXK^{+}17, WLFY21]. Robot [HAAM12, XW23, TBK^{+}10]. Robots [CCZ^{+}23, DZY^{+}22, ZS18, CKS10]. Robust [CSA^{+}23, DYQ^{+}23, EMF12, FXR^{+}17, HBK^{+}16, HNA20, HCJM15, LNYV22, WHR13, XJS^{+}21, YCL^{+}21, YLY^{+}19, ZYT^{+}15, ZYSL12, ZGP^{+}18, ZWZZ23].

Robustness [WCFH23]. Role [DJS16, DHF22, HGE17, MPA13]. Romagna [CGMC11]. Rotational [HG21]. Rotorcraft [MJVL16]. Rough [CP23]. Route [DLY^{+}21, DCF^{+}18, GXYZ21, LH22, WLW^{+}22]. Routes [HLL14, WPL13]. Routine [LCLN18]. routines [FGP11]. Routing [CGZ18]. Rover [EBG^{+}12]. Rule [LLL^{+}16]. Rules [CDR19, CFG13]. Rumor [MGJW20].

S [HRCT16]. S-SMART [HRCT16]. S3 [CYC^{+}21]. S3-Net [CYC^{+}21]. Saddle [WZZ^{+}21]. Safely [LXBW20]. SafeRoute [LXBW20]. Safety [YLS15]. SAGE [WYC^{+}17]. Saliency [DLLT23, LJLZ19, ZFH^{+}18, LLF^{+}19]. Salient [ZLZ15, ZZL^{+}19]. Sample [YGY^{+}23]. Samples [TPM23]. Sampling [LYZ^{+}23, PDR23, YCSH23, ZZZ^{+}22, ZYY^{+}23]. Sampling-Based [ZZZ^{+}22]. SAT [LHC^{+}13, MMPS23]. SAT-based [LHC^{+}13]. Satellite [LPR19, LXW^{+}23]. Scalable [ACC21, BMTT16, CMR15, CCK^{+}18, LRD^{+}22, PFS17, RXK^{+}17, SSL^{+}18, VNL^{+}11, ZZC^{+}22, ZLC^{+}20]. Scale [ASW^{+}19, CQZ^{+}12, HDTG15, HG21, HKMN20, LWLG22, LJC^{+}11, MFLP14, PCC17, PCL18, ZQP^{+}15, ZL19, DLGT19, FGP11, Hsu11, HYL^{+}21, WTK^{+}19, WFX^{+}21, WZM^{+}22]. Scarcity [GB22]. Scenario [HMG^{+}23]. Scenarios [NVDMFD22]. Scene [CYC^{+}21, SSZ^{+}13]. Scenes [LXM^{+}18, ZRX^{+}22]. Scheduling [DGJ^{+}23, GJSC16, GJC17, SMGMC^{+}15, TAL^{+}19, WFX^{+}21, WM21, ZAK^{+}23, RYS10]. Schematization [RHF16]. Scheme [TRZ^{+}19, YTH17]. Schemes [MP23, TCK20]. Science [BTL20, EBG^{+}12, HCTC12, WCBK11]. Science-guided [BTL20]. Scientific [CWR^{+}16]. Score [RSCOVCMM17]. Scoring [HKMN20, NYBG17]. Screenshots [LHO23]. Sea [YLH^{+}23]. Seam [LC15].

Seam-Carved [LC15]. Search [AGP17, AC15, CWCK15, CLL23, CCC^{+}12, DSS^{+}22, HWZL20, JYL^{+}23, KCTT16, LCV17, LCM^{+}12, LTW^{+}16, LJC^{+}11, MOC^{+}11, QCL15, RGH19, SNL^{+}16, SSL^{+}18, TLB^{+}21, WH11a, XXL^{+}17, JPL13, WLH10, WPL13, YSJ13]. Searching [JTZ^{+}11, LRD^{+}22]. second [ESNN13]. second-language [ESNN13]. Secret [XLL^{+}22]. Section [CC12, CCC^{+}12, GST12, HJTZ12, HTDJ12, LZCQ12, SY12, ZCWY14a, CABD13, Edi13, FS13, GCZ13, KN13, LLWC13, RY13, WDSZ13, YNS13, YZEC13]. Sections [DCM15]. Secure [JJKZ22, LLLC19, ST20, YTH17]. Securely [KP17]. Security [HGE17, WMA20, ZGF^{+}23]. Seeding [VS11]. Segmentation [BZW^{+}22, CYC^{+}21, HSJ^{+}22, HCJM15, JHK^{+}22, LCN^{+}16, LWWX20, MWY^{+}23, TS17, WLL^{+}22, WHW^{+}21, YLX^{+}20, YCZ^{+}23, ZZS^{+}21, ZHW^{+}21]. Segmentations [MTC^{+}20]. Selected [CCL15]. Selecting [WL23, OSM^{+}13]. Selection [DPB20, GLL^{+}17, HCRF21, HYL^{+}18, JLL18, JZG22, LYW^{+}19, SLZ^{+}23, YGY^{+}23, YLY^{+}23, FT10, LHG11]. Selective [CCZ^{+}15]. Self [CASR22, HHJ22, YCL^{+}21, YWZ^{+}23, ZCZ^{+}23, ZXY^{+}23]. Self-Adaptive [HHJ22]. Self-attention-based [ZCZ^{+}23]. Self-supervised [CASR22, YWZ^{+}23]. Self-training [ZXY^{+}23]. Self-weighted [YCL^{+}21]. seller [ZC13]. Semantic [CDW^{+}19, DOTD16, HRCT16, HYC^{+}16, LFY^{+}22, SGY^{+}22, WLL^{+}21, YCP^{+}13, YLY^{+}23, ZW19, ZWC23, BGMS13b, CZLS13, CBP13, LKD13, SSZ^{+}13, YLT13, CDLV13]. Semantically [GLL^{+}17, LSW23]. Semantics [HNV14, SC17, Siz12]. Semi [CZG^{+}23, HSJ^{+}22, JYT^{+}12, KLL22, MFI19, STA22, ZY12, ZW19]. Semi-Boosted [MFI19]. Semi-Local [JYT^{+}12].

Semi-structured [ZW19]. Semi-Supervised [CZG^{+}23, ZY12, HSJ^{+}22, KLL22]. Semi-Synchronous [STA22]. Semiparametric [CDGZ16]. Sensed [HCTC12]. Sensing [CTY^{+}19, GCY^{+}15, LCCT12, LYKS23, MRW^{+}12, THL^{+}15, ZSL^{+}15, ZYW^{+}15, ZZS^{+}21]. Sensitive [HLL14, KCSW23, VKLY18, EFMRK^{+}20, LHC^{+}13, SLZ^{+}23, WCBK11]. Sensitivity [MFI19]. Sensitivity-Based [MFI19]. Sensor [SMX15]. Sensors [HYZ15, LCCT12]. Sensory [WHRGC22]. Sentence [XXL^{+}23, CL13]. sentential [BMV13]. Sentiment [PT12, RHD^{+}12, YZZ23, HLGW13]. sentiment-topic [HLGW13]. Separation [ZYH^{+}17]. Sequence [LLX^{+}22, RGC^{+}22, ZCL^{+}21, ZPP^{+}21]. Sequences [ASW^{+}19, GBBD22, LAsO^{+}19, LJC^{+}11, YZL^{+}19]. Sequencing [CLBM15]. Sequential [HLL^{+}23, LHZ22a, WH18, ZC15]. Series [EL14, KBM^{+}21, LCN^{+}21, MTC^{+}20, WC12, WHRGC22]. Serum [RFJ16]. Server [SLZ^{+}23]. Service [BSRSS16, FLLX18, JPS^{+}16, NZS^{+}22, XLZ^{+}22]. Services [SMGMC^{+}15, TPG^{+}19, WLWJ21, BGMS13a, KN13]. Servoing [XW23]. Session [ADM^{+}21]. Session-based [ADM^{+}21]. Set [LPM20]. Sets [LRD^{+}22]. Settings [HDTG15, STA22, WZYM19]. Shallow [ZZKT20]. shape [TDVC13]. Shapelet [DPB20]. Shapelet-transformed [DPB20]. Shapley [LCY^{+}22, SRMW19]. Shared [CZJL15, PMR^{+}17, QHH^{+}21, WYZ23]. Shared-Bike [QHH^{+}21]. Shared-private [WYZ23]. Sharing [EL14, LCM^{+}12, TRZ^{+}19, WLZ^{+}23, XLL^{+}22, CLL^{+}21, dMFA^{+}13]. Sharp [SDXG16]. Shifting [CCGP22]. Shilling [WGL^{+}22]. Shopping [BRSG20]. Shops [WM21]. Short [CASR22, CDW^{+}19, WCP^{+}23, ZWXZ12]. Short-term [WCP^{+}23]. Short-Text [ZWXZ12, CASR22]. Shot [CYC^{+}21, LCY^{+}18, WZYM19, WCP^{+}23, HCWH22, LZW^{+}23, MPS23]. Siamese [LRD^{+}22, SC17]. side [WHJ^{+}11]. Sided [PKCC18]. Sign [BLNN20, JJ14, JZG22, SZX15, ZZZ20a]. Sign-based [JZG22]. Signal [QTM23]. Signatures [BC19]. SignDS [JZG22]. SignDS-FL [JZG22]. Signed [BTL20]. Significance [XZS20]. Significant [CPHL15, LYWW18, XJS^{+}21]. Silo [DCWP22, STA22]. Similar [CCL15, DSS^{+}22, DPSS19, TLB^{+}21]. Similarities [XZR12]. Similarity [BQF^{+}23, EK15, GHZ^{+}17, LYWH20, LAS20, SNL^{+}16, TRH16, SLH13]. Similarly [TRH16]. Simple [CMR15, WTK^{+}19]. Simplifying [HTM15]. Simulating [KF18]. Simulation [HMG^{+}23, HM19, MFLP14, SHX^{+}23, SZS^{+}17, XWC^{+}19, ZLBZ23, FK13]. Simulations [CRRH11]. Simulator [RXK^{+}17]. Simultaneous [HRCT16, ZRX^{+}22]. Single [CYC^{+}21, DDZ^{+}21, HJCK20, WHW^{+}21, ZSS^{+}15]. Single-Shot [CYC^{+}21]. Sites [ZZZ^{+}11]. Size [SLR^{+}16]. Skeleton [BLNN20]. Sketch [MCEG23]. Sketching [SPDR15]. Skills [YFJ^{+}18]. Skin [SLC23]. Sliding [BYK^{+}21]. small [Dha11]. Smart [ASSR18, Bha21, CPHL15, DC21, FMdA^{+}23, HLC^{+}21, KHNB15, MMDY15, YLS15, RVRJ11, HRCT16]. Smartphone [XLF^{+}20]. smartphones [SGD13]. SmartPM [MMS17]. SMARTS [RXK^{+}17]. SmartTransfer [DCF^{+}18]. Smog [XWC^{+}19]. Smooth [LLDT16]. Smoothness [ZCS^{+}12]. SMP [BTVY17]. SNAP [LS16, LCY^{+}15]. Snippets [CHHH18]. Snow [HJCK20]. Soccer [PCF^{+}19, TLWZ11]. Social [ABTS15, ABO17, BTL20, BGMS13b, BCGJ11, BTVY17, CDLV13, CZP^{+}14, CHP17, CCW^{+}19, CYKL16, CCWS17, CDW^{+}21, DJS16, DSB^{+}18, FS17, FZX15, GST12, GZZY17, GW17, GDC19, GTM^{+}14, Goo10, GGC21, HQY^{+}22, HLY^{+}14, HTSC^{+}17, HL19, JJ14, JGL^{+}15, JLX^{+}17, LCV17, LC12, LH12, LCM^{+}12, LWH^{+}20, LGL^{+}22, LBP19, LHZ13, LYF^{+}22, MOC^{+}11, NZW^{+}17, PT12, PCL18, PEK^{+}16, QSRGDAJD13, RHT^{+}18, SFX17, Siz12, STP^{+}18, SZL^{+}23, TWL11, TY12, TY14, TRDD12, TLLS17, WXLY12, YZQ16, ZQP^{+}15, ZWH17, ZL19, ZZZ^{+}11, ZLL^{+}22, ZJSY21, ZRX^{+}22, ZMH^{+}15, BCD^{+}13, BGMS13a, CBP13, EvdHW13, FTCP^{+}13, FK13, Gin13, GCZ13, HCB13, HKO13, LCCS13, LN10, MKL11, SKOM13, SRM^{+}13, WCBK11, YNS13, YJHL11, dMFA^{+}13, BBGG13]. Social-Attribute [GTM^{+}14]. Social-Mobile [SFX17]. socially [YZEC13]. SocialWave [STP^{+}18]. sociotechnical [Sin13]. Sockpuppet [LLPS20]. Sockpuppet-Based [LLPS20]. Soft [WZH16]. solar [RFI^{+}11]. Solder [ZPL^{+}20]. Solving [EBS^{+}22, ZSY^{+}12]. Soter [YLS15]. Sound [OOD^{+}17]. SOUP [ZFH^{+}22]. Source [RBG22, TPM23, ZZS^{+}21, WYNW20]. Source-free [TPM23]. Sources [LNYV22, Min16]. Sourcing [YMLM16]. Souvenir [WST^{+}15]. Space [BC19, BD23, BQF^{+}23, BLAK19, CC12, JCH14, PMR^{+}17, RFI^{+}11, CCG^{+}13, WZY^{+}18, FC15]. SPACE-TA [WZY^{+}18]. Spaces [FC15, FDE15, SCLZ17, YZY^{+}17]. Spammer [FXR^{+}17]. Spammers [WXLY12]. Sparse [DYQ^{+}23, FWZ17, GHZ^{+}17, HG21, HJCK20, LTS^{+}15, LSZH18, LFL^{+}20, SLM^{+}23, TTLG17, WYC^{+}17, WZY^{+}18, ZYSL12, ZLZ15, THY^{+}11, YJHL11]. Sparseness [CSN^{+}17]. Sparsity [XTW17]. Spatial [CRRH11, DSS^{+}22, FXHM16, GMX^{+}21, HYC^{+}16, JWJC16, JSL^{+}19, JHK^{+}22, LLDT16, RHF16, SST^{+}15, TTFS18, WYC^{+}17, WFZ^{+}18, WYD^{+}18, WLL^{+}20, WLW^{+}22, WLW^{+}23, XHZ^{+}23, ZWL^{+}15, ZLZ^{+}22, ZPP^{+}21, DL13]. Spatial-aware [DSS^{+}22]. Spatial-Temporal [FXHM16, HYC^{+}16, WYC^{+}17, WLW^{+}23, WFZ^{+}18, WLW^{+}22, ZLZ^{+}22, ZPP^{+}21]. Spatio [BZX^{+}22, CCZ^{+}23, DDZ^{+}21, GXS^{+}22, HS19, LLS^{+}22, SC22, STP^{+}18, TLB^{+}21, WZFL21, WZFL22, WZM^{+}22, YXL^{+}23]. Spatio-Temporal [DDZ^{+}21, SC22, TLB^{+}21, WZFL21, WZFL22, BZX^{+}22, CCZ^{+}23, GXS^{+}22, HS19, LLS^{+}22, STP^{+}18, WZM^{+}22, YXL^{+}23]. Spatiotemporal [DYQ^{+}23, DCF^{+}18, LLZ^{+}23, LWLG22, LGJ^{+}22, Pat15, TEP11, XZS20, ZC15]. Speak [LBC^{+}22]. Special [AJL18, BTVY17, CKW19, CWLZ15, CALK16, CC12, CL15, CCC^{+}12, CSTZ16, GCY^{+}15, HJTZ12, HYZ15, JLX^{+}17, LZCQ12, SA15, SY12, WZFL21, WZFL22, YTL^{+}22a, ZLB^{+}16, ZCWY14a, ZLSY22a, ZLSY22b, ZWGW17, BBGG13, Che10, CABD13, Edi13, FS13, GY11, GCZ13, Hsu11, HTDJ11, KN13, LLWC13, Lin11, LN10, RY13, WDSZ13, YNS13, YZEC13, ZPY11, GST12, HLY^{+}14, HTDJ12]. Species [SLR^{+}16]. Specific [EK15, LBP19, GJ13, SSV19]. Specification [RGH19]. specifications [BBMP13]. Specified [LCLG19]. Spectral [RSCOVCMM17, WHW^{+}21]. Speech [JTP^{+}21, ZGP^{+}18, ZCZ^{+}23]. Speed [KP17, LLX^{+}22]. Splines [KBM^{+}21]. Sponsored [AC15, CWCK15, QCL15]. Spoof [ZCZ^{+}23]. Spoofing [TMZ^{+}20]. Sports [Bha21, ZDL^{+}12]. Spotting [WLF^{+}18]. SPrank [DOTD16]. Squares [LHS18]. ST [WYC^{+}17]. ST-SAGE [WYC^{+}17]. Stability [ZCWZ18]. Stable [RGH19]. Stage [ZZS^{+}21, SLZ^{+}23].

Staleness [DGK^{+}22]. StarFL [HLC^{+}21]. STARS [LLPS20]. Start [LHS18]. State [BC19, PCC10, RCN10]. state-dependent [RCN10]. Statechart [KW17]. Statechart-Based [KW17]. Static [TAL^{+}19]. Station [HCRF21, WFX^{+}21, RFI^{+}11]. Stations [CLL^{+}21]. Statistical [LC12, XZS20, Mar13]. Statistically [XJS^{+}21]. Status [LCCT12]. STCAPLRS [FXHM16]. Steering [HASS22]. Steering-by-example [HASS22]. Steganalysis [LC15, LSQ11]. Stereotypes [SR17]. Stereotypical [BNS13]. STExplorer [CCZ^{+}23]. Stochastic [CZJL15, JD15, LH12, SC22, TY14, VKLY18]. Stock [YLWX20]. Stop [KL23]. Stopping [ZCWZ18]. Storage [MBR^{+}14]. Store [LGZ^{+}21, WXLY12]. Story [CQZ^{+}12]. storytelling [PCC10]. Strangers [LMC^{+}15]. Strategic [ARGK15, WMA20]. Strategies [YCSH23, ERR13]. Strategy [CCZ^{+}23, DZY^{+}22, HYL^{+}18, OLY^{+}17, RV18, HLT11]. Stream [GGC21, XLZ21, YCY23]. Streamed [DCM15]. Streaming [LWH12, ZFWL17]. Streamlined [CCL15]. Streams [ABTS15, BYK^{+}21, CQZ^{+}12, MLJZ21, RHD^{+}12, YLD^{+}22, ZLY^{+}18, TZY^{+}13]. Street [LPR19, MIS20, ZYY^{+}23]. Streets [LXBW20]. Stroke [LMAP16, WCP^{+}23, ZLZ15]. Structural [AKR^{+}18, DWKP16, PS11, RGH19, TS17, CBP13]. Structure [ASK^{+}21, FZX15, HDPH16, LTS^{+}15, VKA^{+}19, WZHL14, WLH17, YL14, ZPP^{+}21]. Structure-Based [LTS^{+}15]. Structured [FT10, HCWH22, Min16, XTW17, MGJW20, ZW19]. Structures [TWL11, WFZ^{+}18, EvdHW13]. Students [YLC^{+}19]. Studies [LLL^{+}16]. Study [EL14, HBL16, MOC^{+}11, PHL^{+}20, WZS^{+}20, EvdHW13]. Style [SZZ^{+}21]. Styles [HWL^{+}17]. Sub [POM20, BMV13]. Sub-Optimal [POM20]. sub-sentential [BMV13]. Subgraph [PDR23]. Subgraphs [SRMW19]. Subject [ZS18]. Subjective [ACC21]. Subkilometer [DSM^{+}11]. Submodularity [KG11]. Subpaths [XZS20]. Subspace [AWSF21, LCD18, LFL^{+}23]. Subtraction [CCH15]. Subtrajectory [DSS^{+}22]. Suggestion [LJC^{+}11, WFJY12, SLWW13, YJHL11]. suitable [OSM^{+}13]. Suite [HLL^{+}22]. Sulfur [MRW^{+}12]. Summarization [ABG^{+}11, LTW^{+}16, LZP^{+}12, SWZ^{+}21]. Summary [YMC16]. Super [DDZ^{+}21]. Superpixel [LWWX20]. Supervised [BQF^{+}23, CZG^{+}23, DC21, JHK^{+}22, PCC17, TWC^{+}23, WLL^{+}22, YHF21, ZY12, ZCL^{+}18, CASR22, HSJ^{+}22, KLL22, YWZ^{+}23]. Supervision [SP16]. Supply [WM21, ZMH^{+}22]. Supply-Demand-aware [ZMH^{+}22]. Support [AKR^{+}18, NVDMFD22, SZX15, VKA^{+}19, CL11, ESNN13, KDC13]. Surface [HSJ^{+}22, MRW^{+}12, WPA^{+}12, YLH^{+}23]. Surrounding [XWW^{+}21]. Surveillance [JJKZ22, SSZ^{+}13]. Survey [CMPR21, DT16, GXS^{+}22, HCWH22, JGL^{+}15, LPL^{+}22, LZW^{+}23, NCG21, RAK23, SQJ^{+}19, VDL^{+}19, WZYM19, WWD^{+}21, WC20, XWW^{+}21, YLX^{+}20, ZSAL20, ZB20, ZGL^{+}17, JPL13, LHS^{+}13, SRM^{+}13, WH11b]. Survival [KZL^{+}21]. Survivor [SDXG16]. Susceptibility [ODP^{+}17]. sustainability [GY11]. Sustainable [CGMC11, CRRH11, HMS^{+}14, PMSR11]. SVANN [GMX^{+}21]. SVM [CGZ23]. Swarm [HJCK20]. Switching [ZZZ^{+}20b]. Synchronous [STA22]. Synthesis [WLL^{+}20, WLL^{+}21, ZLL^{+}22]. Synthetic [LMAP16]. System [ADM^{+}21, Bha21, CLL^{+}21, CSTZ16, DGJ^{+}23, DJI^{+}16, EL14, FXHM16, FLF^{+}20, FTE21, FSS15, JCW^{+}22, KAH12, LLLC19, LH22, LDFL23, MZL12, MP23, MMS17, NDW^{+}19, SNL^{+}16, SSG^{+}20, SMGMC^{+}15, TLW^{+}15, WST^{+}15, WYG^{+}22, YMLM16, ZLT15, ZWL^{+}15, ZW19, ZYH^{+}20, ERR13, KDC13, SSZ^{+}13, SZC^{+}13]. system-wide [ERR13]. Systematic [AdCK^{+}22, ZDL^{+}12]. Systems [AT15, BC19, BLL^{+}14, CWLZ15, CVCL22, CALK16, CZJL15, CRRH11, CYW^{+}21, DJS16, DPSS19, GDC19, GJS23, HGE17, HJTZ12, HLL^{+}22, HTDJ12, HCJM15, IVS^{+}16, KW17, LCN^{+}21, LLPS20, MBM21, MWS^{+}18, ODL^{+}20, PMR^{+}17, QTM23, RZS^{+}15, SYHB17, TAL^{+}19, TL23, YTL^{+}22a, YTL^{+}22b, ZDW19, BNS13, Edi13, FS13, FKSS13, GCZ13, HBSC13, HTDJ11, LLWC13, QSRGDAJD13, RC13, RY13, Sin13, WW13, YZEC13, ZPY11].

Table [ZB20]. Tag [BBS^{+}16, FSS15, LM11, TC19, ZWH17, CFG13, CZLS13, GJ13, FGL17]. Tagged [TRDD12, ZWH17, ZLY^{+}18, THY^{+}11]. tagging [WHJ^{+}11]. Tags [SC17]. Tail [WZS^{+}15]. Tailed [SLM^{+}23]. Tailored [BQF^{+}23]. Take [LPR19, WLZ^{+}23]. Takeaway [HLF^{+}21]. Taken [ZSLC19]. Taking [WMR17]. TAML [XXZ^{+}21]. Tangled [MWY^{+}23]. TARA [HLF^{+}21]. TARA-Net [HLF^{+}21]. Target [CHP17, TTL^{+}21]. Targeted [LLPS20, WYG^{+}22, MD13, RBK^{+}13]. Targeting [EBG^{+}12]. targets [SZC^{+}13]. Task [BN21a, DPC16, GBC^{+}22, HWZL20, LCKY14, LHZ22a, LGZ^{+}21, NOZ20, TAL^{+}19, TTFS18, WZY^{+}18, XLF^{+}20, ZPL^{+}20, CKS10, DLLT21, LHZ^{+}22b, XXZ^{+}21]. Task-adaptative [LGZ^{+}21]. Taxi [PHL^{+}20, WLF^{+}18, WZS^{+}20, ZQP^{+}15, ZYW^{+}15]. Taxonomy [MPS23, SRM^{+}13, ZZH^{+}22]. TCP [HBL16]. Team [Zhu19]. teaming [TBK^{+}10]. Teams [DHF22, SS11]. Techniques [DCM15, GH18, SQJ^{+}19, ZSY^{+}12]. Technologies [XWW^{+}21]. Technology [HTDJ12, HTDJ11, WW13]. Telco [LZY^{+}16]. Tele [KXZG15]. Tele-Health [KXZG15]. Telepresence [HCJM15]. Temperature [YLH^{+}23]. Tempo [RSCOVCMM17]. Temporal [DDZ^{+}21, EHG21, FXHM16, GBBD22, HYC^{+}16, HLH^{+}21, JWJC16, LLDT16, PMSR11, PEK^{+}16, SC22, TY14, TLB^{+}21, WYC^{+}17, WYD^{+}18, WCS^{+}20, WZFL21, WZFL22, WLW^{+}23, YY15, ZHLL21, BZX^{+}22, BVCH13, BK11, CCZ^{+}23, GXS^{+}22, HS19, LLS^{+}22, LHZ13, STP^{+}18, WFZ^{+}18, WZM^{+}22, WLW^{+}22, YLT13, YXL^{+}23, ZLZ^{+}22, ZPP^{+}21]. Temporal-Spatial [JWJC16]. Temporal-Spatial-Smooth [LLDT16]. temporally [LHC^{+}13]. Tensor [DLLT21, LWLG22, LYLX23, SLZ^{+}23, WYM17]. Tensor-based [SLZ^{+}23]. TensorBeat [WYM17]. TensorFlow [AHJB20]. Tensors [PFS17]. Term [CUG^{+}12, TJL^{+}21, WZS^{+}20, CDS13, SLWW13, WCP^{+}23]. term-suggestion [SLWW13]. TerraFly [ZWL^{+}15]. Terrain [HSJ^{+}22]. Test [DGJ^{+}23, ZWL^{+}19]. testbed [EvdHW13]. Tests [FNS16, BCC^{+}13]. Text [CDS12, CLL23, CDW^{+}19, GOB^{+}12, GWD^{+}21, KCSW23, KCS18, LPL^{+}22, LZCQ12, LZP^{+}12, LCY^{+}18, RHD^{+}12, SDD^{+}16, SZZ^{+}21, TRH12, WLL^{+}20, ZWXZ12, CASR22, HCB13]. Text-to-Image [WLL^{+}20]. Texts [KXZG15]. Textual [NAPI14, YLY^{+}23, YZZ23]. Texture [DTL15]. TextWheel [CQZ^{+}12]. Their [CCWS17, LCLG19, WMWR22]. Theorem [ZGF^{+}23]. Theoretic [GG23]. Theoretical [WZZ^{+}21]. Theory [Bha21, Gin13, JCH14, LMAP16, MRJ16]. Thermal [GRR^{+}15, LGJ^{+}21]. Thermography [GJS23]. Thermography-based [GJS23]. Thing [YSN^{+}17]. Things [YSN^{+}17]. Threaded [KW17]. Three [BLNN20].

Three-dimensional [BLNN20]. TIARA [LZP^{+}12]. Ticket [GG15]. Tier [DCWP22]. Ties [HL19]. Time [ABG^{+}11, BYK^{+}21, BB15, CGZ18, GBBD22, HSBRM22, HNA20, HLL14, JYL^{+}23, KBM^{+}21, LCN^{+}21, MWY^{+}23, MTC^{+}20, RSCOVCMM17, TLW^{+}15, TTFS18, VDL^{+}19, WC12, WTK^{+}19, WCS^{+}20, WLW^{+}23, WHRGC22, XXZ^{+}21, ZLT15, ZLY^{+}18, ZHZ18, BKB10, GXT^{+}23, WFX^{+}21]. Time-Sensitive [HLL14]. Timing [GG15]. TIST [AJL18, BTVY17, CKW19, CL15, LN10, WZFL21, Yan10, ZLB^{+}16, ZWGW17]. TLDS [HCRF21]. Tomographic [CWR^{+}16]. Tools [SPDR15]. Top [DOTD16, EK15, MG16]. Top- [EK15, DOTD16, MG16]. Topic [GLL^{+}17, GOB^{+}12, GZ21, GZ23, HYC^{+}16, JTS^{+}21, LZP^{+}12, PEK^{+}16, WZS^{+}15, WHW^{+}21, YMC16, YCGH12, CDS13, FGP11, HLGW13]. Topic-Aware [PEK^{+}16]. Topic-Based [LZP^{+}12]. TopicNets [GOB^{+}12]. Topics [GLL^{+}17, HL17]. Topological [WLH17]. Total [CGZ18]. Tour [ZWX^{+}22]. Tourists [MNSB15, ZNWC14]. Tournaments [VS11]. TPM [WYD^{+}18]. Traces [WLWJ21, YFJ^{+}18, ZQP^{+}15, HY11, ZX11]. Track [WLZ^{+}23]. Tracking [GZZY17, HRCT16, SSZ^{+}13, YL17, YLX^{+}20, ZYSL12, LHS^{+}13, SZC^{+}13, TLWZ11]. Tractable [FKSS13]. Trading [SWA23, ZKC^{+}23, HLT11]. Traditional [LPL^{+}22, SPDR15]. Traffic [ABTS15, CP23, DLY^{+}21, GLJ^{+}14, HQW22, HTL^{+}20, HLH^{+}21, KL23, KZL^{+}17, LLS^{+}22, LLX^{+}22, LGJ^{+}22, LBC^{+}22, MFLP14, QTM23, RXK^{+}17, WZM^{+}22, WYY^{+}23, XWC^{+}19, XXZ^{+}21, ZLBZ23, ZLZ^{+}22]. Traffic-aware [XXZ^{+}21]. Training [HSJ^{+}22, KLL22, KCS18, KSL^{+}15, LCJ^{+}19, STA22, YCZY21, ZLD^{+}23, TWL^{+}22, ZXY^{+}23]. Trajectories [CLH^{+}22, LRD^{+}22, LBC^{+}22, MJVL16, SGY^{+}22, WLF^{+}18, XLB23, YCP^{+}13, YLT13]. Trajectory [BZW^{+}22, BTTT19, CTC^{+}22, DYQ^{+}23, FL20, HQY^{+}22, HHL^{+}22, MLJZ21, SHX^{+}23, SS22, SLM^{+}23, TLB^{+}21, YBZ^{+}20, YCH^{+}22, Zhe15, ZLSY22a, ZLSY22b, ZLL^{+}22, ZWX^{+}22, ZRX^{+}22, TZY^{+}13, TLWZ11, WPL13]. Transactional [BN21b]. Transductive [EMF12]. Transfer [BN21a, DJI^{+}16, DGL^{+}22, FC15, GB22, HCRF21, JTP^{+}21, LCN^{+}21, LCY^{+}18, LGZ^{+}21, PYD^{+}17, SZZ^{+}21, WCF^{+}20, WYY^{+}23, WYY^{+}19, ZY12, ZCL^{+}18, DSM^{+}11]. Transfer-Learning-Based [HCRF21]. Transferability [XLG^{+}23]. Transferability-Based [XLG^{+}23]. Transferred [DDZ^{+}21]. Transfers [DCF^{+}18]. Transform [DDZ^{+}21]. Transformation [HHJ22]. Transformations [MBM21]. transformed [DPB20]. Transformer [SZZ^{+}21]. Transformer-Based [SZZ^{+}21]. Transition [WYY^{+}19]. Translation [LGJ^{+}21, MD13, Mar13, RBK^{+}13]. Transmission [YZL^{+}19]. Transportation [FLF^{+}20, HSBRM22, LH22, ZDW19]. Transporter [LH22]. Traps [FMdA^{+}23]. Travel [CGZ18, CWC^{+}20, JYL^{+}23, LLL^{+}18, NZS^{+}22, TLC^{+}14, WTK^{+}19, XXZ^{+}21, ZZC12, WPL13, ZX11]. Traveling [LH22, TZY^{+}13]. Treatment [LG16, KDC13]. Treatments [ABB^{+}15]. Tree [MGJW20, MCEG23, DL13]. Tree-structured [MGJW20]. Trees [BFHP12, LNO^{+}18]. TreeSketchNet [MCEG23]. Trembr [FL20]. Trend [WLZ^{+}23, ZZC^{+}20]. Trends [SBS^{+}23, ZZH^{+}22]. Tri [YCZY21]. Tri-Training [YCZY21]. trial [LKD13]. Trip [VKLY18, WLF^{+}18, WTK^{+}19]. Trump [YLWX20]. Truncated [SDXG16]. Trust [OSM^{+}13, BNS13, FS13, ZC13]. Trusted [TPM23]. trusting [FPVC13]. Trustworthy [LWF^{+}23]. Truth [FSW^{+}20, RFJ16, YL14]. TSK [DJI^{+}16]. tuned [EFMRK^{+}20]. Tuning [BN21a, XW23]. tutoring [FKSS13, LLWC13]. Tweet [ZLY^{+}18]. Tweets [SP16, LWZZ13, SWZ^{+}13]. TWIST [DDZ^{+}21]. TWIST-GAN [DDZ^{+}21]. Twitter [AAX13, CDK^{+}13, KN13, LMC^{+}15, MND14, PT12]. Two [HSBRM22, LWWL11, PKCC18, SSG^{+}20]. Two-Dimensional [PKCC18]. Two-Level [HSBRM22]. Two-Word [LWWL11]. Type [LCV17, WZCJ21].

Ubiquitous [ZZZ20a]. Ubiquity [XYS^{+}23]. UI [LHO23]. UMCR [YWZ^{+}17]. uncertain [WHJ^{+}11]. Uncertainties [KHNB15]. Uncertainty [ACC21, BZW^{+}22, CCK^{+}18, LCC^{+}20, WHR13]. Uncertainty-based [BZW^{+}22]. Uncovering [ZJSY21]. Underlying [ZZKT20]. Understand [ZLH18, ZZZ^{+}20b]. Understanding [CYC^{+}21, HYZ15, JTZ^{+}11, LJLZ19, LFW23, ODP^{+}17, RHT^{+}18, SWZ^{+}21, TWL11, WZS^{+}20, ZYT^{+}15, ZL12, ZL19, ZWL^{+}19]. Unexpected [AT15, LT20]. Unexpectedness [AT15]. Unfold [CQZ^{+}12]. Unified [CYKL16, HCCY15, HRCT16, LXJ^{+}20, SLM^{+}21, SZC^{+}14, WLC^{+}16]. Unifying [WCFH23]. Universal [XLG^{+}23]. unlabeled [CCG^{+}13]. Unordered [LRD^{+}22]. Unsupervised [CZG^{+}23, HWT17, LSW23, PT12, SZL^{+}23, TTL^{+}21, TPM23, WC20, XLZ21, ZXY^{+}23, SSZ^{+}13]. Unveiling [YSN^{+}17]. Update [LLLC19]. Upper [ODF17]. Urban [AJL18, CXW^{+}19, CCK^{+}18, DSB^{+}18, HLC^{+}21, KF18, KZL^{+}21, LXBW20, LH22, LDFL23, LGJ^{+}22, SCLZ17, VKLY18, WFZ^{+}18, YKTL14, ZYW^{+}15, ZLZ^{+}22, ZYY^{+}23, ZCWY14a, ZCWY14b]. UrbanKG [LDFL23]. URLs [MSSV11]. Usage [EL14, FLLX18, LZ18, LLZ^{+}23, SMX15, XLF^{+}20, YFJ^{+}18]. Use [HLNL18, LLZ^{+}23, XHZ^{+}23]. User [BLL^{+}14, CRYT12, CTY^{+}19, CDR19, CCC^{+}12, EK15, FXR^{+}17, LCKY14, LPM20, LWH^{+}20, LLZ^{+}23, LZY^{+}16, NZW^{+}17, SDD^{+}16, TY12, TL23, WZHL14, WLWJ21, WWL^{+}22, YFJ^{+}18, YWZ^{+}17, YKTL14, ZCX^{+}15, SWZ^{+}13, ZX11, dMFA^{+}13]. User-Generated [CRYT12, CCC^{+}12, SDD^{+}16, ZX11]. User-Specific [EK15]. User/Item [LPM20]. User/Item-Set [LPM20]. Users [ADJ^{+}20, CHP17, LH12, LHS18, MND14, NYBG17, CCL13]. Using [ASK^{+}21, BC19, BMTT16, CLBM15, CCZ^{+}15, CWR^{+}16, CLL^{+}21, DB16, DOTD16, EFMRK^{+}20, FSW^{+}20, FK13, GJSC16, GDC19, GTM^{+}14, GWDJ15, HCTC12, HM19, HWT17, HJCK20, JV20, JD15, KCTT16, LPR19, LMAP16, LH12, LCM^{+}12, LLDT16, LWZ^{+}23, LWWL11, MFLP14, PCL18, POM20, PS11, RFJ16, RK15, RBK^{+}13, RSCOVCMM17, SDHS15, SRB15, SLR^{+}16, SGTK20, SSV19, SRJP12, TCK20, TY14, TLW^{+}15, WYZ23, XZXM19, XLL^{+}22, YGU15, YY15, YL17, ZWXZ12, ZL12, ZYSL12, ZYH^{+}17, ZLH18, ZZZ20a, CDK^{+}13, CCG^{+}13, EvdHW13, FGP11, GBBD22, KDC13, Min16, MGB^{+}11, PG13, PCC10, SSHL13, SGD13, TDVC13, VNL^{+}11, WTK^{+}19]. Utility [BYK^{+}21, WH18, ZGF^{+}23, ZKC^{+}23, ZLL^{+}22, WCS^{+}20]. Utility-aware [ZLL^{+}22]. Utilizing [AHJB20].

Valid [GLL^{+}17]. Validating [CTY^{+}19]. Validation [BCR21, KDC13]. Value [AKA^{+}21, ZZL^{+}19, KDC13]. Value-based [AKA^{+}21]. Values [SRMW19, SSV19]. VAR [DWKP16]. Variability [GMX^{+}21]. Variable [WHRGC22]. Variables [RGH19]. Variational [GZ23, SZC11]. variety [FTCP^{+}13]. Various [GJS23]. varying [ZC13]. VDI [ZFH^{+}22]. Vector [AKR^{+}18, GZ21, NVDMFD22, SZX15, CL11].

Vector-Quantization-Based [GZ21]. Vehicle [CGZ18, CLH^{+}22]. Vehicles [KL23, MBR^{+}14, QTM23]. Vehicular [NDW^{+}19, XWW^{+}21]. Vélib’ [EL14]. Velocity [ZS18]. Venues [ZNWC14]. Verbal [LJLZ19]. Verification [BCR21, XWC^{+}19]. verifying [GS13]. Vertical [DCWP22, KLL22, KCS18, RYC22, YCZY21]. VFI [TAL^{+}19]. VFI-based [TAL^{+}19]. Via [CCL15, HQW22, PKCC18, SC17, AWSF21, BL16, BPS13, CSA^{+}23, CZLS13, CWC^{+}20, DGK^{+}22, EHG21, EMF12, FC15, GZZY17, GXYZ21, GB22, GZ23, HLT11, LZ18, LPM20, LWLG22, LLZW17, LCY^{+}18, LFL^{+}23, MZL12, MRW^{+}12, NZW^{+}17, PCF^{+}19, QWC^{+}23, RP23, SLM^{+}21, WXLY12, WZHL14, WCBL18, WYC^{+}22, WYY^{+}23, WCP^{+}23, WLL^{+}22, XLF^{+}20, YSN^{+}17, ZYH^{+}20, ZNWC14, ZJSY21, ZZC^{+}22]. Video [BLNN20, EMF12, FZ16, HXY^{+}22, JJKZ22, LLDT16, LYKS23, MWY^{+}23, OLY^{+}17, SRJP12, SWZ^{+}21, WLL^{+}22, XTW17, YLX^{+}20, YCZ^{+}23, ZFWL17, TLWZ11]. Video-Based [XTW17, JJKZ22]. Videos [SZC^{+}14, ZDL^{+}12]. View [CDW^{+}21, DMGSD23, FLLX18, LPR19, LYLX23, MIS20, WJY^{+}18, ZHZ18, ZWZ^{+}19, ZYY^{+}23, CKP^{+}22, KLL22]. Viewpoint [ST19]. village [HKO13]. Virtual [Bai10, BL16, KSL^{+}15, WCBL18, ZDC^{+}13]. Vision [JGL^{+}15, LGJ^{+}21, SZC^{+}13]. Vision-Based [JGL^{+}15]. Visual [CKW19, CCW^{+}19, CXW^{+}19, DPSS19, GOB^{+}12, HNV14, HASS22, HYZ15, HWL^{+}17, LJLZ19, LZCQ12, LZP^{+}12, LLZW17, LYWH20, RHD^{+}12, RAZE18, SR17, SST^{+}15, STP^{+}18, SWZ^{+}21, TRH12, XZXM19, XWC^{+}19, XW23, YZZ23, ZYSL12, ZSS^{+}15, ZW19, ZWL^{+}19, LHS^{+}13]. Visual-textual [YZZ23]. Visual-Verbal [LJLZ19]. Visualisation [VKA^{+}19]. Visualization [CDS12, CQZ^{+}12, KSL^{+}15, SFX17, ST19, ZL12, ZWL^{+}15, CBP13].

Visualizing [LCJ^{+}19]. Voice [PSRL12]. Volume [KKG18, ST19, WYZ23]. Voting [MFB^{+}20]. VSRank [WSGM14]. VSumVis [SWZ^{+}21].

Waiting [HSBRM22, WLW^{+}23]. Walk [LWZ^{+}23, YKTL14, CLSL13]. walks [CST13]. Warping [LLDT16, RSCOVCMM17]. Wasserstein [WYNW20, ZFQX20]. Watch [CQZ^{+}12]. Water [LAsO^{+}19, WCBL18]. Waterway [HSBRM22]. Wavelet [DDZ^{+}21]. Weakly [JHK^{+}22, TWC^{+}23, WLL^{+}22]. Wearable [WHRGC22, YLD^{+}22]. Wearable-Sensory [WHRGC22]. Wearables [ZZD^{+}17]. web [THY^{+}11, AGP17, ABG^{+}11, BLL^{+}14, BBGG13, CDLV13, CCL15, CZLS13, ESNN13, GST12, HDPH16, HWZL20, JPL13, LC12, LH12, MOC^{+}11, NYBG17, SZC^{+}14, TRDD12, YSN^{+}17, ZWXZ12, ZB20, ZZZ^{+}11]. Webpages [NYBG17]. Weighted [MFB^{+}20, MFI19, WZH16, WYC^{+}22, YTH17, LNYV22, YCL^{+}21]. Weighting [WYZ23]. Weights [XSJW23]. Well [SBS^{+}23]. Well-Informed [SBS^{+}23]. Where2Stand [WST^{+}15]. White [YZY^{+}17]. Who [LMC^{+}15, ZZZ^{+}11]. Wi [SCLZ17, ZZZ20a]. Wi-Fi [SCLZ17, ZZZ20a]. wide [ERR13]. WiFi [WYM17]. Wiki [WFJY12]. Wikipedia [LLY12, SRB15, ZWXZ12]. Will [LMC^{+}15]. Willing [WYP22]. Willing-to-Pay [WYP22]. Window [BYK^{+}21]. Window-based [BYK^{+}21]. Wireless [DCM15, ZWGW17]. WiSign [ZZZ20a]. Within [XYS^{+}23]. Within-Basket [XYS^{+}23]. Without [KL23, SLR^{+}16, FSW^{+}20, WLZ^{+}23]. Word [CUG^{+}12, LWWL11, ME13, ZT11]. word-of-mouth [ZT11]. Words [HCCY15]. Worker [CCK^{+}18]. workflow [WHR13]. workflows [HY11]. Workload [ZFH^{+}22, BCC^{+}13]. World [BFC^{+}17, FSS15, WCFH23]. worlds [Bai10, TBK^{+}10].

X [LNO^{+}18]. X-CLEaVER [LNO^{+}18]. XGBoost [XLL^{+}22]. XLearn [YDZ20].

York [JCH14]. YouTube [DJNC21].

Zero [WZYM19]. Zero-Shot [WZYM19].

References

Arias:2013:FTD

[AAX13] Marta Arias, Argimiro Arratia, and Ramon Xuriguera. Forecasting with Twitter data. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1): 8:1–8:??, December 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Antonelli:2015:MCM

[ABB^{+}15] Dario Antonelli, Elena Baralis, Giulia Bruno, Luca Cagliero, Tania Cerquitelli, Silvia Chiusano, Paolo Garza, and Naeem A. Mahoto. MeTA: Characterization of medical treatments at different abstraction levels. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4): 57:1–57:??, August 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Anagnostopoulos:2011:WPS

[ABG^{+}11] Aris Anagnostopoulos, Andrei Z. Broder, Evgeniy Gabrilovich, Vanja Josifovski, and Lance Riedel. Web page summarization for just-in-time contextual advertising. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):14:1–14:??, October 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Assem:2017:RRC

[ABO17] Haytham Assem, Teodora Sandra Buda, and Declan O’Sullivan. RCMC: Recognizing crowd-mobility patterns in cities based on location based social networks data. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (5):70:1–70:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Anantharam:2015:ECT

[ABTS15] Pramod Anantharam, Payam Barnaghi, Krishnaprasad Thirunarayan, and Amit Sheth. Extracting city traffic events from social streams. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4):43:1–43:??, August 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ashkan:2015:LQA

[AC15] Azin Ashkan and Charles L. A. Clarke. Locationand query-aware modeling of browsing and click behavior in sponsored search. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):59:1– 59:??, January 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Alim:2021:CSC

[ACC21] Adil Alim, Jin-Hee Cho, and Feng Chen. CSL+: Scalable collective subjective logic under multidimensional uncertainty. ACM Transactions on Intelligent Systems and Technology (TIST), 12(1):7:1– 7:26, February 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3426193.

An:2017:DDF

[ACPS17] Bo An, Haipeng Chen, Noseong Park, and V. S. Subrahmanian. Datadriven frequency-based airline profit maximization. ACM Transactions on Intelligent Systems and Technology (TIST), 8(4):61:1– 61:??, July 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Antunes:2022:FLH

[AdCK^{+}22] Rodolfo Stoffel Antunes, Cristiano André da Costa, Arne Küderle, Imrana Abdullahi Yari, and Björn Eskofier. Federated learning for healthcare: Systematic review and architecture proposal. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):54:1–54:??, August 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3501813.

Arora:2020:ADC

[ADJ^{+}20] Udit Arora, Hridoy Sankar Dutta, Brihi Joshi, Aditya Chetan, and Tanmoy Chakraborty. Analyzing and detecting collusive users involved in blackmarket retweeting activities. ACM Transactions on Intelligent Systems and Technology (TIST), 11(3): 35:1–35:24, May 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3380537.

Adamczak:2021:SBH

[ADM^{+}21] Jens Adamczak, Yashar Deldjoo, Farshad Bakhshandegan Moghaddam, Peter Knees, Gerard-Paul Leyson, and Philipp Monreal. Session-based hotel recommendations dataset:

As part of the ACM Recommender System Challenge 2019. ACM Transactions on Intelligent Systems and Technology (TIST), 12(1):1:1–1:20, February 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/doi/
10.1145/3412379.

Agrawal:2017:HWS

[AGP17] Rakesh Agrawal, Behzad Golshan, and Evangelos E. Papalexakis. Homogeneity in Web search results: Diagnosis and mitigation. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (5):66:1–66:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Albaqsami:2020:AHM

[AHJB20] Ahmad Albaqsami, Maryam S.

Hosseini, Masoomeh Jasemi, and Nader Bagherzadeh. Adaptive HTF-MPR: an adaptive heterogeneous TensorFlow mapper utilizing Bayesian optimization and genetic algorithms. ACM Transactions on Intelligent Systems and Technology (TIST), 11(5):55:1–55:25, September 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3396949.

An:2018:ATS

[AJL18] Bo An, Nick Jennings, and Zhenhui Jessie Li. ACM TIST special issue on urban intelligence. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3):23:1–23:??, February 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Aydogan:2021:NVB

[AKA^{+}21] Reyhan Aydogan, Özgür Kafali, Furkan Arslan, Catholijn M. Jonker, and Munindar P. Singh. Nova: Value-based negotiation of norms. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4): 45:1–45:29, August 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3465054.

Anaissi:2018:AOO

[AKR^{+}18] Ali Anaissi, Nguyen Lu Dang Khoa, Thierry Rakotoarivelo, Mehrisadat Makki Alamdari, and Yang Wang. Adaptive online one-class support vector machines with applications in structural health monitoring. ACM Transactions on Intelligent Systems and Technology (TIST), 9(6):64:1– 64:??, November 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Azaria:2015:SID

[ARGK15] Amos Azaria, Zinovi Rabinovich, Claudia V. Goldman, and Sarit Kraus. Strategic information disclosure to people with multiple alternatives. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):64:1– 64:??, January 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ali:2021:PAN

[ASK^{+}21] Sarwan Ali, Muhammad Haroon Shakeel, Imdadullah Khan, Safiullah Faizullah, and Muhammad Asad Khan. Predicting attributes of nodes using network structure. ACM Transactions on Intelligent Systems and Technology (TIST), 12(2):21:1– 21:23, March 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3442390.

Auffenberg:2018:CBA

[ASSR18] Frederik Auffenberg, Stephen Snow, Sebastian Stein, and Alex Rogers. A comfortbased approach to smart heating and air conditioning. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3): 28:1–28:??, February 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ao:2019:LSF

[ASW^{+}19] Xiang Ao, Haoran Shi, Jin Wang, Luo Zuo, Hongwei Li, and Qing He. Largescale frequent episode mining from complex event sequences with hierarchies. ACM Transactions on Intelligent Systems and Technology (TIST), 10(4):36:1– 36:??, August 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3326163.

Adamopoulos:2015:URS

[AT15] Panagiotis Adamopoulos and Alexander Tuzhilin. On unexpectedness in recommender systems: Or how to better expect the unexpected. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):54:1–54:??, January 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Abhadiomhen:2021:MCS

[AWSF21] Stanley Ebhohimhen Abhadiomhen, Zhiyang Wang, Xiangjun Shen, and Jianping Fan. Multiview common subspace clustering via coupled low rank representation. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4):44:1–44:25, August 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3465056.

Bainbridge:2010:VWC

[Bai10] William Sims Bainbridge.

Virtual worlds as cultural models. ACM Transactions on Intelligent Systems and Technology (TIST), 1 (1):3:1–3:??, October 2010. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Balakrishnan:2015:RTB

[BB15] Raju Balakrishnan and Rushi P. Bhatt. Real-time bid optimization for groupbuying ads. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):62:1–62:??, January 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Bonchi:2013:ISI

[BBGG13] Francesco Bonchi, Wray Buntine, Ricard Gavaldá, and Shengbo Guo. Introduction to the special issue on Social Web mining. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1): 5:1–5:??, December 2013.

CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Barbieri:2017:EMI

[BBM17] Nicola Barbieri, Francesco Bonchi, and Giuseppe Manco. Efficient methods for influence-based network-oblivious community detection. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2):32:1– 32:??, January 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Baldoni:2013:CRS

[BBMP13] Matteo Baldoni, Cristina Baroglio, Elisa Marengo, and Viviana Patti. Constitutive and regulative specifications of commitment protocols: a decoupled approach. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (2):22:1–22:??, March 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Belem:2016:BRE

[BBS^{+}16] Fabiano M. Belém, Carolina S. Batista, Rodrygo L. T. Santos, Jussara M. Almeida, and Marcos A. Gonçalves. Beyond relevance: Explicitly promoting novelty and diversity in tag recommendation.

ACM Transactions on Intelligent Systems and Technology (TIST), 7(3):26:1– 26:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Banerjee:2019:AAR

[BC19] Suvadeep Banerjee and Abhijit Chatterjee. ALERA: Accelerated reinforcement learning driven adaptation to electro-mechanical degradation in nonlinear control systems using encoded state space error signatures. ACM Transactions on Intelligent Systems and Technology (TIST), 10(4):44:1–44:??, August 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3338123.

Baralis:2013:EPH

[BCC^{+}13] Elena Baralis, Tania Cerquitelli, Silvia Chiusano, Vincenzo D’Elia, Riccardo Molinari, and Davide Susta. Early prediction of the highest workload in incremental cardiopulmonary tests. ACM Transactions on Intelligent Systems and Technology (TIST), 4(4):70:1– 70:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Bellogin:2013:ECS

[BCD^{+}13] Alejandro Bellogín, Iván Cantador, Fernando Díez, Pablo Castells, and Enrique Chavarriaga. An empirical comparison of social, collaborative filtering, and hybrid recommenders. ACM Transactions on Intelligent Systems and Technology (TIST), 4(1):14:1– 14:??, January 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Bonchi:2011:SNA

[BCGJ11] Francesco Bonchi, Carlos Castillo, Aristides Gionis, and Alejandro Jaimes. Social network analysis and mining for business applications. ACM Transactions on Intelligent Systems and Technology (TIST), 2(3): 22:1–22:??, April 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Bozzano:2021:CAB

[BCR21] Marco Bozzano, Alessandro Cimatti, and Marco Roveri. A comprehensive approach to onboard autonomy verification and validation. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4):46:1–46:29, August 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3472715.

Bao:2011:FRC

[BD11] Xinlong Bao and Thomas G.

Dietterich. FolderPredictor: Reducing the cost of reaching the right folder. ACM Transactions on Intelligent Systems and Technology (TIST), 2(1):8:1– 8:??, January 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Baradaaji:2023:JLS

[BD23] A. Baradaaji and F. Dornaika. Joint latent space and label inference estimation with adaptive fused data and label graphs. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):62:1– 62:??, August 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3590172.

Berretti:2012:DFF

[BDP12] Stefano Berretti, Alberto Del Bimbo, and Pietro Pala. Distinguishing facial features for ethnicitybased 3D face recognition. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3):45:1– 45:??, May 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Bistaffa:2017:AGC

[BFC^{+}17] Filippo Bistaffa, Alessandro Farinelli, Jesús Cerquides, Juan Rodríguez-Aguilar, and Sarvapali D. Ramchurn. Algorithms for graph-constrained coalition formation in the real world. ACM Transactions on Intelligent Systems and Technology (TIST), 8(4):60:1– 60:??, July 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Bifet:2012:ERH

[BFHP12] Albert Bifet, Eibe Frank, Geoff Holmes, and Bernhard Pfahringer. Ensembles of restricted Hoeffding trees. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2): 30:1–30:??, February 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Biancalana:2013:ASR

[BGMS13a] Claudio Biancalana, Fabio Gasparetti, Alessandro Micarelli, and Giuseppe Sansonetti. An approach to social recommendation for context-aware mobile services. ACM Transactions on Intelligent Systems and Technology (TIST), 4(1): 10:1–10:??, January 2013.

CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Biancalana:2013:SSQ

[BGMS13b] Claudio Biancalana, Fabio Gasparetti, Alessandro Micarelli, and Giuseppe Sansonetti. Social semantic query expansion. ACM Transactions on Intelligent Systems and Technology (TIST), 4(4):60:1–60:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Berry:2011:PPA

[BGPYS11] Pauline M. Berry, Melinda Gervasio, Bart Peintner, and Neil Yorke-Smith. PTIME: Personalized assistance for calendaring. ACM Transactions on Intelligent Systems and Technology (TIST), 2(4):40:1– 40:??, July 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Bhatia:2021:ISG

[Bha21]  Munish Bhatia.  Intel-

ligent system of gametheory-based decision making in smart sports industry. ACM Transactions on Intelligent Systems and Technology (TIST), 12(3): 29:1–29:23, July 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3447986.

Ben-Israel:2017:LPM

[BI17] Isaac Ben-Israel. The letter from Prof. Maj. Gen. (Ret.) Isaac Ben-Israel. ACM Transactions on Intelligent Systems and Technology (TIST), 8(4):49:1– 49:??, July 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Bhatt:2011:PTM

[BK11] Chidansh Bhatt and Mohan Kankanhalli. Probabilistic temporal multimedia data mining. ACM Transactions on Intelligent Systems and Technology (TIST), 2(2):17:1– 17:??, February 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Benaskeur:2010:CRT

[BKB10] Abder Rezak Benaskeur, Froduald Kabanza, and Eric Beaudry. CORALS: a real-time planner for anti-air defense operations. ACM Transactions on Intelligent Systems and Technology (TIST), 1(2):13:1– 13:??, November 2010. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Borish:2016:RLC

[BL16] Michael Borish and Benjamin Lok. Rapid lowcost virtual human bootstrapping via the crowd. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):47:1– 47:??, July 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Braytee:2019:CML

[BLAK19] Ali Braytee, Wei Liu, Ali Anaissi, and Paul J. Kennedy. Correlated multilabel classification with incomplete label space and class imbalance. ACM Transactions on Intelligent Systems and Technology (TIST), 10(5):56:1–56:??, October 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Bian:2014:EUP

[BLL^{+}14] Jiang Bian, Bo Long, Lihong Li, Taesup Moon, Anlei Dong, and Yi Chang. Exploiting user preference for online learning in Web content optimization systems. ACM Transactions on Intelligent Systems and Technology (TIST), 5(2): 33:1–33:??, April 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Brock:2020:LTD

[BLNN20] Heike Brock, Felix Law, Kazuhiro Nakadai, and Yuji Nagashima. Learning three-dimensional skeleton data from sign language video. ACM Transactions on Intelligent Systems and Technology (TIST), 11(3): 30:1–30:24, May 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3377552.

Belcastro:2016:USD

[BMTT16] Loris Belcastro, Fabrizio Marozzo, Domenico Talia, and Paolo Trunfio. Using scalable data mining for predicting flight delays. ACM Transactions on Intelligent Systems and Technology (TIST), 8(1):5:1– 5:??, October 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Bouamor:2013:MPA

[BMV13] Houda Bouamor, Auréelien Max, and Anne Vilnat. Multitechnique paraphrase alignment: a contribution to pinpointing sub-sentential paraphrases. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3):44:1– 44:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Bashar:2021:ALE

[BN21a] Md Abul Bashar and Richi Nayak. Active learning for effectively finetuning transfer learning to downstream task. ACM Transactions on Intelligent Systems and Technology (TIST), 12(2):24:1– 24:24, March 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3446343.

Bouguessa:2021:BBN

[BN21b] Mohamed Bouguessa and Khaled Nouri. BiNeT-Clus: Bipartite network community detection based on transactional clustering. ACM Transactions on Intelligent Systems and Technology (TIST), 12(1):6:1– 6:26, February 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3423067.

Burnett:2013:STB

[BNS13] Chris Burnett, Timothy J.

Norman, and Katia Sycara. Stereotypical trust and bias in dynamic multiagent systems. ACM Transactions on Intelligent Systems and Technology (TIST), 4(2): 26:1–26:??, March 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Burrows:2013:PAC

[BPS13] Steven Burrows, Martin Potthast, and Benno Stein. Paraphrase acquisition via crowdsourcing and machine learning. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (3):43:1–43:??, June 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Barros:2023:NSS

[BQF^{+}23] Pedro Barros, Fabiane Queiroz, Flávio Figueiredo, Jefersson A. Dos Santos, and Heitor Ramos. A new similarity space tailored for supervised deep metric learning. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):16:1–16:??, February 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3559766.

Bouguessa:2015:IAO

[BR15] Mohamed Bouguessa and Lotfi Ben Romdhane. Identifying authorities in online communities. ACM Transactions on Intelligent Systems and Technology (TIST), 6(3):30:1– 30:??, May 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Banerjee:2020:BRB

[BRSG20] Debopriyo Banerjee, Krothapalli Sreenivasa Rao, Shamik Sural, and Niloy Ganguly. BOXREC: Recommending a Box of preferred outfits in online shopping. ACM Transactions on Intelligent Systems and Technology (TIST), 11(6):69:1–69:28, November 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3408890.

Ben-Shimon:2016:AAR

[BSRSS16] David Ben-Shimon, Lior Rokach, Guy Shani, and Bracha Shapira. Anytime algorithms for recommendation service providers. ACM Transactions on Intelligent Systems and Technology (TIST), 7(3):43:1– 43:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Bi:2013:MLA

[BSW^{+}13] Jinbo Bi, Jiangwen Sun, Yu Wu, Howard Tennen, and Stephen Armeli. A machine learning approach to college drinking prediction and risk factor identification. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (4):72:1–72:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Beigi:2020:SSG

[BTL20] Ghazaleh Beigi, Jiliang Tang, and Huan Liu. Social science-guided feature engineering: a novel approach to signed link analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 11(1):11:1– 11:27, February 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3364222.

Bian:2019:TDC

[BTTT19] Jiang Bian, Dayong Tian, Yuanyan Tang, and Dacheng Tao. Trajectory data classification: a review. ACM Transactions on Intelligent Systems and Technology (TIST), 10(4):33:1–33:??, August 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3330138.

Burt:2017:ISI

[BTVY17] Ronald Burt, Jie Tang, Michalis Vazirgiannis, and Shuang Yang. Introduction to special issue on social media processing (TIST — SMP). ACM Transactions on Intelligent Systems and Technology (TIST), 8 (6):76:1–76:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Batal:2013:TPM

[BVCH13] Iyad Batal, Hamed Valizadegan, Gregory F. Cooper, and Milos Hauskrecht. A temporal pattern mining approach for classifying electronic health record data. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (4):63:1–63:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Bryce:2010:PIB

[BVK10] Daniel Bryce, Michael Verdicchio, and Seungchan Kim. Planning interventions in biological networks. ACM Transactions on Intelligent Systems and Technology (TIST), 1(2):11:1– 11:??, November 2010. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Bai:2015:OPL

[BWC15] Aijun Bai, Feng Wu, and Xiaoping Chen. Online planning for large Markov decision processes with hierarchical decomposition. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4):45:1– 45:??, August 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Baek:2021:RMR

[BYK^{+}21] Yoonji Baek, Unil Yun, Heonho Kim, Hyoju Nam, Hyunsoo Kim, Jerry Chun-Wei Lin, Bay Vo, and Witold Pedrycz. RHUPS: Mining recent high utility patterns with sliding window-based arrival time control over data streams. ACM Transactions on Intelligent Systems and Technology (TIST), 12(2):16:1– 16:27, March 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3430767.

Bi:2022:UBN

[BZW^{+}22] Xin Bi, Chao Zhang, Fangtong Wang, Zhixun Liu, Xiangguo Zhao, Ye Yuan, and Guoren Wang. An uncertainty-based neural network for explainable trajectory segmentation. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):11:1–11:18, February 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3467978.

Bao:2022:CGE

[BZX^{+}22] Han Bao, Xun Zhou, Yiqun Xie, Yingxue Zhang, and Yanhua Li. COVID-GAN+: Estimating human mobility responses to COVID-19 through spatiotemporal generative adversarial networks with enhanced features. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):27:1– 27:23, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3481617.

Chopra:2013:RDA

[CAB^{+}13] Amit K. Chopra, Alexander Artikis, Jamal Bentahar, Marco Colombetti, Frank Dignum, Nicoletta Fornara, Andrew J. I. Jones, Munindar P. Singh, and Pinar Yolum. Research directions in agent communication. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (2):20:1–20:??, March 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chopra:2013:ISS

[CABD13] Amit K. Chopra, Alexander Artikis, Jamal Bentahar, and Frank Dignum. Introduction to the special section on agent communication. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (2):19:1–19:??, March 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2016:ISI

[CALK16] Kuan-Ta Chen, Omar Alonso, Martha Larson, and Irwin King. Introduction to the special issue on crowd in intelligent systems. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4): 44:1–44:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Choudhary:2022:SSS

[CASR22] Nurendra Choudhary, Charu C. Aggarwal, Karthik Subbian, and Chandan K. Reddy. Self-supervised short-text modeling through auxiliary context generation. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3): 51:1–51:21, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3511712.

Cruz:2013:CDV

[CBP13] Juan David Cruz, Cécile Bothorel, and François Poulet. Community detection and visualization in social networks: Integrating structural and semantic information. ACM Transactions on Intelligent Systems and Technology (TIST), 5 (1):11:1–11:??, December 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2022:DAP

[CBPG22] Chien-Lun Chen, Sara Babakniya, Marco Paolieri, and Leana Golubchik. Defending against poisoning backdoor attacks on federated meta-learning. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):76:1–76:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3523062.

Chien:2012:ISS

[CC12] Steve Chien and Amedeo Cesta. Introduction to the special section on artificial intelligence in space. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3):48:1– 48:??, May 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cortizo:2012:ISS

[CCC^{+}12] José Carlos Cortizo, Francisco Carrero, Iván Cantador, José Antonio Troyano, and Paolo Rosso. Introduction to the special section on search and mining usergenerated content. ACM Transactions on Intelligent Systems and Technology (TIST), 3(4):65:1–65:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2013:EEM

[CCG^{+}13] Tianshi Chen, Yunji Chen, Qi Guo, Zhi-Hua Zhou, Ling Li, and Zhiwei Xu. Effective and efficient microprocessor design space exploration using unlabeled design configurations. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1):20:1– 20:??, December 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Costa:2022:SCN

[CCGP22] Miguel Costa, Diogo Costa, Tiago Gomes, and Sandro Pinto. Shifting capsule networks from the cloud to the deep edge. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):105:1–105:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3544562.

Cheng:2015:HBS

[CCH15] Fan-Chieh Cheng, Bo-Hao Chen, and Shih-Chia Huang. A hybrid background subtraction method with background and foreground candidates detection. ACM Transactions on Intelligent Systems and Technology (TIST), 7(1): 7:1–7:??, October 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cheng:2018:SUM

[CCK^{+}18] Shih-Fen Cheng, Cen Chen, Thivya Kandappu, Hoong Chuin Lau, Archan Misra, Nikita Jaiman, Randy Tandriansyah, and Desmond Koh. Scalable urban mobile crowdsourcing: Handling uncertainty in worker movement. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3):26:1– 26:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cheng:2013:CDF

[CCL13] Zhiyuan Cheng, James Caverlee, and Kyumin Lee. A content-driven framework for geolocating microblog users. ACM Transactions on Intelligent Systems and Technology (TIST), 4(1):2:1–2:??, January 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Champaign:2015:EPC

[CCL15] John Champaign, Robin Cohen, and Disney Yan Lam. Empowering patients and caregivers to manage healthcare via streamlined presentation of Web objects selected by modeling learning benefits obtained by similar peers. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4):54:1– 54:??, August 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2019:DMI

[CCW^{+}19] Siming Chen, Shuai Chen, Zhenhuang Wang, Jie Liang, Yadong Wu, and Xiaoru Yuan. D-Map+: Interactive visual analysis and exploration of egocentric and event-centric information diffusion patterns in social media. ACM Transactions on Intelligent Systems and Technology (TIST), 10(1):11:1–11:??, January 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3183347.

Chikhaoui:2017:DCA

[CCWS17] Belkacem Chikhaoui, Mauricio Chiazzaro, Shengrui Wang, and Martin Sotir. Detecting communities of authority and analyzing their influence in dynamic social networks. ACM Transactions on Intelligent Systems and Technology (TIST), 8(6):82:1–82:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2015:KDR

[CCZ^{+}15] Chongyu Chen, Jianfei Cai, Jianmin Zheng, Tat Jen Cham, and Guangming Shi. Kinect depth recovery using a colorguided, region-adaptive, and depth-selective framework. ACM Transactions on Intelligent Systems and Technology (TIST), 6(2): 12:1–12:??, April 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2023:SHA

[CCZ^{+}23] Bolei Chen, Yongzheng Cui, Ping Zhong, Wang Yang, Yixiong Liang, and Jianxin Wang. STExplorer: a hierarchical autonomous exploration strategy with spatio-temporal awareness for aerial robots. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):99:1–99:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3595184.

Chen:2016:SIC

[CDGZ16] Hua Chen, Peng Ding, Zhi Geng, and Xiao-Hua Zhou. Semiparametric inference of the complier average causal effect with nonignorable missing outcomes. ACM Transactions on Intelligent Systems and Technology (TIST), 7(2): 19:1–19:??, January 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chang:2013:IRR

[CDK^{+}13] Yi Chang, Anlei Dong, Pranam Kolari, Ruiqiang Zhang, Yoshiyuki Inagaki, Fernanodo Diaz, Hongyuan Zha, and Yan Liu. Improving recency ranking using Twitter data. ACM Transactions on Intelligent Systems and Technology (TIST), 4(1):4:1–4:??, January 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cena:2013:PSA

[CDLV13] Federica Cena, Antonina Dattolo, Pasquale Lops, and Julita Vassileva. Perspectives in Semantic Adaptive Social Web. ACM Transactions on Intelligent Systems and Technology (TIST), 4(4):59:1–59:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Corno:2019:RRI

[CDR19] Fulvio Corno, Luigi De Russis, and Alberto Monge Roffarello. RecRules: Recommending IF–THEN rules for end-user development. ACM Transactions on Intelligent Systems and Technology (TIST), 10(5): 58:1–58:??, October 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Candan:2012:PMV

[CDS12] K. Selçuk Candan, Luigi Di Caro, and Maria Luisa Sapino. PhC: Multiresolution visualization and exploration of text corpora with parallel hierarchical coordinates. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):22:1– 22:??, February 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cataldi:2013:PET

[CDS13] Mario Cataldi, Luigi Di Caro, and Claudio Schifanella. Personalized emerging topic detection based on a term aging model. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1):

7:1–7:??, December 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cui:2019:STA

[CDW^{+}19] Wanqiu Cui, Junping Du, Dawei Wang, Xunpu Yuan, Feifei Kou, Liyan Zhou, and Nan Zhou. Short text analysis based on dual semantic extension and deep hashing in microblog. ACM Transactions on Intelligent Systems and Technology (TIST), 10(4):38:1–38:??, August 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3326166.

Cui:2021:MMV

[CDW^{+}21] Wanqiu Cui, Junping Du, Dawei Wang, Feifei Kou, and Zhe Xue. MVGAN: Multi-view graph attention network for social event detection. ACM Transactions on Intelligent Systems and Technology (TIST), 12(3): 27:1–27:24, July 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3447270.

Cagliero:2013:PTR

[CFG13] Luca Cagliero, Alessandro Fiori, and Luigi Grimaudo. Personalized tag recommendation based on generalized rules. ACM Transactions on Intelligent Systems and Technology (TIST), 5 (1):12:1–12:??, December 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cattafi:2011:SBP

[CGMC11] Massimiliano Cattafi, Marco Gavanelli, Michela Milano, and Paolo Cagnoli. Sustainable biomass power plant location in the Italian Emilia-Romagna region. ACM Transactions on Intelligent Systems and Technology (TIST), 2(4): 33:1–33:??, July 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cao:2018:MBA

[CGZ18] Zhiguang Cao, Hongliang Guo, and Jie Zhang. A multiagent-based approach for vehicle routing by considering both arriving on time and total travel time. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3):25:1– 25:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2023:IDM

[CGZ23] Shaohan Chen, Chuanhou Gao, and Ping Zhang. Incorporation of data-mined knowledge into black-box SVM for interpretability. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):6:1– 6:??, February 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3548775.

Chen:2010:PSI

[Che10] Yixin Chen. Preface to special issue on applications of automated planning. ACM Transactions on Intelligent Systems and Technology (TIST), 1(2): 9:1–9:??, November 2010. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2018:MQC

[CHHH18] Qin Chen, Qinmin Hu, Jimmy Xiangji Huang, and Liang He. Modeling queries with contextual snippets for information retrieval. ACM Transactions on Intelligent Systems and Technology (TIST), 9(4):47:1– 47:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2017:ECB

[CHP17] Chien-Cheng Chen, Kuo-Wei Hsu, and Wen-Chih Peng. Exploring communication behaviors of users to target potential users in mobile social networks. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (6):79:1–79:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2015:HIR

[CHY15] Bo-Hao Chen, Shih-Chia Huang, and Jian Hui Ye. Hazy image restoration by bi-histogram modification. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4):50:1– 50:??, August 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Che:2022:FMV

[CKP^{+}22] Sicong Che, Zhaoming Kong, Hao Peng, Lichao Sun, Alex Leow, Yong Chen, and Lifang He. Federated multi-view learning for private medical data integration and analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):61:1– 61:??, August 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3501816.

Cirillo:2010:HAT

[CKS10] Marcello Cirillo, Lars Karlsson, and Alessandro Saffiotti. Human-aware task planning: An application to mobile robots. ACM Transactions on Intelligent Systems and Technology (TIST), 1(2):15:1– 15:??, November 2010. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cao:2019:ATS

[CKW19] Nan Cao, Steffen Koch, and David Gotz / Yingcai Wu. ACM TIST special issue on visual analytics. ACM Transactions on Intelligent Systems and Technology (TIST), 10(1):1:1– 1:??, January 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3277019.

Chang:2011:LLS

[CL11] Chih-Chung Chang and Chih-Jen Lin. LIBSVM: a library for support vector machines. ACM Transactions on Intelligent Systems and Technology (TIST), 2 (3):27:1–27:??, April 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cohn:2013:AAS

[CL13] Trevor Cohn and Mirella Lapata. An abstractive approach to sentence compression. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (3):41:1–41:??, June 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Combi:2015:IAT

[CL15] Carlo Combi and Jiming Liu. Introduction to the ACM TIST special issue on intelligent healthcare informatics. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4): 51:1–51:??, August 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Changuel:2015:RSU

[CLBM15] Sahar Changuel, Nicolas Labroche, and Bernadette Bouchon-Meunier. Resources sequencing using automatic prerequisite– outcome annotation. ACM Transactions on Intelligent Systems and Technology (TIST), 6(1):6:1–6:??, March 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2022:OAL

[CLH^{+}22] Meng Chen, Qingjie Liu, Weiming Huang, Teng Zhang, Yixuan Zuo, and Xiaohui Yu. Originaware location prediction based on historical vehicle trajectories. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):5:1–5:18, February 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3462675.

Chen:2021:DPB

[CLL^{+}21] Jianguo Chen, Kenli Li, Keqin Li, Philip S. Yu, and Zeng Zeng. Dynamic planning of bicycle stations in dockless public bicycle-sharing system using gated graph neural network. ACM Transactions on Intelligent Systems and Technology (TIST), 12(2):25:1– 25:22, March 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3446342.

Chen:2023:DDD

[CLL23] Kuan-Chun Chen, Cheng-Te Li, and Kuo-Jung Lee. DDNAS: Discretized differentiable neural architecture search for text classification. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):88:1–88:??, October 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3610299.

Chen:2013:MRW

[CLSL13] Yu-Chih Chen, Yu-Shi Lin, Yu-Chun Shen, and Shou-De Lin. A modified random walk framework for handling negative ratings and generating explanations. ACM Transactions on Intelligent Systems and Technology (TIST), 4(1): 12:1–12:??, January 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chaudhari:2021:ASA

[CMPR21] Sneha Chaudhari, Varun Mithal, Gungor Polatkan, and Rohan Ramanath. An attentive survey of attention models. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):53:1–53:32, October 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3465055.

Chapelle:2015:SSR

[CMR15] Olivier Chapelle, Eren Manavoglu, and Romer Rosales. Simple and scalable response prediction for display advertising. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):61:1– 61:??, January 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Choi:2023:GNR

[CP23] Jeongwhan Choi and Noseong Park. Graph neural rough differential equations for traffic forecasting. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):74:1–74:??, August 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3604808.

Chen:2015:SCP

[CPHL15] Yi-Cheng Chen, Wen-Chih Peng, Jiun-Long Huang, and Wang-Chien Lee. Significant correlation pattern mining in smart homes. ACM Transactions on Intelligent Systems and Technology (TIST), 6(3):35:1– 35:??, May 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cui:2012:WSU

[CQZ^{+}12] Weiwei Cui, Huamin Qu, Hong Zhou, Wenbin Zhang, and Steve Skiena. Watch the story unfold with TextWheel: Visualization of large-scale news streams. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):20:1– 20:??, February 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cioffi-Revilla:2011:GIS

[CRRH11] Claudio Cioffi-Revilla, J. Daniel Rogers, and Atesmachew Hailegiorgis. Geographic information systems and spatial agent-based model simulations for sustainable development. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):10:1– 10:??, October 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Carmel:2012:RBN

[CRYT12] David Carmel, Haggai Roitman, and Elad Yom-Tov. On the relationship between novelty and popularity of user-generated content. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (4):69:1–69:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cai:2023:RDR

[CSA^{+}23] Mingjian Cai, Xiangjun Shen, Stanley Ebhohimhen Abhadiomhen, Yingfeng Cai, and Sirui Tian. Robust dimensionality reduction via low-rank Laplacian graph learning. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3):47:1– 47:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3582698.

Chen:2021:QAE

[CSHL21] Congliang Chen, Li Shen, Haozhi Huang, and Wei Liu. Quantized Adam with error feedback. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):56:1–56:26, October 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3470890.

Cui:2017:ACF

[CSN^{+}17] Chaoran Cui, Jialie Shen, Liqiang Nie, Richang Hong, and Jun Ma. Augmented collaborative filtering for sparseness reduction in personalized POI recommendation. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (5):71:1–71:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2013:RWM

[CST13] Bin Chen, Jian Su, and Chew Lim Tan. Random walks down the mention graphs for event coreference resolution. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (4):74:1–74:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cremonesi:2016:ISI

[CSTZ16] Paolo Cremonesi, Alan Said, Domonkos Tikk, and Michelle X. Zhou. Introduction to the special issue on recommender system benchmarking. ACM Transactions on Intelligent Systems and Technology (TIST), 7(3):38:1– 38:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2022:IBD

[CTC^{+}22] Wen-Cheng Chen, Wan-Lun Tsai, Huan-Hua Chang, Min-Chun Hu, and Wei-Ta Chu. Instant basketball defensive trajectory generation. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):3:1–3:20, February 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3460619.

Cong:2019:EUG

[CTY^{+}19] Phan Thanh Cong, Nguyen Thanh Tam, Hongzhi Yin, Bolong Zheng, Bela Stantic, and Nguyen Quoc Viet Hung. Efficient user guidance for validating participatory sensing data. ACM Transactions on Intelligent Systems and Technology (TIST), 10(4):37:1–37:??, August 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3326164.

Carmel:2012:FBT

[CUG^{+}12] David Carmel, Erel Uziel, Ido Guy, Yosi Mass, and Haggai Roitman. Folksonomybased term extraction for word cloud generation. ACM Transactions on Intelligent Systems and Technology (TIST), 3(4):60:1– 60:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ceh-Varela:2022:PEA

[CVCL22] Edgar Ceh-Varela, Huiping Cao, and Hady W. Lauw. Performance evaluation of aggregation-based group recommender systems for ephemeral groups. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):98:1– 98:??, December 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3542804.

Chen:2020:TRF

[CWC^{+}20] Lei Chen, Zhiang Wu, Jie Cao, Guixiang Zhu, and Yong Ge. Travel recommendation via fusing multi-auxiliary information into matrix factorization. ACM Transactions on Intelligent Systems and Technology (TIST), 11(2):22:1– 22:24, March 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3372118.

Chen:2015:MKM

[CWCK15] Bowei Chen, Jun Wang, Ingemar J. Cox, and Mohan S. Kankanhalli. Multikeyword multi-click advertisement option contracts for sponsored search. ACM Transactions on Intelligent Systems and Technology (TIST), 7(1):5:1–5:??, October 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chai:2022:EFM

[CWCY22] Di Chai, Leye Wang, Kai Chen, and Qiang Yang. Efficient federated matrix factorization against inference attacks. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):59:1–59:??, August 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3501812.

Castells:2015:ISI

[CWLZ15] Pablo Castells, Jun Wang, Rubén Lara, and Dell Zhang. Introduction to the special issue on diversity and discovery in recommender systems. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):52:1– 52:??, January 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2016:UCS

[CWR^{+}16] Chen Chen, Paweł W. Woźniak, Andrzej Romanowski, Mohammad Obaid, Tomasz Jaworski, Jacek Kucharski, Krzysztof Grudzień, Shengdong Zhao, and Morten Fjeld. Using crowdsourcing for scientific analysis of industrial tomographic images. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):52:1– 52:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chin:2013:CPT

[CXW^{+}13] Alvin Chin, Bin Xu, Hao Wang, Lele Chang, Hao Wang, and Lijun Zhu. Connecting people through physical proximity and physical resources at a conference. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3): 50:1–50:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2019:RVR

[CXW^{+}19] Wei Chen, Jing Xia, Xumeng Wang, Yi Wang, Jun Chen, and Liang Chang. RelationLines: Visual reasoning of egocentric relations from heterogeneous urban data. ACM Transactions on Intelligent Systems and Technology (TIST), 10 (1):2:1–2:??, January 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3200766.

Cheng:2021:SNF

[CYC^{+}21] Yuan Cheng, Yuchao Yang, Hai-Bao Chen, Ngai Wong, and Hao Yu. S3-Net: a fast scene understanding network by singleshot segmentation for autonomous driving. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):58:1–58:19, October 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3470660.

Cheong:2023:AIC

[CYC^{+}23] Chin Wang Cheong, Kejing Yin, William K. Cheung, Benjamin C. M. Fung, and Jonathan Poon. Adaptive integration of categorical and multi-relational ontologies with EHR data for medical concept embedding. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):111:1–111:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3625224.

Cheng:2016:UPI

[CYKL16] Chen Cheng, Haiqin Yang, Irwin King, and Michael R. Lyu. A unified pointof-interest recommendation framework in locationbased social networks. ACM Transactions on Intelligent Systems and Technology (TIST), 8(1):10:1– 10:??, October 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Cui:2021:DIR

[CYW^{+}21] Zeyu Cui, Feng Yu, Shu Wu, Qiang Liu, and Liang Wang. Disentangled item representation for recommender systems. ACM Transactions on Intelligent Systems and Technology (TIST), 12(2):20:1– 20:20, March 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3445811.

Chin:2018:EAN

[CYYL18] Wei-Sheng Chin, Bo-Wen Yuan, Meng-Yuan Yang, and Chih-Jen Lin. An efficient alternating Newton method for learning factorization machines. ACM Transactions on Intelligent Systems and Technology (TIST), 9(6):72:1– 72:??, November 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3230710.

Cai:2023:FLG

[CZG^{+}23] Yaoming Cai, Zijia Zhang, Pedram Ghamisi, Zhihua Cai, Xiaobo Liu, and Yao Ding. Fully linear graph convolutional networks for semi-supervised and unsupervised classification. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3):40:1– 40:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3579828.

Chin:2015:FPS

[CZJL15] Wei-Sheng Chin, Yong Zhuang, Yu-Chin Juan, and Chih-Jen Lin. A fast parallel stochastic gradient method for matrix factorization in shared memory systems. ACM Transactions on Intelligent Systems and Technology (TIST), 6 (1):2:1–2:??, March 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2022:IIN

[CZKJ22] Xiaoyu Chen, Yingyan Zeng, Sungku Kang, and Ran Jin. INN: an interpretable neural network for AI incubation in manufacturing. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):85:1–85:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3519313.

Chen:2013:WMS

[CZLS13] Chao Chen, Qiusha Zhu, Lin Lin, and Mei-Ling Shyu. Web media semantic concept retrieval via tag removal and model fusion. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (4):61:1–61:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2014:CCB

[CZP^{+}14] Yi-Cheng Chen, Wen-Yuan Zhu, Wen-Chih Peng, Wang-Chien Lee, and Suh-Yin Lee. CIM: Communitybased influence maximization in social networks.

ACM Transactions on Intelligent Systems and Technology (TIST), 5(2):25:1– 25:??, April 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Chen:2020:PPP

[CZW^{+}20] Chaochao Chen, Jun Zhou, Bingzhe Wu, Wenjing Fang, Li Wang, Yuan Qi, and Xiaolin Zheng. Practical privacy preserving POI recommendation. ACM Transactions on Intelligent Systems and Technology (TIST), 11(5):52:1–52:20, September 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3394138.

DeBoer:2016:PTA

[DB16] Patrick M. De Boer and Abraham Bernstein. PPLib: Toward the automated generation of crowd computing programs using process recombination and autoexperimentation. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):49:1– 49:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Dooms:2016:FDB

[DBDM16] Simon Dooms, Alejandro Bellogín, Toon De Pessemier, and Luc Martens. A framework for dataset benchmarking and its application to a new movie rating dataset. ACM Transactions on Intelligent Systems and Technology (TIST), 7(3):41:1– 41:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Dahmen:2021:ISA

[DC21] Jessamyn Dahmen and Diane J. Cook. Indirectly supervised anomaly detection of clinically meaningful health events from smart home data. ACM Transactions on Intelligent Systems and Technology (TIST), 12(2):18:1– 18:18, March 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3439870.

Du:2018:SMS

[DCF^{+}18] Bowen Du, Yifeng Cui, Yanjie Fu, Runxing Zhong, and Hui Xiong. Smart-Transfer: Modeling the spatiotemporal dynamics of passenger transfers for crowdedness-aware route recommendations. ACM Transactions on Intelligent Systems and Technology (TIST), 9(6):70:1– 70:??, November 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3232229.

Doherty:2015:PMT

[DCM15] Jonathan Doherty, Kevin Curran, and Paul McKevitt. Pattern matching techniques for replacing missing sections of audio streamed across wireless networks. ACM Transactions on Intelligent Systems and Technology (TIST), 6 (2):25:1–25:??, April 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Das:2022:CSF

[DCWP22] Anirban Das, Timothy Castiglia, Shiqiang Wang, and Stacy Patterson. Crosssilo federated learning for multi-tier networks with vertical and horizontal data partitioning. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):99:1–99:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3543433.

Dharejo:2021:TGT

[DDZ^{+}21] Fayaz Ali Dharejo, Farah Deeba, Yuanchun Zhou, Bhagwan Das, Munsif Ali Jatoi, Muhammad Zawish, Yi Du, and Xuezhi Wang. TWIST-GAN: Towards wavelet transform and transferred GAN for spatio-temporal single image super resolution. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):71:1–71:20, December 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3456726.

Danzinger:2023:SAI

[DGJ^{+}23] Philipp Danzinger, Tobias Geibinger, David Janneau, Florian Mischek, Nysret Musliu, and Christian Poschalko. A system for automated industrial test laboratory scheduling. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):3:1– 3:??, February 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3546871.

Damaskinos:2022:FOF

[DGK^{+}22] Georgios Damaskinos, Rachid Guerraoui, Anne-Marie Kermarrec, Vlad Nitu, Rhicheek Patra, and François Taiani. FLeet: Online federated learning via staleness awareness and performance prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):79:1–79:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3527621.

Ding:2022:MME

[DGL^{+}22] Yasan Ding, Bin Guo, Yan Liu, Yunji Liang, Haocheng Shen, and Zhiwen Yu. MetaDetector: Meta event knowledge transfer for fake news detection. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):93:1–93:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3532851.

Ding:2015:LRN

[DGZ15] Wenkui Ding, Xiubo Geng, and Xu-Dong Zhang. Learning to rank from noisy data. ACM Transactions on Intelligent Systems and Technology (TIST), 7(1):1:1– 1:??, October 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Dhar:2011:PFM

[Dha11] Vasant Dhar. Prediction in financial markets: The case for small disjuncts. ACM Transactions on Intelligent Systems and Technology (TIST), 2(3):19:1– 19:??, April 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Dong:2022:HAR

[DHF22] Junyi Dong, Qingze Huo, and Silvia Ferrari. A holistic approach for role inference and action anticipation in human teams. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):95:1– 95:??, December 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3531230.

Deng:2016:EKL

[DJI^{+}16] Zhaohong Deng, Yizhang Jiang, Hisao Ishibuchi, Kup-Sze Choi, and Shitong Wang. Enhanced knowledge-leverage-based TSK fuzzy system modeling for inductive transfer learning. ACM Transactions on Intelligent Systems and Technology (TIST), 8(1):11:1–11:??, October 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Dutta:2021:DAC

[DJNC21] Hridoy Sankar Dutta, Mayank Jobanputra, Himani Negi, and Tanmoy Chakraborty. Detecting and analyzing collusive entities on YouTube. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):64:1–64:28, October 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3477300.

Doerfel:2016:RCR

[DJS16] Stephan Doerfel, Robert Jäschke, and Gerd Stumme. The role of cores in recommender benchmarking for social bookmarking systems. ACM Transactions on Intelligent Systems and Technology (TIST), 7(3): 40:1–40:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Doo:2013:MTF

[DL13] Myungcheol Doo and Ling Liu. Mondrian tree: a fast index for spatial alarm processing. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1): 4:1–4:??, December 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Deng:2019:DMS

[DLGT19] Cheng Deng, Zhao Li, Xinbo Gao, and Dacheng Tao. Deep multi-scale discriminative networks for double JPEG compression forensics. ACM Transactions on Intelligent Systems and Technology (TIST), 10(2):20:1–20:??, February 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3301274.

Duan:2021:NMT

[DLLT21] Mingxing Duan, Kenli Li, Keqin Li, and Qi Tian. A novel multi-task tensor correlation neural network for facial attribute prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 12(1):3:1– 3:22, February 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3418285.

Dai:2023:SAT

[DLLT23] Zeyu Dai, Shengcai Liu, Qing Li, and Ke Tang. Saliency attack: Towards imperceptible black-box adversarial attack. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3):45:1– 45:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3582563.

Dang:2022:FLE

[DLWF22] Trung Kien Dang, Xiang Lan, Jianshu Weng, and Mengling Feng. Federated learning for electronic health records. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):72:1–72:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3514500.

Dai:2021:PPT

[DLY^{+}21] Tianlun Dai, Bohan Li, Ziqiang Yu, Xiangrong Tong, Meng Chen, and Gang Chen. PARP: a parallel traffic condition driven route planning model on dynamic road networks. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):73:1– 73:24, December 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3459099.

deMeo:2013:AUB

[dMFA^{+}13] Pasquale de Meo, Emilio Ferrara, Fabian Abel, Lora Aroyo, and Geert-Jan Houben. Analyzing user behavior across social sharing environments. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1):14:1– 14:??, December 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Dominguez-Martin:2023:NAF

[DMGSD23] Javier Domínguez-Martín, María J. Gómez-Silva, and Arturo De la Escalera. Neural architectures for feature embedding in person re-identification: a comparative view. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):91:1–91:??, October 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3610298.

DiNoia:2016:SSP

[DOTD16] Tommaso Di Noia, Vito Claudio Ostuni, Paolo Tomeo, and Eugenio Di Sciascio. SPrank: Semantic pathbased ranking for top-N recommendations using linked open data. ACM Transactions on Intelligent Systems and Technology (TIST), 8(1):9:1–9:??, October 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Dai:2020:STM

[DPB20] Chenglong Dai, Dechang Pi, and Stefanie I. Becker. Shapelet-transformed multichannel EEG channel selection. ACM Transactions on Intelligent Systems and Technology (TIST), 11(5):58:1–58:27, September 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3397850.

Doucette:2016:MRA

[DPC16] John A. Doucette, Graham Pinhey, and Robin Cohen. Multiagent resource allocation for dynamic task arrivals with preemption. ACM Transactions on Intelligent Systems and Technology (TIST), 8(1):3:1– 3:??, October 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Du:2019:VIR

[DPSS19] Fan Du, Catherine Plaisant, Neil Spring, and Ben Shneiderman. Visual interfaces for recommendation systems: Finding similar and dissimilar peers. ACM Transactions on Intelligent Systems and Technology (TIST), 10(1):9:1– 9:??, January 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3200490.

Dong:2018:SBU

[DSB^{+}18] Xiaowen Dong, Yoshihiko Suhara, Burçin Bozkaya, Vivek K. Singh, Bruno Lepri, and Alex ‘Sandy’ Pentland. Social bridges in urban purchase behavior. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3):33:1– 33:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ding:2011:SCD

[DSM^{+}11] Wei Ding, Tomasz F. Stepinski, Yang Mu, Lourenco Bandeira, Ricardo Ricardo, Youxi Wu, Zhenyu Lu, Tianyu Cao, and Xindong Wu. Subkilometer crater discovery with boosting and transfer learning. ACM Transactions on Intelligent Systems and Technology (TIST), 2(4): 39:1–39:??, July 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Deng:2022:EES

[DSS^{+}22] Liwei Deng, Hao Sun, Rui Sun, Yan Zhao, and Han Su. Efficient and effective similar subtrajectory search: a spatial-aware comprehension approach. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3):35:1– 35:22, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3456723.

Ding:2016:CSP

[DT16] Changxing Ding and Dacheng Tao. A comprehensive survey on pose-invariant face recognition. ACM Transactions on Intelligent Systems and Technology (TIST), 7 (3):37:1–37:??, April 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Dong:2015:NMR

[DTL15] Yongsheng Dong, Dacheng Tao, and Xuelong Li. Nonnegative multiresolution representation-based texture image classification. ACM Transactions on Intelligent Systems and Technology (TIST), 7(1):4:1– 4:??, October 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Demeshko:2016:NCS

[DWKP16] Marina Demeshko, Takashi Washio, Yoshinobu Kawahara, and Yuriy Pepyolyshev. A novel continuous and structural VAR modeling approach and its application to reactor noise analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 7(2): 24:1–24:??, January 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Deng:2023:RLP

[DYQ^{+}23] Bangchao Deng, Dingqi Yang, Bingqing Qu, Benjamin Fankhauser, and Philippe Cudre-Mauroux. Robust location prediction over sparse spatiotemporal trajectory data: Flashback to the right moment! ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):90:1– 90:??, October 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3616541.

Di:2022:FSR

[DZY^{+}22] Kai Di, Yifeng Zhou, Fuhan Yan, Jiuchuan Jiang, Shaofu Yang, and Yichuan Jiang. A foraging strategy with risk response for individual robots in adversarial environments. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):83:1– 83:??, October 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3514499.

Estlin:2012:AAS

[EBG^{+}12] Tara A. Estlin, Benjamin J. Bornstein, Daniel M. Gaines, Robert C. Anderson, David R. Thompson, Michael Burl, Rebecca Castaño, and Michele Judd. AEGIS automated science targeting for the MER Opportunity Rover. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3): 50:1–50:??, May 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Elhamod:2022:CPL

[EBS^{+}22] Mohannad Elhamod, Jie Bu, Christopher Singh, Matthew Redell, Abantika Ghosh, Viktor Podolskiy, Wei-Cheng Lee, and Anuj Karpatne. CoPhy-PGNN: Learning physicsguided neural networks with competing loss functions for solving eigenvalue problems. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):92:1–92:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3530911.

Editors:2013:ISS

[Edi13] Editors. Introduction to special section on intelligent mobile knowledge discovery and management systems. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1):1:1–1:??, December 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Eiras-Franco:2020:FDN

[EFMRK^{+}20] Carlos Eiras-Franco, David Martínez-Rego, Leslie Kanthan, César Piñeiro, Antonio Bahamonde, Bertha Guijarro-Berdiñas, and Amparo Alonso-Betanzos. Fast distributed k NN graph construction using auto-tuned localitysensitive hashing. ACM Transactions on Intelligent Systems and Technology (TIST), 11(6):71:1–71:18, November 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3408889.

Elmadany:2021:IAR

[EHG21] Nour Eldin Elmadany, Yifeng He, and Ling Guan. Improving action recognition via temporal and complementary learning. ACM Transactions on Intelligent Systems and Technology (TIST), 12(3):31:1– 31:24, July 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3447686.

Elbadrawy:2015:USF

[EK15] Asmaa Elbadrawy and George Karypis. Userspecific feature-based similarity models for top-n recommendation of new items. ACM Transactions on Intelligent Systems and Technology (TIST), 6(3):33:1– 33:??, May 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Etienne:2014:MBC

[EL14] Côme Etienne and Oukhellou Latifa. Model-based count series clustering for bike sharing system usage mining: a case study with the Vélib’ system of Paris. ACM Transactions on Intelligent Systems and Technology (TIST), 5 (3):39:1–39:??, September 2014. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ewerth:2012:RVC

[EMF12] Ralph Ewerth, Markus Mühling, and Bernd Freisleben. Robust video content analysis via transductive learning. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3): 41:1–41:??, May 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Elahi:2013:ALS

[ERR13] Mehdi Elahi, Francesco Ricci, and Neil Rubens. Active learning strategies for rating elicitation in collaborative filtering: a systemwide perspective. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1):13:1– 13:??, December 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ehara:2013:PRS

[ESNN13] Yo Ehara, Nobuyuki Shimizu, Takashi Ninomiya, and Hiroshi Nakagawa. Personalized reading support for second-language Web documents. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (2):31:1–31:??, March 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Erriquez:2013:BUS

[EvdHW13] Elisabetta Erriquez, Wiebe van der Hoek, and Michael Wooldridge. Building and using social structures: a case study using the agent ART testbed. ACM Transactions on Intelligent Systems and Technology (TIST), 4(2):25:1– 25:??, March 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Feuz:2015:TLA

[FC15] Kyle D. Feuz and Diane J. Cook. Transfer learning across feature-rich heterogeneous feature spaces via Feature-Space Remapping (FSR). ACM Transactions on Intelligent Systems and Technology (TIST), 6(1): 3:1–3:??, March 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Figueroa:2015:CAT

[FDE15] Nadia Figueroa, Haiwei Dong, and Abdulmotaleb El Saddik. A combined approach toward consistent reconstructions of indoor spaces based on 6D RGB-D odometry and KinectFusion. ACM Transactions on Intelligent Systems and Technology (TIST), 6(2): 14:1–14:??, April 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Fire:2015:DMO

[FE15] Michael Fire and Yuval Elovici. Data mining of online genealogy datasets for revealing lifespan patterns in human population. ACM Transactions on Intelligent Systems and Technology (TIST), 6(2):28:1– 28:??, April 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Fan:2017:RMA

[FGL17] Xiaoyi Fan, Wei Gong, and Jiangchuan Liu. i^{2} tag: RFID mobility and activity identification through intelligent profiling. ACM Transactions on Intelligent Systems and Technology (TIST), 9(1):5:1–5:??, October 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Farrahi:2011:DRL

[FGP11] Katayoun Farrahi and Daniel Gatica-Perez. Discovering routines from large-scale human locations using probabilistic topic models. ACM Transactions on Intelligent Systems and Technology (TIST), 2 (1):3:1–3:??, January 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Fridman:2013:UQR

[FK13] Natalie Fridman and Gal A. Kaminka. Using qualitative reasoning for social simulation of crowds. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3):54:1– 54:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Folsom-Kovarik:2013:TPR

[FKSS13] Jeremiah T. Folsom-Kovarik, Gita Sukthankar, and Sae Schatz. Tractable POMDP representations for intelligent tutoring systems. ACM Transactions on Intelligent Systems and Technology (TIST), 4(2):29:1– 29:??, March 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Fu:2020:TER

[FL20] Tao-Yang Fu and Wang-Chien Lee. Trembr: Exploring road networks for trajectory representation learning. ACM Transactions on Intelligent Systems and Technology (TIST), 11(1):10:1–10:25, February 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3361741.

Feygin:2020:BBI

[FLF^{+}20] Sidney A. Feygin, Jessica R. Lazarus, Edward H. Forscher, Valentine Golfier-Vetterli, Jonathan W. Lee, Abhishek Gupta, Rashid A. Waraich, Colin J. R. Sheppard, and Alexandre M. Bayen. BISTRO: Berkeley Integrated System for Transportation Optimization. ACM Transactions on Intelligent Systems and Technology (TIST), 11(4): 38:1–38:27, July 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3384344.

Fu:2018:MLM

[FLLX18] Yanjie Fu, Junming Liu, Xiaolin Li, and Hui Xiong. A multi-label multi-view learning framework for inapp service usage analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 9(4):40:1– 40:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Fang:2023:RAG

[FLY^{+}23] Yujie Fang, Xin Li, Rui Ye, Xiaoyan Tan, Peiyao Zhao, and Mingzhong Wang. Relation-aware graph convolutional networks for multi-relational network alignment. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2):37:1– 37:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3579827.

Freitas:2023:DLE

[FMdA^{+}23] Lucas Freitas, Valter Martins, Marilton de Aguiar, Lisane de Brisolara, and Paulo Ferreira. Deep learning embedded into smart traps for fruit insect pests detection. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):10:1–10:??, February 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3552435.

Flaxman:2016:GPI

[FNS16] Seth R. Flaxman, Daniel B. Neill, and Alexander J. Smola. Gaussian processes for independence tests with non-iid data in causal inference. ACM Transactions on Intelligent Systems and Technology (TIST), 7(2): 22:1–22:??, January 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Falcone:2013:MKR

[FPVC13] Rino Falcone, Michele Piunti, Matteo Venanzi, and Cristiano Castelfranchi. From manifesta to krypta: The relevance of categories for trusting others. ACM Transactions on Intelligent Systems and Technology (TIST), 4(2):27:1– 27:??, March 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Falcone:2013:ISS

[FS13] Rino Falcone and Munindar P. Singh. Introduction to special section on trust in multiagent systems. ACM Transactions on Intelligent Systems and Technology (TIST), 4(2): 23:1–23:??, March 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Feyisetan:2017:SIP

[FS17] Oluwaseyi Feyisetan and Elena Simperl. Social incentives in paid collaborative crowdsourcing. ACM Transactions on Intelligent Systems and Technology (TIST), 8(6):73:1–73:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Font:2015:AIT

[FSS15] Frederic Font, Joan Serrà, and Xavier Serra. Analysis of the impact of a tag recommendation system in a real-world folksonomy. ACM Transactions on Intelligent Systems and Technology (TIST), 7 (1):6:1–6:??, October 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Fang:2020:AEC

[FSW^{+}20] Xiu Susie Fang, Quan Z. Sheng, Xianzhi Wang, Wei Emma Zhang, Anne H. H. Ngu, and Jian Yang. From appearance to essence: Comparing truth discovery methods without using ground truth. ACM Transactions on Intelligent Systems and Technology (TIST), 11(6):74:1– 74:24, November 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3411749.

Feldman:2010:SCR

[FT10] Michal Feldman and Moshe Tennenholtz. Structured coalitions in resource selection games. ACM Transactions on Intelligent Systems and Technology (TIST), 1 (1):4:1–4:??, October 2010. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Fire:2013:CEL

[FTCP^{+}13] Michael Fire, Lena Tenenboim-Chekina, Rami Puzis, Ofrit Lesser, Lior Rokach, and Yuval Elovici. Computationally efficient link prediction in a variety of social networks. ACM Transactions on Intelligent Systems and Technology (TIST), 5 (1):10:1–10:??, December 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Firdaus:2021:AAR

[FTE21] Mauajama Firdaus, Nidhi Thakur, and Asif Ekbal. Aspect-aware response generation for multimodal dialogue system. ACM Transactions on Intelligent Systems and Technology (TIST), 12(2):15:1– 15:33, March 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3430752.

Fang:2022:GFC

[FWYX22] Chenglong Fang, Feng Wang, Bin Yao, and Jianqiu Xu. GPSClean: a framework for cleaning and repairing GPS data. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3):40:1– 40:22, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3469088.

Feng:2017:MHC

[FWZ17] Xiaodong Feng, Sen Wu, and Wenjun Zhou. Multihypergraph consistent sparse coding. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (6):75:1–75:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Fang:2016:SST

[FXHM16] Quan Fang, Changsheng Xu, M. Shamim Hossain, and G. Muhammad. STCAPLRS: a spatial-temporal contextaware personalized location recommendation system. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4): 59:1–59:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Fu:2017:RSD

[FXR^{+}17] Hao Fu, Xing Xie, Yong Rui, Neil Zhenqiang Gong, Guangzhong Sun, and Enhong Chen. Robust spammer detection in microblogs: Leveraging user carefulness. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (6):83:1–83:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Fire:2016:LPC

[FZ16] Amy Fire and Song-Chun Zhu. Learning perceptual causality from video. ACM Transactions on Intelligent Systems and Technology (TIST), 7(2):23:1– 23:??, January 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Fu:2021:MCE

[FZH^{+}21] Teng Fu, Guido Zampieri, David Hodgson, Claudio Angione, and Yifeng Zeng. Modeling customer experience in a contact center through process log mining. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4):48:1– 48:21, August 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3468269.

Fu:2015:ESG

[FZX15] Hao Fu, Aston Zhang, and Xing Xie. Effective social graph deanonymization based on graph structure and descriptive information. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4): 49:1–49:??, August 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gupta:2022:DML

[GB22] Vinayak Gupta and Srikanta Bedathur. Doing more with less: Overcoming data scarcity for POI recommendation via crossregion transfer. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3):50:1– 50:24, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3511711.

Gupta:2022:MCT

[GBBD22] Vinayak Gupta, Srikanta Bedathur, Sourangshu Bhattacharya, and Abir De. Modeling continuous time sequences with intermittent observations using marked temporal point processes. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):103:1–103:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3545118.

Gao:2022:LCH

[GBC^{+}22] Guangliang Gao, Zhifeng Bao, Jie Cao, A. K. Qin, and Timos Sellis. Locationcentered house price prediction: a multi-task learning approach. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):32:1– 32:25, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3501806.

Guo:2015:ISI

[GCY^{+}15] Bin Guo, Alvin Chin, Zhiwen Yu, Runhe Huang, and Daqing Zhang. An introduction to the special issue on participatory sensing and crowd intelligence. ACM Transactions on Intelligent Systems and Technology (TIST), 6(3):36:1– 36:??, May 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Guy:2011:I

[GCZ11] Ido Guy, Li Chen, and Michelle X. Zhou. Introduction. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (1):1:1–1:??, October 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Guy:2013:ISS

[GCZ13] Ido Guy, Li Chen, and Michelle X. Zhou. Introduction to the special section on social recommender systems. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (1):7:1–7:??, January 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Golpayegani:2019:USD

[GDC19] Fatemeh Golpayegani, Ivana Dusparic, and Siobhan Clarke. Using social dependence to enable neighbourly behaviour in open multi-agent systems. ACM Transactions on Intelligent Systems and Technology (TIST), 10(3):31:1– 31:??, May 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3319402.

Groves:2015:OAT

[GG15] William Groves and Maria Gini. On optimizing airline ticket purchase timing. ACM Transactions on Intelligent Systems and Technology (TIST), 7(1):3:1– 3:??, October 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gao:2023:DIT

[GG23] Lei Gao and Ling Guan.

A discriminant information theoretic learning framework for multi-modal feature representation. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3):55:1– 55:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3587253.

Guo:2021:DDH

[GGC21] Jinjin Guo, Zhiguo Gong, and Longbing Cao. dhCM: Dynamic and hierarchical event categorization and discovery for social media stream. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):57:1–57:25, October 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3470888.

Guo:2023:FCP

[GGY^{+}23] Kun Guo, Wenzhong Guo, Enjie Ye, Yutong Fang, Jiachen Zheng, Ximeng Liu, and Kai Chen. Federated clique percolation for privacy-preserving overlapping community detection. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):76:1– 76:??, August 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3604807.

Goodwin:2018:KRI

[GH18] Travis R. Goodwin and Sanda M. Harabagiu. Knowledge representations and inference techniques for medical question answering. ACM Transactions on Intelligent Systems and Technology (TIST), 9(2): 14:1–14:??, January 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gao:2017:SOL

[GHZ^{+}17] Xingyu Gao, Steven C. H. Hoi, Yongdong Zhang, Jianshe Zhou, Ji Wan, Zhenyu Chen, Jintao Li, and Jianke Zhu. Sparse online learning of image similarity. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (5):64:1–64:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gintis:2013:MMS

[Gin13] Herbert Gintis. Markov models of social dynamics: Theory and applications. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3):53:1– 53:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gedikli:2013:IRA

[GJ13] Fatih Gedikli and Dietmar Jannach. Improving recommendation accuracy based on item-specific tag preferences. ACM Transactions on Intelligent Systems and Technology (TIST), 4(1): 11:1–11:??, January 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ganesan:2017:OSC

[GJC17] Rajesh Ganesan, Sushil Jajodia, and Hasan Cam. Optimal scheduling of cybersecurity analysts for minimizing risk. ACM Transactions on Intelligent Systems and Technology (TIST), 8(4): 52:1–52:??, July 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gupta:2023:ERV

[GJS23] Trasha Gupta, Rajni Jindal, and Indu Sreedevi. Empirical review of various thermography-based computer-aided diagnostic systems for multiple diseases. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3): 56:1–56:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3583778.

Ganesan:2016:DSC

[GJSC16] Rajesh Ganesan, Sushil Jajodia, Ankit Shah, and Hasan Cam. Dynamic scheduling of cybersecurity analysts for minimizing risk using reinforcement learning. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (1):4:1–4:??, October 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gal:2011:AAN

[GKG^{+}11] Ya’akov Gal, Sarit Kraus, Michele Gelfand, Hilal Khashan, and Elizabeth Salmon. An adaptive agent for negotiating with people in different cultures. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):8:1– 8:??, October 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gurung:2014:TIP

[GLJ^{+}14] Sashi Gurung, Dan Lin, Wei Jiang, Ali Hurson, and Rui Zhang. Traffic information publication with privacy preservation. ACM Transactions on Intelligent Systems and Technology (TIST), 5(3):44:1–44:??, September 2014. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gao:2017:FSV

[GLL^{+}17] Yang Gao, Yuefeng Li, Raymond Y. K. Lau, Yue Xu, and Md Abul Bashar. Finding semantically valid and relevant topics by association-based topic selection model. ACM Transactions on Intelligent Systems and Technology (TIST), 9(1):3:1–3:??, October 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Guri:2017:BAG

[GME17] Mordechai Guri, Matan Monitz, and Yuval Elovici. Bridging the air gap between isolated networks and mobile phones in a practical cyber-attack. ACM Transactions on Intelligent Systems and Technology (TIST), 8(4):50:1– 50:??, July 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gupta:2021:SVA

[GMX^{+}21] Jayant Gupta, Carl Molnar, Yiqun Xie, Joe Knight, and Shashi Shekhar. Spatial variability aware deep neural networks (SVANN): a general approach. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):76:1–76:21, December 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3466688.

Gretarsson:2012:TVA

[GOB^{+}12] Brynjar Gretarsson, John O’Donovan, Svetlin Bostandjiev, Tobias Höllerer, Arthur Asuncion, David Newman, and Padhraic Smyth. TopicNets: Visual analysis of large text corpora with topic modeling. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):23:1– 23:??, February 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Goolsby:2010:SMC

[Goo10]  Rebecca Goolsby.  So-

cial media as crisis platform: The future of community maps/crisis maps. ACM Transactions on Intelligent Systems and Technology (TIST), 1(1):7:1– 7:??, October 2010. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gasso:2011:BOL

[GPSB11] Gilles Gasso, Aristidis Pappaioannou, Marina Spivak, and Léon Bottou. Batch and online learning algorithms for nonconvex Neyman–Pearson classification. ACM Transactions on Intelligent Systems and Technology (TIST), 2(3): 28:1–28:??, April 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gou:2019:LMR

[GQY^{+}19] Jianping Gou, Wenmo Qiu, Zhang Yi, Yong Xu, Qirong Mao, and Yongzhao Zhan. A local mean representation-based Knearest neighbor classifier. ACM Transactions on Intelligent Systems and Technology (TIST), 10(3):29:1– 29:??, May 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3319532.

Ghosh:2015:MTD

[GRR^{+}15] Siddhartha Ghosh, Steve Reece, Alex Rogers, Stephen Roberts, Areej Malibari, and Nicholas R. Jennings. Modeling the thermal dynamics of buildings: a latent-force- model-based approach. ACM Transactions on Intelligent Systems and Technology (TIST), 6 (1):7:1–7:??, March 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gerard:2013:FVP

[GS13] Scott N. Gerard and Munindar P. Singh. Formalizing and verifying protocol refinements. ACM Transactions on Intelligent Systems and Technology (TIST), 4(2):21:1– 21:??, March 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Goethals:2023:PIC

[GSM23] Sofie Goethals, Kenneth Sörensen, and David Martens. The privacy issue of counterfactual explanations: Explanation linkage attacks. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):83:1–83:??, October 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3608482.

Gabrilovich:2012:ISS

[GST12] Evgeniy Gabrilovich, Zhong Su, and Jie Tang. Introduction to the Special Section on Computational Models of Collective Intelligence in the Social Web. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (4):58:1–58:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gong:2014:JLP

[GTM^{+}14] Neil Zhenqiang Gong, Ameet Talwalkar, Lester Mackey, Ling Huang, Eui Chul Richard Shin, Emil Stefanov, Elaine (Runting) Shi, and Dawn Song. Joint link prediction and attribute inference using a social-attribute network. ACM Transactions on Intelligent Systems and Technology (TIST), 5(2):27:1– 27:??, April 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Glenski:2017:RES

[GW17] Maria Glenski and Tim Weninger. Rating effects on social news posts and comments. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (6):78:1–78:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Guo:2021:CTG

[GWD^{+}21] Bin Guo, Hao Wang, Yasan Ding, Wei Wu, Shaoyang Hao, Yueqi Sun, and Zhiwen Yu. Conditional text generation for harmonious humanmachine interaction. ACM Transactions on Intelligent Systems and Technology (TIST), 12(2):14:1– 14:50, March 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3439816.

Guan:2015:DML

[GWDJ15] Tao Guan, Yuesong Wang, Liya Duan, and Rongrong Ji. On-device mobile landmark recognition using binarized descriptor with multifeature fusion. ACM Transactions on Intelligent Systems and Technology (TIST), 7(1):12:1– 12:??, October 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gao:2023:CYF

[GWL^{+}23] Yujia Gao, Pengfei Wang, Liang Liu, Chi Zhang, and Huadong Ma. Configure your federation: Hierarchical attention-enhanced meta-learning network for personalized federated learning. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4): 63:1–63:??, August 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3591362.

Gao:2022:GAN

[GXS^{+}22] Nan Gao, Hao Xue, Wei Shao, Sichen Zhao, Kyle Kai Qin, Arian Prabowo, Mohammad Saiedur Rahaman, and Flora D.

Salim. Generative adversarial networks for spatiotemporal data: a survey. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):22:1– 22:25, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3474838.

Gunarathna:2023:RTR

[GXT^{+}23] Udesh Gunarathna, Hairuo Xie, Egemen Tanin, Shanika Karunasekera, and Renata Borovica-Gajic. Real-time road network optimization with coordinated reinforcement learning. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):72:1–72:??, August 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3603379.

Guo:2021:ROE

[GXYZ21] Pengzhan Guo, Keli Xiao, Zeyang Ye, and Wei Zhu. Route optimization via environment-aware deep network and reinforcement learning. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):74:1–74:21, December 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3461645.

Ge:2011:MLC

[GXZ^{+}11] Yong Ge, Hui Xiong, Wenjun Zhou, Siming Li, and Ramendra Sahoo. Multifocal learning for customer problem analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 2(3):24:1– 24:??, April 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Gomes:2011:ISI

[GY11] Carla Gomes and Qiang Yang. Introduction to special issue on computational sustainability. ACM Transactions on Intelligent Systems and Technology (TIST), 2(4):31:1– 31:??, July 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Guo:2022:FLP

[GYL^{+}22] Xu Guo, Han Yu, Boyang Li, Hao Wang, Pengwei Xing, Siwei Feng, Zaiqing Nie, and Chunyan Miao. Federated learning for personalized humor recognition. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):68:1–68:??, August 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3511710.

Gong:2019:MMC

[GYT19] Chen Gong, Jian Yang, and Dacheng Tao. Multi-modal curriculum learning over graphs. ACM Transactions on Intelligent Systems and Technology (TIST), 10(4): 35:1–35:??, August 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3322122.

Gupta:2021:VQB

[GZ21] Amulya Gupta and Zhu Zhang. Vector-quantization-based topic modeling. ACM Transactions on Intelligent Systems and Technology (TIST), 12(3):34:1– 34:30, July 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3450946.

Gupta:2023:NTM

[GZ23] Amulya Gupta and Zhu Zhang. Neural topic modeling via discrete variational inference. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2):23:1– 23:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3570509.

Gao:2017:ECM

[GZZY17] Yue Gao, Hanwang Zhang, Xibin Zhao, and Shuicheng Yan. Event classification in microblogs via social tracking. ACM Transactions on Intelligent Systems and Technology (TIST), 8(3): 35:1–35:??, April 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hajimirsadeghi:2012:CIL

[HAAM12] Hossein Hajimirsadeghi, Majid Nili Ahmadabadi, Babak Nadjar Araabi, and Hadi Moradi. Conceptual imitation learning in a human-robot interaction paradigm. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):28:1–28:??, February 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hografer:2022:SEP

[HASS22] Marius Hogräfer, Marco Angelini, Giuseppe Santucci, and Hans-Jörg Schulz. Steering-by-example for progressive visual analytics. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):96:1–96:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3531229.

Herdagdelen:2012:BGP

[HB12] Amaç Herdagdelen and Marco Baroni. Bootstrapping a game with a purpose for commonsense collection. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (4):59:1–59:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hegedus:2016:RDL

[HBK^{+}16] István Hegedűs, Árpád Berta, Levente Kocsis, András A. Benczúr, and Márk Jelasity. Robust decentralized low-rank matrix decomposition. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):62:1– 62:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hours:2016:CAS

[HBL16] Hadrien Hours, Ernst Biersack, and Patrick Loiseau. A causal approach to the study of TCP performance. ACM Transactions on Intelligent Systems and Technology (TIST), 7(2):25:1– 25:??, January 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hoens:2013:RMR

[HBSC13] T. Ryan Hoens, Marina Blanton, Aaron Steele, and Nitesh V. Chawla. Reliable medical recommendation systems with patient privacy. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (4):67:1–67:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Han:2013:LNS

[HCB13] Bo Han, Paul Cook, and Timothy Baldwin. Lexical normalization for social media text. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (1):5:1–5:??, January 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hai:2015:ABU

[HCCY15] Zhen Hai, Kuiyu Chang, Gao Cong, and Christopher C. Yang. An association-based unified framework for mining features and opinion words. ACM Transactions on Intelligent Systems and Technology (TIST), 6(2):26:1– 26:??, April 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Huang:2015:ARM

[HCJM15] Meiyu Huang, Yiqiang Chen, Wen Ji, and Chunyan Miao. Accurate and robust moving-object segmentation for telepresence systems. ACM Transactions on Intelligent Systems and Technology (TIST), 6 (2):17:1–17:??, April 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hou:2021:TTL

[HCRF21] Chenyu Hou, Bin Cao, Sijie Ruan, and Jing Fan. TLDS: a transfer-learning-based delivery station location selection pipeline. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4):50:1– 50:24, August 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3469084.

Hayden:2012:UCM

[HCTC12] David S. Hayden, Steve Chien, David R. Thompson, and Rebecca Castaño. Using clustering and metric learning to improve science return of remote sensed imagery. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3): 51:1–51:??, May 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hu:2022:WCK

[HCWH22] Yang Hu, Adriane Chapman, Guihua Wen, and Dame Wendy Hall. What can knowledge bring to machine learning? — a survey of low-shot learning for structured data. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3):48:1– 48:45, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3510030.

Han:2016:CHA

[HDPH16] Shuguang Han, Peng Dai, Praveen Paritosh, and David Huynh. Crowdsourcing human annotation on Web page structure: Infrastructure design and behavior-based quality control. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4): 56:1–56:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hennes:2015:MLS

[HDTG15] Daniel Hennes, Steven De Jong, Karl Tuyls, and Ya’akov (Kobi) Gal. Metastrategies in large-scale bargaining settings. ACM Transactions on Intelligent Systems and Technology (TIST), 7(1):10:1– 10:??, October 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Heckerman:2019:TAH

[Hec19] David Heckerman. Toward accounting for hidden common causes when inferring cause and effect from observational data. ACM Transactions on Intelligent Systems and Technology (TIST), 10(5):51:1– 51:??, October 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hong:2021:SRI

[HG21] Thanh Phuoc Hong and Ling Guan. A scale and rotational invariant key-point detector based on sparse coding. ACM Transactions on Intelligent Systems and Technology (TIST), 12(3): 36:1–36:19, July 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3452009.

Harel:2017:CSR

[HGE17] Yaniv Harel, Irad Ben Gal, and Yuval Elovici. Cyber security and the role of intelligent systems in addressing its challenges. ACM Transactions on Intelligent Systems and Technology (TIST), 8(4):49:1– 49:??, July 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Huang:2022:SAF

[HHJ22] Shih-Chia Huang, Quoc-Viet Hoang, and Da-Wei Jaw. Self-adaptive feature transformation networks for object detection in low luminance images. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):13:1– 13:11, February 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3480973.

Huang:2022:DRL

[HHL^{+}22] Jianbin Huang, Longji Huang, Meijuan Liu, He Li, Qinglin Tan, Xiaoke Ma, Jiangtao Cui, and De-Shuang Huang. Deep reinforcement learning-based trajectory pricing on ridehailing platforms. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3):41:1– 41:19, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3474841.

Huang:2020:SIS

[HJCK20] Shih-Chia Huang, Da-Wei Jaw, Bo-Hao Chen, and Sy-Yen Kuo. Single image snow removal using sparse representation and particle swarm optimizer. ACM Transactions on Intelligent Systems and Technology (TIST), 11(2):20:1– 20:15, March 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3372116.

Hoi:2012:ISS

[HJTZ12] Steven C. H. Hoi, Rong Jin, Jinhui Tang, and Zhi-Hua Zhou. Introduction to the special section on distance metric learning in intelligent systems. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3):52:1– 52:??, May 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Horvath:2020:CBA

[HKMN20] Gábor Horváth, Edith Kovács, Roland Molontay, and Szabolcs Nováczki. Copula-based anomaly scoring and localization for large-scale, high-dimensional continuous data. ACM Transactions on Intelligent Systems and Technology (TIST), 11(3):26:1– 26:26, May 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3372274.

Hung:2013:OBI

[HKO13] Benjamin W. K. Hung, Stephan E. Kolitz, and Asuman Ozdaglar. Optimizationbased influencing of village social networks in a counterinsurgency. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3):52:1– 52:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hoang:2017:MTB

[HL17] Tuan-Anh Hoang and Ee-Peng Lim. Modeling topics and behavior of microbloggers: an integrated approach. ACM Transactions on Intelligent Systems and Technology (TIST), 8(3): 44:1–44:??, April 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hsieh:2019:IOS

[HL19] Hsun-Ping Hsieh and Cheng-Te Li. Inferring online social ties from offline geographical activities. ACM Transactions on Intelligent Systems and Technology (TIST), 10(2):17:1–17:??, February 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3293319.

Huang:2021:SHF

[HLC^{+}21] Anbu Huang, Yang Liu, Tianjian Chen, Yongkai Zhou, Quan Sun, Hongfeng Chai, and Qiang Yang. StarFL: Hybrid federated learning architecture for smart urban computing. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4):43:1– 43:23, August 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3467956.

He:2021:TNF

[HLF^{+}21] Yifan He, Zhao Li, Lei Fu, Anhui Wang, Peng Zhang, Shuigeng Zhou, Ji Zhang, and Ting Yu. TARA-Net: a fusion network for detecting takeaway rider accidents. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):72:1–72:19, December 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3457218.

He:2013:DJS

[HLGW13] Yulan He, Chenghua Lin, Wei Gao, and Kam-Fai Wong. Dynamic joint sentiment-topic model. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1):6:1–6:??, December 2013. CODEN ????

ISSN 2157-6904 (print), 2157-6912 (electronic).

Huang:2021:THG

[HLH^{+}21] Ling Huang, Xing-Xing Liu, Shu-Qiang Huang, Chang-Dong Wang, Wei Tu, Jia-Meng Xie, Shuai Tang, and Wendi Xie. Temporal hierarchical graph attention network for traffic prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):68:1–68:21, December 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3446430.

Hsu:2011:PMC

[HLJ11] Jane Yung-Jen Hsu, Chia-Chun Lian, and Wan-Rong Jih. Probabilistic models for concurrent chatting activity recognition. ACM Transactions on Intelligent Systems and Technology (TIST), 2(1):4:1– 4:??, January 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hsieh:2014:MRT

[HLL14] Hsun-Ping Hsieh, Cheng-Te Li, and Shou-De Lin. Measuring and recommending time-sensitive routes from location-based data. ACM Transactions on Intelligent Systems and Technology (TIST), 5(3):45:1– 45:??, September 2014. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hu:2022:OBS

[HLL^{+}22] Sixu Hu, Yuan Li, Xu Liu, Qinbin Li, Zhaomin Wu, and Bingsheng He. The OARF benchmark suite: Characterization and implications for federated learning systems. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):63:1–63:??, August 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3510540.

He:2023:FFA

[HLL^{+}23] Mingkai He, Jing Lin, Jinwei Luo, Weike Pan, and Zhong Ming. FLAG: a feedback-aware local and global model for heterogeneous sequential recommendation. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):14:1–14:??, February 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3557046.

Huang:2018:QBP

[HLNL18] Michael Xuelin Huang, Jiajia Li, Grace Ngai, and Hong Va Leong. Quick bootstrapping of a personalized gaze model from realuse interactions. ACM Transactions on Intelligent Systems and Technology (TIST), 9(4):43:1– 43:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Huang:2011:LBC

[HLT11] Szu-Hao Huang, Shang-Hong Lai, and Shih-Hsien Tai. A learning-based contrarian trading strategy via a dual-classifier model. ACM Transactions on Intelligent Systems and Technology (TIST), 2(3):20:1– 20:??, April 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

He:2014:ISI

[HLY^{+}14] Qi He, Juanzi Li, Rong Yan, John Yen, and Haizheng Zhang. Introduction to the Special Issue on Linking Social Granularity and Functions. ACM Transactions on Intelligent Systems and Technology (TIST), 5(2):22:1– 22:??, April 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Herd:2019:DCR

[HM19] Benjamin C. Herd and Simon Miles. Detecting causal relationships in simulation models using intervention-based counterfactual analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 10(5):47:1–47:??, October 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Huang:2015:HMC

[HMCW15] Shanshan Huang, Jun Ma, Peizhe Cheng, and Shuaiqiang Wang. A Hybrid Multigroup CoClustering recommendation framework based on information fusion. ACM Transactions on Intelligent Systems and Technology (TIST), 6(2): 27:1–27:??, April 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Han:2023:DSA

[HMG^{+}23] Zhenyu Han, Siran Ma, Changzheng Gao, Erzhuo Shao, Yulai Xie, Yang Zhang, Lu Geng, and Yong Li. Disease simulation in airport scenario based on individual mobility model. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):84:1– 84:??, October 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3593589.

Hossain:2014:AFS

[HMS^{+}14] M. Shahriar Hossain, Manish Marwah, Amip Shah, Layne T. Watson, and Naren Ramakrishnan. AutoLCA: a framework for sustainable redesign and assessment of products. ACM Transactions on Intelligent Systems and Technology (TIST), 5(2):34:1– 34:??, April 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Horne:2020:RFN

[HNA20] Benjamin D. Horne, Jeppe Nørregaard, and Sibel Adali. Robust fake news detection over time and attack. ACM Transactions on Intelligent Systems and Technology (TIST), 11(1): 7:1–7:23, February 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:

                                        //dl.acm.org/doi/abs/
                                        10.1145/3363818.

Heath:2014:CST

[HNV14] Derrall Heath, David Norton, and Dan Ventura. Conveying semantics through visual metaphor. ACM Transactions on Intelligent Systems and Technology (TIST), 5(2):31:1– 31:??, April 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hu:2022:DET

[HQW22] Yue Hu, Ao Qu, and Dan Work. Detecting extreme traffic events via a context augmented graph autoencoder. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):101:1–101:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3539735.

Han:2022:ATP

[HQY^{+}22] Nan Han, Shaojie Qiao, Kun Yue, Jianbin Huang, Qiang He, Tingting Tang, Faliang Huang, Chunlin He, and Chang-An Yuan. Algorithms for trajectory points clustering in location-based social networks. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3): 43:1–43:29, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3480972.

Hardegger:2016:SUB

[HRCT16] Michael Hardegger, Daniel Roggen, Alberto Calatroni, and Gerhard Tröster. S-SMART: a unified Bayesian framework for simultaneous semantic mapping, activity recognition, and tracking. ACM Transactions on Intelligent Systems and Technology (TIST), 7(3):34:1– 34:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

He:2019:STA

[HS19] Suining He and Kang G. Shin. Spatio-temporal adaptive pricing for balancing mobility-on-demand networks. ACM Transactions on Intelligent Systems and Technology (TIST), 10(4):39:1–39:??, August 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3331450.

Hammedi:2022:TLO

[HSBRM22] Wided Hammedi, Sidi Mohammed Senouci, Philippe Brunet, and Metzli Ramirez-Martinez. Two-level optimization to reduce waiting time at locks in inland waterway transportation. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):91:1– 91:??, December 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3527822.

He:2022:EIS

[HSJ^{+}22] Wenchong He, Arpan Man Sainju, Zhe Jiang, Da Yan, and Yang Zhou. Earth imagery segmentation on terrain surface with limited training labels: a semi-supervised approach based on physics-guided graph co-training. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):26:1– 26:22, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3481043.

Hsu:2011:ISI

[Hsu11] Chun-Nan Hsu. Introduction to special issue on large-scale machine learning. ACM Transactions on Intelligent Systems and Technology (TIST), 2(3): 25:1–25:??, April 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hua:2011:ISI

[HTDJ11] Xian-Sheng Hua, Qi Tian, Alberto Del Bimbo, and Ramesh Jain. Introduction to the special issue on intelligent multimedia systems and technology. ACM Transactions on Intelligent Systems and Technology (TIST), 2(2):9:1– 9:??, February 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hua:2012:ISS

[HTDJ12] Xian-Sheng Hua, Qi Tian, Alberto Del Bimbo, and Ramesh Jain. Introduction to the Special Section on Intelligent Multimedia Systems and Technology Part II. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (3):39:1–39:??, May 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Huang:2020:TDE

[HTL^{+}20] Yapei Huang, Yun Tian, Zhijie Liu, Xiaowei Jin, Yanan Liu, Shifeng Zhao, and Daxin Tian. A traffic density estimation model based on crowdsourcing privacy protection. ACM Transactions on Intelligent Systems and Technology (TIST), 11(4):46:1– 46:18, July 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3391707.

Hirschprung:2015:SDD

[HTM15] Ron Hirschprung, Eran Toch, and Oded Maimon. Simplifying data disclosure configurations in a cloud computing environment. ACM Transactions on Intelligent Systems and Technology (TIST), 6(3): 32:1–32:??, May 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hirschprung:2017:AOA

[HTSC^{+}17] Ron Hirschprung, Eran Toch, Hadas Schwartz-Chassidim, Tamir Mendel, and Oded Maimon. Analyzing and optimizing access control choice architectures in online social networks. ACM Transactions on Intelligent Systems and Technology (TIST), 8(4): 57:1–57:??, July 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hu:2017:COM

[HWCL17] Han Hu, Yonggang Wen, Tat-Seng Chua, and Xuelong Li. Cost-optimized microblog distribution over geo-distributed data centers: Insights from crossmedia analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 8(3):40:1– 40:??, April 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hu:2017:VCF

[HWL^{+}17] Zhenhen Hu, Yonggang Wen, Luoqi Liu, Jianguo Jiang, Richang Hong, Meng Wang, and Shuicheng Yan. Visual classification of furniture styles. ACM Transactions on Intelligent Systems and Technology (TIST), 8(5):67:1–67:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Huang:2017:UAI

[HWT17] Chao Huang, Dong Wang, and Jun Tao. An unsupervised approach to inferring the localness of people using incomplete geotemporal online checkin data. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (6):80:1–80:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Huang:2020:MTL

[HWZL20] Jizhou Huang, Haifeng Wang, Wei Zhang, and Ting Liu. Multi-task learning for entity recommendation and document ranking in Web search. ACM Transactions on Intelligent Systems and Technology (TIST), 11(5):54:1–54:24, September 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3396501.

Hao:2023:HMA

[HXC^{+}23] Qianyue Hao, Fengli Xu, Lin Chen, Pan Hui, and Yong Li. Hierarchical multi-agent model for reinforced medical resource allocation with imperfect information. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):8:1–8:??, February 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3552436.

Hu:2022:DAF

[HXY^{+}22] Ziheng Hu, Hongtao Xie, Lingyun Yu, Xingyu Gao, Zhihua Shang, and Yongdong Zhang. Dynamicaware federated learning for face forgery video detection. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4): 57:1–57:??, August 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3501814.

Haigh:2011:RLL

[HY11] Karen Zita Haigh and Fusun Yaman. RECY-CLE: Learning looping workflows from annotated traces. ACM Transactions on Intelligent Systems and Technology (TIST), 2(4): 42:1–42:??, July 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

He:2016:STT

[HYC^{+}16] Tieke He, Hongzhi Yin, Zhenyu Chen, Xiaofang Zhou, Shazia Sadiq, and Bin Luo. A spatialtemporal topic model for the semantic annotation of POIs in LBSNs. ACM Transactions on Intelligent Systems and Technology (TIST), 8(1):12:1– 12:??, October 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

He:2022:BAB

[HYHFV22] Yulin He, Xuan Ye, Joshua Zhexue Huang, and Philippe Fournier-Viger. Bayesian attribute bagging-based extreme learning machine for high-dimensional classification and regression. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):30:1– 30:26, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3495164.

Huang:2018:CFR

[HYL^{+}18] Dingjiang Huang, Shunchang Yu, Bin Li, Steven C. H. Hoi, and Shuigeng Zhou. Combination forecasting reversion strategy for online portfolio selection. ACM Transactions on Intelligent Systems and Technology (TIST), 9(5): 58:1–58:??, July 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Hu:2021:IID

[HYL^{+}21] Chuanbo Hu, Minglei Yin, Bin Liu, Xin Li, and Yanfang Ye. Identifying illicit drug dealers on Instagram with large-scale multimodal data fusion. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):59:1– 59:23, October 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3472713.

Hong:2015:VUR

[HYZ15] Richang Hong, Shuicheng Yan, and Zhengyou Zhang. Visual understanding with RGB-D sensors: an introduction to the special issue. ACM Transactions on Intelligent Systems and Technology (TIST), 6(2):11:1– 11:??, April 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ibrahim:2016:IEM

[IVS^{+}16] Azhar Mohd Ibrahim, Ibrahim Venkat, K. G. Subramanian, Ahamad Tajudin Khader, and Philippe De Wilde. Intelligent evacuation management systems: a review. ACM Transactions on Intelligent Systems and Technology (TIST), 7 (3):36:1–36:??, April 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ji:2021:LGE

[JBLW21] Shengwei Ji, Chenyang Bu, Lei Li, and Xindong Wu. Local graph edge partitioning. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):61:1–61:25, October 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3466685.

Joseph:2014:CIB

[JCH14] Kenneth Joseph, Kathleen M. Carley, and Jason I. Hong. Check-ins in “Blau Space”: Applying Blau’s macrosociological theory to foursquare check-ins from new York city. ACM Transactions on Intelligent Systems and Technology (TIST), 5(3):46:1–46:??, September 2014. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Jiang:2022:PCC

[JCW^{+}22] Renhe Jiang, Zekun Cai, Zhaonan Wang, Chuang Yang, Zipei Fan, Quanjun Chen, Xuan Song, and Ryosuke Shibasaki. Predicting citywide crowd dynamics at big events: a deep learning system. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):21:1– 21:24, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3472300.

Jumadinova:2015:APM

[JD15] Janyl Jumadinova and Prithviraj Dasgupta. Automated pricing in a multiagent prediction market using a partially observable stochastic game. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4):48:1– 48:??, August 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ji:2015:WLM

[JGL^{+}15] Rongrong Ji, Yue Gao, Wei Liu, Xing Xie, Qi Tian, and Xuelong Li. When location meets social multimedia: a survey on vision-based recognition and mining for geo-social multimedia analytics. ACM Transactions on Intelligent Systems and Technology (TIST), 6(1): 1:1–1:??, March 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Jiang:2022:WSS

[JHK^{+}22] Zhe Jiang, Wenchong He, Marcus Stephen Kirby, Arpan Man Sainju, Shaowen Wang, Lawrence V. Stanislawski, Ethan J. Shavers, and E. Lynn Usery. Weakly supervised spatial deep learning for Earth image segmentation based on imperfect polyline labels. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):25:1– 25:20, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3480970.

Javari:2014:CBC

[JJ14] Amin Javari and Mahdi Jalili. Cluster-based collaborative filtering for sign prediction in social networks with positive and negative links. ACM Transactions on Intelligent Systems and Technology (TIST), 5(2):24:1– 24:??, April 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Javari:2015:ANR

[JJ15] Amin Javari and Mahdi Jalili. Accurate and novel recommendations: an algorithm based on popularity forecasting. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):56:1–56:??, January 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Jiang:2022:FDG

[JJKZ22] Meng Jiang, Taeho Jung, Ryan Karl, and Tong Zhao. Federated dynamic graph neural networks with secure aggregation for video-based distributed surveillance. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):56:1– 56:??, August 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3501808.

Jin:2019:LFE

[JLH19] Hai Jin, Yuanfeng Lian, and Jing Hua. Learning facial expressions with 3D mesh convolutional neural network. ACM Transactions on Intelligent Systems and Technology (TIST), 10 (1):7:1–7:??, January 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3200572.

Jin:2020:MNP

[JLJ^{+}20] Di Jin, Bingyi Li, Pengfei Jiao, Dongxiao He, Hongyu Shan, and Weixiong Zhang. Modeling with node popularities for autonomous overlapping community detection. ACM Transactions on Intelligent Systems and Technology (TIST), 11(3): 27:1–27:23, May 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3373760.

Jian:2018:EMI

[JLL18] Ling Jian, Jundong Li, and Huan Liu. Exploiting multilabel information for noise-resilient feature selection. ACM Transactions on Intelligent Systems and Technology (TIST), 9(5): 52:1–52:??, July 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Jiang:2023:REK

[JLW^{+}23] Lu Jiang, Kunpeng Liu, Yibin Wang, Dongjie Wang, Pengyang Wang, Yanjie Fu, and Minghao Yin. Reinforced explainable knowledge concept recommendation in MOOCs. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3):43:1– 43:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3579991.

Ji:2017:MSM

[JLX^{+}17] Rongrong Ji, Wei Liu, Xing Xie, Yiqiang Chen, and Jiebo Luo. Mobile social multimedia analytics in the big data era: an introduction to the special issue. ACM Transactions on Intelligent Systems and Technology (TIST), 8(3):34:1– 34:??, April 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Jiang:2013:MSB

[JPL13] Daxin Jiang, Jian Pei, and Hang Li. Mining search and browse logs for Web search: a survey. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (4):57:1–57:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Jiang:2016:CCS

[JPS^{+}16] Yexi Jiang, Chang-Shing Perng, Anca Sailer, Ignacio Silva-Lepe, Yang Zhou, and Tao Li. CSM: a cloud service marketplace for complex service acquisition. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (1):8:1–8:??, October 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Jiang:2019:SEL

[JSL^{+}19] Zhe Jiang, Arpan Man Sainju, Yan Li, Shashi Shekhar, and Joseph Knight. Spatial ensemble learning for heterogeneous geographic data with class ambiguity. ACM Transactions on Intelligent Systems and Technology (TIST), 10(4): 43:1–43:??, August 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3337798.

Jiang:2021:GCE

[JTP^{+}21] Di Jiang, Conghui Tan, Jinhua Peng, Chaotao Chen, Xueyang Wu, Weiwei Zhao, Yuanfeng Song, Yongxin Tong, Chang Liu, Qian Xu, Qiang Yang, and Li Deng. A GDPR-compliant ecosystem for speech recognition with transfer, federated, and evolutionary learning. ACM Transactions on Intelligent Systems and Technology (TIST), 12(3):30:1– 30:19, July 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3447687.

Jiang:2021:IFT

[JTS^{+}21] Di Jiang, Yongxin Tong, Yuanfeng Song, Xueyang Wu, Weiwei Zhao, Jinhua Peng, Rongzhong Lian, Qian Xu, and Qiang Yang. Industrial federated topic modeling. ACM Transactions on Intelligent Systems and Technology (TIST), 12(1):2:1–2:22, February 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3418283.

Jiang:2011:UMS

[JTZ^{+}11] Yingying Jiang, Feng Tian, Xiaolong (Luke) Zhang, Guozhong Dai, and Hongan Wang. Understanding, manipulating and searching hand-drawn concept maps. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):11:1– 11:??, October 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Jan:2020:MEB

[JV20] Zohaib Md. Jan and Brijesh Verma. Multiple elimination of base classifiers in ensemble learning using accuracy and diversity comparisons. ACM Transactions on Intelligent Systems and Technology (TIST), 11(6):67:1–67:17, November 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3405790.

Jia:2016:LPT

[JWJC16] Yantao Jia, Yuanzhuo Wang, Xiaolong Jin, and Xueqi Cheng. Location prediction: a temporalspatial Bayesian model. ACM Transactions on Intelligent Systems and Technology (TIST), 7(3):31:1– 31:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Jin:2023:DGC

[JYL^{+}23] Guangyin Jin, Huan Yan, Fuxian Li, Yong Li, and Jincai Huang. Dual graph convolution architecture search for travel time estimation. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):64:1–64:??, August 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3591361.

Ji:2012:CAS

[JYT^{+}12] Rongrong Ji, Hongxun Yao, Qi Tian, Pengfei Xu, Xiaoshuai Sun, and Xianming Liu. Context-aware semi-local feature detector. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3):44:1– 44:??, May 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Jiang:2022:SFL

[JZG22] Xue Jiang, Xuebing Zhou, and Jens Grossklags. SignDS-FL: Local differentially private federated learning with sign-based dimension selection. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):74:1–74:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3517820.

Kolomvatsos:2012:FLS

[KAH12] Kostas Kolomvatsos, Christos Anagnostopoulos, and Stathes Hadjiefthymiades. A fuzzy logic system for bargaining in information markets. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):32:1–32:??, February 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Katsimerou:2016:CEI

[KAH^{+}16] Christina Katsimerou, Joris Albeda, Alina Huldtgren, Ingrid Heynderickx, and Judith A. Redi. Crowdsourcing empathetic intelligence: The case of the annotation of EMMA database for emotion and mood recognition. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):51:1– 51:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Koutroulis:2021:KCC

[KBM^{+}21] Georgios Koutroulis, Leo Botler, Belgin Mutlu, Konrad Diwold, Kay Römer, and Roman Kern. KOM-POS: Connecting causal knots in large nonlinear time series with nonparametric regression splines. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):66:1– 66:27, October 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3480971.

Katz:2018:VEC

[KCS18] Gilad Katz, Cornelia Caragea, and Asaf Shabtai. Vertical ensemble co-training for text classification. ACM Transactions on Intelligent Systems and Technology (TIST), 9(2):21:1– 21:??, January 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Kasthuriarachchy:2023:MST

[KCSW23] Buddhika Kasthuriarachchy, Madhu Chetty, Adrian Shatte, and Darren Walls. Meaning-sensitive text data augmentation with intelligent masking. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):104:1– 104:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3623403.

Kim:2016:UCI

[KCTT16] Yubin Kim, Kevyn Collins-Thompson, and Jaime Teevan. Using the crowd to improve search result ranking and the search experience. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):50:1– 50:??, July 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Khan:2013:VOM

[KDC13] Atif Khan, John A. Doucette, and Robin Cohen. Validation of an ontological medical decision support system for patient treatment using a repository of patient data: Insights into the value of machine learning. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (4):68:1–68:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Kaminka:2018:SUP

[KF18] Gal A. Kaminka and Natalie Fridman. Simulating urban pedestrian crowds of different cultures. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3):27:1– 27:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Krause:2011:SAO

[KG11] Andreas Krause and Carlos Guestrin. Submodularity and its applications in optimized information gathering. ACM Transactions on Intelligent Systems and Technology (TIST), 2(4): 32:1–32:??, July 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Kljucaric:2023:DLI

[KG23] Luke Kljucaric and Alan D. George. Deep learning inferencing with highperformance hardware accelerators. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):68:1–68:??, August 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3594221.

Kim:2015:AAR

[KHNB15] Eunju Kim, Sumi Helal, Chris Nugent, and Mark Beattie. Analyzing activity recognition uncertainties in smart home environments. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4):52:1– 52:??, August 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Khan:2018:NIC

[KKG18] Naimul Mefraz Khan, Riadh Ksantini, and Ling Guan. A novel imagecentric approach toward direct volume rendering. ACM Transactions on Intelligent Systems and Technology (TIST), 9(4):42:1– 42:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Kang:2023:HET

[KL23] Jian Kang and Dan Lin.

Highly efficient traffic planning for autonomous vehicles to cross intersections without a stop. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2):29:1– 29:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3572034.

Kim:2017:DFN

[KLL17] Jungeun Kim, Jae-Gil Lee, and Sungsu Lim. Differential flattening: a novel framework for community detection in multi-layer graphs. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2): 27:1–27:??, January 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Kang:2022:FSS

[KLL22] Yan Kang, Yang Liu, and Xinle Liang. FedCVT: Semi-supervised vertical federated learning with cross-view training. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):64:1–64:??, August 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3510031.

Kim:2020:GCC

[KLLL20] Jungeun Kim, Jae-Gil Lee, Byung Suk Lee, and Jiajun Liu. Geosocial coclustering: a novel framework for geosocial community detection. ACM Transactions on Intelligent Systems and Technology (TIST), 11(4):45:1– 45:26, July 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3391708.

Kim:2022:NNE

[KMH22] Cheolhyeong Kim, Haeseong Moon, and Hyung Ju Hwang. NEAR: Neighborhood edge AggregatoR for graph classification. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3):45:1– 45:17, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3506714.

King:2013:ISS

[KN13] Irwin King and Wolfgang Nejdl. Introduction to the special section on Twitter and microblogging services. ACM Transactions on Intelligent Systems and Technology (TIST), 4(1):1:1– 1:??, January 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Kolman:2017:SCG

[KP17] Eyal Kolman and Benny Pinkas. Securely computing a ground speed model. ACM Transactions on Intelligent Systems and Technology (TIST), 8(4):54:1– 54:??, July 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Kulev:2018:BAI

[KPF18] Igor Kulev, Pearl Pu, and Boi Faltings. A Bayesian approach to interventionbased clustering. ACM Transactions on Intelligent Systems and Technology (TIST), 9(4):44:1– 44:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Kucuktunc:2015:DCR

[KSKÇ15] Onur Küçüktunç, Erik Saule, Kamer Kaya, and Ümit V. Çatalyürek. Diversifying citation recommendations. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4): 55:1–55:??, January 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Kyan:2015:ABD

[KSL^{+}15] Matthew Kyan, Guoyu Sun, Haiyan Li, Ling Zhong, Paisarn Muneesawang, Nan Dong, Bruce Elder, and Ling Guan. An approach to ballet dance training through MS Kinect and visualization in a CAVE virtual reality environment. ACM Transactions on Intelligent Systems and Technology (TIST), 6 (2):23:1–23:??, April 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Kleinmann:2017:ACS

[KW17] Amit Kleinmann and Avishai Wool. Automatic construction of statechart-based anomaly detection models for multi-threaded industrial control systems. ACM Transactions on Intelligent Systems and Technology (TIST), 8(4):55:1– 55:??, July 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Kim:2015:RPR

[KXZG15] Mi-Young Kim, Ying Xu, Osmar R. Zaiane, and Randy Goebel. Recognition of patient-related named entities in noisy tele-health texts. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4): 59:1–59:??, August 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Khezerlou:2017:TFA

[KZL^{+}17] Amin Vahedian Khezerlou, Xun Zhou, Lufan Li, Zubair Shafiq, Alex X. Liu, and Fan Zhang. A traffic flow approach to early detection of gathering events: Comprehensive results. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (6):74:1–74:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Khezerlou:2021:DPU

[KZL^{+}21] Amin Vahedian Khezerlou, Xun Zhou, Xinyi Li, W. Nick Street, and Yanhua Li. DILSA+: Predicting urban dispersal events through deep survival analysis with enhanced urban features. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4):49:1– 49:25, August 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3469085.

Lopes:2020:GBR

[LAS20] Ramon Lopes, Renato Assunção, and Rodrygo L. T. Santos. Graph-based recommendation meets Bayes and similarity measures. ACM Transactions on Intelligent Systems and Technology (TIST), 11(1):3:1– 3:26, February 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3356882.

Leyli-Abadi:2019:MJN

[LAsO^{+}19] Milad Leyli-Abadi, Allou samé, Latifa Oukhellou, Nicolas Cheifetz, Pierre Mandel, Cédric Féliers, and Olivier Chesneau. Mixture of joint nonhomogeneous Markov chains to cluster and model water consumption behavior sequences. ACM Transactions on Intelligent Systems and Technology (TIST), 10(6):71:1– 71:??, December 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3347452.

Luo:2022:LTS

[LBC^{+}22] Hui Luo, Zhifeng Bao, Gao Cong, J. Shane Culpepper, and Nguyen Lu Dang Khoa. Let trajectories speak out the traffic bottlenecks. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):8:1–8:21, February 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3465058.

Likhyani:2019:LSI

[LBP19] Ankita Likhyani, Srikanta Bedathur, and Deepak P. Location-specific influence quantification in location-based social networks. ACM Transactions on Intelligent Systems and Technology (TIST), 10(3): 23:1–23:??, May 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3300199.

Li:2023:ALC

[LBT23] Pan Li, Brian Brost, and Alexander Tuzhilin. Adversarial learning for cross domain recommendations. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):5:1– 5:??, February 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3548776.

Lampos:2012:NES

[LC12] Vasileios Lampos and Nello Cristianini. Nowcasting events from the social Web with statistical learning. ACM Transactions on Intelligent Systems and Technology (TIST), 3(4):72:1– 72:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2015:IAC

[LC15] Qingzhong Liu and Zhongxue Chen. Improved approaches with calibrated neighboring joint density to steganalysis and seamcarved forgery detection in JPEG images. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):63:1– 63:??, January 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2016:CDD

[LC16] Furui Liu and Laiwan Chan. Causal discovery on discrete data with extensions to mixture model. ACM Transactions on Intelligent Systems and Technology (TIST), 7(2):21:1– 21:??, January 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2020:PBU

[LCC^{+}20] Wenhe Liu, Xiaojun Chang, Ling Chen, Dinh Phung, Xiaoqin Zhang, Yi Yang, and Alexander G. Hauptmann. Pair-based uncertainty and diversity promoting early active learning for person reidentification. ACM Transactions on Intelligent Systems and Technology (TIST), 11(2):21:1–21:15, March 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3372121.

Lee:2013:CES

[LCCS13] Kyumin Lee, James Caverlee, Zhiyuan Cheng, and Daniel Z. Sui. Campaign extraction from social media. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1): 9:1–9:??, December 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2012:LIS

[LCCT12] Xueying Li, Huanhuan Cao, Enhong Chen, and Jilei Tian. Learning to infer the status of heavyduty sensors for energyefficient context-sensing. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):35:1– 35:??, February 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2017:RGR

[LCD17] Xuelong Li, Guosheng Cui, and Yongsheng Dong. Refined-graph regularizationbased nonnegative matrix factorization. ACM Transactions on Intelligent Systems and Technology (TIST), 9(1):1:1–1:??, October 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2018:DOS

[LCD18] Xuelong Li, Guosheng Cui, and Yongsheng Dong. Discriminative and orthogonal subspace constraints-based nonnegative matrix factorization. ACM Transactions on Intelligent Systems and Technology (TIST), 9 (6):65:1–65:??, November 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3229051.

Liu:2019:DVT

[LCJ^{+}19] Dongyu Liu, Weiwei Cui, Kai Jin, Yuxiao Guo, and Huamin Qu. DeepTracker: Visualizing the training process of convolutional neural networks. ACM Transactions on Intelligent Systems and Technology (TIST), 10(1):6:1– 6:??, January 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3200489.

Li:2014:LPH

[LCKY14] Nan Li, William Cushing, Subbarao Kambhampati, and Sungwook Yoon. Learning probabilistic hierarchical task networks as probabilistic context-free grammars to capture user preferences. ACM Transactions on Intelligent Systems and Technology (TIST), 5 (2):29:1–29:??, April 2014. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2019:LLA

[LCLG19] Yue Liu, Zheng Cai, Chunchen Liu, and Zhi Geng. Local learning approaches for finding effects of a specified cause and their causal paths. ACM Transactions on Intelligent Systems and Technology (TIST), 10(5):49:1–49:??, October 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2018:AEB

[LCLN18] Chen Li, William K. Cheung, Jiming Liu, and Joseph K. Ng. Automatic extraction of behavioral patterns for elderly mobility and daily routine analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 9(5): 54:1–54:??, July 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Leung:2012:ISM

[LCM^{+}12] Clement H. C. Leung, Alice W. S. Chan, Alfredo Milani, Jiming Liu, and Yuanxi Li. Intelligent social media indexing and sharing using an adaptive indexing search engine. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3):47:1– 47:??, May 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2016:MLR

[LCN^{+}16] Teng Li, Bin Cheng, Bingbing Ni, Guangchan Liu, and Shuicheng Yan. Multitask low-rank affinity graph for image segmentation and image annotation. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):65:1– 65:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2021:CMT

[LCN^{+}21] Zijian Li, Ruichu Cai, Hong Wei Ng, Marianne Winslett, Tom Z. J. Fu, Boyan Xu, Xiaoyan Yang, and Zhenjie Zhang. Causal mechanism transfer network for time series domain adaptation in mechanical systems. ACM Transactions on Intelligent Systems and Technology (TIST), 12(2):23:1– 23:21, March 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3445033.

Lagree:2017:YTS

[LCV17] Paul Lagrée, Bogdan Cautis, and Hossein Vahabi. Asyou-type social aware search. ACM Transactions on Intelligent Systems and Technology (TIST), 8(5):63:1– 63:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2023:FRF

[LCX^{+}23] Yunqi Li, Hanxiong Chen, Shuyuan Xu, Yingqiang Ge, Juntao Tan, Shuchang Liu, and Yongfeng Zhang. Fairness in recommendation: Foundations, methods, and applications. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):95:1– 95:??, October 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3610302.

Liu:2015:SPA

[LCY^{+}15] Si Liu, Qiang Chen, Shuicheng Yan, Changsheng Xu, and Hanqing Lu. Snap & Play: Auto-generated personalized find-the-difference game. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4): 65:1–65:??, January 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2018:FST

[LCY^{+}18] Wenhe Liu, Xiaojun Chang, Yan Yan, Yi Yang, and Alexander G. Hauptmann. Few-shot text and image classification via analogical transfer learning. ACM Transactions on Intelligent Systems and Technology (TIST), 9(6):71:1– 71:??, November 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3230709.

Liu:2022:GSE

[LCY^{+}22] Zelei Liu, Yuanyuan Chen, Han Yu, Yang Liu, and Lizhen Cui. GTG-Shapley: Efficient and accurate participant contribution evaluation in federated learning. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):60:1– 60:??, August 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3501811.

Liu:2023:UUK

[LDFL23] Yu Liu, Jingtao Ding, Yanjie Fu, and Yong Li. UrbanKG: an urban knowledge graph system. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):60:1–60:??, August 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3588577.

Luo:2016:IMD

[LDTX16] Tie Luo, Sajal K. Das, Hwee Pink Tan, and Lirong Xia. Incentive mechanism design for crowdsourcing: an all-pay auction approach. ACM Transactions on Intelligent Systems and Technology (TIST), 7(3): 35:1–35:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Lin:2023:CCD

[LFG^{+}23] Zhuoyi Lin, Lei Feng, Xingzhi Guo, Yu Zhang, Rui Yin, Chee Keong Kwoh, and Chi Xu. COMET: Convolutional dimension interaction for collaborative filtering. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):59:1–59:??, August 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3588576.

Lyu:2020:HPL

[LFL^{+}20] Gengyu Lyu, Songhe Feng, Yidong Li, Yi Jin, Guojun Dai, and Congyan Lang. HERA: Partial label learning by combining heterogeneous loss with sparse and low-rank regularization. ACM Transactions on Intelligent Systems and Technology (TIST), 11(3): 34:1–34:19, May 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3379501.

Lyu:2023:RLL

[LFL^{+}23] Gengyu Lyu, Songhe Feng, Wei Liu, Shuoyan Liu, and Congyan Lang. Redundant label learning via subspace representation and global disambiguation. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):15:1–15:??, February 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3558547.

Lin:2023:MAU

[LFW23] Rui Lin, Jing Fan, and Haifeng Wu. Multi-aspect understanding with cooperative graph attention networks for medical dialogue information extraction. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):103:1–103:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3620675.

Lyu:2023:PKC

[LFWY23] Gengyu Lyu, Songhe Feng, Shaokai Wang, and Zhen Yang. Prior knowledge constrained adaptive graph framework for partial label learning. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2):25:1– 25:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3569421.

Li:2022:CFP

[LFY^{+}22] Fuxian Li, Jie Feng, Huan Yan, Depeng Jin, and Yong Li. Crowd flow prediction for irregular regions with semantic graph attention network. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):81:1–81:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3501805.

Luo:2016:BDI

[LG16] Peng Luo and Zhi Geng. Bounds on direct and indirect effects of treatment on a continuous endpoint. ACM Transactions on Intelligent Systems and Technology (TIST), 7(2):20:1– 20:??, January 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2021:DLT

[LGJ^{+}21] Shuo Liu, Mingliang Gao, Vijay John, Zheng Liu, and Erik Blasch. Deep learning thermal image translation for night vision perception. ACM Transactions on Intelligent Systems and Technology (TIST), 12(1):9:1– 9:18, February 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3426239.

Lu:2022:MMC

[LGJ^{+}22] Bin Lu, Xiaoying Gan, Haiming Jin, Luoyi Fu, Xinbing Wang, and Haisong Zhang. Make more connections: Urban traffic flow forecasting with spatiotemporal adaptive gated graph convolution network. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):28:1– 28:25, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3488902.

Li:2016:MCA

[LGL^{+}16] Zhifeng Li, Dihong Gong, Qiang Li, Dacheng Tao, and Xuelong Li. Mutual component analysis for heterogeneous face recognition. ACM Transactions on Intelligent Systems and Technology (TIST), 7(3): 28:1–28:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2022:DPG

[LGL^{+}22] Ke Li, Bin Guo, Jiaqi Liu, Jiangtao Wang, Haoyang Ren, Fei Yi, and Zhiwen Yu. Dynamic probabilistic graphical model for progressive fake news detection on social media platform. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):86:1– 86:??, October 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3523060.

Li:2017:MAO

[LGZ^{+}17] Zhifeng Li, Dihong Gong, Kai Zhu, Dacheng Tao, and Xuelong Li. Multifeature anisotropic orthogonal Gaussian process for automatic age estimation. ACM Transactions on Intelligent Systems and Technology (TIST), 9(1):2:1– 2:??, October 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2021:MTA

[LGZ^{+}21] Yan Liu, Bin Guo, Daqing Zhang, Djamal Zeghlache, Jingmin Chen, Sizhe Zhang, Dan Zhou, Xinlei Shi, and Zhiwen Yu. Meta-Store: a task-adaptative meta-learning model for optimal store placement with multi-city knowledge transfer. ACM Transactions on Intelligent Systems and Technology (TIST), 12(3): 28:1–28:23, July 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3447271.

Lerman:2012:USM

[LH12] Kristina Lerman and Tad Hogg. Using stochastic models to describe and predict social dynamics of Web users. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (4):62:1–62:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Lin:2022:TTP

[LH22] Fandel Lin and Hsun-Ping Hsieh. Traveling transporter problem: Arranging a new circular route in a public transportation system based on heterogeneous nonmonotonic urban data. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3):49:1– 49:25, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3510034.

Lu:2013:SBA

[LHC^{+}13] Qiang Lu, Ruoyun Huang, Yixin Chen, You Xu, Weixiong Zhang, and Guoliang Chen. A SAT-based approach to cost-sensitive temporally expressive planning. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1): 18:1–18:??, December 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2011:CCD

[LHG11] Bin Li, Steven C. H. Hoi, and Vivekanand Gopalkrishnan. CORN: Correlationdriven nonparametric learning approach for portfolio selection. ACM Transactions on Intelligent Systems and Technology (TIST), 2 (3):21:1–21:??, April 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2011:MMM

[LHJ^{+}11] Zhenhui Li, Jiawei Han, Ming Ji, Lu-An Tang, Yintao Yu, Bolin Ding, Jae-Gil Lee, and Roland Kays. MoveMine: Mining moving object data for discovery of animal movement patterns. ACM Transactions on Intelligent Systems and Technology (TIST), 2(4):37:1– 37:??, July 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Leiva:2023:DUS

[LHO23] Luis A. Leiva, Asutosh Hota, and Antti Oulasvirta. Describing UI screenshots in natural language. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):19:1–19:??, February 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3564702.

Li:2013:SAM

[LHS^{+}13] Xi Li, Weiming Hu, Chunhua Shen, Zhongfei Zhang, Anthony Dick, and Anton Van Den Hengel. A survey of appearance models in visual object tracking. ACM Transactions on Intelligent Systems and Technology (TIST), 4(4):58:1–58:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2018:CDR

[LHS18] Cheng-Te Li, Chia-Tai Hsu, and Man-Kwan Shan. A cross-domain recommendation mechanism for coldstart users based on partial least squares regression. ACM Transactions on Intelligent Systems and Technology (TIST), 9(6):67:1– 67:??, November 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3231601.

Liu:2013:STC

[LHZ13] Nathan N. Liu, Luheng He, and Min Zhao. Social temporal collaborative ranking for context aware movie recommendation. ACM Transactions on Intelligent Systems and Technology (TIST), 4(1):15:1– 15:??, January 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2022:FFA

[LHZ22a] Cheng-Te Li, Cheng Hsu, and Yang Zhang. FairSR: Fairness-aware sequential recommendation through multi-task learning with preference graph embeddings. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):16:1–16:21, February 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3495163.

Liu:2022:FMT

[LHZ^{+}22b] Yijing Liu, Dongming Han, Jianwei Zhang, Haiyang Zhu, Mingliang Xu, and Wei Chen. Federated multitask graph learning. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):80:1–80:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3527622.

Ling:2011:ISI

[Lin11]  Charles X. Ling.  In-

troduction to special issue on machine learning for business applications. ACM Transactions on Intelligent Systems and Technology (TIST), 2(3):18:1– 18:??, April 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liao:2011:MCS

[LJC^{+}11] Zhen Liao, Daxin Jiang, Enhong Chen, Jian Pei, Huanhuan Cao, and Hang Li. Mining concept sequences from large-scale search logs for contextaware query suggestion. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):17:1– 17:??, October 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2017:PMR

[LJL^{+}17] Yang Li, Jing Jiang, Ting Liu, Minghui Qiu, and Xiaofei Sun. Personalized microtopic recommendation on microblogs. ACM Transactions on Intelligent Systems and Technology (TIST), 8(6):77:1–77:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liang:2019:CTB

[LJLZ19] Haoran Liang, Ming Jiang, Ronghua Liang, and Qi Zhao. CapVis: Toward better understanding of visualverbal saliency consistency. ACM Transactions on Intelligent Systems and Technology (TIST), 10(1):10:1– 10:??, January 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3200767.

Lee:2013:SFI

[LKD13] Yugyung Lee, Saranya Krishnamoorthy, and Deendayal Dinakarpandian. A semantic framework for intelligent matchmaking for clinical trial eligibility criteria. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (4):71:1–71:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2016:VFE

[LLDT16] Xiaoyan Li, Tongliang Liu, Jiankang Deng, and Dacheng Tao. Video face editing using temporalspatial-smooth warping. ACM Transactions on Intelligent Systems and Technology (TIST), 7(3):32:1– 32:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2019:CSD

[LLF^{+}19] Zun Li, Congyan Lang, Jiashi Feng, Yidong Li, Tao Wang, and Songhe Feng. Co-saliency detection with graph matching.

ACM Transactions on Intelligent Systems and Technology (TIST), 10(3):22:1– 22:??, May 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3313874.

Li:2016:OSC

[LLL^{+}16] Jiuyong Li, Thuc Duy Le, Lin Liu, Jixue Liu, Zhou Jin, Bingyu Sun, and Saisai Ma. From observational studies to causal rule mining. ACM Transactions on Intelligent Systems and Technology (TIST), 7(2): 14:1–14:??, January 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2018:PAT

[LLL^{+}18] Jie Liu, Bin Liu, Yanchi Liu, Huipeng Chen, Lina Feng, Hui Xiong, and Yalou Huang. Personalized air travel prediction: a multifactor perspective. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3):30:1– 30:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Lai:2021:DEF

[LLL21] Chih-Te Lai, Cheng-Te Li, and Shou-De Lin. Deep energy factorization model for demographic prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 12(1):8:1– 8:16, February 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3426240.

Li:2023:LPP

[LLL23] Chu-Chen Li, Cheng-Te Li, and Shou-De Lin. Learning privacy-preserving embeddings for image data to be published. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):105:1–105:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3623404.

Li:2019:SDS

[LLLC19] Jin Li, Tong Li, Zheli Liu, and Xiaofeng Chen. Secure deduplication system with active key update and its application in IoT. ACM Transactions on Intelligent Systems and Technology (TIST), 10(6):69:1– 69:??, December 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3356468.

Liu:2020:SDA

[LLPS20] Rui Liu, Runze Liu, Andrea Pugliese, and V. S.

Subrahmanian. STARS: Defending against sockpuppetbased targeted attacks on reviewing systems. ACM Transactions on Intelligent Systems and Technology (TIST), 11(5):56:1–56:25, September 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3397463.

Li:2021:PPC

[LLS^{+}21] Shilei Li, Meng Li, Jiongming Su, Shaofei Chen, Zhimin Yuan, and Qing Ye. PP-PG: Combining parameter perturbation with policy gradient methods for effective and efficient explorations in deep reinforcement learning. ACM Transactions on Intelligent Systems and Technology (TIST), 12(3):35:1– 35:21, July 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3452008.

Li:2022:DST

[LLS^{+}22] He Li, Xuejiao Li, Liangcai Su, Duo Jin, Jianbin Huang, and Deshuang Huang. Deep spatiotemporal adaptive 3D convolutional neural networks for traffic flow prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):19:1– 19:21, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3510829.

Li:2013:ISS

[LLWC13] Qing Li, Xiangfeng Luo, Liu Wenyin, and Cristina Conati. Introduction to the special section on intelligent tutoring and coaching systems. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (2):28:1–28:??, March 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Lin:2020:CDM

[LLX^{+}20] Adi Lin, Jie Lu, Junyu Xuan, Fujin Zhu, and Guangquan Zhang. A causal Dirichlet mixture model for causal inference from observational data. ACM Transactions on Intelligent Systems and Technology (TIST), 11(3):33:1– 33:29, May 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3379500.

Lu:2022:GSN

[LLX^{+}22] Zhilong Lu, Weifeng Lv, Zhipu Xie, Bowen Du, Guixi Xiong, Leilei Sun, and Haiquan Wang. Graph sequence neural network with an attention mechanism for traffic speed prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 13 (2):20:1–20:24, April 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3470889.

Li:2012:ERQ

[LLY12] Xiaonan Li, Chengkai Li, and Cong Yu. Entityrelationship queries over Wikipedia. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (4):70:1–70:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2023:YHY

[LLZ^{+}23] Tong Li, Yong Li, Mingyang Zhang, Sasu Tarkoma, and Pan Hui. You are how you use apps: User profiling based on spatiotemporal app usage behavior. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):71:1– 71:??, August 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3597212.

Liu:2017:IVL

[LLZW17] Yan Liu, Yang Liu, Shenghua Zhong, and Songtao Wu.

Implicit visual learning: Image recognition via dissipative learning model. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2):31:1– 31:??, January 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Lipczak:2011:ETR

[LM11] Marek Lipczak and Evangelos Milios. Efficient tag recommendation for reallife data. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (1):2:1–2:??, October 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Leiva:2016:GGG

[LMAP16] Luis A. Leiva, Daniel Martín-Albo, and Réjean Plamondon. Gestures à go go: Authoring synthetic human-like stroke gestures using the kinematic theory of rapid movements. ACM Transactions on Intelligent Systems and Technology (TIST), 7(2):15:1– 15:??, January 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Lee:2015:WWR

[LMC^{+}15] Kyumin Lee, Jalal Mahmud, Jilin Chen, Michelle Zhou, and Jeffrey Nichols.

Who will retweet this? Detecting strangers from Twitter to retweet information. ACM Transactions on Intelligent Systems and Technology (TIST), 6(3): 31:1–31:??, May 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Lee:2013:CPR

[LMWS13] Suk Jin Lee, Yuichi Motai, Elisabeth Weiss, and Shumei S. Sun. Customized prediction of respiratory motion with clustering from multiple patient interaction. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (4):69:1–69:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2010:IAT

[LN10] Huan Liu and Dana Nau.

Introduction to the ACM TIST special issue AI in social computing and cultural modeling. ACM Transactions on Intelligent Systems and Technology (TIST), 1 (1):2:1–2:??, October 2010. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2011:I

[LN11] Huan Liu and Dana Nau.

Introduction. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):6:1–6:??, October 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Lucchese:2018:XCL

[LNO^{+}18] Claudio Lucchese, Franco Maria Nardini, Salvatore Orlando, Raffaele Perego, Fabrizio Silvestri, and Salvatore Trani. X-CLEaVER: Learning ranking ensembles by growing and pruning trees. ACM Transactions on Intelligent Systems and Technology (TIST), 9 (6):62:1–62:??, November 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3205453.

Li:2022:AWR

[LNYV22] Shenghui Li, Edith Ngai, Fanghua Ye, and Thiemo Voigt. Auto-weighted robust federated learning with corrupted data sources. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):73:1–73:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3517821.

Li:2022:STC

[LPL^{+}22] Qian Li, Hao Peng, Jianxin Li, Congying Xia, Renyu Yang, Lichao Sun, Philip S. Yu, and Lifang He. A survey on text classification: From traditional to deep learning. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):31:1– 31:41, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3495162.

Li:2020:CPC

[LPM20] Lin Li, Weike Pan, and Zhong Ming. CoFi-points: Collaborative filtering via pointwise preference learning on user/item-set. ACM Transactions on Intelligent Systems and Technology (TIST), 11(4):41:1– 41:24, July 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3389127.

Law:2019:TLA

[LPR19] Stephen Law, Brooks Paige, and Chris Russell. Take a look around: Using street view and satellite images to estimate house prices. ACM Transactions on Intelligent Systems and Technology (TIST), 10(5): 54:1–54:??, October 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Loffler:2022:DSM

[LRD^{+}22] Christoffer Löffler, Luca Reeb, Daniel Dzibela, Robert Marzilger, Nicolas Witt, Björn M. Eskofier, and Christopher Mutschler. Deep Siamese metric learning: a highly scalable approach to searching unordered sets of trajectories. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):6:1– 6:23, February 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3465057.

Leskovec:2016:SGP

[LS16] Jure Leskovec and Rok Sosic. SNAP: a generalpurpose network analysis and graph-mining library. ACM Transactions on Intelligent Systems and Technology (TIST), 8(1):1:1– 1:??, October 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2011:NJD

[LSQ11] Qingzhong Liu, Andrew H. Sung, and Mengyu Qiao. Neighboring joint densitybased JPEG steganalysis. ACM Transactions on Intelligent Systems and Technology (TIST), 2(2):16:1– 16:??, February 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Luo:2020:ECN

[LSW^{+}20] Ping Luo, Kai Shu, Junjie Wu, Li Wan, and Yong Tan. Exploring correlation network for cheating detection. ACM Transactions on Intelligent Systems and Technology (TIST), 11(1): 12:1–12:23, February 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:

//dl.acm.org/doi/abs/
10.1145/3364221.

Li:2023:SDH

[LSW23] Jia Li, Dandan Song, and Zhijing Wu. A semantically driven hybrid network for unsupervised entity alignment. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2): 20:1–20:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3567829.

Lu:2018:SPA

[LSZH18] Jing Lu, Doyen Sahoo, Peilin Zhao, and Steven C. H. Hoi. Sparse passive-aggressive learning for bounded online kernel methods. ACM Transactions on Intelligent Systems and Technology (TIST), 9(4):45:1–45:??, February 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2020:LUR

[LT20] Pan Li and Alexander Tuzhilin. Latent unexpected recommendations. ACM Transactions on Intelligent Systems and Technology (TIST), 11(6):70:1– 70:25, November 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3404855.

Liu:2015:LSB

[LTS^{+}15] Fan Liu, Jinhui Tang, Yan Song, Liyan Zhang, and Zhenmin Tang. Local structure-based sparse representation for face recognition. ACM Transactions on Intelligent Systems and Technology (TIST), 7 (1):2:1–2:??, October 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2016:MNS

[LTW^{+}16] Zechao Li, Jinhui Tang, Xueming Wang, Jing Liu, and Hanqing Lu. Multimedia news summarization in search. ACM Transactions on Intelligent Systems and Technology (TIST), 7(3): 33:1–33:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2018:FCD

[LWC^{+}18] Qi Liu, Runze Wu, Enhong Chen, Guandong Xu, Yu Su, Zhigang Chen, and Guoping Hu. Fuzzy cognitive diagnosis for modelling examinee performance. ACM Transactions on Intelligent Systems and Technology (TIST), 9(4): 48:1–48:??, February 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2023:TAC

[LWF^{+}23] Haochen Liu, Yiqi Wang, Wenqi Fan, Xiaorui Liu, Yaxin Li, Shaili Jain, Yunhao Liu, Anil Jain, and Jiliang Tang. Trustworthy AI: a computational perspective. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):4:1–4:??, February 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3546872.

Li:2012:MRC

[LWH12] Peipei Li, Xindong Wu, and Xuegang Hu. Mining recurring concept drifts with limited labeled streaming data. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2): 29:1–29:??, February 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2020:JNM

[LWH^{+}20] Junwei Li, Le Wu, Richang Hong, Kun Zhang, Yong Ge, and Yan Li. A joint neural model for user behavior prediction on social networking platforms. ACM Transactions on Intelligent Systems and Technology (TIST), 11(6):72:1– 72:25, November 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3406540.

Liang:2022:CFS

[LWLG22] Weichao Liang, Zhiang Wu, Zhe Li, and Yong Ge. CrimeTensor: Fine-scale crime prediction via tensor learning with spatiotemporal consistency. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):33:1– 33:24, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3501807.

u:2023:GBA

[LWW^{+}23] Shuaiyi L(y)u, Kai Wang, Yuliang Wei, Hongri Liu, Qilin Fan, and Bailing Wang. GNN-based advanced feature integration for ICS anomaly detection. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):106:1–106:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3620676.

Liu:2011:TWC

[LWWL11] Zhanyi Liu, Haifeng Wang, Hua Wu, and Sheng Li. Two-word collocation extraction using monolingual word alignment method. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):16:1– 16:??, October 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2020:SRM

[LWWX20] Hui Liu, Haiou Wang, Yan Wu, and Lei Xing. Superpixel region merging based on deep network for medical image segmentation. ACM Transactions on Intelligent Systems and Technology (TIST), 11(4):39:1– 39:22, July 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3386090.

Li:2023:RLE

[LWZ^{+}23] Xing Li, Wei Wei, Ruizhi Zhang, Zhenyu Shi, Zhiming Zheng, and Xiangnan Feng. Representation learning of enhanced graphs using random walk graph convolutional network. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3): 46:1–46:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3582841.

Liu:2013:NER

[LWZZ13] Xiaohua Liu, Furu Wei, Shaodian Zhang, and Ming Zhou. Named entity recognition for tweets. ACM Transactions on Intelligent Systems and Technology (TIST), 4(1):3:1–3:??, January 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Lian:2014:MCH

[LX14] Defu Lian and Xing Xie. Mining check-in history for personalized location naming. ACM Transactions on Intelligent Systems and Technology (TIST), 5(2): 32:1–32:??, April 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Levy:2020:SLN

[LXBW20] Sharon Levy, Wenhan Xiong, Elizabeth Belding, and William Yang Wang. SafeRoute: Learning to navigate streets safely in an urban environment. ACM Transactions on Intelligent Systems and Technology (TIST), 11(6):66:1–66:17, November 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3402818.

Li:2021:NED

[LXC^{+}21] Xingjian Li, Haoyi Xiong, Zeyu Chen, Jun Huan, Cheng-Zhong Xu, and Dejing Dou. “In-Network Ensemble”: Deep ensemble learning with diversified knowledge distillation. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):63:1– 63:19, October 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3473464.

Liu:2020:UGA

[LXJ^{+}20] Zhuang Liu, Keli Xiao, Bo Jin, Kaiyu Huang, Degen Huang, and Yunxia Zhang. Unified generative adversarial networks for multiple-choice oriented machine comprehension. ACM Transactions on Intelligent Systems and Technology (TIST), 11(3):25:1– 25:20, May 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3372120.

Liu:2018:HPC

[LXM^{+}18] Xiaobai Liu, Qian Xu, Yadong Mu, Jiadi Yang, Liang Lin, and Shuicheng Yan. High-precision camera localization in scenes with repetitive patterns. ACM Transactions on Intelligent Systems and Technology (TIST), 9(6):66:1– 66:??, November 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2023:LRS

[LXW^{+}23] Tong Li, Yanxin Xi, Huandong Wang, Yong Li, Sasu Tarkoma, and Pan Hui. Learning representations of satellite imagery by leveraging point-of-interests. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):61:1– 61:??, August 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3589344.

Liu:2022:FSR

[LYF^{+}22] Zhiwei Liu, Liangwei Yang, Ziwei Fan, Hao Peng, and Philip S. Yu. Federated social recommendation with graph neural network. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):55:1– 55:??, August 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3501815.

Lu:2023:RLA

[LYKS23] Sidi Lu, Xin Yuan, Aggelos K. Katsaggelos, and Weisong Shi. Reinforcement learning for adaptive video compressive sensing. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):81:1– 81:??, October 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3608479.

Luo:2023:HLR

[LYLX23] Qilun Luo, Ming Yang, Wen Li, and Mingqing Xiao. Hyper-Laplacian regularized multi-view clustering with exclusive L21 regularization and tensor log-determinant minimization approach. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3):53:1– 53:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3587034.

Ling:2019:BBM

[LYW^{+}19] Zhaolong Ling, Kui Yu, Hao Wang, Lin Liu, Wei Ding, and Xindong Wu.

BAMB: a balanced Markov blanket discovery approach to feature selection. ACM Transactions on Intelligent Systems and Technology (TIST), 10(5):52:1–52:??, October 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2020:DNC

[LYWH20] Xueliang Liu, Xun Yang, Meng Wang, and Richang Hong. Deep neighborhood component analysis for visual similarity modeling. ACM Transactions on Intelligent Systems and Technology (TIST), 11(3):29:1– 29:15, May 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3375787.

Liu:2018:MSM

[LYWW18] Qiang Liu, Feng Yu, Shu Wu, and Liang Wang. Mining significant microblogs for misinformation identification: an attention-based approach. ACM Transactions on Intelligent Systems and Technology (TIST), 9 (5):50:1–50:??, July 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2023:JER

[LYZ^{+}23] Qibin Li, Nianmin Yao, Nai Zhou, Jian Zhao, and Yanan Zhang. A joint entity and relation extraction model based on efficient sampling and explicit interaction. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):77:1–77:??, October 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3604811.

Li:2018:EUB

[LZ18] Liangda Li and Hongyuan Zha. Energy usage behavior modeling in energy disaggregation via Hawkes processes. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3): 36:1–36:??, February 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Li:2023:RBE

[LZC23] Lei Li, Yongfeng Zhang, and Li Chen. On the relationship between explanation and recommendation: Learning to rank explanations for improved performance. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2): 21:1–21:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3569423.

Liu:2012:ISS

[LZCQ12] Shixia Liu, Michelle X. Zhou, Giuseppe Carenini, and Huamin Qu. Introduction to the special section on intelligent visual interfaces for text analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):19:1– 19:??, February 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2011:PPL

[LZCS11] Zhiyuan Liu, Yuzhou Zhang, Edward Y. Chang, and Maosong Sun. PLDA+: Parallel latent Dirichlet allocation with data placement and pipeline processing. ACM Transactions on Intelligent Systems and Technology (TIST), 2(3): 26:1–26:??, April 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2012:TIT

[LZP^{+}12] Shixia Liu, Michelle X. Zhou, Shimei Pan, Yangqiu Song, Weihong Qian, Weijia Cai, and Xiaoxiao Lian. TIARA: Interactive, topicbased visual text summarization and analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):25:1– 25:??, February 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Liu:2023:RFS

[LZW^{+}23] Tianying Liu, Lu Zhang, Yang Wang, Jihong Guan, Yanwei Fu, Jiajia Zhao, and Shuigeng Zhou. Recent few-shot object detection algorithms: a survey with performance comparison. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):66:1– 66:??, August 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3593588.

Luo:2016:TUA

[LZY^{+}16] Chen Luo, Jia Zeng, Mingxuan Yuan, Wenyuan Dai, and Qiang Yang. Telco user activity level prediction with massive mobile broadband data. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):63:1– 63:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Marton:2013:DPP

[Mar13] Yuval Marton. Distributional phrasal paraphrase generation for statistical machine translation. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3):39:1– 39:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Mansoury:2021:FBP

[MBM21] Masoud Mansoury, Robin Burke, and Bamshad Mobasher. Flatter is better: Percentile transformations for recommender systems. ACM Transactions on Intelligent Systems and Technology (TIST), 12(2):19:1– 19:16, March 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3437910.

Momtazpour:2014:CSI

[MBR^{+}14] Marjan Momtazpour, Patrick Butler, Naren Ramakrishnan, M. Shahriar Hossain, Mohammad C. Bozchalui, and Ratnesh Sharma. Charging and storage infrastructure design for electric vehicles. ACM Transactions on Intelligent Systems and Technology (TIST), 5(3):42:1–42:??, September 2014. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Manfredi:2023:TST

[MCEG23] Gilda Manfredi, Nicola Capece, Ugo Erra, and Monica Gruosso. TreeSketch-Net: From sketch to 3D tree parameters generation.

ACM Transactions on Intelligent Systems and Technology (TIST), 14(3):41:1– 41:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3579831.

Madnani:2013:GTP

[MD13] Nitin Madnani and Bonnie J. Dorr. Generating targeted paraphrases for improved translation. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3):40:1– 40:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Moon:2013:IBM

[ME13] Taesun Moon and Katrin Erk. An inference-based model of word meaning in context as a paraphrase distribution. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (3):42:1–42:??, June 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Mash:2020:HCC

[MFB^{+}20] Moshe Mash, Roy Fairstein, Yoram Bachrach, Kobi Gal, and Yair Zick. Humancomputer coalition formation in weighted voting games. ACM Transactions on Intelligent Systems and Technology (TIST), 11(6):73:1–73:20, November 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3408294.

Mikhail:2019:SBN

[MFI19] Joseph W. Mikhail, John M. Fossaceca, and Ronald Iammartino. A semiboosted nested model with sensitivity-based weighted binarization for multidomain network intrusion detection. ACM Transactions on Intelligent Systems and Technology (TIST), 10 (3):28:1–28:??, May 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3313778.

Mcardle:2014:UDF

[MFLP14] Gavin Mcardle, Eoghan Furey, Aonghus Lawlor, and Alexei Pozdnoukhov. Using digital footprints for a city-scale traffic simulation. ACM Transactions on Intelligent Systems and Technology (TIST), 5 (3):41:1–41:??, September 2014. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Moody:2016:NCF

[MG16] Jennifer Moody and David H. Glass. A novel classification framework for evaluating individual and aggregate diversity in top-N recommendations. ACM Transactions on Intelligent Systems and Technology (TIST), 7(3):42:1– 42:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Mithal:2011:MGF

[MGB^{+}11] Varun Mithal, Ashish Garg, Shyam Boriah, Michael Steinbach, Vipin Kumar, Christopher Potter, Steven Klooster, and Juan Carlos Castilla-Rubio. Monitoring global forest cover using data mining. ACM Transactions on Intelligent Systems and Technology (TIST), 2(4):36:1– 36:??, July 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ma:2020:ABR

[MGJW20] Jing Ma, Wei Gao, Shafiq Joty, and Kam-Fai Wong. An attention-based rumor detection model with treestructured recursive neural networks. ACM Transactions on Intelligent Systems and Technology (TIST), 11 (4):42:1–42:28, July 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:

                                        //dl.acm.org/doi/abs/
                                        10.1145/3391250.

Maltinsky:2017:NNM

[MGS17a] Alex Maltinsky, Ran Giladi, and Yuval Shavitt. On network neutrality measurements. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (4):56:1–56:??, July 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Mirsky:2017:COP

[MGS17b] Reuth Mirsky, Ya’akov (Kobi) Gal, and Stuart M. Shieber. CRADLE: an online plan recognition algorithm for exploratory domains. ACM Transactions on Intelligent Systems and Technology (TIST), 8(3):45:1– 45:??, April 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Mirsky:2019:GPR

[MGSK19] Reuth Mirsky, Kobi Gal, Roni Stern, and Meir Kalech. Goal and plan recognition design for plan libraries. ACM Transactions on Intelligent Systems and Technology (TIST), 10(2):14:1–14:??, February 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3234464.

Minkov:2016:EEU

[Min16] Einat Minkov. Event extraction using structured learning and rich domain knowledge: Application across domains and data sources. ACM Transactions on Intelligent Systems and Technology (TIST), 7(2): 16:1–16:??, January 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Maddalena:2023:QEP

[MIRS23] Eddy Maddalena, Luis-Daniel Ibáñez, Neal Reeves, and Elena Simperl. Qrowdsmith: Enhancing paid microtask crowdsourcing with gamification and furtherance incentives. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):86:1–86:??, October 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3604940.

Maddalena:2020:MPI

[MIS20] Eddy Maddalena, Luis-Daniel Ibáñez, and Elena Simperl. Mapping points of interest through street view imagery and paid crowdsourcing. ACM Transactions on Intelligent Systems and Technology (TIST), 11(5):63:1–63:28, September 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3403931.

Morris:2016:DNM

[MJVL16] Robert Morris, Matthew Johnson, K. Brent Venable, and James Lindsey. Designing noise-minimal rotorcraft approach trajectories. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4): 58:1–58:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ma:2011:LRE

[MKL11] Hao Ma, Irwin King, and Michael R. Lyu. Learning to recommend with explicit and implicit social relations. ACM Transactions on Intelligent Systems and Technology (TIST), 2(3): 29:1–29:??, April 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Mao:2021:FGB

[MLJZ21] Jiali Mao, Jiaye Liu, Cheqing Jin, and Aoying Zhou. Feature groupingbased trajectory outlier detection over distributed streams. ACM Transactions on Intelligent Systems and Technology (TIST), 12(2):22:1–22:23, March 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/doi/
10.1145/3444753.

Mullner:2023:RNR

[MLSK23] Peter Müllner, Elisabeth Lex, Markus Schedl, and Dominik Kowald. ReuseKNN: Neighborhood reuse for differentially private KNN-Based recommendations. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):80:1– 80:??, October 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3608481.

Montali:2013:MBC

[MMC^{+}13] Marco Montali, Fabrizio M.

Maggi, Federico Chesani, Paola Mello, and Wil M. P. van der Aalst. Monitoring business constraints with the event calculus. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1):17:1– 17:??, December 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Motai:2015:SCD

[MMDY15] Yuichi Motai, Dingkun Ma, Alen Docef, and Hiroyuki Yoshida. Smart colonography for distributed medical databases with group kernel feature analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4):58:1– 58:??, August 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Maggi:2023:DAD

[MMPS23] Fabrizio Maria Maggi, Andrea Marrella, Fabio Patrizi, and Vasyl Skydanienko. Data-aware declarative process mining with SAT. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):75:1–75:??, August 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3600106.

Marrella:2017:IPA

[MMS17] Andrea Marrella, Massimo Mecella, and Sebastian Sardina. Intelligent process adaptation in the SmartPM system. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2): 25:1–25:??, January 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Mahmud:2014:HLI

[MND14] Jalal Mahmud, Jeffrey Nichols, and Clemens Drews. Home location identification of Twitter users. ACM Transactions on Intelligent Systems and Technology (TIST), 5 (3):47:1–47:??, September 2014. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Muntean:2015:LPM

[MNSB15] Cristina Ioana Muntean, Franco Maria Nardini, Fabrizio Silvestri, and Ranieri Baraglia. On learning prediction models for tourists paths. ACM Transactions on Intelligent Systems and Technology (TIST), 7 (1):8:1–8:??, October 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

McNally:2011:CSC

[MOC^{+}11] Kevin McNally, Michael P.

O’Mahony, Maurice Coyle, Peter Briggs, and Barry Smyth. A case study of collaboration and reputation in social Web search. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):4:1– 4:??, October 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Mahesar:2023:ASD

[MP23] Quratul-Ain Mahesar and Simon Parsons. Argument schemes and a dialogue system for explainable planning. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):89:1–89:??, October 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3610301.

Marathe:2013:AFN

[MPA13] Achla Marathe, Zhengzheng Pan, and Andrea Apolloni. Analysis of friendship network and its role in explaining obesity. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (3):56:1–56:??, June 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Moscato:2023:FSN

[MPS23] Vincenzo Moscato, Marco Postiglione, and Giancarlo Sperlí. Few-shot named entity recognition: Definition, taxonomy and research directions. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):94:1–94:??, October 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3609483.

Moshfeghi:2016:GTA

[MRJ16] Yashar Moshfeghi, Alvaro Francisco Huertas Rosero, and Joemon M. Jose. A game-theory approach for effective crowdsource-based relevance assessment. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):55:1– 55:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Mandrake:2012:SSD

[MRW^{+}12] Lukas Mandrake, Umaa Rebbapragada, Kiri L. Wagstaff, David Thompson, Steve Chien, Daniel Tran, Robert T. Pappalardo, Damhnait Gleeson, and Rebecca Castaño. Surface sulfur detection via remote sensing and onboard classification. ACM Transactions on Intelligent Systems and Technology (TIST), 3(4):77:1–77:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ma:2011:LDM

[MSSV11] Justin Ma, Lawrence K.

Saul, Stefan Savage, and Geoffrey M. Voelker. Learning to detect malicious URLs. ACM Transactions on Intelligent Systems and Technology (TIST), 2(3): 30:1–30:??, April 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Muralidhar:2020:CRT

[MTC^{+}20] Nikhil Muralidhar, Anika Tabassum, Liangzhe Chen, Supriya Chinthavali, Naren Ramakrishnan, and B. Aditya Prakash. Cut-n-Reveal: Time series segmentations with explanations. ACM Transactions on Intelligent Systems and Technology (TIST), 11(5):53:1–53:26, September 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3394118.

Muralidhar:2018:III

[MWS^{+}18] Nikhil Muralidhar, Chen Wang, Nathan Self, Marjan Momtazpour, Kiyoshi Nakayama, Ratnesh Sharma, and Naren Ramakrishnan. illiad: InteLLigent invariant and anomaly detection in cyber-physical systems. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3): 35:1–35:??, February 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Mei:2023:FRT

[MWY^{+}23] Jianbiao Mei, Mengmeng Wang, Yu Yang, Yanjun Li, and Yong Liu. Fast real-time video object segmentation with a tangled memory network. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3):51:1– 51:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3585076.

Ma:2012:RPC

[MZL12] Huadong Ma, Chengbin Zeng, and Charles X. Ling. A reliable people counting system via multiple cameras. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2): 31:1–31:??, February 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Mao:2022:CEF

[MZY^{+}22] Yuzhu Mao, Zihao Zhao, Guangfeng Yan, Yang Liu, Tian Lan, Linqi Song, and Wenbo Ding. Communication-efficient federated learning with adaptive quantization. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):67:1–67:??, August 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3510587.

Neviarouskaya:2014:IIT

[NAPI14] Alena Neviarouskaya, Masaki Aono, Helmut Prendinger, and Mitsuru Ishizuka. Intelligent interface for textual attitude analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 5(3):48:1– 48:??, September 2014. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Notaro:2021:SAM

[NCG21] Paolo Notaro, Jorge Cardoso, and Michael Gerndt. A survey of AIOps methods for failure management. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):81:1– 81:45, December 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3483424.

Ning:2019:DRL

[NDW^{+}19] Zhaolong Ning, Peiran Dong, Xiaojie Wang, Joel J. P. C. Rodrigues, and Feng Xia. Deep reinforcement learning for vehicular edge computing: an intelligent offloading system. ACM Transactions on Intelligent Systems and Technology (TIST), 10(6):60:1– 60:??, October 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Nelke:2020:MCB

[NOZ20] Sofia Amador Nelke, Steven Okamoto, and Roie Zivan. Market clearingbased dynamic multi-agent task allocation. ACM Transactions on Intelligent Systems and Technology (TIST), 11(1):4:1–4:25, February 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3356467.

Nanni:2016:DPC

[NTM^{+}16] Mirco Nanni, Roberto Trasarti, Anna Monreale, Valerio Grossi, and Dino Pedreschi. Driving profiles computation and monitoring for car insurance CRM. ACM Transactions on Intelligent Systems and Technology (TIST), 8(1):14:1– 14:??, October 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Navia-Vazquez:2022:BDS

[NVDMFD22] A. Navia-Vázquez, R. Díaz-Morales, and M. Fernández-Díaz. Budget distributed support vector machine for non-ID federated learning scenarios. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):100:1–100:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3539734.

Neria:2017:RSF

[NYBG17] Michal Ben Neria, Nancy-Sarah Yacovzada, and Irad Ben-Gal. A riskscoring feedback model for Webpages and Web users based on browsing behavior. ACM Transactions on Intelligent Systems and Technology (TIST), 8(4): 53:1–53:??, July 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Niu:2022:ERT

[NZS^{+}22] Hongting Niu, Hengshu Zhu, Ying Sun, Xinjiang Lu, Jing Sun, Zhiyuan Zhao, Hui Xiong, and Bo Lang. Exploring the risky travel area and behavior of car-hailing service. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):9:1– 9:22, February 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3465059.

Nie:2017:LUA

[NZW^{+}17] Liqiang Nie, Luming Zhang, Meng Wang, Richang Hong, Aleksandr Farseev, and Tat-Seng Chua. Learning user attributes via mobile social multimedia analytics. ACM Transactions on Intelligent Systems and Technology (TIST), 8(3): 36:1–36:??, April 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ottens:2017:DUC

[ODF17] Brammert Ottens, Christos Dimitrakakis, and Boi Faltings. DUCT: an upper confidence bound approach to distributed constraint optimization problems. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (5):69:1–69:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Oliveira:2020:RAE

[ODL^{+}20] Samuel E. L. Oliveira, Victor Diniz, Anisio Lacerda, Luiz Merschmanm, and Gisele L. Pappa. Is rank aggregation effective in recommender systems? An experimental analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 11(2):16:1– 16:26, March 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3365375.

Ovelgonne:2017:URB

[ODP^{+}17] Michael Ovelgönne, Tudor Dumitras, B. Aditya Prakash, V. S. Subrahmanian, and Benjamin Wang. Understanding the relationship between human behavior and susceptibility to cyber attacks: a datadriven approach. ACM Transactions on Intelligent Systems and Technology (TIST), 8(4):51:1– 51:??, July 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ou:2017:AIV

[OLY^{+}17] Xinyu Ou, Hefei Ling, Han Yu, Ping Li, Fuhao Zou, and Si Liu. Adult image and video recognition by a deep multicontext network and fine-to-coarse strategy. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (5):68:1–68:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Oramas:2017:SMR

[OOD^{+}17] Sergio Oramas, Vito Claudio Ostuni, Tommaso Di Noia, Xavier Serra, and Eugenio Di Sciascio. Sound and music recommendation with knowledge graphs. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2):21:1– 21:??, January 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Osman:2013:TMA

[OSM^{+}13] Nardine Osman, Carles Sierra, Fiona Mcneill, Juan Pane, and John Debenham. Trust and matching algorithms for selecting suitable agents. ACM Transactions on Intelligent Systems and Technology (TIST), 5 (1):16:1–16:??, December 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ou:2022:AAE

[OSW^{+}22] Jinxiang Ou, Yunheng Shen, Feng Wang, Qiao Liu, Xuegong Zhang, and Hairong Lv. AggEnhance: Aggregation enhancement by class interior points in federated learning with non-IID data. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):104:1–104:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3544495.

Okada:2013:MDA

[OY13] Isamu Okada and Hitoshi Yamamoto. Mathematical description and analysis of adaptive risk choice behavior. ACM Transactions on Intelligent Systems and Technology (TIST), 4(1): 17:1–17:??, January 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Pachet:2017:JOA

[Pac17] François Pachet. A joyful ode to automatic orchestration. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2): 18:1–18:??, January 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Paik:2016:PDM

[Pai16] Jiaul H. Paik. Parameterized decay model for information retrieval. ACM Transactions on Intelligent Systems and Technology (TIST), 7(3):27:1– 27:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Patel:2015:DSI

[Pat15] Dhaval Patel. On discovery of spatiotemporal influence-based moving clusters. ACM Transactions on Intelligent Systems and Technology (TIST), 6 (1):4:1–4:??, March 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Pool:2014:DDC

[PBvL14] Simon Pool, Francesco Bonchi, and Matthijs van Leeuwen. Descriptiondriven community detection. ACM Transactions on Intelligent Systems and Technology (TIST), 5(2): 28:1–28:??, April 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Porteous:2010:API

[PCC10] Julie Porteous, Marc Cavazza, and Fred Charles. Applying planning to interactive storytelling: Narrative control using state constraints. ACM Transactions on Intelligent Systems and Technology (TIST), 1(2):10:1– 10:??, November 2010. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Peng:2017:SLM

[PCC17] Chong Peng, Jie Cheng, and Qiang Cheng. A supervised learning model for high-dimensional and largescale data. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2):30:1–30:??, January 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Pappalardo:2019:PDD

[PCF^{+}19] Luca Pappalardo, Paolo Cintia, Paolo Ferragina, Emanuele Massucco, Dino Pedreschi, and Fosca Giannotti. PlayeRank: Datadriven performance evaluation and player ranking in soccer via a machine learning approach. ACM Transactions on Intelligent Systems and Technology (TIST), 10(5):59:1– 59:??, October 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Peng:2018:EPH

[PCL18] Xuefeng Peng, Li-Kai Chi, and Jiebo Luo. The effect of pets on happiness: a large-scale multi-factor analysis using social multimedia. ACM Transactions on Intelligent Systems and Technology (TIST), 9(5): 60:1–60:??, July 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Preti:2023:MAM

[PDR23] Giulia Preti, Gianmarco De Francisci Morales, and Matteo Riondato. MaNI-ACS: Approximate mining of frequent subgraph patterns through sampling. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3):54:1– 54:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3587254.

Phan:2016:TAP

[PEK^{+}16] Nhathai Phan, Javid Ebrahimi, David Kil, Brigitte Piniewski, and Dejing Dou. Topicaware physical activity propagation with temporal dynamics in a health social network. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (1):2:1–2:??, October 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Papalexakis:2017:TDM

[PFS17] Evangelos E. Papalexakis, Christos Faloutsos, and Nicholas D. Sidiropoulos. Tensors for data mining and data fusion: Models, applications, and scalable algorithms. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2):16:1–16:??, January 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Park:2013:CPC

[PG13] Yubin Park and Joydeep Ghosh. CUDIA: Probabilistic cross-level imputation using individual auxiliary information. ACM Transactions on Intelligent Systems and Technology (TIST), 4(4):66:1–66:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Pan:2020:DDH

[PHL^{+}20] Menghai Pan, Weixiao Huang, Yanhua Li, Xun Zhou, Zhenming Liu, Rui Song, Hui Lu, Zhihong Tian, and Jun Luo. DHPA: Dynamic human preference analytics framework: a case study on taxi drivers’ learning curve analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 11(1):8:1–8:19, February 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3360312.

Peng:2018:ICD

[PKCC18] Chong Peng, Zhao Kang, Shuting Cai, and Qiang Cheng. Integrate and conquer: Double-sided twodimensional k-means via integrating of projection and manifold construction. ACM Transactions on Intelligent Systems and Technology (TIST), 9(5):57:1– 57:??, July 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Peng:2017:NMF

[PKH^{+}17] Chong Peng, Zhao Kang, Yunhong Hu, Jie Cheng, and Qiang Cheng. Nonnegative matrix factorization with integrated graph and feature learning. ACM Transactions on Intelligent Systems and Technology (TIST), 8(3):42:1– 42:??, April 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Panagopoulos:2017:AEC

[PMR^{+}17] Athanasios Aris Panagopoulos, Sasan Maleki, Alex Rogers, Matteo Venanzi, and Nicholas R. Jennings. Advanced economic control of electricity-based space heating systems in domestic coalitions with shared intermittent energy resources. ACM Transactions on Intelligent Systems and Technology (TIST), 8(4): 59:1–59:??, July 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Patnaik:2011:TDM

[PMSR11] Debprakash Patnaik, Manish Marwah, Ratnesh K. Sharma, and Naren Ramakrishnan. Temporal data mining approaches for sustainable chiller management in data centers. ACM Transactions on Intelligent Systems and Technology (TIST), 2(4):34:1– 34:??, July 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Pereira:2020:USO

[POM20] Ramon Fraga Pereira, Nir Oren, and Felipe Meneguzzi. Using suboptimal plan detection to identify commitment abandonment in discrete environments. ACM Transactions on Intelligent Systems and Technology (TIST), 11(2):23:1– 23:26, March 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3372119.

Pellungrini:2018:DMA

[PPPM18] Roberto Pellungrini, Luca Pappalardo, Francesca Pratesi, and Anna Monreale. A data mining approach to assess privacy risk in human mobility data. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3):31:1– 31:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Prettenhofer:2011:CLA

[PS11] Peter Prettenhofer and Benno Stein. Cross-lingual adaptation using structural correspondence learning. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):13:1– 13:??, October 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Potthast:2012:IRC

[PSLB12] Martin Potthast, Benno Stein, Fabian Loose, and Steffen Becker. Information retrieval in the Commentsphere. ACM Transactions on Intelligent Systems and Technology (TIST), 3(4):68:1–68:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Peng:2012:MVC

[PSRL12] Wei Peng, Tong Sun, Shriram Revankar, and Tao Li. Mining the “Voice of the customer” for business prioritization. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):38:1–38:??, February 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Paltoglou:2012:TMD

[PT12] Georgios Paltoglou and Mike Thelwall. Twitter, MySpace, Digg: Unsupervised sentiment analysis in social media. ACM Transactions on Intelligent Systems and Technology (TIST), 3(4):66:1–66:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Pan:2017:TLB

[PYD^{+}17] Weike Pan, Qiang Yang, Yuchao Duan, Ben Tan, and Zhong Ming. Transfer learning for behavior ranking. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (5):65:1–65:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Qin:2015:SSA

[QCL15] Tao Qin, Wei Chen, and Tie-Yan Liu. Sponsored search auctions: Recent advances and future directions. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4): 60:1–60:??, January 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Qiao:2021:CDC

[QCZ^{+}21] Jie Qiao, Ruichu Cai, Kun Zhang, Zhenjie Zhang, and Zhifeng Hao. Causal discovery with confounding cascade nonlinear additive noise models. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):80:1–80:28, December 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3482879.

Qiao:2021:DCN

[QHH^{+}21] Shaojie Qiao, Nan Han, Jianbin Huang, Kun Yue, Rui Mao, Hongping Shu, Qiang He, and Xindong Wu. A dynamic convolutional neural network based shared-bike demand forecasting model. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):70:1–70:24, December 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3447988.

Quijano-Sanchez:2013:SFG

[QSRGDAJD13] Lara Quijano-Sanchez, Juan A.

Recio-Garcia, Belen Diaz-Agudo, and Guillermo Jimenez-Diaz. Social factors in group recommender systems. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (1):8:1–8:??, January 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Qu:2023:AAD

[QTM23] Ao Qu, Yihong Tang, and Wei Ma. Adversarial attacks on deep reinforcement learning-based traffic signal control systems with colluding vehicles. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):113:1–113:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3625236.

Qin:2023:DGA

[QWC^{+}23] Xin Qin, Jindong Wang, Yiqiang Chen, Wang Lu, and Xinlong Jiang. Domain generalization for activity recognition via adaptive feature fusion. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):9:1–9:??, February 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3552434.

Rokh:2023:CSM

[RAK23] Babak Rokh, Ali Azarpeyvand, and Alireza Khanteymoori. A comprehensive survey on model quantization for deep neural networks in image classification. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):97:1–97:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3623402.

Rossi:2018:IVG

[RAZE18] Ryan A. Rossi, Nesreen K. Ahmed, Rong Zhou, and Hoda Eldardiry. Interactive visual graph mining and learning. ACM Transactions on Intelligent Systems and Technology (TIST), 9 (5):59:1–59:??, July 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Rai:2022:RSC

[RBG22] Sawan Rai, Ramesh Chandra Belwal, and Atul Gupta. A review on source code documentation. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):84:1–84:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3519312.

Resnik:2013:UTP

[RBK^{+}13] Philip Resnik, Olivia Buzek, Yakov Kronrod, Chang Hu, Alexander J. Quinn, and Benjamin B. Bederson. Using targeted paraphrasing and monolingual crowdsourcing to improve translation. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3): 38:1–38:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Rashidi:2013:CMM

[RC13] Parisa Rashidi and Diane J.

Cook. COM: a method for mining and monitoring human activity patterns in home-based health monitoring systems. ACM Transactions on Intelligent Systems and Technology (TIST), 4(4):64:1–64:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Roos:2010:ESD

[RCN10] Patrick Roos, J. Ryan Carr, and Dana S. Nau. Evolution of state-dependent risk preferences. ACM Transactions on Intelligent Systems and Technology (TIST), 1 (1):6:1–6:??, October 2010. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ren:2022:GGR

[RDX22] Hanchi Ren, Jingjing Deng, and Xianghua Xie. GRNN: Generative regression neural network — a data leakage attack for federated learning. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):65:1–65:??, August 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3510032.

Rendle:2012:FML

[Ren12] Steffen Rendle. Factorization machines with libFM. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3):57:1– 57:??, May 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Reddy:2011:PSA

[RFI^{+}11] Sudhakar Y. Reddy, Jeremy D.

Frank, Michael J. Iatauro, Matthew E. Boyce, Elif Kürklü, Mitchell Ai-Chang, and Ari K. Jónsson. Planning solar array operations on the International Space Station. ACM Transactions on Intelligent Systems and Technology (TIST), 2(4): 41:1–41:??, July 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Radanovic:2016:IEC

[RFJ16] Goran Radanovic, Boi Faltings, and Radu Jurca. Incentives for effort in crowdsourcing using the peer truth serum. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):48:1– 48:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ren:2022:DHC

[RGC^{+}22] Siyuan Ren, Bin Guo, Longbing Cao, Ke Li, Jiaqi Liu, and Zhiwen Yu. DeepExpress: Heterogeneous and coupled sequence modeling for express delivery prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):89:1–89:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3526087.

Rahmadi:2019:SSS

[RGH19] Ridho Rahmadi, Perry Groot, and Tom Heskes. Stable specification search in structural equation models with latent variables. ACM Transactions on Intelligent Systems and Technology (TIST), 10(5):48:1– 48:??, October 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Rohrdantz:2012:FBV

[RHD^{+}12] Christian Rohrdantz, Ming C. Hao, Umeshwar Dayal, Lars-Erik Haug, and Daniel A. Keim. Feature-based visual sentiment analysis of text document streams. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):26:1– 26:??, February 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Rao:2016:LHC

[RHF16] Huaming Rao, Shih-Wen Huang, and Wai-Tat Fu. Leveraging human computations to improve schematization of spatial relations from imagery. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):54:1– 54:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ranganath:2018:UIR

[RHT^{+}18] Suhas Ranganath, Xia Hu, Jiliang Tang, Suhang Wang, and Huan Liu. Understanding and identifying rhetorical questions in social media. ACM Transactions on Intelligent Systems and Technology (TIST), 9(2):17:1–17:??, January 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Reches:2015:CCU

[RK15] Shulamit Reches and Meir Kalech. Choosing a candidate using efficient allocation of biased information. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):66:1– 66:??, January 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Reches:2014:FEC

[RKH14] Shulamit Reches, Meir Kalech, and Philip Hendrix. A framework for effectively choosing between alternative candidate partners. ACM Transactions on Intelligent Systems and Technology (TIST), 5(2): 30:1–30:??, April 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Reyes:2023:PCD

[RP23] Óscar Reyes and Eduardo Pérez. Performing cancer diagnosis via an isoform expression rankingbased LSTM model. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):110:1– 110:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3625237.

Rodriguez-Serrano:2017:TDA

[RSCOVCMM17] Francisco Jose Rodriguez-Serrano, Julio Jose Carabias-Orti, Pedro Vera-Candeas, and Damian Martinez-Munoz. Tempo driven audio-to-score alignment using spectral decomposition and online dynamic time warping. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2):22:1– 22:??, January 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Reyes:2018:ESP

[RV18] Oscar Reyes and Sebastián Ventura. Evolutionary strategy to perform batchmode active learning on multi-label data. ACM Transactions on Intelligent Systems and Technology (TIST), 9(4):46:1– 46:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ramchurn:2011:ABH

[RVRJ11] Sarvapali D. Ramchurn, Perukrishnen Vytelingum, Alex Rogers, and Nicholas R. Jennings. Agent-based homeostatic control for green energy in the smart grid. ACM Transactions on Intelligent Systems and Technology (TIST), 2(4): 35:1–35:??, July 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ramamohanarao:2017:SSM

[RXK^{+}17] Kotagiri Ramamohanarao, Hairuo Xie, Lars Kulik, Shanika Karunasekera, Egemen Tanin, Rui Zhang, and Eman Bin Khunayn. SMARTS: Scalable microscopic adaptive road traffic simulator. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2):26:1–26:??, January 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ren:2023:GLA

[RXL^{+}23] Jing Ren, Feng Xia, Ivan Lee, Azadeh Noori Hoshyar, and Charu Aggarwal.

Graph learning for anomaly analytics: Algorithms, applications, and challenges. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2):28:1– 28:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3570906.

Reddy:2013:ISS

[RY13] Chandan K. Reddy and Cristopher C. Yang. Introduction to the special section on intelligent systems for health informatics. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (4):62:1–62:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ren:2022:IAV

[RYC22] Zhenghang Ren, Liu Yang, and Kai Chen. Improving availability of vertical federated learning: Relaxing inference on nonoverlapping data. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):58:1–58:??, August 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3501817.

Refanidis:2010:CBA

[RYS10] Ioannis Refanidis and Neil Yorke-Smith. A constraintbased approach to scheduling an individual’s activities. ACM Transactions on Intelligent Systems and Technology (TIST), 1 (2):12:1–12:??, November 2010. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ribeiro:2015:MPE

[RZS^{+}15] Marco Tulio Ribeiro, Nivio Ziviani, Edleno Silva De Moura, Itamar Hata, Anisio Lacerda, and Adriano Veloso. Multiobjective Pareto-efficient approaches for recommender systems. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):53:1– 53:??, January 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shen:2015:ISI

[SA15] Dou Shen and Deepak Agarwal. Introduction to the special issue on online advertising. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):57:1–57:??, January 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Said:2013:MRC

[SBD13] Alan Said, Shlomo Berkovsky, and Ernesto W. De Luca. Movie recommendation in context. ACM Transactions on Intelligent Systems and Technology (TIST), 4(1):13:1–13:??, January 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Sarkar:2023:AHM

[SBS^{+}23] Souvika Sarkar, Biddut Sarker Bijoy, Syeda Jannatus Saba, Dongji Feng, Yash Mahajan, Mohammad Ruhul Amin, Sheikh Rabiul Islam, and Shubhra Kanti Karmaker (“Santu”). Ad-hoc monitoring of COVID-19 global research trends for well-informed policy making. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2): 26:1–26:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3576901.

Sandouk:2017:LCM

[SC17] Ubai Sandouk and Ke Chen. Learning contextualized music semantics from tags via a Siamese neural network. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2): 24:1–24:??, January 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Saxena:2022:MST

[SC22] Divya Saxena and Jiannong Cao. Multimodal spatio-temporal prediction with stochastic adversarial networks. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):18:1– 18:23, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3458025.

Sun:2023:WYN

[SCC^{+}23] Heli Sun, Chen Cao, Xuguang Chu, Tingting Hu, Junzhi Lu, Liang He, Zhi Wang, Hui He, and Hui Xiong. What your next check-in might look like: Next check-in behavior prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):112:1–112:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3625234.

Shen:2017:DDD

[SCLZ17] Jiaxing Shen, Jiannong Cao, Xuefeng Liu, and Chisheng Zhang. DMAD: Data-driven measuring of Wi-Fi access point deployment in urban spaces.

ACM Transactions on Intelligent Systems and Technology (TIST), 9(1):11:1– 11:??, October 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Schulz:2016:MTN

[SDD^{+}16] Sarah Schulz, Guy De Pauw, Orphée De Clercq, Bart Desmet, Véronique Hoste, Walter Daelemans, and Lieve Macken. Multimodular text normalization of Dutch user-generated content. ACM Transactions on Intelligent Systems and Technology (TIST), 7 (4):61:1–61:??, July 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Sawant:2015:AGC

[SDHS15] Anshul Sawant, John P. Dickerson, Mohammad T. Hajiaghayi, and V. S. Subrahmanian. Automated generation of counterterrorism policies using multiexpert input. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4):44:1– 44:??, August 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shakarian:2012:AGA

[SDS12] Paulo Shakarian, John P. Dickerson, and V. S. Subrahmanian. Adversarial geospatial abduction problems. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2): 34:1–34:??, February 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shan:2016:SBS

[SDXG16] Na Shan, Xiaogang Dong, Pingfeng Xu, and Jianhua Guo. Sharp bounds on survivor average causal effects when the outcome is binary and truncated by death. ACM Transactions on Intelligent Systems and Technology (TIST), 7(2):18:1– 18:??, January 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Sang:2017:ESM

[SFX17] Jitao Sang, Quan Fang, and Changsheng Xu. Exploiting social-mobile information for location visualization. ACM Transactions on Intelligent Systems and Technology (TIST), 8(3): 39:1–39:??, April 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Subbu:2013:LMF

[SGD13] Kalyan Pathapati Subbu, Brandon Gozick, and Ram Dantu. LocateMe: Magnetic-fields-based indoor localization using smartphones. ACM Transactions on Intelligent Systems and Technology (TIST), 4(4):73:1–73:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shah:2018:DOL

[SGJC18] Ankit Shah, Rajesh Ganesan, Sushil Jajodia, and Hasan Cam. Dynamic optimization of the level of operational effectiveness of a CSOC under adverse conditions. ACM Transactions on Intelligent Systems and Technology (TIST), 9(5): 51:1–51:??, July 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Singhal:2020:CBM

[SGTK20] Divya Singhal, Abhinav Gupta, Anurag Tripathi, and Ravi Kothari. CNNbased multiple manipulation detector using frequency domain features of image residuals. ACM Transactions on Intelligent Systems and Technology (TIST), 11(4):40:1– 40:26, July 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3388634.

Sun:2022:PFL

[SGY^{+}22] Heli Sun, Xianglan Guo, Zhou Yang, Xuguang Chu, Xinwang Liu, and Liang He. Predicting future locations with semantic trajectories. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):7:1–7:20, February 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3465060.

Strohmaier:2012:EFI

[SHB^{+}12] Markus Strohmaier, Denis Helic, Dominik Benz, Christian Körner, and Roman Kern. Evaluation of folksonomy induction algorithms. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (4):74:1–74:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shao:2023:IIT

[SHX^{+}23] Erzhuo Shao, Zhenyu Han, Yulai Xie, Yang Zhang, Lu Geng, and Yong Li. Interior individual trajectory simulation with population distribution constraint. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):2:1– 2:??, February 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3529108.

Shi:2013:ACL

[SHZ13] Ziqiang Shi, Jiqing Han, and Tieran Zheng. Audio classification with low-rank matrix representation features. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1): 15:1–15:??, December 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Singh:2013:NBG

[Sin13] Munindar P. Singh. Norms as a basis for governing sociotechnical systems. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1):21:1– 21:??, December 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Sizov:2012:LGS

[Siz12]  Sergej Sizov.  Latent

geospatial semantics of social media. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (4):64:1–64:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Sun:2023:MBR

[SJCM23] Wei Sun, Shaoxiong Ji, Erik Cambria, and Pekka Marttinen. Multitask balanced and recalibrated network for medical code prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):17:1–17:??, February 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3563041.

Shi:2014:MLC

[SKF^{+}14] Chuan Shi, Xiangnan Kong, Di Fu, Philip S. Yu, and Bin Wu. Multilabel classification based on multi-objective optimization. ACM Transactions on Intelligent Systems and Technology (TIST), 5(2): 35:1–35:??, April 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Saito:2013:DCI

[SKOM13] Kazumi Saito, Masahiro Kimura, Kouzou Ohara, and Hiroshi Motoda. Detecting changes in information diffusion patterns over social networks. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3):55:1– 55:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shi:2023:SLI

[SLC23] Yanhang Shi, Xue Li, and Siguang Chen. Skin lesion intelligent diagnosis in edge computing networks: an FCL approach. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):69:1– 69:??, August 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3595186.

Shi:2013:MCM

[SLH13] Yue Shi, Martha Larson, and Alan Hanjalic. Mining contextual movie similarity with matrix factorization for context-aware recommendation. ACM Transactions on Intelligent Systems and Technology (TIST), 4(1):16:1–16:??, January 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shen:2021:MMK

[SLM^{+}21] Xiangjun Shen, Kou Lu, Sumet Mehta, Jianming Zhang, Weifeng Liu, Jianping Fan, and Zhengjun Zha. MKEL: Multiple kernel ensemble learning via unified ensemble loss for image classification. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4):40:1–40:21, August 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3457217.

Shi:2023:MIL

[SLM^{+}23] Lei Shi, Yuankai Luo, Shuai Ma, Hanghang Tong, Zhetao Li, Xiatian Zhang, and Zhiguang Shan. Mobility inference on long-tailed sparse trajectory. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):18:1–18:??, February 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3563457.

Siddharthan:2016:CCR

[SLR^{+}16] Advaith Siddharthan, Christopher Lambin, Anne-Marie Robinson, Nirwan Sharma, Richard Comont, Elaine O’Mahony, Chris Mellish, and René Van Der Wal. Crowdsourcing without a crowd: Reliable online species identification using Bayesian models to minimize crowd size. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):45:1– 45:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shieh:2013:RTS

[SLWW13] Jyh-Ren Shieh, Ching-Yung Lin, Shun-Xuan Wang, and Ja-Ling Wu. Relational term-suggestion graphs incorporating multipartite concept and expertise networks. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1):19:1– 19:??, December 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shen:2023:CST

[SLZ^{+}23] Ziyu Shen, Binghui Liu, Qing Zhou, Zheng Liu, Bin Xia, and Yun Li. Costsensitive tensor-based dualstage attention LSTM with feature selection for data center server power forecasting. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2): 24:1–24:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3569422.

Soto-Mendoza:2015:DPS

[SMGMC^{+}15] Valeria Soto-Mendoza, J. Antonio García-Macías, Edgar Chávez, Ana I. Martínez-García, Jesús Favela, Patricia Serrano-Alvarado, and Maythé R. Zúñiga Rojas. Design of a predictive scheduling system to improve assisted living services for elders. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4):53:1– 53:??, August 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Sang:2015:ASC

[SMX15] Jitao Sang, Tao Mei, and Changsheng Xu. Activity sensor: Check-in usage mining for local recommendation. ACM Transactions on Intelligent Systems and Technology (TIST), 6(3): 41:1–41:??, May 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Semertzidis:2016:CPS

[SNL^{+}16]  Theodoros  Semertzidis,

Jasminko Novak, Michalis Lazaridis, Mark Melenhorst, Isabel Micheel, Dimitrios Michalopoulos, Martin Böckle, Michael G. Strintzis, and Petros Daras. A crowd-powered system for fashion similarity search. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4): 46:1–46:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Sintsova:2016:DDS

[SP16] Valentina Sintsova and Pearl Pu. Dystemo: Distant supervision method for multi-category emotion recognition in tweets.

ACM Transactions on Intelligent Systems and Technology (TIST), 8(1):13:1– 13:??, October 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Stapleton:2015:CST

[SPDR15] Gem Stapleton, Beryl Plimmer, Aidan Delaney, and Peter Rodgers. Combining sketching and traditional diagram editing tools. ACM Transactions on Intelligent Systems and Technology (TIST), 6(1): 10:1–10:??, March 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Sharma:2019:CFN

[SQJ^{+}19] Karishma Sharma, Feng Qian, He Jiang, Natali Ruchansky, Ming Zhang, and Yan Liu. Combating fake news: a survey on identification and mitigation techniques. ACM Transactions on Intelligent Systems and Technology (TIST), 10(3):21:1– 21:??, May 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3305260.

Schindler:2017:HMR

[SR17] Alexander Schindler and Andreas Rauber. Harnessing music-related visual stereotypes for music information retrieval. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2):20:1– 20:??, January 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Sepehri-Rad:2015:ICW

[SRB15] Hoda Sepehri-Rad and Denilson Barbosa. Identifying controversial Wikipedia articles using editor collaboration networks. ACM Transactions on Intelligent Systems and Technology (TIST), 6(1):5:1–5:??, March 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Suk:2012:VHM

[SRJP12] Myunghoon Suk, Ashok Ramadass, Yohan Jin, and B. Prabhakaran. Video human motion recognition using a knowledge-based hybrid method based on a hidden Markov model. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3):42:1– 42:??, May 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Schuster:2013:PSC

[SRM^{+}13] Daniel Schuster, Alberto Rosi, Marco Mamei, Thomas Springer, Markus Endler, and Franco Zambonelli. Pervasive social context: Taxonomy and survey. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3):46:1– 46:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Skibski:2019:ECS

[SRMW19] Oskar Skibski, Talal Rahwan, Tomasz P. Michalak, and Michael Wooldridge. Enumerating connected subgraphs and computing the Myerson and Shapley values in graph-restricted games. ACM Transactions on Intelligent Systems and Technology (TIST), 10(2):15:1–15:??, February 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3235026.

Sukthankar:2011:ARD

[SS11] Gita Sukthankar and Katia Sycara. Activity recognition for dynamic multi-agent teams. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):18:1– 18:??, October 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Spurlock:2015:EGD

[SS15] Scott Spurlock and Richard Souvenir. An evaluation of gamesourced data for human pose estimation. ACM Transactions on Intelligent Systems and Technology (TIST), 6(2):19:1– 19:??, April 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Sharma:2022:ATG

[SS22] Arun Sharma and Shashi Shekhar. Analyzing trajectory gaps to find possible rendezvous region. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3):36:1– 36:23, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3467977.

Shah:2020:TCP

[SSG^{+}20] Ankit Shah, Arunesh Sinha, Rajesh Ganesan, Sushil Jajodia, and Hasan Cam. Two can play that game: an adversarial evaluation of a cyber-alert inspection system. ACM Transactions on Intelligent Systems and Technology (TIST), 11(3): 32:1–32:20, May 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3377554.

Shi:2013:NLR

[SSHL13] Yue Shi, Pavel Serdyukov, Alan Hanjalic, and Martha Larson. Nontrivial landmark recommendation using geotagged photos. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3):47:1– 47:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shen:2018:MDH

[SSL^{+}18] Xiaobo Shen, Fumin Shen, Li Liu, Yun-Hao Yuan, Weiwei Liu, and Quan-Sen Sun. Multiview discrete hashing for scalable multimedia search. ACM Transactions on Intelligent Systems and Technology (TIST), 9(5):53:1– 53:??, July 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Strohmeier:2021:CFI

[SSLM21] Martin Strohmeier, Matthew Smith, Vincent Lenders, and Ivan Martinovic. Classi-Fly: Inferring aircraft categories from open data. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):79:1– 79:23, December 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3480969.

Shakarian:2011:GGA

[SSS11] Paulo Shakarian, V. S. Subrahmanian, and Maria Luisa Sapino. GAPs: Geospatial abduction problems. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):7:1– 7:??, October 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shi:2015:ESC

[SST^{+}15] Miaojing Shi, Xinghai Sun, Dacheng Tao, Chao Xu, George Baciu, and Hong Liu. Exploring spatial correlation for visual object retrieval. ACM Transactions on Intelligent Systems and Technology (TIST), 6(2): 24:1–24:??, April 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Strobl:2019:ECF

[SSV19] Eric V. Strobl, Peter L.

Spirtes, and Shyam Visweswaran. Estimating and controlling the false discovery rate of the PC algorithm using edge-specific P -values. ACM Transactions on Intelligent Systems and Technology (TIST), 10(5):46:1– 46:??, October 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Song:2013:FOU

[SSZ^{+}13] Xuan Song, Xiaowei Shao, Quanshi Zhang, Ryosuke Shibasaki, Huijing Zhao, Jinshi Cui, and Hongbin Zha. A fully online and unsupervised system for large and high-density area surveillance: Tracking, semantic scene learning and abnormality detection. ACM Transactions on Intelligent Systems and Technology (TIST), 4(2):35:1– 35:??, March 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shi:2019:CBV

[ST19] Neng Shi and Yubo Tao. CNNs based viewpoint estimation for volume visualization. ACM Transactions on Intelligent Systems and Technology (TIST), 10(3): 27:1–27:??, May 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3309993.

Shmueli:2020:MSM

[ST20] Erez Shmueli and Tamir Tassa. Mediated secure multi-party protocols for collaborative filtering. ACM Transactions on Intelligent Systems and Technology (TIST), 11(2):15:1– 15:25, March 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3375402.

Stripelis:2022:SSF

[STA22] Dimitris Stripelis, Paul M.

Thompson, and José Luis Ambite. Semi-synchronous federated learning for energyefficient training and accelerated convergence in cross-silo settings. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):78:1–78:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3524885.

Sun:2018:SVA

[STP^{+}18] Guodao Sun, Tan Tang, Tai-Quan Peng, Ronghua Liang, and Yingcai Wu. SocialWave: Visual analysis of spatio-temporal diffusion of information on social media. ACM Transactions on Intelligent Systems and Technology (TIST), 9(2): 15:1–15:??, January 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Sun:2023:RLQ

[SWA23] Shuo Sun, Rundong Wang, and Bo An. Reinforcement learning for quantitative trading. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3): 44:1–44:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3582560.

Shen:2013:RUT

[SWZ^{+}13] Keyi Shen, Jianmin Wu, Ya Zhang, Yiping Han, Xiaokang Yang, Li Song, and Xiao Gu. Reorder user’s tweets. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (1):6:1–6:??, January 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Sun:2021:VIV

[SWZ^{+}21] Guodao Sun, Hao Wu, Lin Zhu, Chaoqing Xu, Haoran Liang, Binwei Xu, and Ronghua Liang. VSumVis: Interactive visual understanding and diagnosis of video summarization model. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4):41:1–41:28, August 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3458928.

Sugiyama:2012:ISS

[SY12] Masashi Sugiyama and Qiang Yang. Introduction to the special section on the 2nd Asia Conference on Machine Learning (ACML 2010). ACM Transactions on Intelligent Systems and Technology (TIST), 3(2): 27:1–27:??, February 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Schedl:2017:IIM

[SYHB17] Markus Schedl, Yi-Hsuan Yang, and Perfecto Herrera-Boyer. Introduction to intelligent music systems and applications. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2):17:1– 17:??, January 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shao:2011:VIG

[SZC11] Yuanlong Shao, Yuan Zhou, and Deng Cai. Variational inference with graph regularization for image annotation. ACM Transactions on Intelligent Systems and Technology (TIST), 2(2):11:1–11:??, February 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Song:2013:OSM

[SZC^{+}13] Xuan Song, Huijing Zhao, Jinshi Cui, Xiaowei Shao, Ryosuke Shibasaki, and Hongbin Zha. An online system for multiple interacting targets tracking: Fusion of laser and vision, tracking and learning. ACM Transactions on Intelligent Systems and Technology (TIST), 4(1):18:1– 18:??, January 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Song:2014:UGF

[SZC^{+}14] Yicheng Song, Yongdong Zhang, Juan Cao, Jinhui Tang, Xingyu Gao, and Jintao Li. A unified geolocation framework for Web videos. ACM Transactions on Intelligent Systems and Technology (TIST), 5(3): 49:1–49:??, July 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Sun:2023:MUM

[SZL^{+}23] Li Sun, Zhongbao Zhang, Gen Li, Pengxin Ji, Sen Su, and Philip S. Yu. MC^{2}: Unsupervised multiple social network alignment. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):70:1–70:??, August 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3596514.

Song:2017:PSH

[SZS^{+}17] Xuan Song, Quanshi Zhang, Yoshihide Sekimoto, Ryosuke Shibasaki, Nicholas Jing Yuan, and Xing Xie. Prediction and simulation of human mobility following natural disasters. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2):29:1– 29:??, January 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shi:2012:BMA

[SZT12] Lixin Shi, Yuhang Zhao, and Jie Tang. Batch mode active learning for networked data. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):33:1– 33:??, February 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Sun:2015:LSV

[SZX15] Chao Sun, Tianzhu Zhang, and Changsheng Xu. Latent support vector machine modeling for sign language recognition with Kinect. ACM Transactions on Intelligent Systems and Technology (TIST), 6(2): 20:1–20:??, April 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Shi:2021:GGT

[SZZ^{+}21] Yukai Shi, Sen Zhang, Chenxing Zhou, Xiaodan Liang, Xiaojun Yang, and Liang Lin. GTAE: Graph transformer-based autoencoders for linguisticconstrained text style transfer. ACM Transactions on Intelligent Systems and Technology (TIST), 12 (3):32:1–32:16, July 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3448733.

Tariq:2019:EES

[TAL^{+}19] Umair Ullah Tariq, Haider Ali, Lu Liu, John Panneerselvam, and Xiaojun Zhai. Energy-efficient static task scheduling on VFI-based NoC–HMPSoCs for intelligent edge devices in cyberphysical systems. ACM Transactions on Intelligent Systems and Technology (TIST), 10(6):66:1–66:??, October 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tassa:2022:PPC

[TB22] Tamir Tassa and Alon Ben Horin. Privacypreserving collaborative filtering by distributed mediation. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):102:1–102:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3542950.

Talamadupula:2010:PHR

[TBK^{+}10] Kartik Talamadupula, J. Benton, Subbarao Kambhampati, Paul Schermerhorn, and Matthias Scheutz. Planning for human-robot teaming in open worlds. ACM Transactions on Intelligent Systems and Technology (TIST), 1(2):14:1– 14:??, November 2010. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tajeuna:2021:MCC

[TBW21] Etienne Gael Tajeuna, Mohamed Bouguessa, and Shengrui Wang. Mining customers’ changeable electricity consumption for effective load forecasting. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4):47:1– 47:26, August 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3466684.

Tonge:2019:PAT

[TC19] Ashwini Tonge and Cornelia Caragea. Privacyaware tag recommendation for accurate image privacy prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 10(4):40:1–40:??, August 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3335054.

Tama:2020:EID

[TCK20] Bayu Adhi Tama, Marco Comuzzi, and Jonghyeon Ko. An empirical investigation of different classifiers, encoding, and ensemble schemes for next event prediction using business process event logs. ACM Transactions on Intelligent Systems and Technology (TIST), 11(6):68:1– 68:34, November 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3406541.

Tabia:2013:PBA

[TDVC13] Hedi Tabia, Mohamed Daoudi, Jean-Philippe Vandeborre, and Olivier Colot. A parts-based approach for automatic 3D shape categorization using belief functions. ACM Transactions on Intelligent Systems and Technology (TIST), 4(2): 33:1–33:??, March 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Teng:2023:IHF

[Ten23] Shang-Hua Teng. “Intelligent heuristics are the future of computing”. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):96:1–96:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3627708.

Toole:2011:SCC

[TEP11] Jameson L. Toole, Nathan Eagle, and Joshua B. Plotkin. Spatiotemporal correlations in criminal offense records. ACM Transactions on Intelligent Systems and Technology (TIST), 2(4):38:1– 38:??, July 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tangmunarunkit:2015:OGE

[THL^{+}15] H. Tangmunarunkit, C. K.

Hsieh, B. Longstaff, S. Nolen, J. Jenkins, C. Ketcham, J. Selsky, F. Alquaddoomi, D. George, J. Kang, Z. Khalapyan, J. Ooms, N. Ramanathan, and D. Estrin. Ohmage: a general and extensible end-to-end participatory sensing platform. ACM Transactions on Intelligent Systems and Technology (TIST), 6(3): 38:1–38:??, May 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tang:2011:IAK

[THY^{+}11] Jinhui Tang, Richang Hong, Shuicheng Yan, Tat-Seng Chua, Guo-Jun Qi, and Ramesh Jain. Image annotation by kNN-sparse graph-based label propagation over noisily tagged web images. ACM Transactions on Intelligent Systems and Technology (TIST), 2(2):14:1–14:??, February 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tao:2021:PHM

[TJL^{+}21] Shuo Tao, Jingang Jiang, Defu Lian, Kai Zheng, and Enhong Chen. Predicting human mobility with reinforcement-learning-based long-term periodicity modeling. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):78:1–78:23, December 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3469860.

Tran:2023:MNB

[TL23] Nhu-Thuat Tran and Hady W. Lauw. Memory networkbased interpreter of user preferences in contentaware recommender systems. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):108:1–108:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3625239.

Tedjopurnomo:2021:STS

[TLB^{+}21] David Alexander Tedjopurnomo, Xiucheng Li, Zhifeng Bao, Gao Cong, Farhana Choudhury, and A. K. Qin. Similar trajectory search with spatiotemporal deep representation learning. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):77:1–77:26, December 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3466687.

Tan:2014:OOT

[TLC^{+}14] Chang Tan, Qi Liu, Enhong Chen, Hui Xiong, and Xiang Wu. Object-oriented travel package recommendation. ACM Transactions on Intelligent Systems and Technology (TIST), 5 (3):43:1–43:??, September 2014. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tu:2017:PPI

[TLLS17] Cunchao Tu, Zhiyuan Liu, Huanbo Luan, and Maosong Sun. PRISM: Profession identification in social media. ACM Transactions on Intelligent Systems and Technology (TIST), 8(6):81:1–81:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tang:2015:RTH

[TLW^{+}15] Ao Tang, Ke Lu, Yufei Wang, Jie Huang, and Houqiang Li. A real-time hand posture recognition system using deep neural networks. ACM Transactions on Intelligent Systems and Technology (TIST), 6 (2):21:1–21:??, April 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tong:2011:APL

[TLWZ11] Xiaofeng Tong, Jia Liu, Tao Wang, and Yimin Zhang. Automatic player labeling, tracking and field registration and trajectory mapping in broadcast soccer video. ACM Transactions on Intelligent Systems and Technology (TIST), 2(2):15:1–15:??, February 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tu:2020:LGI

[TMZ^{+}20] Xiaoguang Tu, Zheng Ma, Jian Zhao, Guodong Du, Mei Xie, and Jiashi Feng. Learning generalizable and identity-discriminative representations for face antispoofing. ACM Transactions on Intelligent Systems and Technology (TIST), 11(5):60:1–60:19, September 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3402446.

Tran:2013:CPB

[TNSP13] Vien Tran, Khoi Nguyen, Tran Cao Son, and Enrico Pontelli. A conformant planner based on approximation: CpA(H). ACM Transactions on Intelligent Systems and Technology (TIST), 4(2):36:1– 36:??, March 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Thukral:2019:DER

[TPG^{+}19] Deepak Thukral, Adesh Pandey, Rishabh Gupta, Vikram Goyal, and Tanmoy Chakraborty. DiffQue: Estimating relative difficulty of questions in community question answering services. ACM Transactions on Intelligent Systems and Technology (TIST), 10(4):42:1–42:??, August 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3337799.

Tian:2023:SFU

[TPM23] Qing Tian, Shun Peng, and Tinghuai Ma. Sourcefree unsupervised domain adaptation with trusted pseudo samples. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2):30:1– 30:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3570510.

Trivedi:2012:LSB

[TRDD12] Anusua Trivedi, Piyush Rai, Hal Daumé III, and Scott L. Duvall. Leveraging social bookmarks from partially tagged corpus for improved Web page clustering. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (4):67:1–67:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Thai:2012:VAO

[TRH12] Vinhtuan Thai, Pierre-Yves Rouille, and Siegfried Handschuh. Visual abstraction and ordering in faceted browsing of text collections. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):21:1– 21:??, February 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Towne:2016:MSS

[TRH16] W. Ben Towne, Carolyn P.

Rosé, and James D. Herbsleb. Measuring similarity similarly: LDA and human perception. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (1):7:1–7:??, October 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tang:2019:EPP

[TRZ^{+}19] Wenjuan Tang, Ju Ren, Kuan Zhang, Deyu Zhang, Yaoxue Zhang, and Xuemin (Sherman) Shen. Efficient and privacy-preserving fogassisted health data sharing scheme. ACM Transactions on Intelligent Systems and Technology (TIST), 10 (6):68:1–68:??, December 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3341104.

Tian:2017:TMS

[TS17] Mi Tian and Mark B. Sandler. Towards music structural segmentation across genres: Features, structural hypotheses, and annotation principles. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2):23:1–23:??, January 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tran:2018:RTF

[TTFS18] Luan Tran, Hien To, Liyue Fan, and Cyrus Shahabi. A real-time framework for task assignment in hyperlocal spatial crowdsourcing. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3):37:1– 37:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tian:2021:CIG

[TTL^{+}21] Jiajie Tian, Qihao Tang, Rui Li, Zhu Teng, Baopeng Zhang, and Jianping Fan. A camera identity-guided distribution consistency method for unsupervised multi-target domain person re-identification. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4):38:1–38:18, August 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3454130.

Tao:2017:LSC

[TTLG17] Dapeng Tao, Dacheng Tao, Xuelong Li, and Xinbo Gao. Large sparse cone non-negative matrix factorization for image annotation. ACM Transactions on Intelligent Systems and Technology (TIST), 8(3): 37:1–37:??, April 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tian:2023:DLD

[TWC^{+}23] Jieru Tian, Yongxin Wang, Zhenduo Chen, Xin Luo, and Xinshun Xu. Diagnose like doctors: Weakly supervised fine-grained classification of breast cancer. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2):34:1– 34:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3572033.

Tang:2011:GPU

[TWL11] Lei Tang, Xufei Wang, and Huan Liu. Group profiling for understanding social structures. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):15:1–15:??, October 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tian:2022:FWF

[TWL^{+}22] Yuanyishu Tian, Yao Wan, Lingjuan Lyu, Dezhong Yao, Hai Jin, and Lichao Sun. FedBERT: When federated learning meets pretraining. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):66:1–66:??, August 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3510033.

Tang:2012:RUI

[TY12] Xuning Tang and Christopher C. Yang. Ranking user influence in healthcare social media. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (4):73:1–73:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tang:2014:DSM

[TY14] Xuning Tang and Christopher C. Yang. Detecting social media hidden communities using dynamic stochastic blockmodel with temporal Dirichlet process. ACM Transactions on Intelligent Systems and Technology (TIST), 5(2):36:1– 36:??, April 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Tian:2020:MGD

[TZC^{+}20] Qing Tian, Wenqiang Zhang, Meng Cao, Liping Wang, Songcan Chen, and Hujun Yin. Momentguided discriminative manifold correlation learning on ordinal data. ACM Transactions on Intelligent Systems and Technology (TIST), 11(5):61:1–61:18, September 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3402445.

Tang:2013:FTC

[TZY^{+}13] Lu-An Tang, Yu Zheng, Jing Yuan, Jiawei Han, Alice Leung, Wen-Chih Peng, and Thomas La Porta. A framework of traveling companion discovery on trajectory data streams. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1):3:1– 3:??, December 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ullah:2015:ERL

[UAS15] Md Zia Ullah, Masaki Aono, and Md Hanif Seddiqui. Estimating a ranked list of human genetic diseases by associating phenotype-gene with genedisease bipartite graphs. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4):56:1– 56:??, August 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Verenich:2019:SCB

[VDL^{+}19] Ilya Verenich, Marlon Dumas, Marcello La Rosa, Fabrizio Maria Maggi, and Irene Teinemaa. Survey and cross-benchmark comparison of remaining time prediction methods in business process monitoring. ACM Transactions on Intelligent Systems and Technology (TIST), 10(4):34:1– 34:??, August 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3331449.

Vogogias:2019:BVS

[VKA^{+}19] Athanasios Vogogias, Jessie Kennedy, Daniel Archambault, Benjamin Bach, V. Anne Smith, and Hannah Currant. BayesPiles: Visualisation support for Bayesian network structure learning. ACM Transactions on Intelligent Systems and Technology (TIST), 10 (1):5:1–5:??, January 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3230623.

Varakantham:2018:RSS

[VKLY18] Pradeep Varakantham, Akshat Kumar, Hoong Chuin Lau, and William Yeoh. Risk-sensitive stochastic orienteering problems for trip optimization in urban environments. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3):24:1– 24:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Vasuki:2011:SAR

[VNL^{+}11] Vishvas Vasuki, Nagarajan Natarajan, Zhengdong Lu, Berkant Savas, and Inderjit Dhillon. Scalable affiliation recommendation using auxiliary networks. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):3:1– 3:??, October 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Verma:2022:IDB

[VPD^{+}22] Rohit Verma, Sugandh Pargal, Debasree Das, Tanusree Parbat, Sai Shankar Kambalapalli, Bivas Mitra, and Sandip Chakraborty. Impact of driving behavior on Commuter’s comfort during cab rides: Towards a new perspective of driver rating. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):87:1–87:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3523063.

Vu:2011:FSK

[VS11] Thuc Vu and Yoav Shoham.

Fair seeding in knockout tournaments. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):9:1–9:??, October 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2018:RFI

[WAL18] Suhang Wang, Charu Aggarwal, and Huan Liu. Random-forest-inspired neural networks. ACM Transactions on Intelligent Systems and Technology (TIST), 9(6):69:1– 69:??, November 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3232230.

Wang:2012:LCR

[WC12] Zhenxing Wang and Laiwan Chan. Learning causal relations in multivariate time series data. ACM Transactions on Intelligent Systems and Technology (TIST), 3(4):76:1–76:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wilson:2020:SUD

[WC20] Garrett Wilson and Diane J. Cook. A survey of unsupervised deep domain adaptation. ACM Transactions on Intelligent Systems and Technology (TIST), 11(5):51:1–51:46, September 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3400066.

Wyatt:2011:ICC

[WCBK11] Danny Wyatt, Tanzeem Choudhury, Jeff Bilmes, and James A. Kitts. Inferring colocation and conversation networks from privacy-sensitive audio with implications for computational social science. ACM Transactions on Intelligent Systems and Technology (TIST), 2(1):7:1–7:??, January 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2018:VME

[WCBL18] Bingsheng Wang, Zhiqian Chen, Arnold P. Boedihardjo, and Chang-Tien Lu. Virtual metering: an efficient water disaggregation algorithm via nonintrusive load monitoring. ACM Transactions on Intelligent Systems and Technology (TIST), 9(4):39:1– 39:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2020:TLD

[WCF^{+}20] Jindong Wang, Yiqiang Chen, Wenjie Feng, Han Yu, Meiyu Huang, and Qiang Yang. Transfer learning with dynamic distribution adaptation. ACM Transactions on Intelligent Systems and Technology (TIST), 11(1):6:1–6:25, February 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3360309.

Wu:2023:UGI

[WCFH23] Yingwen Wu, Sizhe Chen, Kun Fang, and Xiaolin Huang. Unifying gradients to improve real-world robustness for deep networks. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):101:1–101:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3617895.

Wang:2023:HSI

[WCP^{+}23] Wei-Yao Wang, Teng-Fong Chan, Wen-Chih Peng, Hui-Kuo Yang, Chih-Chuan Wang, and Yao-Chung Fan. How is the stroke? Inferring shot influence in badminton matches via long shortterm dependencies. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):7:1–7:??, February 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3551391.

Wang:2020:MHU

[WCS^{+}20] Jun-Zhe Wang, Yi-Cheng Chen, Wen-Yueh Shih, Lin Yang, Yu-Shao Liu, and Jiun-Long Huang. Mining high-utility temporal patterns on time intervalbased data. ACM Transactions on Intelligent Systems and Technology (TIST), 11 (4):43:1–43:31, July 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:

//dl.acm.org/doi/abs/
10.1145/3391230.

Wang:2013:ISS

[WDSZ13] Haifeng Wang, Bill Dolan, Idan Szpektor, and Shiqi Zhao. Introduction to special section on paraphrasing. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3): 37:1–37:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2012:EFW

[WFJY12] Haofen Wang, Linyun Fu, Wei Jin, and Yong Yu. EachWiki: Facilitating wiki authoring by annotation suggestion. ACM Transactions on Intelligent Systems and Technology (TIST), 3(4):71:1–71:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2021:PAR

[WFX^{+}21] Guang Wang, Zhihan Fang, Xiaoyang Xie, Shuai Wang, Huijun Sun, Fan Zhang, Yunhuai Liu, and Desheng Zhang. Pricing-aware real-time charging scheduling and charging station expansion for large-scale electric buses. ACM Transactions on Intelligent Systems and Technology (TIST), 12(1):13:1–13:26, February 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3428080.

Wang:2018:LUC

[WFZ^{+}18] Pengyang Wang, Yanjie Fu, Jiawei Zhang, Xiaolin Li, and Dan Lin. Learning urban community structures: a collective embedding perspective with periodic spatial-temporal mobility graphs. ACM Transactions on Intelligent Systems and Technology (TIST), 9(6):63:1– 63:??, November 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wu:2023:GFF

[WGF^{+}23] Hao Wu, Jianyang Gu, Xiaojin Fan, He Li, Lidong Xie, and Jian Zhao. 3Dguided frontal face generation for pose-invariant recognition. ACM Transactions on Intelligent Systems and Technology (TIST), 14 (2):31:1–31:??, April 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3572035.

Wang:2022:GBS

[WGL^{+}22] Zongwei Wang, Min Gao, Jundong Li, Junwei Zhang, and Jiang Zhong. Graybox shilling attack: an adversarial learning approach. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):82:1– 82:??, October 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3512352.

Wu:2010:OFU

[WH10] Fang Wu and Bernardo A.

Huberman. Opinion formation under costly expression. ACM Transactions on Intelligent Systems and Technology (TIST), 1 (1):5:1–5:??, October 2010. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2011:IIS

[WH11a] Jingdong Wang and Xian-Sheng Hua. Interactive image search by color map.

ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):12:1– 12:??, October 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2011:ALM

[WH11b] Meng Wang and Xian-Sheng Hua. Active learning in multimedia annotation and retrieval: a survey. ACM Transactions on Intelligent Systems and Technology (TIST), 2(2):10:1– 10:??, February 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2018:IHU

[WH18] Jun-Zhe Wang and Jiun-Long Huang. On incremental high utility sequential pattern mining. ACM Transactions on Intelligent Systems and Technology (TIST), 9(5):55:1– 55:??, July 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2013:LIC

[WHC13] Zhengxiang Wang, Yiqun Hu, and Liang-Tien Chia. Learning image-to-class distance metric for image classification. ACM Transactions on Intelligent Systems and Technology (TIST), 4(2):34:1– 34:??, March 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wu:2011:DML

[WHJ^{+}11] Lei Wu, Steven C. H. Hoi, Rong Jin, Jianke Zhu, and Nenghai Yu. Distance metric learning from uncertain side information for automated photo tagging. ACM Transactions on Intelligent Systems and Technology (TIST), 2(2):13:1– 13:??, February 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wolf:2013:DUR

[WHR13] Hannes Wolf, Klaus Herrmann, and Kurt Rothermel. Dealing with uncertainty: Robust workflow navigation in the healthcare domain. ACM Transactions on Intelligent Systems and Technology (TIST), 4(4):65:1–65:??, September 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wu:2022:RLV

[WHRGC22] Xian Wu, Chao Huang, Pablo Robles-Granda, and Nitesh V. Chawla. Representation learning on variable length and incomplete wearable-sensory time series. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):97:1–97:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3531228.

Wu:2021:BSB

[WHW^{+}21] Qiong Wu, Adam Hare, Sirui Wang, Yuwei Tu, Zhenming Liu, Christopher G. Brinton, and Yanhua Li. BATS: a spectral biclustering approach to single document topic modeling and segmentation. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):54:1– 54:29, October 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3468268.

Widmer:2017:GCE

[Wid17] Gerhard Widmer. Getting closer to the essence of music: The con espressione manifesto. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2):19:1–19:??, January 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2018:CAB

[WJY^{+}18] Pengwei Wang, Lei Ji, Jun Yan, Dejing Dou, Nisansa De Silva, Yong Zhang, and Lianwen Jin. Concept and attention-based CNN for question retrieval in multi-view learning. ACM Transactions on Intelligent Systems and Technology (TIST), 9(4):41:1– 41:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wu:2023:SCL

[WL23] Yanzhao Wu and Ling Liu. Selecting and composing learning rate policies for deep neural networks. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2): 22:1–22:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3570508.

Wu:2016:RMC

[WLC^{+}16] Le Wu, Qi Liu, Enhong Chen, Nicholas Jing Yuan, Guangming Guo, and Xing Xie. Relevance meets coverage: a unified framework to generate diversified recommendations. ACM Transactions on Intelligent Systems and Technology (TIST), 7(3):39:1– 39:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2020:FMH

[WLC^{+}20] Meng Wang, Hui Li, Jiangtao Cui, Sourav S. Bhowmick, and Ping Liu. FROST: Movement historyconscious facility relocation. ACM Transactions on Intelligent Systems and Technology (TIST), 11(1): 9:1–9:26, February 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:

                                        //dl.acm.org/doi/abs/
                                        10.1145/3361740.

Wang:2018:STP

[WLF^{+}18] Pengfei Wang, Guannan Liu, Yanjie Fu, Yuanchun Zhou, and Jianhui Li. Spotting trip purposes from taxi trajectories: a general probabilistic model. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3):29:1– 29:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2021:RLR

[WLFY21] Meng-Xiang Wang, Wang-Chien Lee, Tao-Yang Fu, and Ge Yu. On representation learning for road networks. ACM Transactions on Intelligent Systems and Technology (TIST), 12(1):11:1–11:27, February 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3424346.

Ward:2011:PMA

[WLG11] Jamie A. Ward, Paul Lukowicz, and Hans W. Gellersen. Performance metrics for activity recognition. ACM Transactions on Intelligent Systems and Technology (TIST), 2 (1):6:1–6:??, January 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2010:AIS

[WLH10] Meng Wang, Bo Liu, and Xian-Sheng Hua. Accessible image search for colorblindness. ACM Transactions on Intelligent Systems and Technology (TIST), 1 (1):8:1–8:??, October 2010. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wu:2017:CDT

[WLH17] Zhonggang Wu, Zhao Lu, and Shan-Yuan Ho. Community detection with topological structure and attributes in information networks. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2):33:1–33:??, January 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2020:EET

[WLL^{+}20] Min Wang, Congyan Lang, Liqian Liang, Songhe Feng, Tao Wang, and Yutong Gao. End-to-end textto-image synthesis with spatial constrains. ACM Transactions on Intelligent Systems and Technology (TIST), 11(4):47:1– 47:19, July 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3391709.

Wang:2021:FGS

[WLL^{+}21] Min Wang, Congyan Lang, Liqian Liang, Songhe Feng, Tao Wang, and Yutong Gao. Fine-grained semantic image synthesis with object-attention generative adversarial network. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):60:1–60:18, October 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3470008.

Wei:2022:WSV

[WLL^{+}22] Lili Wei, Congyan Lang, Liqian Liang, Songhe Feng, Tao Wang, and Shidi Chen. Weakly supervised video object segmentation via dual-attention cross-branch fusion. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3): 46:1–46:20, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3506716.

Wen:2022:DMC

[WLW^{+}22] Haomin Wen, Youfang Lin, Huaiyu Wan, Shengnan Guo, Fan Wu, Lixia Wu, Chao Song, and Yinghui Xu. Deep-Route+: Modeling couriers’ spatial-temporal behaviors and decision preferences for package pickup route prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):24:1– 24:23, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3481006.

Wen:2023:EWC

[WLW^{+}23] Haomin Wen, Youfang Lin, Fan Wu, Huaiyu Wan, Zhongxiang Sun, Tianyue Cai, Hongyu Liu, Shengnan Guo, Jianbin Zheng, Chao Song, and Lixia Wu. Enough waiting for the couriers: Learning to estimate package pick-up arrival time from couriers’ spatial-temporal behaviors. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3):50:1– 50:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3582561.

Wang:2023:TBE

[WLWC23] Hu Wang, Hui Li, Meng Wang, and Jiangtao Cui. Toward balancing the efficiency and effectiveness in k-facility relocation problem. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3): 52:1–52:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3587039.

Wang:2021:LMU

[WLWJ21] Huandong Wang, Yong Li, Gang Wang, and Depeng Jin. Linking multiple user identities of multiple services from massive mobility traces. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4):39:1–39:28, August 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3439817.

Wu:2023:DOL

[WLZ^{+}23] Wendi Wu, Zongren Li, Yawei Zhao, Chen Yu, Peilin Zhao, Ji Liu, and Kunlun He. Decentralized online learning: Take benefits from others’ data without sharing your own to track global trend. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):13:1–13:??, February 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3559765.

Winter:2021:CBS

[WM21] Felix Winter and Nysret Musliu. Constraintbased scheduling for paint shops in the automotive supply industry. ACM Transactions on Intelligent Systems and Technology (TIST), 12(2):17:1– 17:25, March 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3430710.

Waniek:2020:SAA

[WMA20] Marcin Waniek, Tomasz P.

Michalak, and Aamena Alshamsi. Strategic attack & defense in security diffusion games. ACM Transactions on Intelligent Systems and Technology (TIST), 11(1):5:1–5:35, February 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3357605.

Wu:2018:IDC

[WMH18] Ou Wu, Xue Mao, and Weiming Hu. Iteratively divide-and-conquer learning for nonlinear classification and ranking. ACM Transactions on Intelligent Systems and Technology (TIST), 9(2):18:1– 18:??, January 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wu:2017:TPU

[WMR17] Yanqiu Wu, Tehila Minkus, and Keith W. Ross. Taking the pulse of US college campuses with location-based anonymous mobile apps. ACM Transactions on Intelligent Systems and Technology (TIST), 9(1):12:1– 12:??, October 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Waniek:2022:HMC

[WMWR22] Marcin Waniek, Tomasz P. Michalak, Michael Wooldridge, and Talal Rahwan. How members of covert networks conceal the identities of their leaders. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):12:1–12:29, February 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3490462.

Wagstaff:2012:DLS

[WPA^{+}12] Kiri L. Wagstaff, Julian Panetta, Adnan Ansar, Ronald Greeley, Mary Pendleton Hoffer, Melissa Bunte, and Norbert Schörghofer. Dynamic landmarking for surface feature identification and change detection. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3):49:1– 49:??, May 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wei:2013:EPA

[WPL13] Ling-Yin Wei, Wen-Chih Peng, and Wang-Chien Lee. Exploring patternaware travel routes for trajectory search. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3):48:1– 48:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2014:VNF

[WSGM14] Shuaiqiang Wang, Jiankai Sun, Byron J. Gao, and Jun Ma. VSRank: a novel framework for rankingbased collaborative filtering. ACM Transactions on Intelligent Systems and Technology (TIST), 5(3): 51:1–51:??, July 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2015:WHP

[WST^{+}15] Yinting Wang, Mingli Song, Dacheng Tao, Yong Rui, Jiajun Bu, Ah Chung Tsoi, Shaojie Zhuo, and Ping Tan. Where2Stand: a human position recommendation system for souvenir photography. ACM Transactions on Intelligent Systems and Technology (TIST), 7(1):9:1–9:??, October 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2019:SBT

[WTK^{+}19] Hongjian Wang, Xianfeng Tang, Yu-Hsuan Kuo, Daniel Kifer, and Zhenhui Li. A simple baseline for travel time estimation using large-scale trip data. ACM Transactions on Intelligent Systems and Technology (TIST), 10(2):19:1– 19:??, February 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3293317.

Wang:2020:NMC

[WTL20] Chien-Chih Wang, Kent Loong Tan, and Chih-Jen Lin. Newton methods for convolutional neural networks. ACM Transactions on Intelligent Systems and Technology (TIST), 11(2):19:1– 19:30, March 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3368271.

Wang:2013:RCI

[WW13] Fei-Yue Wang and Pak Kin Wong. Research commentary: Intelligent systems and technology for integrative and predictive medicine: an ACP approach. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (2):32:1–32:??, March 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2021:CSG

[WWD^{+}21] Yu Wang, Yuelin Wang, Kai Dang, Jie Liu, and Zhuo Liu. A comprehensive survey of grammatical error correction. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):65:1–65:51, October 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3474840.

Wu:2022:FFN

[WWL^{+}22] Chuhan Wu, Fangzhao Wu, Lingjuan Lyu, Yongfeng Huang, and Xing Xie. FedCTR: Federated native ad CTR prediction with cross-platform user behavior data. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):62:1–62:??, August 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3506715.

Wang:2016:RPG

[WWZ^{+}16] Tianben Wang, Zhu Wang, Daqing Zhang, Tao Gu, Hongbo Ni, Jiangbo Jia, Xingshe Zhou, and Jing Lv. Recognizing Parkinsonian gait pattern by exploiting fine-grained movement function features. ACM Transactions on Intelligent Systems and Technology (TIST), 8(1):6:1–6:??, October 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2012:IOS

[WXLY12] Guan Wang, Sihong Xie, Bing Liu, and Philip S. Yu. Identify online store review spammers via social review graph. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (4):61:1–61:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2016:CFI

[WXZ^{+}16] Senzhang Wang, Sihong Xie, Xiaoming Zhang, Zhoujun Li, Philip S. Yu, and Yueying He. Coranking the future influence of multiobjects in bibliographic network through mutual reinforcement. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):64:1– 64:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2017:SSS

[WYC^{+}17] Weiqing Wang, Hongzhi Yin, Ling Chen, Yizhou Sun, Shazia Sadiq, and Xiaofang Zhou. ST-SAGE: a spatial-temporal sparse additive generative model for spatial item recommendation. ACM Transactions on Intelligent Systems and Technology (TIST), 8(3): 48:1–48:??, April 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2022:PMP

[WYC^{+}22] Yuandong Wang, Hongzhi Yin, Tong Chen, Chunyang Liu, Ben Wang, Tianyu Wo, and Jie Xu. Passenger mobility prediction via representation learning for dynamic directed and weighted graphs. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):2:1–2:25, February 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3446344.

Wang:2018:TTP

[WYD^{+}18] Weiqing Wang, Hongzhi Yin, Xingzhong Du, Quoc Viet Hung Nguyen, and Xiaofang Zhou. TPM: a temporal personalized model for spatial item recommendation. ACM Transactions on Intelligent Systems and Technology (TIST), 9 (6):61:1–61:??, November 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3230706.

Wang:2022:DDT

[WYG^{+}22] Liang Wang, Zhiwen Yu, Bin Guo, Dingqi Yang, Lianbo Ma, Zhidan Liu, and Fei Xiong. Datadriven targeted advertising recommendation system for outdoor billboard. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):29:1– 29:23, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3495159.

Wang:2017:TTD

[WYM17] Xuyu Wang, Chao Yang, and Shiwen Mao. Tensor-Beat: Tensor decomposition for monitoring multiperson breathing beats with commodity WiFi. ACM Transactions on Intelligent Systems and Technology (TIST), 9(1):8:1– 8:??, October 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wu:2020:DAC

[WYNW20] Hanrui Wu, Yuguang Yan, Michael K. Ng, and Qingyao Wu. Domainattention conditional Wasserstein distance for multisource domain adaptation. ACM Transactions on Intelligent Systems and Technology (TIST), 11(4):44:1– 44:19, July 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3391229.

Wen:2022:MWP

[WYP22] Yu-Ting Wen, Hui-Kuo Yang, and Wen-Chih Peng. Mining willing-to-pay behavior patterns from payment datasets. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):14:1–14:19, February 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3485848.

Wu:2019:OHT

[WYY^{+}19] Hanrui Wu, Yuguang Yan, Yuzhong Ye, Huaqing Min, Michael K. Ng, and Qingyao Wu. Online heterogeneous transfer learning by knowledge transition. ACM Transactions on Intelligent Systems and Technology (TIST), 10(3): 26:1–26:??, May 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3309537.

Wang:2023:DCT

[WYY^{+}23] Mudan Wang, Yuan Yuan, Huan Yan, Hongjie Sui, Fan Zuo, Yue Liu, Yong Li, and Depeng Jin. Discovering causes of traffic congestion via deep transfer clustering. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):79:1–79:??, October 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3604810.

Wang:2023:CVP

[WYZ23] Wenshan Wang, Su Yang, and Weishan Zhang. Customer volume prediction using fusion of sharedprivate dynamic weighting over multiple modalities. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3):42:1– 42:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3579826.

Wang:2021:MCB

[WZCJ21] Daheng Wang, Qingkai Zeng, Nitesh V. Chawla, and Meng Jiang. Modeling complementarity in behavior data with multitype itemset embedding. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4):42:1– 42:25, August 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3458724.

Wang:2021:ATS

[WZFL21] Senzhang Wang, Junbo Zhang, Yanjie Fu, and Yong Li. ACM TIST special issue on deep learning for spatio-temporal data: Part 1. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):67:1– 67:3, December 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3495188.

Wang:2022:ISI

[WZFL22] Senzhang Wang, Junbo Zhang, Yanjie Fu, and Yong Li. Introduction to the special issue on deep learning for spatiotemporal data:Part 2. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):17:1– 17:4, April 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3510023.

Wang:2016:SCW

[WZH16] Jialei Wang, Peilin Zhao, and Steven C. H. Hoi. Soft confidence-weighted learning. ACM Transactions on Intelligent Systems and Technology (TIST), 8(1): 15:1–15:??, October 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2014:IUI

[WZHL14] Jinpeng Wang, Wayne Xin Zhao, Yulan He, and Xiaoming Li. Infer user interests via link structure regularization. ACM Transactions on Intelligent Systems and Technology (TIST), 5 (2):23:1–23:??, April 2014. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2022:MCA

[WZM^{+}22] Senzhang Wang, Meiyue Zhang, Hao Miao, Zhaohui Peng, and Philip S. Yu. Multivariate correlationaware spatio-temporal graph convolutional networks for multi-scale traffic prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3): 38:1–38:22, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3469087.

Wang:2015:PLL

[WZS^{+}15] Yi Wang, Xuemin Zhao, Zhenlong Sun, Hao Yan, Lifeng Wang, Zhihui Jin, Liubin Wang, Yang Gao, Ching Law, and Jia Zeng. Peacock: Learning longtail topic features for industrial applications. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4):47:1– 47:??, August 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2020:ULT

[WZS^{+}20] Guang Wang, Fan Zhang, Huijun Sun, Yang Wang, and Desheng Zhang. Understanding the long-term evolution of electric taxi networks: a longitudinal measurement study on mobility and charging patterns. ACM Transactions on Intelligent Systems and Technology (TIST), 11(4): 48:1–48:27, July 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3393671.

Wang:2018:SCE

[WZY^{+}18] Leye Wang, Daqing Zhang, Dingqi Yang, Animesh Pathak, Chao Chen, Xiao Han, Haoyi Xiong, and Yasha Wang. SPACE-TA: Cost-effective task allocation exploiting intradata and interdata correlations in sparse crowdsensing. ACM Transactions on Intelligent Systems and Technology (TIST), 9(2): 20:1–20:??, January 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wang:2019:SZS

[WZYM19] Wei Wang, Vincent W. Zheng, Han Yu, and Chunyan Miao. A survey of zeroshot learning: Settings, methods, and applications. ACM Transactions on Intelligent Systems and Technology (TIST), 10(2):13:1– 13:??, February 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3293318.

Wei:2016:MDC

[WZZ^{+}16] Yunchao Wei, Yao Zhao, Zhenfeng Zhu, Shikui Wei, Yanhui Xiao, Jiashi Feng, and Shuicheng Yan.

Modality-dependent crossmedia retrieval. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):57:1– 57:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Wu:2021:TRL

[WZZ^{+}21] Wendi Wu, Yawei Zhao, En Zhu, Xinwang Liu, Xingxing Zhang, Lailong Luo, Shixiong Wang, and Jianping Yin. A theoretical revisit to linear convergence for saddle point problems. ACM Transactions on Intelligent Systems and Technology (TIST), 12(1):10:1–10:17, February 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3420035.

Xu:2023:SAR

[XHZ^{+}23] Ronghui Xu, Weiming Huang, Jun Zhao, Meng Chen, and Liqiang Nie. A spatial and adversarial representation learning approach for land use classification with POIs. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):114:1–114:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3627824.

Xie:2021:SDS

[XJS^{+}21] Yiqun Xie, Xiaowei Jia, Shashi Shekhar, Han Bao, and Xun Zhou. Significant DBSCAN+: Statistically robust densitybased clustering. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):62:1–62:26, October 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3474842.

Xin:2016:EGF

[XKW^{+}16] Bo Xin, Yoshinobu Kawahara, Yizhou Wang, Lingjing Hu, and Wen Gao. Efficient generalized fused lasso and its applications. ACM Transactions on Intelligent Systems and Technology (TIST), 7(4):60:1– 60:??, July 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Xu:2023:QOR

[XLB23] Jianqiu Xu, Hua Lu, and Zhifeng Bao. A query optimizer for range queries over multi-attribute trajectories. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):12:1–12:??, February 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3555811.

Xia:2020:DPP

[XLF^{+}20] Tong Xia, Yong Li, Jie Feng, Depeng Jin, Qing Zhang, Hengliang Luo, and Qingmin Liao. Deep-App: Predicting personalized smartphone app usage via context-aware multi-task learning. ACM Transactions on Intelligent Systems and Technology (TIST), 11(6):64:1–64:12, November 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3408325.

Xiang:2023:TQE

[XLG^{+}23] Tao Xiang, Hangcheng Liu, Shangwei Guo, Yan Gan, Wenjian He, and Xiaofeng Liao. Towards query-efficient black-Box attacks: a universal dual transferabilitybased framework. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):58:1–58:??, August 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3583777.

Xie:2022:ELF

[XLL^{+}22] Lunchen Xie, Jiaqi Liu, Songtao Lu, Tsung-Hui Chang, and Qingjiang Shi. An efficient learning framework for federated XG-Boost using secret sharing and distributed optimization. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):77:1–77:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3523061.

Xuan:2021:BNU

[XLZ21] Junyu Xuan, Jie Lu, and Guangquan Zhang. Bayesian nonparametric unsupervised concept drift detection for data stream mining. ACM Transactions on Intelligent Systems and Technology (TIST), 12(1): 5:1–5:22, February 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3420034.

Xu:2022:PPA

[XLZ^{+}22] Xiaolong Xu, Wentao Liu, Yulan Zhang, Xuyun Zhang, Wanchun Dou, Lianyong Qi, and Md Zakirul Alam Bhuiyan. PSDF: Privacy-aware IoV service deployment with federated learning in cloudedge computing. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):70:1–70:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3501810.

Xu:2023:DWP

[XSJW23] Meng Xu, Yechao She, Yang Jin, and Jianping Wang. Dynamic weights and prior reward in policy fusion for compound agent learning. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):107:1–107:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3623405.

Xie:2017:JSS

[XTW17] Liping Xie, Dacheng Tao, and Haikun Wei. Joint structured sparsity regularized multiview dimension reduction for video-based facial expression recognition. ACM Transactions on Intelligent Systems and Technology (TIST), 8(2): 28:1–28:??, January 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Xu:2023:DRL

[XW23] Meng Xu and Jianping Wang. Deep reinforcement learning for parameter tuning of robot visual servoing.

ACM Transactions on Intelligent Systems and Technology (TIST), 14(2):33:1– 33:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3579829.

Xu:2019:TSV

[XWC^{+}19] Mingliang Xu, Hua Wang, Shili Chu, Yong Gan, Xiaoheng Jiang, Yafei Li, and Bing Zhou. Traffic simulation and visual verification in smog. ACM Transactions on Intelligent Systems and Technology (TIST), 10 (1):3:1–3:??, January 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3200491.

Xia:2021:CSK

[XWW^{+}21] Zhenchang Xia, Jia Wu, Libing Wu, Yanjiao Chen, Jian Yang, and Philip S. Yu. A comprehensive survey of the key technologies and challenges surrounding vehicular ad hoc networks. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4):37:1– 37:30, August 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3451984.

Xu:2017:DOD

[XXL^{+}17] Jun Xu, Long Xia, Yanyan Lan, Jiafeng Guo, and Xueqi Cheng. Directly optimize diversity evaluation measures: a new approach to search result diversification. ACM Transactions on Intelligent Systems and Technology (TIST), 8(3): 41:1–41:??, April 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Xu:2023:CLM

[XXL^{+}23] Lingling Xu, Haoran Xie, Zongxi Li, Fu Lee Wang, Weiming Wang, and Qing Li. Contrastive learning models for sentence representations. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):67:1–67:??, August 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3593590.

Xu:2021:TTA

[XXZ^{+}21] Jiajie Xu, Saijun Xu, Rui Zhou, Chengfei Liu, An Liu, and Lei Zhao. TAML: a traffic-aware multi-task learning model for estimating travel time. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):75:1– 75:14, December 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3466686.

Xu:2023:MWB

[XYS^{+}23] En Xu, Zhiwen Yu, Zhuo Sun, Bin Guo, and Lina Yao. Modeling withinbasket auxiliary item recommendation with matchability and ubiquity. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3):49:1– 49:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3574157.

Xiong:2017:DDA

[XZH^{+}17] Haoyi Xiong, Jinghe Zhang, Yu Huang, Kevin Leach, and Laura E. Barnes. Daehr: a discriminant analysis framework for electronic health record data and an application to early detection of mental health disorders. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (3):47:1–47:??, April 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Xu:2012:MLE

[XZR12] Jun-Ming Xu, Xiaojin Zhu, and Timothy T. Rogers. Metric learning for estimating psychological similarities. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3): 55:1–55:??, May 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Xie:2020:DIS

[XZS20] Yiqun Xie, Xun Zhou, and Shashi Shekhar. Discovering interesting subpaths with statistical significance from spatiotemporal datasets. ACM Transactions on Intelligent Systems and Technology (TIST), 11(1):2:1–2:24, February 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3354189.

Xiong:2015:EEE

[XZW^{+}15] Haoyi Xiong, Daqing Zhang, Leye Wang, J. Paul Gibson, and Jie Zhu. EEMC: Enabling energy-efficient mobile crowdsensing with anonymous participants. ACM Transactions on Intelligent Systems and Technology (TIST), 6(3):39:1– 39:??, May 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Xie:2019:VAH

[XZXM19] Cong Xie, Wen Zhong, Wei Xu, and Klaus Mueller. Visual analytics of heterogeneous data using hypergraph learning. ACM Transactions on Intelligent Systems and Technology (TIST), 10(1):4:1– 4:??, January 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/ft_gateway.
                                        cfm?id=3200765.

Yang:2010:IAT

[Yan10] Qiang Yang. Introduction to ACM TIST. ACM Transactions on Intelligent Systems and Technology (TIST), 1(1):1:1–1:??, October 2010. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yadamjav:2020:QRC

[YBZ^{+}20] Munkh-Erdene Yadamjav, Zhifeng Bao, Baihua Zheng, Farhana M. Choudhury, and Hanan Samet. Querying recurrent convoys over trajectory data. ACM Transactions on Intelligent Systems and Technology (TIST), 11(5):59:1–59:24, September 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3400730.

Yang:2012:MRM

[YC12] Yi-Hsuan Yang and Homer H. Chen. Machine recognition of music emotion: a review. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3): 40:1–40:??, May 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yin:2012:LCT

[YCGH12] Zhijun Yin, Liangliang Cao, Quanquan Gu, and Jiawei Han. Latent community topic analysis: Integration of community discovery with topic modeling. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (4):63:1–63:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yao:2022:PPT

[YCH^{+}22] Lin Yao, Zhenyu Chen, Haibo Hu, Guowei Wu, and Bin Wu. Privacy preservation for trajectory publication based on differential privacy. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3): 42:1–42:21, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3474839.

Yan:2021:SWR

[YCL^{+}21] Caixia Yan, Xiaojun Chang, Minnan Luo, Qinghua Zheng, Xiaoqin Zhang, Zhihui Li, and Feiping Nie. Self-weighted robust LDA for multiclass classification with edge classes. ACM Transactions on Intelligent Systems and Technology (TIST), 12(1):4:1–4:19, February 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3418284.

Yan:2013:STM

[YCP^{+}13] Zhixian Yan, Dipanjan Chakraborty, Christine Parent, Stefano Spaccapietra, and Karl Aberer. Semantic trajectories: Mobility data computation and annotation. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (3):49:1–49:??, June 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yang:2023:RPL

[YCSH23] Wun-Ting Yang, Chiao-Ting Chen, Chuan-Yun Sang, and Szu-Hao Huang. Reinforced PU-learning with hybrid negative sampling strategies for recommendation. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3):57:1– 57:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3582562.

Yin:2023:CBA

[YCY23] Chunyong Yin, Shuangshuang Chen, and Zhichao Yin. Clustering-based active learning classification towards data stream. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2):38:1– 38:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3579830.

Yao:2023:AGA

[YCZ^{+}23] Rui Yao, Ying Chen, Yong Zhou, Fuyuan Hu, Jiaqi Zhao, Bing Liu, and Zhiwen Shao. Attentionguided adversarial attack for video object segmentation. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):102:1–102:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3617067.

Yin:2021:IFR

[YCZY21] Chunyong Yin, Haoqi Cuan, Yuhang Zhu, and Zhichao Yin. Improved fake reviews detection model based on vertical ensemble tri-training and active learning. ACM Transactions on Intelligent Systems and Technology (TIST), 12(3):33:1– 33:19, July 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3450285.

Ye:2020:XLA

[YDZ20] Juan Ye, Simon Dobson, and Franco Zambonelli. XLearn: Learning activity labels across heterogeneous datasets. ACM Transactions on Intelligent Systems and Technology (TIST), 11(2):17:1– 17:28, March 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3368272.

Yang:2018:CUS

[YFJ^{+}18] Longqi Yang, Chen Fang, Hailin Jin, Matthew D. Hoffman, and Deborah Estrin. Characterizing user skills from application usage traces with hierarchical attention recurrent networks. ACM Transactions on Intelligent Systems and Technology (TIST), 9 (6):68:1–68:??, November 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3232231.

Yang:2015:IPI

[YGU15] Yiyang Yang, Zhiguo Gong, and Leong Hou U. Identifying points of interest using heterogeneous features. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):68:1– 68:??, January 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yang:2023:CFS

[YGY^{+}23] Shuai Yang, Xianjie Guo, Kui Yu, Xiaoling Huang, Tingting Jiang, Jin He, and Lichuan Gu. Causal feature selection in the presence of sample selection bias. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):78:1– 78:??, October 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3604809.

Yuan:2021:MGC

[YHF21] Changsen Yuan, Heyan Huang, and Chong Feng. Multi-graph cooperative learning towards distant supervised relation extraction. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):52:1–52:21, October 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3466560.

Yu:2011:CBS

[YJHL11] Jie Yu, Xin Jin, Jiawei Han, and Jiebo Luo. Collectionbased sparse label propagation and its application on social group suggestion from photos. ACM Transactions on Intelligent Systems and Technology (TIST), 2(2):12:1– 12:??, February 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ying:2014:MUC

[YKTL14] Josh Jia-Ching Ying, Wen-Ning Kuo, Vincent S. Tseng, and Eric Hsueh-Chan Lu. Mining user check-in behavior with a random walk for urban point-of-interest recommendations. ACM Transactions on Intelligent Systems and Technology (TIST), 5(3):40:1–40:??, September 2014. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yang:2014:SOG

[YL14] Jaewon Yang and Jure Leskovec. Structure and overlaps of ground-truth communities in networks. ACM Transactions on Intelligent Systems and Technology (TIST), 5(2):26:1– 26:??, April 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yang:2017:TID

[YL17] Xitong Yang and Jiebo Luo. Tracking illicit drug dealing and abuse on Instagram using multimodal analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (4):58:1–58:??, July 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yao:2019:PAP

[YLC^{+}19] Huaxiu Yao, Defu Lian, Yi Cao, Yifan Wu, and Tao Zhou. Predicting academic performance for college students: a campus behavior perspective. ACM Transactions on Intelligent Systems and Technology (TIST), 10 (3):24:1–24:??, May 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3299087.

Yang:2019:FML

[YLCT19] Qiang Yang, Yang Liu, Tianjian Chen, and Yongxin Tong. Federated machine learning: Concept and applications. ACM Transactions on Intelligent Systems and Technology (TIST), 10(2):12:1–12:??, February 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3298981.

Yan:2022:CFM

[YLD^{+}22] Runze Yan, Xinwen Liu, Janine Dutcher, Michael Tumminia, Daniella Villalba, Sheldon Cohen, David Creswell, Kasey Creswell, Jennifer Mankoff, Anind Dey, and Afsaneh Doryab. A computational framework for modeling biobehavioral rhythms from mobile and wearable data streams. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3):47:1– 47:27, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3510029.

Yang:2023:HHG

[YLH^{+}23] Hanchen Yang, Wengen Li, Siyun Hou, Jihong Guan, and Shuigeng Zhou. HiGRN: a hierarchical graph recurrent network for global sea surface temperature prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):73:1–73:??, August 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3597937.

Ye:2015:SSB

[YLS15] Yanfang Ye, Tao Li, and Haiyin Shen. Soter: Smart bracelets for children’s safety. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4): 46:1–46:??, August 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ying:2013:MGT

[YLT13] Josh Jia-Ching Ying, Wang-Chien Lee, and Vincent S. Tseng. Mining geographictemporal-semantic patterns in trajectories for location prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 5(1):2:1–2:??, December 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yuan:2020:DTS

[YLWX20] Kun Yuan, Guannan Liu, Junjie Wu, and Hui Xiong. Dancing with Trump in the stock market: a deep information echoing model. ACM Transactions on Intelligent Systems and Technology (TIST), 11(5):62:1–62:22, September 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3403578.

Yao:2020:VOS

[YLX^{+}20] Rui Yao, Guosheng Lin, Shixiong Xia, Jiaqi Zhao, and Yong Zhou. Video object segmentation and tracking: a survey. ACM Transactions on Intelligent Systems and Technology (TIST), 11(4):36:1– 36:47, July 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3391743.

Yu:2019:RHR

[YLY^{+}19] Zeng Yu, Tianrui Li, Ning Yu, Yi Pan, Hongmei Chen, and Bing Liu. Reconstruction of hidden representation for robust feature extraction. ACM Transactions on Intelligent Systems and Technology (TIST), 10(2):18:1–18:??, February 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3284174.

Yao:2023:CCD

[YLY^{+}23] Jing Yao, Zheng Liu, Junhan Yang, Zhicheng Dou, Xing Xie, and Ji-Rong Wen. CDSM: Cascaded deep semantic matching on textual graphs leveraging ad-hoc neighbor selection. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2):32:1– 32:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3573204.

Ye:2016:GIL

[YMC16] Jintao Ye, Zhao Yan Ming, and Tat Seng Chua. Generating incremental length summary based on hierarchical topic coverage maximization. ACM Transactions on Intelligent Systems and Technology (TIST), 7 (3):29:1–29:??, April 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

You:2016:CFP

[YMLM16] Linlin You, Gianmario Motta, Kaixu Liu, and Tianyi Ma. CITY FEED: a pilot system of citizensourcing for city issue management. ACM Transactions on Intelligent Systems and Technology (TIST), 7 (4):53:1–53:??, July 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yang:2013:ISS

[YNS13] Shanchieh Jay Yang, Dana Nau, and John Salerno. Introduction to the special section on social computing, behavioral-cultural modeling, and prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3):51:1– 51:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yen:2013:LIS

[YSJ13] Neil Y. Yen, Timothy K. Shih, and Qun Jin. LONET: an interactive search network for intelligent lecture path generation. ACM Transactions on Intelligent Systems and Technology (TIST), 4(2):30:1– 30:??, March 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yao:2017:UCM

[YSN^{+}17] Lina Yao, Quan Z. Sheng, Anne H. H. Ngu, Xue Li, and Boualem Benattalah. Unveiling correlations via mining human-thing interactions in the Web of Things. ACM Transactions on Intelligent Systems and Technology (TIST), 8 (5):62:1–62:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yeh:2017:SIB

[YTH17] Lo-Yao Yeh, Woei-Jiunn Tsaur, and Hsin-Han Huang. Secure IoT-based, incentiveaware emergency personnel dispatching scheme with weighted fine-grained access control. ACM Transactions on Intelligent Systems and Technology (TIST), 9(1):10:1–10:??, October 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yang:2022:ISI

[YTL^{+}22a] Qiang Yang, Yongxin Tong, Yang Liu, Yangqiu Song, Hao Peng, and Boi Faltings. Introduction to the special issue on the federated learning: Algorithms, systems, and applications: Part 1. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):52:1–52:??, August 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3514223.

Yang:2022:PFL

[YTL^{+}22b] Qiang Yang, Yongxin Tong, Yang Liu, Yangqiu Song, Hao Peng, and Boi Faltings. Preface to federated learning: Algorithms, systems, and applications: Part 2. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):69:1–69:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3536420.

Yu:2023:ODI

[YWS^{+}23] Guangsheng Yu, Xu Wang, Caijun Sun, Ping Yu, Wei Ni, and Ren Ping Liu. Obfuscating the dataset: Impacts and applications. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):85:1– 85:??, October 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3597936.

Yin:2017:UUI

[YWZ^{+}17] Hao Yin, Wei Wang, Xu Zhang, Yongqiang Lyu, Geyong Min, and Dongchao Guo. UMCR: User interaction-driven mobile content retrieval. ACM Transactions on Intelligent Systems and Technology (TIST), 9(1):7:1–7:??, October 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yang:2023:SSD

[YWZ^{+}23] Wenlu Yang, Hongjun Wang, Yinghui Zhang, Zehao Liu, and Tianrui Li. Self-supervised discriminative representation learning by fuzzy autoencoder. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):11:1– 11:??, February 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3555777.

Yu:2023:STG

[YXL^{+}23] Shuo Yu, Feng Xia, Shihao Li, Mingliang Hou, and Quan Z. Sheng. Spatiotemporal graph learning for epidemic prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2):36:1– 36:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3579815.

Yang:2015:UHC

[YY15] Haodong Yang and Christopher C. Yang. Using health-consumer-contributed data to detect adverse drug reactions by association mining with temporal analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 6(4): 55:1–55:??, August 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yu:2013:ISS

[YZEC13] Zhiwen Yu, Daqing Zhang, Nathan Eagle, and Diane Cook. Introduction to the special section on intelligent systems for socially aware computing. ACM Transactions on Intelligent Systems and Technology (TIST), 4(3):45:1– 45:??, June 2013. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yang:2019:MAC

[YZL^{+}19] Bailin Yang, Luhong Zhang, Frederick W. B. Li, Xiaoheng Jiang, Zhigang Deng, Meng Wang, and Mingliang Xu. Motion-aware compression and transmission of mesh animation sequences. ACM Transactions on Intelligent Systems and Technology (TIST), 10 (3):25:1–25:??, May 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3300198.

Yang:2016:PCM

[YZQ16] Dingqi Yang, Daqing Zhang, and Bingqing Qu. Participatory cultural mapping based on collective behavior data in location-based social networks. ACM Transactions on Intelligent Systems and Technology (TIST), 7(3):30:1– 30:??, April 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Ying:2017:EIW

[YZY^{+}17] Xuhang Ying, Jincheng Zhang, Lichao Yan, Yu Chen, Guanglin Zhang, Minghua Chen, and Ranveer Chandra. Exploring indoor white spaces in metropolises. ACM Transactions on Intelligent Systems and Technology (TIST), 9(1):9:1– 9:??, October 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Yin:2023:HRD

[YZZ23] Chunyong Yin, Sun Zhang, and Qingkui Zeng. Hybrid representation and decision fusion towards visualtextual sentiment. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3):48:1– 48:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3583076.

Zhadan:2023:MAR

[ZAK^{+}23] Anastasia Zhadan, Alexander Allahverdyan, Ivan Kondratov, Vikenty Mikheev, Ovanes Petrosian, Aleksei Romanovskii, and Vitaliy Kharin. Multi-agent reinforcement learning-based adaptive heterogeneous DAG scheduling. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):87:1–87:??, October 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3610300.

Zhang:2020:WTE

[ZB20] Shuo Zhang and Krisztian Balog. Web table extraction, retrieval, and augmentation: a survey. ACM Transactions on Intelligent Systems and Technology (TIST), 11(2):13:1– 13:35, March 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3372117.

Zhong:2022:FHF

[ZBW^{+}22] Zhengyi Zhong, Weidong Bao, Ji Wang, Xiaomin Zhu, and Xiongtao Zhang. FLEE: a hierarchical federated learning framework for distributed deep neural network over cloud, edge, and end device. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):71:1–71:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3514501.

Zhou:2012:LAD

[ZBZX12] Ke Zhou, Jing Bai, Hongyuan Zha, and Gui-Rong Xue. Leveraging auxiliary data for learning to rank. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):37:1– 37:??, February 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2013:FTM

[ZC13] Jie Zhang and Robin Cohen. A framework for trust modeling in multiagent electronic marketplaces with buying advisors to consider varying seller behavior and the limiting of seller bids. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (2):24:1–24:??, March 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2015:SSI

[ZC15] Jia-Dong Zhang and Chi-Yin Chow. Spatiotemporal sequential influence modeling for location recommendations: a gravity-based approach. ACM Transactions on Intelligent Systems and Technology (TIST), 7(1):11:1–11:??, October 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhu:2015:FMF

[ZCG15] Yu Zhu, Wenbin Chen, and Guodong Guo. Fusing multiple features for depthbased action recognition. ACM Transactions on Intelligent Systems and Technology (TIST), 6(2):18:1– 18:??, April 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhuang:2018:SRL

[ZCL^{+}18] Fuzhen Zhuang, Xiaohu Cheng, Ping Luo, Sinno Jialin Pan, and Qing He. Supervised representation learning with double encodinglayer autoencoder for transfer learning. ACM Transactions on Intelligent Systems and Technology (TIST), 9(2):16:1–16:??, January 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhao:2021:PCL

[ZCL^{+}21] Qi Zhao, Chuqiao Chen, Guangcan Liu, Qingshan Liu, and Shengyong Chen. Parallel connected LSTM for matrix sequence prediction with elusive correlations. ACM Transactions on Intelligent Systems and Technology (TIST), 12(4): 51:1–51:16, August 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3469437.

Zhai:2012:MML

[ZCS^{+}12] Deming Zhai, Hong Chang, Shiguang Shan, Xilin Chen, and Wen Gao. Multiview metric learning with global consistency and local smoothness. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3):53:1– 53:??, May 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zheng:2014:ISS

[ZCWY14a] Yu Zheng, Licia Capra, Ouri Wolfson, and Hai Yang. Introduction to the special section on urban computing. ACM Transactions on Intelligent Systems and Technology (TIST), 5 (3):37:1–37:??, September 2014. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zheng:2014:UCC

[ZCWY14b] Yu Zheng, Licia Capra, Ouri Wolfson, and Hai Yang. Urban computing: Concepts, methodologies, and applications. ACM Transactions on Intelligent Systems and Technology (TIST), 5(3):38:1–38:??, September 2014. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2018:SCA

[ZCWZ18] Yexun Zhang, Wenbin Cai, Wenquan Wang, and Ya Zhang. Stopping criterion for active learning with model stability. ACM Transactions on Intelligent Systems and Technology (TIST), 9(2):19:1– 19:??, January 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhu:2015:MMU

[ZCX^{+}15] Hengshu Zhu, Enhong Chen, Hui Xiong, Kuifei Yu, Huanhuan Cao, and Jilei Tian. Mining mobile user preferences for personalized context-aware recommendation. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):58:1– 58:??, January 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhu:2023:LSA

[ZCZ^{+}23] Yupeng Zhu, Yanxiang Chen, Zuxing Zhao, Xueliang Liu, and Jinlin Guo. Local self-attention-based hybrid multiple instance learning for partial spoof speech detection. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):93:1–93:??, October 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3616540.

Zhang:2013:GVR

[ZDC^{+}13] Weishi Zhang, Guiguang Ding, Li Chen, Chunping Li, and Chengbo Zhang. Generating virtual ratings from Chinese reviews to augment online recommendations. ACM Transactions on Intelligent Systems and Technology (TIST), 4 (1):9:1–9:??, January 2013. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2012:GAS

[ZDL^{+}12] Ning Zhang, Ling-Yu Duan, Lingfang Li, Qingming Huang, Jun Du, Wen Gao, and Ling Guan. A generic approach for systematic analysis of sports videos. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3): 46:1–46:??, May 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhou:2019:LCN

[ZDW19] Junhao Zhou, Hong-Ning Dai, and Hao Wang. Lightweight convolution neural networks for mobile edge computing in transportation cyber physical systems. ACM Transactions on Intelligent Systems and Technology (TIST), 10 (6):67:1–67:??, December 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

https://dl.acm.org/ft_
gateway.cfm?id=3339308.

Zhang:2018:RCS

[ZFH^{+}18] Dingwen Zhang, Huazhu Fu, Junwei Han, Ali Borji, and Xuelong Li. A review of co-saliency detection algorithms: Fundamentals, applications, and challenges. ACM Transactions on Intelligent Systems and Technology (TIST), 9(4):38:1– 38:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2022:CST

[ZFH^{+}22] Yao Zhang, Wenping Fan, Qichen Hao, Xinya Wu, and Min-Ling Zhang. CAFE and SOUP: Toward adaptive VDI workload prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):94:1–94:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3529536.

Zhang:2022:IPI

[ZFLD22] Lin Zhang, Lixin Fan, Yong Luo, and Ling-Yu Duan. Intrinsic performance influence-based participant contribution estimation for horizontal federated learning. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):88:1–88:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3523059.

Zhang:2020:KAA

[ZFQX20] Yingying Zhang, Quan Fang, Shengsheng Qian, and Changsheng Xu. Knowledgeaware attentive Wasserstein adversarial dialogue response generation. ACM Transactions on Intelligent Systems and Technology (TIST), 11(4):37:1– 37:20, July 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3384675.

Zhao:2019:PRG

[ZFS^{+}19] Guoshuai Zhao, Hao Fu, Ruihua Song, Tetsuya Sakai, Zhongxia Chen, Xing Xie, and Xueming Qian. Personalized reason generation for explainable song recommendation. ACM Transactions on Intelligent Systems and Technology (TIST), 10(4):41:1– 41:??, August 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3337967.

Zhang:2017:EEM

[ZFWL17] Wei Zhang, Rui Fan, Yonggang Wen, and Fang Liu. Energy-efficient mobile video streaming: a location-aware approach. ACM Transactions on Intelligent Systems and Technology (TIST), 9(1):6:1– 6:??, October 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2023:NFL

[ZGF^{+}23] Xiaojin Zhang, Hanlin Gu, Lixin Fan, Kai Chen, and Qiang Yang. No free lunch theorem for security and utility in federated learning. ACM Transactions on Intelligent Systems and Technology (TIST), 14(1):1:1–1:??, February 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3563219.

Zhao:2017:PLS

[ZGL^{+}17] Hongke Zhao, Yong Ge, Qi Liu, Guifeng Wang, Enhong Chen, and Hefu Zhang. P2P lending survey: Platforms, recent advances and prospects. ACM Transactions on Intelligent Systems and Technology (TIST), 8(6):72:1–72:??, September 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2018:DLE

[ZGP^{+}18] Zixing Zhang, Jürgen Geiger, Jouni Pohjalainen, Amr El-Desoky Mousa, Wenyu Jin, and Björn Schuller. Deep learning for environmentally robust speech recognition: an overview of recent developments. ACM Transactions on Intelligent Systems and Technology (TIST), 9(5): 49:1–49:??, July 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zheng:2015:TDM

[Zhe15]  Yu Zheng.  Trajectory

data mining: an overview. ACM Transactions on Intelligent Systems and Technology (TIST), 6(3):29:1– 29:??, May 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhu:2021:CLG

[ZHLL21] Yisheng Zhu, Hu Han, Guangcan Liu, and Qingshan Liu. Collaborative local-global learning for temporal action proposal. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):55:1– 55:14, October 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3466181.

Zhuo:2019:RMA

[Zhu19] Hankz Hankui Zhuo. Recognizing multi-agent plans when action models and team plans are both incomplete. ACM Transactions on Intelligent Systems and Technology (TIST), 10(3): 30:1–30:??, May 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3319403.

Zhou:2021:AEA

[ZHW^{+}21] Qianli Zhou, Tianrui Hui, Rong Wang, Haimiao Hu, and Si Liu. Attentive excitation and aggregation for bilingual referring image segmentation. ACM Transactions on Intelligent Systems and Technology (TIST), 12(2):26:1– 26:17, March 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3446345.

Zhang:2018:RTH

[ZHZ18] Desheng Zhang, Tian He, and Fan Zhang. Real-time human mobility modeling with multi-view learning. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3):22:1– 22:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhou:2021:UMB

[ZJSY21] Yiyi Zhou, Rongrong Ji, Jinsong Su, and Jiaquan Yao. Uncovering media bias via social network learning. ACM Transactions on Intelligent Systems and Technology (TIST), 12(1):12:1– 12:12, February 2021. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3422181.

Zhang:2023:TPU

[ZKC^{+}23] Xiaojin Zhang, Yan Kang, Kai Chen, Lixin Fan, and Qiang Yang. Trading off privacy, utility, and efficiency in federated learning. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):98:1–98:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3595185.

Zhang:2012:DFE

[ZL12] Yi Zhang and Tao Li.

DClusterE: a framework for evaluating and understanding document clustering using visualization. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):24:1– 24:??, February 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2019:UEO

[ZL19] Jason Shuo Zhang and Qin Lv. Understanding event organization at scale in event-based social networks. ACM Transactions on Intelligent Systems and Technology (TIST), 10(2): 16:1–16:??, February 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3243227.

Zhang:2016:PAT

[ZLB^{+}16] Kun Zhang, Jiuyong Li, Elias Bareinboim, Bernhard Schölkopf, and Judea Pearl. Preface to the ACM TIST special issue on causal discovery and inference. ACM Transactions on Intelligent Systems and Technology (TIST), 7(2): 17:1–17:??, January 2016. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zaji:2023:OBD

[ZLBZ23] Amirhossein Zaji, Zheng Liu, Takashi Bando, and Lihua Zhao. Ontologybased driving simulation for traffic lights optimization. ACM Transactions on Intelligent Systems and Technology (TIST), 14(3): 39:1–39:??, June 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3579839.

Zhu:2020:FMM

[ZLC^{+}20] Lei Zhu, Xu Lu, Zhiyong Cheng, Jingjing Li, and Huaxiang Zhang. Flexible multi-modal hashing for scalable multimedia retrieval. ACM Transactions on Intelligent Systems and Technology (TIST), 11(2): 14:1–14:20, March 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https:

//dl.acm.org/doi/abs/
10.1145/3365841.

Zhang:2023:NAL

[ZLD^{+}23] Jinghui Zhang, Dingyang Lv, Qiangsheng Dai, Fa Xin, and Fang Dong. Noiseaware local model training mechanism for federated learning. ACM Transactions on Intelligent Systems and Technology (TIST), 14(4):65:1–65:??, August 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3591363.

Zhou:2020:DCN

[ZLG^{+}20] Yuxiang Zhou, Lejian Liao, Yang Gao, Heyan Huang, and Xiaochi Wei. A discriminative convolutional neural network with context-aware attention. ACM Transactions on Intelligent Systems and Technology (TIST), 11(5):57:1–57:21, September 2020. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3397464.

Zhang:2018:UOG

[ZLH18] Yingjie Zhang, Beibei Li, and Jason Hong. Using online geotagged and crowdsourced data to understand human offline behavior in the city: an economic perspective. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3):32:1–32:??, February 2018. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zheng:2022:UAP

[ZLL^{+}22] Zhirun Zheng, Zhetao Li, Jie Li, Hongbo Jiang, Tong Li, and Bin Guo. Utility-aware and privacypreserving trajectory synthesis model that resists social relationship privacy attacks. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3): 44:1–44:28, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3495160.

Zheng:2022:ISIa

[ZLSY22a] Kai Zheng, Yong Li, Cyrus Shahabi, and Hongzhi Yin. Introduction to the special issue on intelligent trajectory analytics: Part I. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):1:1– 1:2, February 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3495230.

Zheng:2022:ISIb

[ZLSY22b] Kai Zheng, Yong Li, Cyrus Shahabi, and Hongzhi Yin. Introduction to the special issue on intelligent trajectory analytics: Part II. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3):34:1– 34:2, June 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3510021.

Zhang:2015:RTS

[ZLT15] Liyan Zhang, Fan Liu, and Jinhui Tang. Real-time system for driver fatigue detection by RGB-D camera. ACM Transactions on Intelligent Systems and Technology (TIST), 6(2):22:1– 22:??, April 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2018:GER

[ZLY^{+}18] Chao Zhang, Dongming Lei, Quan Yuan, Honglei Zhuang, Lance Kaplan, Shaowen Wang, and Jiawei Han. GeoBurst+: Effective and real-time local event detection in geotagged tweet streams. ACM Transactions on Intelligent Systems and Technology (TIST), 9(3):34:1– 34:??, February 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2015:SPL

[ZLZ15] Zhao Zhang, Cheng-Lin Liu, and Ming-Bo Zhao. A sparse projection and lowrank recovery framework for handwriting representation and salient stroke feature extraction. ACM Transactions on Intelligent Systems and Technology (TIST), 6(1):9:1–9:??, March 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2017:LKK

[ZLZ^{+}17] Shichao Zhang, Xuelong Li, Ming Zong, Xiaofeng Zhu, and Debo Cheng. Learning k for kNN classification. ACM Transactions on Intelligent Systems and Technology (TIST), 8(3):43:1– 43:??, April 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2022:UTD

[ZLZ^{+}22] Yingxue Zhang, Yanhua Li, Xun Zhou, Jun Luo, and Zhi-Li Zhang. Urban traffic dynamics prediction — a continuous spatial-temporal metalearning approach. ACM Transactions on Intelligent Systems and Technology (TIST), 13(2):23:1– 23:19, April 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3474837.

Zhao:2023:FCG

[ZLZW23] Tianxiang Zhao, Dongsheng Luo, Xiang Zhang, and Suhang Wang. Faithful and consistent graph neural network explanations with rationale alignment. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):92:1– 92:??, October 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3616542.

Zhuang:2015:CDS

[ZMH^{+}15] Jinfeng Zhuang, Tao Mei, Steven C. H. Hoi, Xian-Sheng Hua, and Yongdong Zhang. Community discovery from social media by low-rank matrix recovery. ACM Transactions on Intelligent Systems and Technology (TIST), 5(4):67:1– 67:??, January 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zheng:2022:SDA

[ZMH^{+}22] Bolong Zheng, Lingfeng Ming, Qi Hu, Zhipeng Lü, Guanfeng Liu, and Xiaofang Zhou. Supplydemand-aware deep reinforcement learning for dynamic fleet management. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3):37:1– 37:19, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3467979.

Zhao:2014:PRL

[ZNWC14] Yi-Liang Zhao, Liqiang Nie, Xiangyu Wang, and Tat-Seng Chua. Personalized recommendations of locally interesting venues to tourists via crossregion community matching. ACM Transactions on Intelligent Systems and Technology (TIST), 5(3): 50:1–50:??, July 2014. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhou:2011:RPA

[ZNYH11] Yue Zhou, Bingbing Ni, Shuicheng Yan, and Thomas S. Huang. Recognizing pairactivities by causality analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 2(1): 5:1–5:??, January 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zheng:2020:CAD

[ZPL^{+}20] Zimu Zheng, Jie Pu, Linghui Liu, Dan Wang, Xiangming Mei, Sen Zhang, and Quanyu Dai. Contextual anomaly detection in solder paste inspection with multi-task learning. ACM Transactions on Intelligent Systems and Technology (TIST), 11(6):65:1– 65:17, November 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3383261.

Zhou:2021:PEL

[ZPP^{+}21] Haoyi Zhou, Hao Peng, Jieqi Peng, Shuai Zhang, and Jianxin Li. POLLA: Enhancing the local structure awareness in long sequence spatial-temporal modeling. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):69:1–69:24, December 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3447987.

Zhang:2011:ISI

[ZPY11] Daqing Zhang, Matthai Philipose, and Qiang Yang. Introduction to the special issue on intelligent systems for activity recognition. ACM Transactions on Intelligent Systems and Technology (TIST), 2(1):1:1– 1:??, January 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2015:CSS

[ZQP^{+}15] Wangsheng Zhang, Guande Qi, Gang Pan, Hua Lu, Shijian Li, and Zhaohui Wu. City-scale social event detection and evaluation with taxi traces. ACM Transactions on Intelligent Systems and Technology (TIST), 6 (3):40:1–40:??, May 2015. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhu:2022:SPC

[ZRX^{+}22] Yanliang Zhu, Dongchun Ren, Yi Xu, Deheng Qian, Mingyu Fan, Xin Li, and Huaxia Xia. Simultaneous past and current social interaction-aware trajectory prediction for multiple intelligent agents in dynamic scenes. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):10:1–10:16, February 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3466182.

Zdesar:2018:OVP

[ZS18] Andrej Zdesar and Igor Skrjanc. Optimum velocity profile of multiple Bernstein–Bézier curves subject to constraints for mobile robots. ACM Transactions on Intelligent Systems and Technology (TIST), 9(5):56:1– 56:??, July 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2020:AAD

[ZSAL20] Wei Emma Zhang, Quan Z.

Sheng, Ahoud Alhazmi, and Chenliang Li. Adversarial attacks on deeplearning models in natural language processing: a survey. ACM Transactions on Intelligent Systems and Technology (TIST), 11(3): 24:1–24:41, May 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3374217.

Zhang:2015:EDQ

[ZSL^{+}15] Bo Zhang, Zheng Song, Chi Harold Liu, Jian Ma, and Wendong Wang. An event-driven QoI-aware participatory sensing framework with energy and budget constraints. ACM Transactions on Intelligent Systems and Technology (TIST), 6(3):42:1– 42:??, May 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zamani:2019:AAT

[ZSLC19] Hamed Zamani, Markus Schedl, Paul Lamere, and Ching-Wei Chen. An analysis of approaches taken in the ACM RecSys Challenge 2018 for automatic music playlist continuation. ACM Transactions on Intelligent Systems and Technology (TIST), 10(5):57:1– 57:??, October 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2015:RDI

[ZSS^{+}15] Quanshi Zhang, Xuan Song, Xiaowei Shao, Huijing Zhao, and Ryosuke Shibasaki. From RGB-D images to RGB images: Single labeling for mining visual models. ACM Transactions on Intelligent Systems and Technology (TIST), 6(2):16:1– 16:??, April 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2012:EAL

[ZSY^{+}12] Xiaoqin Shelley Zhang, Bhavesh Shrestha, Sungwook Yoon, Subbarao Kambhampati, Phillip Di-Bona, Jinhong K. Guo, Daniel McFarlane, Martin O. Hofmann, Kenneth Whitebread, Darren Scott Appling, Elizabeth T. Whitaker, Ethan B. Trewhitt, Li Ding, James R. Michaelis, Deborah L. McGuinness, James A. Hendler, Janardhan Rao Doppa, Charles Parker, Thomas G. Dietterich, Prasad Tadepalli, Weng-Keen Wong, Derek Green, Anton Rebguns, Diana Spears, Ugur Kuter, Geoff Levine, Gerald DeJong, Reid L. Mac-Tavish, Santiago Ontañón, Jainarayan Radhakrishnan, Ashwin Ram, Hala Mostafa, Huzaifa Zafar, Chongjie Zhang, Daniel Corkill, Victor Lesser, and Zhexuan Song. An ensemble architecture for learning complex problem-solving techniques from demonstration. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (4):75:1–75:??, September 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2011:HMF

[ZT11] Richong Zhang and Thomas Tran. A helpfulness modeling framework for electronic word-of-mouth on consumer opinion platforms. ACM Transactions on Intelligent Systems and Technology (TIST), 2(3): 23:1–23:??, April 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2019:RVA

[ZW19] Chen Zhang and Hao Wang. ResumeVis: a visual analytics system to discover semantic information in semi-structured resume data. ACM Transactions on Intelligent Systems and Technology (TIST), 10 (1):8:1–8:??, January 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL

                                        https://dl.acm.org/ft_
                                        gateway.cfm?id=3230707.

Zhang:2023:OOS

[ZWC23] Donglin Zhang, Xiao-Jun Wu, and Guoqing Chen. ONION: Online semantic autoencoder hashing for cross-modal retrieval. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2):27:1– 27:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3572032.

Zhu:2017:ATS

[ZWGW17] Wenwu Zhu, Jean Walrand, Yike Guo, and Zhi Wang. ACM TIST special issue on data-driven intelligence for wireless networking. ACM Transactions on Intelligent Systems and Technology (TIST), 9 (1):4:1–4:??, October 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2017:LBP

[ZWH17] Jiaming Zhang, Shuhui Wang, and Qingming Huang. Location-based parallel tag completion for geo-tagged social image retrieval. ACM Transactions on Intelligent Systems and Technology (TIST), 8(3): 38:1–38:??, April 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zheng:2022:JOE

[ZWH^{+}22] Qian Zheng, Yueming Wang, Zhenfang Hu, Xiaobo Zhang, Zhaohui Wu, and Gang Pan. Jointly optimizing expressional and residual models for 3D facial expression removal. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):90:1– 90:??, December 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3533312.

Zhang:2015:TGO

[ZWL^{+}15] Mingjin Zhang, Huibo Wang, Yun Lu, Tao Li, Yudong Guang, Chang Liu, Erik Edrosa, Hongtai Li, and Naphtali Rishe. TerraFly GeoCloud: an online spatial data analysis and visualization system. ACM Transactions on Intelligent Systems and Technology (TIST), 6(3):34:1– 34:??, May 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhao:2019:VAA

[ZWL^{+}19] Ying Zhao, Lei Wang, Shijie Li, Fangfang Zhou, Xiaoru Lin, Qiang Lu, and Lei Ren. A visual analysis approach for understanding durability test data of automotive products. ACM Transactions on Intelligent Systems and Technology (TIST), 10(6):70:1–70:??, December 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhou:2022:CTL

[ZWX^{+}22] Fan Zhou, Pengyu Wang, Xovee Xu, Wenxin Tai, and Goce Trajcevski. Contrastive trajectory learning for tour recommendation. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):4:1– 4:25, February 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3462331.

Zhang:2012:AKR

[ZWXZ12] Weinan Zhang, Dingquan Wang, Gui-Rong Xue, and Hongyuan Zha. Advertising keywords recommendation for short-text Web pages using Wikipedia. ACM Transactions on Intelligent Systems and Technology (TIST), 3(2):36:1– 36:??, February 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2019:MVF

[ZWZ^{+}19] Yongshan Zhang, Jia Wu, Chuan Zhou, Zhihua Cai, Jian Yang, and Philip S. Yu. Multi-view fusion with extreme learning machine for clustering. ACM Transactions on Intelligent Systems and Technology (TIST), 10(5):53:1–53:??, October 2019. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2016:EFC

[ZWZS16] Kun Zhang, Zhikun Wang, Jiji Zhang, and Bernhard Schölkopf. On estimation of functional causal models: General results and application to the postnonlinear causal model. ACM Transactions on Intelligent Systems and Technology (TIST), 7(2):13:1– 13:??, January 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhou:2023:LEC

[ZWZZ23] Jianhang Zhou, Guancheng Wang, Shaoning Zeng, and Bob Zhang. Learning with Euler collaborative representation for robust pattern analysis. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):109:1–109:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3625235.

Zheng:2011:LTR

[ZX11] Yu Zheng and Xing Xie. Learning travel recommendations from usergenerated GPS traces. ACM Transactions on Intelligent Systems and Technology (TIST), 2(1):2:1– 2:??, January 2011. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhu:2023:UGR

[ZXY^{+}23] Yanqiao Zhu, Yichen Xu, Feng Yu, Qiang Liu, and Shu Wu. Unsupervised graph representation learning with cluster-aware selftraining and refining. ACM Transactions on Intelligent Systems and Technology (TIST), 14(5):82:1–82:??, October 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3608480.

Zhang:2012:TML

[ZY12] Yu Zhang and Dit-Yan Yeung. Transfer metric learning with semi-supervised extension. ACM Transactions on Intelligent Systems and Technology (TIST), 3 (3):54:1–54:??, May 2012. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zeng:2022:CCB

[ZYC^{+}22] Bixiao Zeng, Xiaodong Yang, Yiqiang Chen, Hanchao Yu, and Yingwei Zhang. CLC: a consensusbased label correction approach in federated learning. ACM Transactions on Intelligent Systems and Technology (TIST), 13(5):75:1–75:??, October 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3519311.

Zhang:2017:DSM

[ZYH^{+}17] Peng Zhang, Qian Yu, Yuexian Hou, Dawei Song, Jingfei Li, and Bin Hu. A distribution separation method using irrelevance feedback data for information retrieval. ACM Transactions on Intelligent Systems and Technology (TIST), 8(3):46:1– 46:??, April 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2020:DMB

[ZYH^{+}20] Xiang Zhang, Lina Yao, Chaoran Huang, Tao Gu, Zheng Yang, and Yunhao Liu. DeepKey: a multimodal biometric authentication system via deep decoding gaits and brainwaves. ACM Transactions on Intelligent Systems and Technology (TIST), 11(4): 49:1–49:24, July 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3393619.

Zhang:2012:RVT

[ZYSL12] Shengping Zhang, Hongxun Yao, Xin Sun, and Shaohui Liu. Robust visual tracking using an effective appearance model based on sparse coding. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3): 43:1–43:??, May 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zha:2015:RMF

[ZYT^{+}15] Zheng-Jun Zha, Yang Yang, Jinhui Tang, Meng Wang, and Tat-Seng Chua. Robust multiview feature learning for RGB-D image understanding. ACM Transactions on Intelligent Systems and Technology (TIST), 6(2):15:1– 15:??, April 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2015:SPU

[ZYW^{+}15] Fuzheng Zhang, Nicholas Jing Yuan, David Wilkie, Yu Zheng, and Xing Xie. Sensing the pulse of urban refueling behavior: a perspective from taxi mobility. ACM Transactions on Intelligent Systems and Technology (TIST), 6(3):37:1– 37:??, May 2015. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2023:ESV

[ZYY^{+}23] Guozhen Zhang, Jinhui Yi, Jian Yuan, Yong Li, and Depeng Jin. DAS: Efficient street view image sampling for urban prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 14(2): 35:1–35:??, April 2023. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3576902.

Zheng:2012:MTP

[ZZC12] Yan-Tao Zheng, Zheng-Jun Zha, and Tat-Seng Chua. Mining travel patterns from geotagged photos. ACM Transactions on Intelligent Systems and Technology (TIST), 3(3): 56:1–56:??, May 2012. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhou:2020:FPT

[ZZC^{+}20] Binbin Zhou, Sha Zhao, Longbiao Chen, Shijian Li, Zhaohui Wu, and Gang Pan. Forecasting price trend of bulk commodities leveraging crossdomain open data fusion. ACM Transactions on Intelligent Systems and Technology (TIST), 11(1):1:1– 1:26, February 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3354287.

Zhou:2022:TSP

[ZZC^{+}22] Jun Zhou, Longfei Zheng, Chaochao Chen, Yan Wang, Xiaolin Zheng, Bingzhe Wu, Cen Chen, Li Wang, and Jianwei Yin. Toward scalable and privacypreserving deep neural network via algorithmiccryptographic co-design. ACM Transactions on Intelligent Systems and Technology (TIST), 13(4):53:1– 53:??, August 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3501809.

Zhang:2017:EPE

[ZZD^{+}17] Ruide Zhang, Ning Zhang, Changlai Du, Wenjing Lou, Y. Thomas Hou, and Yuichi Kawamoto. From electromyogram to password: Exploring the privacy impact of wearables in augmented reality. ACM Transactions on Intelligent Systems and Technology (TIST), 9(1):13:1– 13:??, October 2017. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2019:MCI

[ZZGH19] Hao Zhang, Shuigeng Zhou, Jihong Guan, and Jun (Luke) Huan. Measuring conditional independence by independent residuals for causal discovery. ACM Transactions on Intelligent Systems and Technology (TIST), 10(5):50:1– 50:??, October 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhou:2022:GNN

[ZZH^{+}22] Yu Zhou, Haixia Zheng, Xin Huang, Shufeng Hao, Dengao Li, and Jumin Zhao. Graph neural networks: Taxonomy, advances, and trends. ACM Transactions on Intelligent Systems and Technology (TIST), 13(1):15:1–15:54, February 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1145/3495161.

Zhuo:2020:DUP

[ZZKT20] Hankz Hankui Zhuo, Yantian Zha, Subbarao Kambhampati, and Xin Tian. Discovering underlying plans based on shallow models. ACM Transactions on Intelligent Systems and Technology (TIST), 11(2):18:1– 18:30, March 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3368270.

Zhu:2019:EVC

[ZZL^{+}19] Chunbiao Zhu, Wenhao Zhang, Thomas H. Li, Shan Liu, and Ge Li. Exploiting the value of the center-dark channel prior for salient object detection. ACM Transactions on Intelligent Systems and Technology (TIST), 10(3):32:1– 32:??, May 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/ft_gateway.
cfm?id=3319368.

Zhang:2023:AAN

[ZZL^{+}23] Yihao Zhang, Chu Zhao, Weiwen Liao, Wei Zhou, and Meng Yuan. Asymmetrical attention networks fused autoencoder for debiased recommendation. ACM Transactions on Intelligent Systems and Technology (TIST), 14(6):100:1–100:??, December 2023. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3596498.

Zhao:2021:MSF

[ZZS^{+}21] Jiaqi Zhao, Yong Zhou, Boyu Shi, Jingsong Yang, Di Zhang, and Rui Yao. Multi-stage fusion and multi-source attention network for multi-modal remote sensing image segmentation. ACM Transactions on Intelligent Systems and Technology (TIST), 12(6):82:1–82:20, December 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3484440.

Zhao:2011:WDW

[ZZZ^{+}11] Shiwan Zhao, Michelle X.

Zhou, Xiatian Zhang, Quan Yuan, Wentao Zheng, and Rongyao Fu. Who is doing what and when: Social map-based recommendation for content-centric social Web sites. ACM Transactions on Intelligent Systems and Technology (TIST), 3(1):5:1–5:??, October 2011. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2019:DDF

[ZZZ^{+}19] Ya-Lin Zhang, Jun Zhou, Wenhao Zheng, Ji Feng, Longfei Li, Ziqi Liu, Ming Li, Zhiqiang Zhang, Chaochao Chen, Xiaolong Li, Yuan (Alan) Qi, and Zhi-Hua Zhou. Distributed deep forest and its application to automatic detection of cash-out fraud. ACM Transactions on Intelligent Systems and Technology (TIST), 10(5):55:1– 55:??, October 2019. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Zhang:2020:WUA

[ZZZ20a] Lei Zhang, Yixiang Zhang, and Xiaolong Zheng. WiSign: Ubiquitous American Sign Language recognition using commercial Wi-Fi devices. ACM Transactions on Intelligent Systems and Technology (TIST), 11(3): 31:1–31:24, May 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3377553.

Zhao:2020:UDR

[ZZZ^{+}20b] Yawei Zhao, Qian Zhao, Xingxing Zhang, En Zhu, Xinwang Liu, and Jianping Yin. Understand dynamic regret with switching cost for online decision making. ACM Transactions on Intelligent Systems and Technology (TIST), 11(3):28:1– 28:21, May 2020. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3375788.

Zhang:2022:IAS

[ZZZ^{+}22] Yifan Zhang, Jinghuai Zhang, Jindi Zhang, Jianping Wang, Kejie Lu, and Jeff Hong. Integrating algorithmic samplingbased motion planning with learning in autonomous driving. ACM Transactions on Intelligent Systems and Technology (TIST), 13(3): 39:1–39:27, June 2022. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/doi/10.1145/3469086.